A Bibliography of Publications in ACM SIGAda Ada Letters

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

28 August 2024 Version 4.18

Title word cross-reference

# [Dew17, Duf08a, FM09a]. #1 [Duf08b, Shu93]. #100 [Bri12a]. #101 [Obr12a]. #102 [Obr12b]. #103 [Pan12a]. #104 [Kan12a]. #105 [Bri12b]. #106 [Bri12c]. #136 [Puc17]. #137 [Reb17a]. #138 [dev17a]. #139 [dev17b]. #140 [Qui17]. #141 [Dev17c]. #142 [Ano17a]. #143 [Ano17b]. #144 [Ano17c]. #145 [Reb17b]. #146 [Moy17a]. #147 [Moy17b]. #148 [Moy17c]. #149 [Moy17d]. #150 [Dew17]. #151 [Moy17e]. #2 [Duf08c, Hir94c]. #22 [DFGZ09]. #23 [Duf09a]. #24 [Duf09b]. #25 [Bri09a]. #26 [Duf09c]. #27 [Dew09a]. #28 [Dew09b]. #29 [Obr09]. #30 [Bar09b]. #31 [Dew09c]. #32 [Bar09c]. #34 [Bar09d]. #35 [Rog09b]. #36 [Bar09e]. #37 [Rog09c]. #38 [Bar09f]. #39 [Rog09d]. #40 [Bar09g]. #41 [FM09b]. #42 [Bar09h]. #43 [Bar09i]. #44 [Duf09d]. #45 [Bar09j]. #46 [Dew09d]. #47 [Bar09k]. #48 [Och09a]. #49 [Bar09l]. #5 [Hea08a]. #50 [Duf09e]. #51 [Bar09m]. #52 [Bri09b]. #54 [Bri09c]. #55 [Och09b]. #56 [Och09c]. #57 [Och09d]. #58 [Och09e]. #59 [Cha09]. #6 [Hea08b]. #61 [MC09a]. #62 [MC09b]. #63 [Dis09]. #64 [Bri09d]. #65 [Bri11a]. #66 [Bri11b]. #67 [Bri11c]. #68 [Moy11a]. #69 [Moy11b]. #7 [Gas08]. #70 [Rog11b]. #71 [KW11a, KW11b]. #73 [KW11c, KW11d, KW11e, KW11f]. #77 [Bri11d]. #78 [Bri11e]. #79 [Bri11f]. #8 [Hea08c]. #80 [Cha11]. #81 [Rog11c]. #82 [Moy11c]. #83 [Moy11d]. #84 [Qui11a]. #85 [Qui11b]. #86 [Och11]. #87 [Qui11c].

#88 [Och12a]. #89 [Pan12b]. #9 [Hea08d]. #90 [Qui12]. #91 [Och12b]. #92 [Pan12c]. #93 [Rog12a]. #94 [Pan12d]. #95 [Och12c]. #96 [Pan12e]. #97 [Bri12d]. #98 [Rog12b]. #99 [Bri12e].

+ [Nyb07]. 10^{th} [Ano00i]. 2 [Reb17a]. 3 [Reb17b]. 8 [SGW90a]. = [Nyb07]. ^{sm} [Sil98]. st [Ano99a]. ^{th} [Ano02d]. μ [PV98].

-1- [Gor83]. -bit [SGW90a]. -or- [Woo99].

.NET [Bro09, CSH03, HCW04].

/DDS [DRSK23]. /design [San12]. /Java [Och09d, Och09e, Och09b]. /multi [Taf13b]. /multi-threaded [Taf13b].

05 [RC10a].

1 [Moy17a, dev17a]. 1.0 [Fag00b]. 11 [Ano02d, SHLR80]. 11/780 [SHLR80]. 11th [Cra22]. 12C [Che09]. 130J [Con03b]. 14th [MR10]. 16 [McC06a]. 1750A [RM88, Roa88, Roa89]. 178C [Bro11]. 178C/ED [Che09]. 1980 [ACM80]. 1987 [Bar87, Off88c]. 1988 [Puk88]. 1st [Ano91a].

2 [Car06b, LVTL22, Moy17b, dev17b]. 2.0 [Wis99]. 2000 [Ano00k, Ano00v]. 2001 [Ano00j, Ano01b, Ano02b]. 2002 [Ano02a, Ano02c, Ano02e]. 2005 [Bar07b, BW07b, BW07a, Car06a, Car06b, CH06, CR07, CR05, Dew06, Duf08b, Duf08c, Duf08a, Ler03, McC06a, MPV10, MWM10, MS04, MSK05, MC09b, Moo10, Och09a, PdlPH^{+}07, RM07, RT09, Taf06, UPRZ07, WB07a, WB07b, WMAB10, WB10a, Whi10, ZBW07]. 2006 [Ano06f]. 2012 [BT14, Car17, EGC13, HG14, LWB13, Moy17a, Moy17b, Moy17c, Rui13, SC13, Sch10b, SP12, Tro12, WGC17]. 2014 [CAC^{+}13, EH13, HG14]. 2018 [MH20]. 2020 [Bur13b]. 2022 [Cra22, MC22, Mos22, SHT^{+}23, Taf22b]. 202X [Taf21, Cou21]. 2167 [Buc87, FG86, GG87, Ros86a]. 2167A [Ros86b]. 248C/ED [Che09]. 278A/ED109A [Che09].

3 [Moy17c]. 3Cs [LWF91].

4th [Rog09e].

5G [ICS22]. 5th [Ano92a].

6 [Ano99l, Cle86]. 60 [HvKPT87]. 653 [GZdlP15, Tok03]. 6th [Ano93b, BW93b, Ano93k].

780 [SHLR80]. 7th [Ano92b].

802.1AS [TPG21]. ’82 [CF82]. 83

[BT14, Dew09d].

’91 [ACM91b, ACM91a]. ’91/Summer [ACM91b]. ’92 [Ano92f, Ano92n, Ano92o, Ano92m]. ’93 [Ano93n, Ano93o, Ano93p, Ano93l]. ’94 [Ano93m, Ano94h, Gau95, bY94]. 94C [Che09]. ’95 [Ano95m, AR95, And04, Bal95b, Bal97, BHD98, Bar01, BBB98, Bot99b, Bro97, Bro98b, BDT99, BM97, CSH03, Che97, Col99a, CR05, Cra95, DCBM97, Dew09d, DPB^{+}97, Dor99, GD00, Gau96, GSX99, Gib00, Hai00, HCBM98a, HCBM98b, HDHH98, KF98, Kie97, KR01b, Lit97, LKN97, MP98, MY98, Moo97, Mor96a, Mor96b, PV98, PV99a, PS06, Pow97, PDN97, Pri96, Pri01, RW99, RDS98, RLPD98, Ros96, SS97, Taf01a, Taf01c, TNGC05, UKDH97, VGD^{+}97, WWB99, WBP97, WJS^{+}02, Wel03, Whe95, Whi97, Wol97, Wol99, Wol01, Yu98, dB97a, dB97b, dB99]. 95/NT [BBB98]. ’98 [STF98, Lei99b]. ’99 [Ano99i, Ano99j, Ano00w]. 9X [AV93, Bak91c, Bar93, BWD90, Bur90, BE91, BD92, BW92, BW94, Car92, Els90a, GHVVW94, Hir94a, Hir94b, Kam91, Loc91, Moo93, Plo92, Sei91, SC92, VE92, Web93, Wel91, Wre92, Ano93d, Bal95a, Bal94, Bar95, BCF94, Dob90, Els91, LMV93, Ros95, Rym94, Bar14, Räi94].

= [Gon91b, Goo85, Bra99].

AADL [ALB^{+}14, BM23, Buz16, DPP^{+}09, DRSK23, Fei14, FD16, GSP^{+}11, Glu09, HSB^{+}22, HG14, LHFD13, PF20, RD23, SB23, SLNM05, SP07, SHT^{+}23, XWZ^{+}23]. Abnormal [Pap89]. aboard [Ros96]. Abort [BQ90, GL89]. Abstract [BYY86, Car91, CdN16, CBW^{+}21, GES89, Leb82, SHR82, Wei90b, Joh93, Sel99]. Abstraction [Bar00, Coh85, CG87a, HCBM98b, LKH16, Moo18, Yeh82, CG87b]. Abstractions [Ano00w, BWK^{+}01]. Abuse [Mos22]. academic [Car01]. Academy [Gri98, SCFG04]. ACATS [EK11, EK12, Smi04]. Acceleration [JARKS22]. accelerator [MMP13a]. Acceptance [Rog85]. Access [Bel82, Gre90, Gan04]. Access-Before-Elaboration [Bel82]. Accessibility [Bar95, Duf09d, FM09a, FM09b]. Accessing [BW02, GZdlP18]. Account [Bak93a]. accurate [Tan91b]. ACEC [Boe90, Com90, Ano90a, Ano90b]. achieve [And05]. achieved [WMAB10]. Achieving [Aus22, LV23]. Ackermann [Wic86]. ACM [ACM80, Ano93a, Gri95, Har94c, STF98]. ACM-SIGPLAN [ACM80]. ACM/SIGAda [Gri95]. ACPS [BH90, CCC21]. Acquisition [CA89]. acronym [Sha93]. across [VMNM85]. Act [Car96]. action [Sei14]. Actions [BW89, Nae05]. active [CM94]. Activities [Ano92c, Ano92d, Ano93c, Ano94b, Ano94a, Joh94, Vla93, Vla94, Weg82, Whi95]. ADA [Ano88b, ACM80, ACM82, ACM91b, Ano90c, Ano90d, Ano91c, Ano92g, Ano92h, Ano92i, Ano93c, Ano93a, Ano93b, Ano93h, Ano93k, Ano97, Ano00i, Ano02d, Bar87, Con97b, Con97c, Con97d, Cra22, Gro07, Lei02, MR10, Moo85, Mor96a, Mor96b, Obe94, SPS88, Sof88, Wes97a, Wes97b, BBB98, LRS09, SGW90a, ACM87a, ACM91a, ACM87b, ACM89, Abb96, ACP11a, ACP11b, AR95, Age85, AB98, AGG^{+}80, ARPT18, ABGH13, AH01, AID05, AP11, AKM^{+}91, Ad93, AdlPT97, Als83, AS87, And88, And04, And05, Ano87, Ano88a, Ano89b, Ano89a, Ano89c, Ano90a, Ano90b, Ano91b, Ano91a, Ano92c, Ano92d, Ano92j, Ano92m, Ano93c, Ano93a, Ano93d, Ano93f, Ano93g, Ano93l, Ano93m, Ano94a, Ano94c, Ano94d, Ano94h, Ano99b, Ano99i, Ano00a, Ano00b, Ano00j, Ano00l, Ano00m, Ano02a]. Ada [Ano02b, Ano06d, Ano06b, Ano06c, Ano06a, Ano06e, Ano10b, AV93, AD82, AP84, Ard87, AA88, AA89, AC85, AB87, ACWB89, AG88, AdB90, AW01, Bac82, Bac84, Bag98, Bak86, Bak87a, Bak87b, Bak88, Bak90a, Bak90c, Bak90b, Bak91b, Bak91c, Bak93b, BOM97, Bal95a, Bal94, Bal95b, Bal97, BTVC99, BST90, BMNS85, Bar85b, BM85, BT88a, BT88b, BCS89, BHD98, Bar01, Bar09a, Bar88, Bar93, Bar95, Bar07a, Bar07b, BT14, Bar14, BP13, BMW94, BGK^{+}82, BCG^{+}84, BFG85, BD91, BBB97, Bec83, Bei92, Bei97, Bei84, Bel80, Bel82, BCHR12, BBH80, BA82, BA90a, Ben84, BKW82, Ber83, Ber84, BB85, Ber15, Ber05, BD99, BDD^{+}82, BHN20, Bis80, Bis86, Bis91, BCF94, Boe90, Bon84, Boo11, BKWS88, BG90, Bos13, BCD83, BC95, Bot99a, Bot99b, Bot00a, Bot00b, Boy87]. Ada [Boy89, BdlPZ10, BDF^{+}85, Bra85, Bra94, Bra98, Bra99, Bra83a, Bra83b, Bri92a, Bri94, Bri12b, Bri12c, Bri12d, Bri12e, Bri12a, Bro80, Bro82, Bro83, Bro88, Bro96, Bro97, Bro98a, Bro98b, BD01, BA07, BHL^{+}93, Bro04, BDT99, Bru17, BM97, Bru82, Bry90a, Bry90b, Bry88, Buc87, BF99, BK85, Buh85, BKW85, BKC91, BW90a, BW90b, Bun85, BN87, BL86, Bur85b, Bur87b, BW87, BW89, BWD90, Bur90, BW90c, BW90d, BE91, BD92, BW92, BW93b, BW94, BW99, BWK^{+}01, BR01, BB02, BWV03, BW03, BDV04, BW07b, BW07a, BTB^{+}10, BW13a, Bur13b, BWM13, BW16b, BDS81, Bux85a, BH90, Cam92, CVW03, Car00, Car01, CS02, CSH03, Car06a, Car06b, CH06, CB07, Car11, CA89, Car17, Car22a, Car88a, Car88b, Car89a, Car89b, Car90, Car92, Car94, CS94, Car96, Car22b]. Ada [CN96, CS91, Cel97, Cha82, CH97, CLY98, CBW94, CF82, Cha09, CG82, CHHB90a, CHHB90b, CAU88, CU89, Che92, Che97, CR07, Che91b, Chr87a, Chr87b, CSSW09, CSSW10, CM89, CM90a, CM90d, CWW80, Cla97, Cla87b, Cla87c, Cle82, Cle86, Coh81, Coh82, Coh88, Col99a, Col95a, CR97, CG88, Col89, Col87, CR05, Com90, Con03a, Con97b, Con03b, CG87a, Cor83, CSL^{+}87, CS87, Cou21, Cra82a, Cra82b, Cra95, CDM87, CEG23, Cro95, DF84, DGCR^{+}84, DS87, Dav82, DeL88a, DeL88b, DeW86, DCBM97, Deb83, DFS^{+}80, Dew84, Dew01, Dew06, DFGZ09, Dew09d, DZM87, DCC85, DPB^{+}97, DoD87b, Dob90, DRF97, Dob83, Dom87, DD87, DGLM85, Dor99, Dri91c, Dri91a, Dri91b, Dri91d, Dri91e, Duf08b, Duf08c, Duf08a, Dul03, DH80, DH82, Dun98, Ear92, Ehr94, EGC13, Ell83]. Ada [Elr88, Elr89, Els90c, Els90a, Els91, EKPPR04, FHN83, Fag00a, Fag00b, FME01, Fai80, Fal91, Fal82, FGN85, FG82, Fan84, Far82, Fel09, Fel11, FCS83, FMN80, FG86, Fir87a, Fir88, Fir90, Fir87b, Fis84a, Fle86, Fli98, FSS87, FNS^{+}85, FA82, Fra87b, FMG90, Fre86b, Fri98a, Fri98b, Fri83, Fro87, Fro15, Fuj87, FOFY87, Fus91, Gal22, Gal20, GH99, GH01, Gar83, GB87, GGP^{+}90, GST^{+}97, GD00, Gas08, GSP^{+}11, Gau95, Gau96, GSX99, GES89, GHL82, Gib00, Gic90, Gid96, GB94, Gil99a, Gil99b, Gil84, GCM90, GL89, GHVVW94, GBCGDBC97, Gon88, Gon91a, GDAG97, Goo80, Goo85, GS88, GW80, Gra83, GG87, GMO92, Gre16, Gre18, Gri98, Gro86, GR80, GS85, GDHM02, GG99, HPT81, Hag91, Hai00, Hal83, HR07, HD85, Har85]. Ada [HS87, Har88, HMRF97, Har99a, Har87, HB88, HL86, Har82, Har94a, Har94c, Har97, Hek83, HL85a, HL85b, HCBM98a, HCBM98b, HMC88, HHR^{+}86, Hil22, Hil82, Hir92, Hir94a, Hir94b, HLRS80, Hod91a, Hod91b, HNS98, Hof86, HDHH98, Hos89, Hou83, HM03, HM91, HW88a, Huf82, HHBC90, HG14, HvKPT87, HCW04, Hun88, HSW87, HW88b, ISO91a, ISO91b, IMM85, Jam98b, Jam99, Jan88, JF98a, JF98b, JEKC89, Jha90, JA82, KPPÉR06, KF98, Kam83, KGW^{+}85, KJEC87, Kam91, Kam98, Kan12b, KB87, KPR93, Ker99, Ker82, Ker86, Ker88a, Ker89, Ker90a, Ker90b, Ker92a, Ker92b, Ker93a, Ker93b, Ker94a, Ker94b, Ker95, Ker96a, Ker96b, Ker97, Ker98, Khr95, Kie97, KR01b, KB97a, KMS82, KUP^{+}83, KBT84, Kle06, Kle21, Klu87, KU84, Kni87, KR88, Kni90, Kni09, KS84, KM98]. Ada [KT87, KB83, KBL80, KVT88a, KVT88b, Kru90, KETT96, KP86b, KP86a, Lad89, Lah82, LMP90, LHBK87, Lap04, LSH98, Lat09, Lat91, Lav95, Law97, LP85, Lea87a, Lea04, Lea87b, Led95b, LN91, LCN91, LMA94, Lef87, Lei96, LL98, Lei99a, Lei99b, Lei00, LLL03, Lei06, Leo85, Ler03, Lev88, Lev89, Lev97a, Lev05a, Lev09a, Lev82a, Lev82b, Li82, LXY98, LYB^{+}10, LW01, LW02, LWB13, Lin82, Lin83, Lit97, LM83a, LM83b, LBO84, Lla92, LV87, LVM90, Loc91, LMV93, LKN97, Lof93, Lom83, Lop99, LT99, LB80, Low99a, LD87, LP80, LNR87, LA99, MK87, Mac80, Mac86, Mac84, Mac96, MMSN09, Mah11, Mah12a, Mal88, MF04, Mar99, Mar05, ML91, MM21, Mar21, Mar86, MK83, Mat87a, Mat96, Mat87b, MB91, Mat91, MP85, Mau07, MR87a, Maz89b]. Ada [McC87a, McC99, McC00, McC07, McC09, McC10, McC87b, McC90a, McC90b, MR83, McD88a, McD88b, McD89, McE03, MR87b, Mea87, Med91, MP84, MG87, Men87, Men09, MPV10, MKP91a, MK91, MKP91b, Mic07, MWM10, Mid87, ML95a, ML95b, MP98, MS04, MSK05, MC09b, Mog91, Mol83, MY98, Moo97, Moo91, MP91, Moo93, Moo96, Moo98, Moo10, MMP13b, Mor87, Mos20, Mos22, Moy17a, Moy17b, Moy17c, Mud87, Mun96, MH97, MF91, Mur87, Mur90, MH98, MH09, MS87, MP89, NKN93, NMT92, NM92, NIM07, Nie86, NWW82, NW83, NW^{+}84, Not80, O’L07, Off88a, Obr09, Och09d, Och09e, Och09c, Och09a, Och09b, Och09f, Och11, Off87, OW82, Pag82, PV13, PZ97a, PZ97b, PBB^{+}88, PMJPA01, PG94, Pau87, Pau93, Paz90, Per88, PWDD80, PDG83, PB98, Pet10, PS84]. Ada [Pie85, Pie87, Pie90, PV98, PV99b, PV99a, PMM13a, PMMT15, PRQ21, Pio86, PS06, Plo92, Plo98, Plo01, PD82, Pot04, PVV85, PR90, Pow97, PDN97, Pri96, Pri01, Pri82, Puk93, Puk94, PdlPH^{+}07, Pul95, PG91, Pyl84, Qui90c, Qui90d, Rac88, Räi94, RC10a, RW99, RLC01, RM07, RC10b, Reb17a, Reb17b, Ree85, Ree86, Reh87, Rei87, RDS98, RLPD98, RS91, RB85, Rie94, Rie98, RH01, RH02, RH03, RTH15, RT21, Riv17, RM88, Roa88, Roa89, Rog85, Rog88, Rog97, Rog09a, Rog21, Rom01, Rom86, Rom88, Rom05, Ros87b, Ros87c, Ros95, Ros96, Ros09, RT09, Ros11a, Ros11b, RMT11, Ros22, RLHS80, Ros87d, RR90, Ros86a, Ros86c, RTM82, Rou85, Rud83, Rui13, Ryb94, Rym94, Sac89, SGS92, SRC13a, SRC13b, SC13, SRC15, SWR82, San03a, San89, San03b, SW87]. Ada [Sch87a, SSJ85, Sch09, Sch10a, SF82, SS85, Sch10b, SP12, SC87, Seb87, SS91, Sei91, Sei92, SC92, SB99, SHLR80, SB80, SHR82, SAH01, Sho87, Shu87, SN88a, Sil98, Sim82, Sin07, Sma09, Smi84, SCD^{+}85, Sny91, Spi00, Spu86, Squ91a, Squ91b, Squ91c, Sri06a, Sri06b, Sri06d, Sri06c, SSFO86, Sta83, SGJP89, SM92, Ste80, SC01, SYW85, SS97, Sum87, SN88b, SC04a, SCFG04, SC04b, Swa07a, Swa07b, Swa09a, Swa10, Syi95, TTRH85, Taf82, Taf01a, Taf01c, Taf06, Taf13a, TMPM14, TMPM16, Taf21, TBD22, Taf22b, Tai86, Tan91a, Tan91b, TP09, Ter87, TR87, TCRW88, Tha82, The90, Tic82, TG09, TGH10, TGH13, Tin90, Tis83, Toa96, Tv88, TNGC05, Tok15, Tom97, Ton99, Too91, Tro06, Tro12, Trü95, Tuc97, UKDH97, UPRZ07, Van86, Var01b, VW13, VR16]. Ada [Vas91, Vau98, Ver22, VE92, Ves89, VGD^{+}97, Vla93, Vla94, Vok92, VMNM85, Vol87, Vol90, Wai98, WBS97, WWB99, Wal85b, Wal87, Wal91, WFF^{+}87, Wan90, Wan99, WGC17, WA02, WA07, WD93, Wat87, Wau83, Wea10, Web93, Weg82, Wei89, Wel85, WKT84, Wel91, WBP97, WJS^{+}02, Wel03, WT03, WB07a, WB07b, WMAB10, WB10a, WBCS13, WCB16, WGA90b, Wes97a, Wes97b, WQ83, Whe84, Whe86, Whe87, Whe95, Whe97, Whi81, Whi97, WW01, Whi10, Whi82, Wic82, Wic86, Wic98, Wil87, Win84, Win90, Win91, Wol97, Wol99, Wol01, WV01, Wol84, Won90, WL98, Won99, WMM10, Woo88a, Woo88b, WT88, WT89, Woo99, Woo87, WV98, Wre92, WB89, XZ02, XRL^{+}88, Yav85, Yem82, YG80, Yu98, bY93, bY94, ZEdlP13, ZW83, ZBW07, de 87, dB97a, dB97b, dB99, vdL84]. Ada [vdL85, vHLKBO85, Rog11d]. Ada-05 [RC10a]. Ada-2005 [CR07]. Ada-94 [Gau95, bY94]. Ada-95 [Gau96]. Ada-Appropriate [BST90]. Ada-Based [SPS88, Sof88, Che91b, Gal22, Abb96]. Ada-COBOL [Bro96]. Ada-embedded [DD87]. Ada-Europe [Ano99i, NWW82, NW83, NW^{+}84]. Ada-In-Ada [Taf82]. Ada-like [Khr95]. Ada-LINPACK [PG91]. Ada-LISP [DS87]. Ada-related [FG86]. Ada/Linux [SRC15]. Ada/Mindstorms [Fag00b, FME01]. Ada/Tcl [Wes97a, Wes97b]. Ada05 [Hea08b].

Ada2005 [FM09b]. Ada83 [Bak91a, Bak93c, Van94]. Ada95 [Gar09, OB97, Bre97, Due97, Faß01, FM09a, Gan01, Hea04, Hea08b, KFS97, KK03, Lev98a, Lew02, MCS97, Mun96, NDP97, NDM98, NDP99, NDP00, Nyb05, PC05, Rym98, Wis99, Wor97, XCZ04]. Ada95-programmed [Faß01]. Ada95/C [Gar09]. Ada95/DSA [Gan01]. Ada’96 [Rob97]. Ada’97 [ACM97]. Ada9X [GHVVW93, Van94]. Adabase [Tic82]. AdaGIDE [CC98]. AdaHorn [BHN20]. Adaing [PV99b]. AdaJUG [MFD85]. AdaPT [GHVVW93, GHVVW94]. adapted [CXY01]. Adapting [EK12, GGP^{+}90, TGH13, Bis88]. Ada(R) [Fri87]. AdaSlicer [SC04a]. AdaTEC [ACM82, MFD85]. AdaTEC/AdaJUG [MFD85]. AdaWebPack [Rez22]. Add [Gre99a]. Adding [Cla87c, Hal83, Sac89, SRC13a]. Additional [Ano06d, Cla87b, Whi10]. Address [Bux85b, Boe99, Bux85a, Car01, Dew01, McC99, Sel99, Taf01b]. Addressing [RDS98]. ADEPT [GSTV97, SHT^{+}23]. Adjustable [Lea87b]. ADL [Ker88b]. Administration [O’L07]. Administrators [Hos89]. Adoption [CCC21, Mog91]. advanced [LP06]. Advancing [BCF94]. Aegis [Nil12a]. aerial [SG06]. Affinities [SRC15]. affordable [Dav05]. after [Klu87]. Agent [Hai00]. agents [LS98]. aggregates [Duf08b, Duf08c]. AI [BCB^{+}22, SK22]. AI-augmented [BCB^{+}22]. AI-oriented [SK22]. aid [EF01]. AIDA [Maz89a]. AIDOaRt [BCB^{+}22]. AIE [Bra82]. AIM [BF86, Fre86a]. Air [Aus22, Gri98, Hum22, ACW04, Kle06, OWSB08]. Airborne [LT99]. aircraft [Con03b, Swa09a]. AIs [BV03, GHV03]. AJIS [Och09c]. AJPO [Coh81]. Alan [Rog97, Rog09e]. Alf [Sei14]. Algebra [Klu87, DCC85]. Algebraic [LM83a, LM83b, BH14]. Algol [HvKPT87]. Algorithm [Cra98, JF98b, RLPD98, Woo88a, Woo88b, WT89, CXY01, JF98a, NS03, SN04, WT88]. Algorithms [Har87, JARKS22, MS87, SS20, Ste80, Yem82, Bar09a, Hea08d, SGS92]. Alire [MC22]. ALISA [BCMC23]. alive [Mah11, Mah12a]. Allocated [Lef87]. allocating [WB07a]. Allocation [KPP97, WKT84]. allowed [Fos20]. ally [Ano17a]. alone [Pow90]. Alternative [LCN91, AV93, VE92]. Always [Law97]. America [Bar14]. Analogies [HL86]. analysable [BW94]. Analysis [And20, Ano90b, BH90, CCC21, Con97a, FHN83, FD16, FMG90, Gen91, GP93, Had90, HS87, KB87, KBT84, LSH98, LKH16, LKSL19, Mar21, MGF16, MP98, PR98, PG91, RS91, RDP97, Rog88, RG90, Shu91, Wag20, Wal91, WHNB91, XWZ^{+}23, ACP11a, ACP11b, AID05, AD03, BF86, Bla07, CFH^{+}13, CBW94, CH04, CBB^{+}97, Col99b, Com90, Coo97, Cro95, Dew07b, DV01, Ehr94, Fir91a, Fir91b, GSP^{+}11, Glu09, GDHM02, JR10, KK03, KNB08, Lat09, LSRM12, Och12c, Saï08, Shu93, SLNM05, SP07, SN04, SU91, Ven08, WV02, Wha13, WW01, ZdlP02]. analyst [Too91]. analytical [MCS97]. Analyzer [SB80]. analyzers [Bar08]. Analyzing [Har87]. anatomy [San03b]. Android [Fos20]. Andy [Rog97, Rog09e]. ANIARA [JBT^{+}22]. Animation [Cra98, JF98b, JF98a]. ANNA [KBL80, KB83, SRN85]. Annex [Ano10a, Bal97, BW15, ALB^{+}14, AH01, AW01, Ber05, DPB^{+}97, GH01, LHFD13, PT99, Qui11a, Qui11b, Qui11c, Qui12, RH01, Moo97, TBA98, dB97a]. Annex-E [Moo97]. Annotating [KBL80]. Announcement [Ano01b, Ano10b]. Announcements [Ano00c, Ano00d, Ano06e, Ano06f]. Annoying [Far82]. Annual [ACM91b, Ano92a, Ano92k, Ano93a, Ano93i, STF98]. anomaly [RA91]. Anonymous [WGA90b, WGA90a]. ANSI [The90, Fis84a, Moo91, Smi84]. Answer [GA90, Law97]. Any [Gre90]. Anything [Hil22]. Anyway [Fir88]. Aonix [BE02]. APE [HNS98, San89]. APEX [Wai21]. API [KQT^{+}21, Men09]. APIs [BH14, Fli98]. Apparently [Hof86]. Apples [Fir88]. Applets [KFS97]. Applicability [LSRM12, Roa89, RM88, Roa88]. Application [BKW85, Hai00, Kie97, RDP97, RH02, RH03, Wai98, ACW04, BW99, BV13, Col99a, Dav05, HEUV99, LG88, Nyb05, PL07, Ros04, Saï08, Wis99]. Application-defined [RH02, RH03]. Applications [All87, BK22, Che97, Chr87a, Cor83, Cra82a, DH80, DH82, GCM90, HSW87, JBT^{+}22, MR87b, Mid87, NPT97, PS84, SB23, Wei90a, Abb96, BMW94, BWM13, Chr87b, DPB^{+}97, HMC88, McC10, MS11, MKK99, Mos06, PV99a, PV02, Puk94, Rog11a, Rog11d, VC01, Vas91, ZHP06]. Applied [Har22]. Applying [BF99, GP93, Pri96, Sil98]. Approach [BFG85, Col87, DGBMCG97, FOSC23, Fir87b, GCM90, GA90, Gra83, Har82, Hir94c, KR88, KB83, LM83a, LM83b, SC87, VGGS20, Wal91, Woo88a, Woo88b, YQZ^{+}23, HM03, Kni09, Lit97, San12, SS91, Ven08, Wan99, WRL13, Yav85]. Approaches [AC85, Gib00, Whe19, MCS97]. Appropriate [BST90, Hof86]. Approved [Ano89b, Ano99d, KW91]. Approximation [Pag82]. April [Puk88]. APSE [Hou83, Boy86, Bux85b, DGCR^{+}84, Dru82, Fri87, ML86, MB91]. arch [Bar98]. archetypes [Pan12c, Pan12d, Pan12e, Pan12a, PV13]. Architectural [Sel99, Gan03]. Architecture [CBB^{+}97, FG82, Har22, ILMV83, Lah82, Pro20, Sim82, Bar09f, BS13, Edg01, GBC^{+}14, HEUV99, KS01, LRS09, Mor95a, NBZ^{+}20, PV98, SAH01, Spi00, Swa07a, Swa07b, Swa09b, SB11, SB12, Wha13]. architecture-based [Edg01]. Architecture-Level [Pro20].

Architectures

[BM23, Red85, Tok16, Dob00, WMAB10]. Arcturus [Sta83]. Arduino [RT21]. Areas [BW90c, BW90a]. ARG [Bar98]. arguing [Syi95]. Arguments [Gór20]. Aria [GSTV97]. Aria-Java [GSTV97]. ARINC [GZdlP15, Tok03]. ARINC-653 [GZdlP15]. ARINC653 [DPP^{+}09]. Arising [Rob92]. Arithmetic [Fis84b, Fro15, Lea87b]. Arlington [ACM82]. array [Rog09d]. ARTEWG [Ano87, KGW^{+}85, Ano92c, Ano92d, Ano94d, Kam95]. Artificial [Ano94b, Ano94e, Ano95b, Ano95c, BSPK22, JBT^{+}22, SCC22, Wol85, Joh94, Lav95]. ASEET [McD88a, McD88b, McD89]. ASIS [Col95a, CR97, RC01, Vla94, Ano99d, Ano99c, Ano99l, Ano00w, AN05, BRC98, CBB^{+}97, Col99b, Coo97, Dru99, FRS97, Hov00, LSP01, PR98, RT09, Ros21, RSZ96, Vla93, Wis99]. ASIS-Based [PR98, Coo97]. ASISRG [Col95b, Rob97]. ASIStint [FRS97]. ASISWG [Vla94, Ano94a, Col95b, Rob97, Vla93]. ASISWG/ASISRG [Col95b, Rob97]. asked [Col95a, CR97, Mat96]. aspect [PC05]. AspectAda [PC05]. Aspects [LWF91]. Asserting [Moy17d]. Assessing [HCT^{+}98, HG14]. Assessment [Ano93f, BDT99, BN87, Kni90, OWSB08, Rei87, Ros21, Ano89a, Bra99, Bro07]. assessments [Ton99]. Assignment [Rob92, Mor95a]. assist [Low99a]. Associated [BN87]. Assurance [Gór20, Har22, Mol83, SCC22, Fis12, GBC^{+}14, Jar07, Jen09, Lan10, McE03]. AST [LT99]. Asynchronism [BE91, Els90a]. Asynchronous [BHR02, BWD90, CHHB90a, CHHB90b, Els90c, Pow90, Qui90b, Qui90a, Qui90d, Tv88, de 88, AV93, HHBC90]. Atlanta [McC06a]. ATMAda [ML86]. ATmega16 [RC10a]. Atom [Lev82a, Lev82b]. Atomic [BW89, PVF01, SRC13b]. Atool [FNS^{+}85]. Attempting [Mar19]. Attitudes [Gil99a, Gil99b, Rog85]. Attribute [SS89, BW03, Duf09c]. attribute-based [BW03]. attributes [SRC13b, SC13, Win91]. Augmented [Tro20, BCB^{+}22, Wel03]. Auto [Kim21, Zhu90]. Auto-Generated [Kim21]. Automated [FD16, Puk93, BCHR12, BB85, Lit97]. Automatic [Ala13, Car00, Car06a, CEG23, KB87, LZL03, LKH16, ML91, PBB^{+}88, SN94, TRT16, Wal85b, CS02, OS12, LRS09]. Automatically [Nyb10a]. Automating [Rad94, San01b, WG20]. Automation [Buc87, JBT^{+}22, Mye85, Bre97, Coo97]. Automotive [BMGS20, SSB^{+}20]. Autonomous [CCC23, LC22, LV23, ZDM22]. Autopilot [LV23]. availability [Aus22]. available [Ker98]. Aviation [O’L07]. Avionics [SPS88, Sof88, Tok16, Bar08, BCF94, Bro11, CS91, LVM90, Rom05, BRF92]. Avoid [Men88]. avoiding [JR10]. AWA [XRL^{+}88]. Awarded [McC06a]. Awards [Gri95, Har99b, Har00, Har01, McC06a, MH20]. Aware [ZDM22]. awareness [SG06]. aWaypoint [Hum22]. AWING [FC91]. AWS [Obr09].

back [Car11, Cha07a]. Bagatelles [Far82]. Bakar [BCHR12]. Ballistics [Rud83, Tem84]. bare [UPRZ07]. Barriers [BW16a, Led95a]. Base [Dru99, MP91]. Based [Ano92b, AL00, BCB^{+}22, CdN16, Che91b, CG88, Cri01, DeL88a, FOSC23, Gal22, GCM90, Gra83, JF98b, Kru90, Leb82, LNR87, PR98, Riv17, SPS88, Sof88, SWR82, SC87, SB23, TRT16, VRH21, Wal91, Wil87, Abb96, BW03, Bur13a, CM94, Coo97, DeL88b, Dob00, Edg01, Fei14, Gan03, Gór20, Hir94a, Hir94b, KR01b, Kni09, LW07, LYB^{+}10, LW02, MMSN09, Moy11c, Moy11d, PV98, PdlPH^{+}07, RTH15, SAH01, Sny91, Spi00, VGGS20, WA07, Wha13, XZ02, Hea08a, JF98a, PB98]. bases [LSP01]. Basic [Bri94, Hum22, KS84, Reh87, Hod91a, Hod91b, Och11]. Basis [MP84, Mor87, NDP97]. BATCES [Hir94c, Shu93]. Be [Bar85b, Ker82, BH14, Bak93a, Bos12, CS87, Cro14, FBL^{+}10, Lad89, Moo96, Mor95a, Taf06, WMAB10]. beauty [Gas08]. Been [Ano99d]. Before [Bel82, GG16, Bar14, Taf01b]. beginner [Lau07]. beginning [GG16]. Begins [GG16]. Behavior [BKC91, ALB^{+}14, Goo13]. Behaviour [Ber15]. Behind [Lev82b]. being [Har94c]. bench [Wai98]. Benchmark [HF84, PC90, PG91, Wei89, Wei90a, CM90d]. Benchmarking [CC18, UKDH97]. Benchmarks [AW89, CM90f, Ves90a, AW88, SC06, Ves90b]. Beneficial [Rac89, Rac88]. Benefits [GD00]. best [Bar07a, Bar07b]. Better [Bak87a, Har97, BH14, Wel03]. Between [AG88, Dew09d, KETT96, Lei02, Mar05, Pot04]. Beyond [Buc87, LVTL22, LSP01, RM07, WB07a, Kle06, Moo10, Mor95b]. Bibliography [Fir90]. binary [Saï08]. Binding [BM97, Bry88, Moo91, Wes97a, Wes97b]. Bindings [McC90a, McC90b, Puk88, AN05, Bar01, Cha09]. Biography [Spu86]. Birds [CWW80, Dew07a]. Birds-of-a-feather [Dew07a]. Bit [Ano17c, MP89, SGW90a]. BlazeNet [Kam98]. Block [Win84]. Blockchain [TS20]. Blocking [GS88]. Board [Ada88, Off88a, Off88b, Off88c, Tas88, AB98, EF01, ML95a, UPRZ07, Off88a]. Boards [LL98]. Booch [SJ91]. Boogie [Lei12b]. Book [Led92, Rog97, DeW86, Rog09e, Rog11d]. Booleans [Wic93]. Boston [ACM80, ACM87a]. both [Sma09]. Bounded [Cha13, Rog09b, Rog09c]. branch [Lat09]. Breaking [Car96]. breaks [Taf01b]. bridged [LRS09]. Bridging [Qui17]. brief [Oli94]. Bringing [Mos20, Taf13a]. btypes [Moy17a, Moy17b, Moy17c]. Budgets [Gre16, RH07, Sri06a]. buffer [Rog09b, Rog09c]. Build [BT88b, Bod19, Car22a, Sal92]. builder [Boy86]. Building [Arn86, Dob00, Goo13, MVG99, MS11, PVV85, Taf91a, TRT16, TP98, UZ07, Taf91b, Rog11d]. built [Jar07, Moo97]. built-in [Jar07, Moo97]. Burns [Rog97, Rog09e]. Byron [Gor83]. Byte [Bal97, And05]. Bytes [Ano17c].

C [CHGH19, AN05, CB07, Cha09, Con03b, Cro14, Dor99, Gar09, Khr95, Kle21, LT99, Mar05, Mar21, MC09b, MC09a, NKN93, Qui12, Syi95, Toa96, Whe97]. C# [Bro09, KPPÉR06]. C-130J [Con03b]. C/C [Mar05, Mar21]. C2AADL Reverse [YQZ^{+}23]. CAD [BKW^{+}94]. CADA [BK85]. CAEDE [BKW85, WHNB91]. CAIS [CSA^{+}87, How86, Obe85, Orb85, Ree88, Rob86, Wol85]. CAIS/CASWG/SEI [Rob86]. Call [Ano92b, Ano93h, Ano93l, Ano93m, Ano93o, Ano94c, Ano94h, Ano99f, Ano02e, WGA90b]. caller [WGA90b]. calls [GH99, GG99, Och09c]. came [Car11]. Can [Cro14, AAAG21, WMAB10, PVF01]. cannot [Bos12]. Capabilities [BCB^{+}22, NPT97, Bri09b, Bri09c]. Capability [Boe90, Com90, Dob83, Goo80, Moo97, Whi10, Ano90a, Ano90b]. Capstone [BRW97]. Capture [Woo88a, Woo88b]. Case [BA82, BCMC23, CG82, FAT^{+}23, KPP97, NAT20, Rog21, SSB^{+}20, Shu87, Tra89, Var01c, CBW94, Cle86, DPB^{+}97, Fav91, Fre86b, GBC^{+}14, KPPÉR06, KB97a, LVM90, Sch91, Sum87, SCFG04, Var01a, VC01, Wad92, Wek90, Ker92a, Ker92b, Ker93a, Ker93b, Ker94a, Ker94b, Ker95, Ker96a, Ker96b, Ker97, Ker98, KM98, Mat91, PS06, Ric20]. CASWG [Rob86]. catalog [Mar19]. Catalogue [AKM^{+}91]. Catch [MRB06]. CAUWG [Ano92g, Ano92h]. cc [WMAB10]. cc-NUMA [WMAB10]. CDROM [Con97c]. Ceiling [Ano06c, CR07, GS88, LG88, MSM^{+}03, RW99, RLC01, RCWB02]. Center [Ell83, SPS88, Sof88]. Certification [WG20, BBPT12, San01b]. certified [Bar09m]. CFP [Ano06e]. Chair [RH96, Bro99, Bro00a, Bro00b, Bro00c, Bro00d, Bro01, Col01, Col02, Har94a, Har94b, McC06b]. Chairperson [Bri86, PR86, Pla86, Tex86, Bar85a, Fir86, Squ86]. Challenge [ACM87b, Ano87, Lit97]. Challenges [BSPK22, GPZdlP21, Gór20, Kle21, Ric20, Mar19]. change [SRC13a]. Changes [Bro82, BQ90, Har94a, AdlP01, BB02, RCWB02, SC06, WV02]. changing [Dew09a, Dew09b]. channel [Mah12b, Ben94]. Chapter [Ano99h, Bar09c, Bar09d, Bar09e, Bar09f, Bar09g, Bar09h, Bar09i, Bar09j, Bar09k, Bar09l, Bar09m]. Chapters [Ano95a, Ano00h, Ano00r, Ano00s]. Character [Arn86, MP89, SGW90a]. Characteristics [SSFO86, Mah13]. Characters [Ano17c, SGW90b]. Charles [Hea04]. Charrette [RLHS80]. Charter [Ano95c]. Charting [PV13]. Charts [Bec83, Bis86, BL86]. Check [Bro83]. Checking [KB83, LKH16, WQ83, BHR^{+}11, BCHR12, BW99, Cha13, KNB08, LVTL22, RR14, Ros11a, SP12]. checks [CAC^{+}13, Due97, Duf09d, EK12, FM09a, FM09b]. Cheddar [SLNM04]. child [Bal95c]. CHILL [MP84]. China [Rie94]. Chinese [Won90]. choice [Rog11a]. Choosing [Irw96]. CIFO [Pow97]. Cincinnati [LC86]. citizen [Har94c]. Class [Wol01, dB99, dB97a]. Classes [Rom00, Ros95]. Classic [NMT92, NM92]. Classic-Ada [NM92]. Classical [Dav82, SGS92]. Classification [Che90]. Classifying [MK87, Ros86c]. Classwide [Hea08d]. Clause [Men88, Rac89, Rac88, Ros87a]. Clauses [Nyb87, Coh94, Mar99]. CLAW [BM97]. client [Obr12b, Qui11a]. client/server [Qui11a]. Clock [Gre18, GP18, PC90, TPG21]. Clocks [Ano06a, WB10b, dlPZ03]. closed [Wan99]. Closures [Hos90]. Cloud [BK22, SSGH^{+}22]. cluster [AID05]. Clustering [MK87]. CMM [Con03b]. Co [BMGS20, Har22, LKH16, MP98, XWZ^{+}23]. Co-Assurance [Har22]. Co-design [MP98]. Co-Designs [LKH16]. Co-engineering [BMGS20]. Co-modelling [XWZ^{+}23]. COBOL [AB87, Bro96]. COCOON [Wel97a]. Code [AD82, Bal97, BMNS85, BBB97, Col99b, Con97a, Fir88, Fle86, HSB^{+}22, MK87, MP98, Moo18, PDV98, Riv17, RR90, SHLR80, TRT16, Tin90, Tuc97, Win90, WB89, Bar08, CBB^{+}97, Coo97, HG14, KB97b, KNB08, Log13a, Log13b, Mau07, Pan12c, Pan12d, Pan12e, Pan12a, PV13, Puk93, PdlPH^{+}07, Rad94, RA91, WW01]. Code-Level [HSB^{+}22]. coded [SGW90a]. Coding [Ros86b, Van86, Ros11a, Ros86a]. Coherent [Kim21]. Cohesion [Nie86, HD85, XCZ04]. Collection [Coh86]. Columbus [Fal91]. COM [Bot99b]. combinations [ML91]. Combined [RSC16]. Combining [Kie99, KR01a, PQR18, XWZ^{+}23, Kan12b]. Combs [Wal85a]. comm [OS12]. Command [Cra82b, DDJ98, FMS98, Gic90, SSJ85, Whe84, Wil87, dev17a, dev17b, BF99, Faß01, FC91]. commentaries [Ano89b]. Comments [Har88, Hek83, Ree88, Wek90]. Commercial [Cra82a, Gar83, Lei99b, Lei00, Woo99, Ano92g, Ano92h]. Commercializing [Lei96, Lei06]. Commercially [Ker98]. Committee [Ano92e, Ker88b, Pla86, Ano94f, Ano95e, Ano95f, Ano95g, Bar85a]. Common [MB08, ER86, Mar19]. Commonly [Mat96]. communicated [And05]. Communication [AB98, AG88, CAU88, DPB^{+}97, Els90c, GSTV97, ICS22, LC22, Ros87d, Sac89, Van90, dB99, Bar09k, Gan01, ML99, OS12, dB97a].

Communications

[CKF90, GZdlP15, KC90]. Community [Bru17, Dob01a, Mun96, McE03]. Companie [Rog85]. Comparative [CC18, JA82, MP84, Ros21, SN04]. Comparing [Bal95a, KPP97, KPPÉR06]. Comparison [Boy87, Bro97, Bro98a, Bro98b, MH98, Tok16, Ber05, Mah13, Pot04, SC01]. COMPASS [BBB^{+}23]. COMPASTA [BBB^{+}23]. Compatible [Shu91, Fir91b]. Competitiveness [ACM91b, BW91, Wil91]. Compilable [Ker82]. compilation [Bal14, Khr95]. compiled [Mau07]. Compiler [Ano90a, Ano90b, AD82, AP84, Boe90, Bra94, Bro80, EJK89, Fal91, Goo80, GW80, HMC88, Mar20, Mol83, NW83, NW^{+}84, Off87, RS91, RLHS80, SN94, Sim82, SK22, TTRH85, Taf82, TR87, WFF^{+}87, BBPT12, Cle86, Cro90, Dew07b, Fri87, Hos88, JR10, KSD12, KPR93, Kir12, MSK05, NIM07, San03b, Taf01c, ZHP06, Com90]. Compilers [ACWB89, BFG85, Fli98, ML91]. compiling [WA02]. complement [LLL03]. Complementing [TP09]. Complete [Bis86, SJ91]. completing [Mic01, Sri06d]. Completion [Pap89, Och12a, Och12b]. Complex [BC16, CBB^{+}97, Hod91a, Hod91b, Sel99, Squ91a, Squ91b, WRL13]. Complexity [DCBM97]. Compliance [Tom97]. Compliant [GG87]. component [Dav04, LW07]. component-based [LW07]. Components [AdlPT97, BT88a, BT88b, Car90, Dau87, FA82, Gib00, Gon90, Lat91, Lev92b, Lev93b, Lev93c, Lev93e, Lev94b, Lev99a, Lev00, Lev01b, Lev02a, Lev10, Lev15a, LM83a, LM83b, Rob92, Wai98, Yu97, Car92, Car04, Con97b, Fai94, Lev90, Lev92a, Lev93a, Lev93d, Lev94a, Lev94c, Lev95a, Lev95b, Lev95c, Lev95d, Lev96a, Lev96b, Lev97b, Lev97c, Lev98b, Lev98c, Lev99b, Lev02b, Lev04, Lev05d, Lev05b, Lev05c, Lev06, Lev08, Lev09b, Lev11b, Lev11c, Lev13, Rie98, dB97b]. Composable [BT88a, BT88b]. Compositional [MWRH13]. Comprehensive [Elr88]. Compression [JARKS22]. Computation [TMPM16]. computational [Bar09a]. Computer [Ano99f, Bra82, DoD87a, MH98, OW82, Puk88, Whe86, Whe87, Boe99, CC98, DRH98, FME01, Toa96, LC86]. Computers [Bun85, BRF92, BCF94]. Computing [BK22, Cor83, DRSK23, CMWT21, PZ97b]. Con [Dev17c]. Con-figure [Dev17c]. Concept [Bac82, LB80]. Concepts [EHP80, Sho87, Bag98, BS13, Gic91, Swa09b, SB11, SB12]. Conceptual [MK87, Mac84]. Conceptualization [DZM87]. Concerns [Pro20, FG86]. concert [Bei97]. Conclusions [MR10, dlPU07]. Concurrency [Bro98b, Lea87a, NDM98, RK01, Bar09l, BW10a, Kie01, Mic13, dlPM13, Rog97]. Concurrent [BKS87, Car90, Car91, CAU88, Che97, Cla87a, Coh82, Har87, KF98, Kim21, LKN97, MNG16, NMT92, San97, Tai86, TT02, Wel97a, Bar09a, BW99, BWK^{+}01, EKPPR04, GSX99, HM03, Pet10]. concurrently [CXY01]. conditional [LS98]. Conference [ACM82, ACM97, Ano99a, Ano06e, Ano06f, STF98, ACM87a, Ano92b]. confessions [Car01]. confidence [Goo13]. Confidentiality [RSK^{+}19]. Confidentiality-by-Construction [RSK^{+}19]. Configuration [MKP91a, Ter87, Kan12a, MKP91b]. configuring [Bis88]. Conflict [Lev01a]. Conformance [BdlP15]. Conformity [BDT99, BW15, Bra99, Ton99]. Connected [ICS22]. conquer [Taf12]. consideration [dlPP02]. Considerations [Bra83a, Won90]. Considered [Gon91b, Gon91a, Lad89, Duf09a, Duf09b, Moo96, Mor95a]. Consistency [KB83]. consortium [DV01]. constrained [LCB09]. Constraint [Bro83]. Constraints [MMPT16, TCRW88, Bei92]. Construction [BRKS22, Con97a, RSK^{+}19, Bar09h, Cha07a, Cha07b]. constructor [Duf08a]. constructors [MC09b, MC09a]. Constructs [OB97]. Contacts [Ano99g, Ano00f, Ano00g, Ano00p, Ano00q, Ano06g]. Container [MF04, DB09]. Containers [Dro22, Hea08a]. Contemporary [Boy89]. Context [DRSK23, SC06]. Continuous [BCB^{+}22, ALB^{+}14, KS01]. Continuum [BK22, SSGH^{+}22]. Contract [CdN16, HSB^{+}22, BHR^{+}11, BCHR12]. Contract-Based [CdN16]. contractor [Sma09]. contracts [Hir92, Log13a, Log13b, Ree85, Ree86]. Control [Aus22, BW16a, DCBM97, DDJ98, FMS98, Fri98a, Gre16, Lev88, MKP91a, Mor87, Qui90a, Sac89, Sch87a, SSJ85, Tv88, Wil87, WV98, de 88, AV93, BHR02, BR94, BF99, BWD90, CVW03, Elr89, Faß01, Fri98b, Gar09, GS10, Gre13, Lev98a, Lev05a, Lev09a, LSR^{+}88, MKP91b, ML95a, OWSB08, Qui90b, Spi00, TT02, VE92, WP13]. Controlled [Cel97, Kir12]. controller [Bre97, OS12]. controllers [GDAG97, HMRF97]. Controlling [Lev89, Ros87b, Ros87c]. Controls [Elr88]. convention [Ros95]. conventional [Con03a, Joh93]. Conventions [Van86]. convergence [BD01, KSD12]. Conversion [Mar86, SSJ85, Fro87, Wal85b]. Converting [Col99a, Wei90b, Moo93]. Cooperative [CCC23, Lei99a]. coordination [Fer97]. Coq [CAC^{+}13, Hug23]. CORBA [Bal99, Ber05, BF99, CN96, Cla97, Gid96, Ker99, Moo97, PQT99, ZHP06]. CorC [BRKS22]. Core [VRH21, JARKS22, LYB^{+}10, MMP13a, Nyb07, PMM13a, Rog12a, Rog12b, TD03]. cores [ZDM22]. Coroutines [Ves89]. Corporation [KM81, OW82]. correct [NIM07]. Correcting [ZBW07]. Correctness [Bal14, Bar00, BRKS22, Cha07a, MD22]. Correctness-by-Construction [BRKS22].

Cost [HS87, Ver21]. Costs [BKW82, HEUV99]. COUNT [SS89]. Counter [Gol93]. Counter-intuitive [Gol93]. counting [Bri12d, Bri12e, Bri12a]. Coupling [HD85, Nie86]. Course [CH97, JF98b, MH98, Wau83, CC98, JF98a, Lau07, MY98, Ruo05, Taf01c, Yu98]. Courseware [JF98b, JF98a]. CPS [SSB^{+}20, SS20, dNW23]. CPU [BW93a]. CQE [Mar19]. Crash [LV23]. Create [Gal20]. Creating [Cam92, Lei02, Och09c]. Creation [KBT84]. Creator [Wel97a]. Creek [Con97c]. Critical [AL00, Fra87a, LV23, Pro20, WCB16, YQZ^{+}23, Bro07, Car99b, Col99b, Dav04, Gar09, HB96, LHFD13, MGF16, Nil12b, Rog11a, SG06]. Criticality [ZDM22]. critique [PZ97b, VE92]. Cross [Bur87b, Bro03, HSWP12, Och09d]. Cross-Debugging [Bur87b]. cross-domain [HSWP12]. cross-platform [Bro03]. Crossroads [Ano95d]. Crusader [Edg01]. CS [CLY98, Ruo05, SS97]. CS1 [Car06b, MRB06]. CS1/2 [Car06b]. cue [New99]. Culling [RLPD98]. cultural [Oli94]. current [Bal99, GHV03]. curriculum [Rym94]. Cut [Lar22]. CVE [Mar19]. CWE [MB08, Mar19]. Cyber [BCB^{+}22, Bod19, MGB^{+}23, MGF16, Tro20, Whe19, XWZ^{+}23, ALB^{+}14, Fis12]. Cyber-Physical [BCB^{+}22, MGF16, MGB^{+}23, XWZ^{+}23, ALB^{+}14]. Cyber-Resilience [Whe19]. Cycle [MR83, Mur87, BK22, BF86]. Cycles [BMGS20, Ste12]. Cyclic [Ber15, Due97].

D 1 [Sha93]. Dafny [Lei12a]. DAG [ZDM22]. DAIS [BSPK22]. DARK [VBF89, VBF90]. Data [Ano90b, BK22, Bak86, BYY86, CA89, Car91, Dru99, Dun98, GES89, Hof86, JF98b, Kim21, Mar05, Mos22, Nyb10b, Riv17, SHR82, SJ91, SCC22, Wic82, Yeh82, And05, Bal95a, Bar01, Com90, CG87b, Dew09a, Dew09b, DB09, Gan04, JF98a, KETT96, LSP01, Moy11c, OS12]. Data-intensive [BK22]. Data-Types [Hof86, Wic82]. Database [BDD^{+}82, Hal83, OP85b, PVV85, SCD^{+}85, Tic82, FNS^{+}85, Ros04, Vas91]. Databases [McC87b, OP85a]. Dataflow [Jam98a, LKSL19, Jam98b]. DAWG [Pau86]. DBMS [MR87b]. DC [STF98, Ano99l]. DCOM [Bot99b]. DDC [Cle86]. DDS [DRSK23]. Dead [Gre05, MM98, EF01]. Deadline [BW16c, CR18, Sri06c, ABGH13, BW16b]. deadlines [Sri06c]. Deadlock [Che91a, Lev89, Lev98a]. Deadlocks [CAU88, Che90, GHL82, EGC13, TNGC05]. Deadness [HL85a, HL85b]. deal [Woo99]. Dear [Bot99a, Bot00b, Bry90a, Bry90b]. Debate [Ano93p]. Debugger [MP85]. Debugging [Bur87b, Dom87, Fai80, FRS97, GG16, HSW87, LP85, NPT97, Taf91a, Tuc97, BJRW96, DCC85, Taf91b]. decade [Bal14]. December [ACM80, ACM87a, Rob97]. Decentralized [LW02, XZ02]. decides [Fos20]. decision [EF01, Elr89]. deck [EF01]. declarations [Hod91a, Hod91b]. Declarative [Gal20]. Decomposition [BCD83]. Deep [LC22]. default [Ros86a]. Defense [Ada88, Eme83, Moo94, Ros87a, Sma09, Off88b, Off88c, Tas88]. Deferred [SRC13b, SC13]. defined [RH02, RH03, WB10b]. Defining [Con97b, Goo85, TBD22]. Definition [Ano06b, AD82, BBH80, KMS82, WGC17, Win90, Sri06d]. Definition-Use [WGC17]. Definitional [Vol87]. DEGAS [LP06, PL07]. degradation [Lev09a]. delay [BRF92, BW02, LA99]. Delays [RB85]. Delegation [Räi94]. Demo [Gon88]. demonstrably [NIM07]. demonstrably-correct [NIM07]. Demonstration [LD87, MNG16]. Denotational [MP84]. Department [Eme83]. Dependability [Gór20].

Dependence [Che92, Che97, Coh88]. Dependency [LSH98]. depending [Led95a]. Deployment [BK22]. Dereference [Ber86b]. Describing [Tai86, Ano88a]. Description [Bon84, HL85a, HL85b, MMSN09, Car88a]. Descriptions [MP84]. Descriptive [LWF91]. Descriptors [Bis80]. Design [Als83, BKS87, BHD98, Bei84, BYY86, BRW97, Boo82, Boy87, Buc87, BK85, BKW85, CM98, CS94, CG82, Fal82, GG16, GES89, Gor83, GR80, Har85, Har82, KF98, Ker92b, Ker93a, Ker93b, Kie89, Lat91, Lev82b, Lin82, Lin83, MGB^{+}23, MK83, MGF16, MNG16, Mur87, Pri82, Rud83, SPS88, Sof88, SWR82, San97, Shu91, Tem84, WBS97, Wal91, Whe19, WL98, Zhu90, Bag98, Bal95b, BT14, BKW^{+}94, BWK^{+}01, Car94, CM90d, Cro95, DB09, Fir91a, GSP^{+}11, Hos88, IMM85, Ker88a, Ker89, Ker90a, Ker94a, Ker94b, Ker95, Ker96a, Ker96b, Ker97, Ker98, KB97a, KB97b, Kle89, LVM90, MMN09, MP98, Pio86, PL07, Pul95, RDS98, Ros86a, San12, Sch91, Shu93, Sol91b, SU91, Var03, dlPZR^{+}01, Ad93, Ker90b, Ker92a, MNG16]. design/development [Pul95]. Designed [Rom00]. Designing [Che91b, Cla87a, Mos22, Pet10, Ros11a, Wad92, MF04]. Designs [BKC91, KB87, LKH16]. Desk [Sri06f]. Destructive [DM91]. detailed [Mah13, VBF90]. Detecting [BCH^{+}19, CXY01]. Detection [Che91a, FAT^{+}23, HL85a, HL85b]. detector [RA91]. determination [ML91]. Determined [Bar85b]. Deterministic [LMP90, GB94, RC10a]. Develop [Yu97, BC95, ML95b, Trü95]. Developer [Cra22, Ker93a, Whe86, Whe87, Dul03]. Developers [Har82, Ker90b, Ker92b, Ker93b, Lei99a, Ker86, Ker88a, Ker88b, Ker89, Ker90a, Ker92a, Ker94a, Ker94b, Ker95, Ker96a]. Developing [BB85, Col87, Lei12a, Mea87, NS03, Rob92, Ros11b, SG06, dB97b, BMW94, BWK^{+}01, Ros04, Sch09]. Development [Ano92i, Ano93g, BCB^{+}22, Bar85b, BGK^{+}82, BCG^{+}84, Bro03, Buc87, Bun85, Car89a, Fal91, GMO92, Gro07, Har22, Ker88b, Lad89, LNR87, OW82, PBB^{+}88, Reh87, SS87, Ter87, Wal87, Wil87, YQZ^{+}23, de 87, Bar08, Ben94, Bjo13, BdlPZ10, Car99a, Car88a, Car88b, Che92, Dew01, DA13, Edg01, Fir91b, Gar09, GDHM02, Lap04, Low99a, Mat96, MP91, OS12, Pul95, RDS98, Sny91, Spi00, SVK^{+}14, Wha13]. Developments [Bis91]. device [Dor99, LHFD13, MWRH13, NAF05]. Devices [FOSC23]. Devon [Bar87]. devoted [Bow92]. DFP [AB15]. DHACM [Tuc97]. Dhrystone [Wei89]. DIADEM [AG88]. Diagnostic [vdL84]. diagnostics [KPR93]. Diagrams [SJ91]. dialect [Men09]. DIANA [Taf82]. Did [Mor95a, Bri11d, Bri11e, Bri11f]. Difference [EHP80, Led92]. differences [NKN93]. Different [JA82, Mar21]. Difficulties [McC87a, Rob92]. Digital [GPZdlP21, PL07, HDHH98]. Dimensional [GP93, Rog88, Mac96]. dimensionality [SP12]. Dining [Age85]. DIR [BMW94]. DIR/SEE [BMW94]. directions [GST^{+}97]. Directive [DoD87a, DoD87b]. Discipline [Dru82]. disciplines [Bar09a]. discovery [KB97a, KW11a, KW11b, KW11c, KW11d, KW11e, KW11f]. Discrete [AS87, Bru82, Sho87, Wei90b, LP06, PL07]. Discrete-Event [AS87, Sho87]. Discriminants [Cla87c]. Discussion [Bry88]. disk [Nyb05]. dispatchable [ML99]. Dispatching [Ano06b, BA98, WB15, Asp01, Bur01, Och09d, Sri06b]. displays [BC95]. distance [SBH^{+}98]. Distributable [CDM87]. Distributed [AA88, AA89, AC85, BSPK22, Bal97, BKL85, Bis91, CM90c, Cle82, Cor83, CKF90, DGCR^{+}84, DGBMCG97, DZM87, DB09, Dob90, EJK89, Fuj87, GLV97, Gid96, Har99a, HW88a, HSW87, ILMV83, Jam98a, Jan88, JEKC89, KJEC87, KC90, KU84, Kni87, KR88, KVT88a, Mud87, NPT97, Pau87, Ros87d, Sac89, SV99, Taf91a, Vol87, Vol90, WV98, AAAG21, AW01, BTVC99, Ber05, Bro03, CMWT21, Con97b, DPB^{+}97, Gan01, Gan03, GH99, GH01, Gar90, GST^{+}97, GDHM02, GG99, HW88b, IMM85, Jam98b, Jam99, Kam95, KVT88b, LT99, Moo97, MKK99, NDP99, PZ97a, PT99, Qui11a, Qui11b, Qui11c, Qui12, RK99, Sot06, Taf91b, TP98, TGH10, TGH13, TPG21, UKDH97, UZ07, VGD^{+}97, Wel91, Wol97, Wol99, Moo97, TBA98]. Distributing [VMNM85]. Distribution [GGP^{+}90, Mud87, Vol90, AdB90, Bak90d, Bis88, DPB^{+}97, GdlP02, HP01, TG09, VHP10]. Diversely [Rom00]. divide [Taf12]. divide-and-conquer [Taf12]. division [Fro87, WBS97]. DL [Ker86]. Do [Ano99c, Ano99l, Bod19, Lei99b, Lei00, LM94, Bro11, Che09]. DO-178C [Bro11, Che09]. DO-178C/ED-12C [Che09]. DO-248C [Che09]. DO-248C/ED-94C [Che09]. DO-278A [Che09]. DO-278A/ED109A [Che09]. Document [Hov00, LRS09]. document-driven [LRS09]. Documentation [Whe86, Whe87, WB89]. Documenting [LP80]. DOD [Buc87, DoD87a, DoD87b, FG86, Fri83, GG87, Ros86b, Ros86a, Whi95]. DOD-STD-2167 [Buc87, FG86, GG87, Ros86a]. DoD-STD-2167A [Ros86b]. Does [Dru82]. dollars [Low99b]. Domain [RDP97, HSWP12, Jac13]. domain-specific [Jac13]. Domains [WB15]. Dorothy [DeW86]. DOS/PC/Ada [WD93]. Download [RDP97]. DPS [Cle86]. Dr. [Mor96a, Mor96b]. Draft [Lei99a, Ros86b, Ano10a]. Dragoon [AdB90]. dramoletto [Gre05]. Drawing [BL86]. Drift [Lev15b]. DRIP [MNG16]. drive [Nyb05]. Driven [CHHB90a, CHHB90b, MP85, YQZ^{+}23, DA13, HHBC90, Lap04, LRS09, WD93]. drivers [Dor99]. Driving [CCC23]. DRLMS [HDHH98]. DROOPI [QKP01]. DSA [Gan01, Gan04, Ker99, Moo97, PQT99, Qui12]. DSL [HSWP12]. DTD [Nyb10a]. DTD-specific [Nyb10a]. Dual [AW89, AW88, Gar09]. due [Nae05]. during [WGA90b]. Dynamic [Ano06c, Cel97, KT87, Lat09, Lef87, MD90, MSM^{+}03, RW99, Ros87b, Tin90, WW01, BW97a, CR05, Nil12b, Och12c, RLC01, Ros87c, Taf13a]. Dynamics [WBS97].

each [LLL03]. EACM [RA91]. Eagles [Bak91b]. earliest [Sri06c]. Early [Gri98, PDG83, CVW03]. easy [LW01]. Echo [Kni09]. ECLIPSE [Pie85, Gro07]. Ecological [Mur90]. economic [Wil91]. economics [Bar09a, RH91]. Ecosystem [BRKS22, SHT^{+}23]. ED-12C [Che09]. ED-94C [Che09]. ED109A [Che09]. EDF [Bur13a, CC18, TPG21, WB10a, ZBW07]. Edge [JBT^{+}22, SSGH^{+}22, BCHR12, Kan12b]. Edge-Cloud [SSGH^{+}22]. edition [Rog09e]. Editor [Bak92, Sch87b, Bri11b, Don90, MC90, Sri06f]. Editorial [Ano99e, Ano00e, Ano00n, Ano00o, Car02, Fis83, Sri06e]. Education [Ano92e, Ber84, McC00, McD88b, Weg82, LC86, Mac86, McC99, Toa96]. education/training [Mac86]. Educational [Rom88]. effect [Dis09]. Effective [Bai10, Bis80, BQ90]. Effectively [FOFY87]. effectiveness [Smi04]. Efficiency [Ard87, BFG85, EHP80, GS85, JA82, Sac89, Duf09b]. Efficient [AB15, Bur85b, KT87, Qui90c, Ros87d, SF82, Con97b, FSS87, Kir12, Rog09d]. effort [Bow92, EH13]. Eight [MP89]. Eight-Bit [MP89]. Eighth [ACM91b, Ano97]. Elaboration [Bel82, Gal20, Web93]. ELASTIC [NBZ^{+}20, SSGH^{+}22]. Electron [CA89]. Electronic [EF01]. Elementary [Mat87a, Sal92, Dri91c, Dri91a, Dri91d, ISO91a, Squ91a, Squ91b, Squ91c, Tan91b]. Elements [Coh86]. Elimination [Bro83]. Embedded [BTP22, Bra82, Chr87a, Col87, Cor83, DH80, DH82, FOSC23, Gal20, GG16, Glu09, Gre21, JARKS22, Kim21, LL98, Mid87, Mye85, PS84, Rog09a, TR87, TCRW88, Wag85, Whe86, Whe87, BC11, Buh85, Chr87b, DPB^{+}97, DD87, DA13, HMC88, LFT12, LCB09, Low99a, McC10, MS11, Mic02, Mos06, Pet10, Pot04, Rog11d, Spi00, SVK^{+}14, WWB99]. Ember [Gre21]. Empirical [FOFY87, JF98b, JF98a]. ENABLE [VGGS20]. ENABLE-S3 [VGGS20]. Encapsulation [Mat91]. Encoding [Ano17c, Bak93b]. End [BMNS85, Bro80, Bun85, GW80, Sim82, TGH13]. end-to-end [TGH13]. Endian [Coh94, Mar99, And05]. Endian-independent [Coh94]. Endian-safe [Mar99]. Endianness [Qui17]. ends [LW01]. Enforcers [CdN16]. Enforcing [CH04, BW93a]. Engine [Led92]. Engineered [Lat91]. Engineering [Ano92b, Ano99a, Ano99f, Ano00d, Ber83, Har97, Jac13, McC00, McD88b, MNG16, Mye85, Wai98, YQZ^{+}23, Bai10, Boe99, BMGS20, Cha07a, Dav04, Dav05, DA13, Fei14, Glu09, HS98, HCBM98a, Jen09, McC99, MY98, SBH^{+}98, SC04b, Wan99, Wel97b]. engineers [HS98]. English [Ano00c]. enhanced [ML86]. Enhancing [BHR^{+}11, Taf01a]. Ensure [NAT20]. Ensuring [Tro20]. entity [San12]. entity-life [San12]. Entries [Pow90, Led95a]. Entropy [Tro20]. Entropy-Augmented [Tro20]. entry [Led95a]. Enumeration [Mar19, MB08]. enviroments [KM98]. Environment [Ano92c, Ano92d, Ano93c, Ano93a, Ano94d, Ard87, BDD^{+}82, BHL^{+}93, BP94, BK85, BKW85, CSA^{+}87, Cra82b, DeL88a, EJK89, Fal91, Hou83, HW88a, Lev82a, Lev82b, LNR87, MSW85, MB91, McC87a, MR83, Pie85, Red85, Sta83, Wil87, XRL^{+}88, AKM^{+}91, Ano88a, BMW94, Bux85a, CC98, CSH03, DeL88b, Fel86, FSS87, Gar09, HCW04, HW88b, ML86, Mat91, RC10a, WD93]. Environments [ACM87b, All87, Ano91a, Bak87a, BKL85, BDF^{+}85, BDS81, Fai80, Fan84, Leb82, Obe94, Pys85, Wag85, Ano87, HBTW99, KGW^{+}85, PG94]. envy [Woo99]. EPTs [GS02]. Equivalent [SCD92]. ERA [LM94]. ERAM [Aus22, Sch10a]. Eratosthenes [And88, Col98, Dri89a, Dri89b, Hek89]. Erroneous [Coh88]. Error [Fro15, Kru90, PF20, LHFD13]. Errors [DM91, HL85a, PF20]. essence [McE03]. established [Kle21]. Estimation [FAT^{+}23]. Europe [Ano00j, Ano02a, Ano06e, Ano94c, Ano99i, Ano00b, NWW82, NW83, NW^{+}84]. European [ACW04]. Evaluate [SC06]. Evaluating [BFG85, RS91]. Evaluation [Ano90a, Ano90b, Bar08, Boe90, Bra94, Com90, Fal91, Fri87, HR07]. Event [AS87, Bru82, CHHB90a, CHHB90b, LW02, MP85, SRC15, Sho87, XZ02, HHBC90, KGL98, LP06, PG94, PL07]. Event-based [LW02, XZ02]. Event-Driven [CHHB90a, CHHB90b, MP85, HHBC90]. Events [SPS88, WB15, Sof88]. ever [Mor95a]. Everything [Boo11]. Evidence [Gór20]. Evidence-based [Gór20]. Evolution [Ano93d, HR07, Jam98b, KS01, PV13]. Evolve [BR01, Rom01]. Evolving [Mac80, Rym94, Sch91, Kle21]. examinations [Lit97]. Example [BKW85, CHHB90a, CHHB90b, Col89, CHGH19, Shu87, Whe86, Whe87, CN96, HHBC90, Spi00, Sum87, Car88b]. examples [Led95a]. Except [RS01]. Exception [Ano17a, BS01, BR01, Gau95, HM91, Li82, RdlPZFM01, San01a, WV01, AC03, Och09e, RS01, Rom01, SC01, Taf01a, Var01b]. Exception-ally [Ano17a]. Exceptions [Kie01, Ler01, MBW01, Qui90d, RK01, Var01c, Wol01, KR01b, PMJPA01, Var01a]. Excerpts [Off88b]. exchange [DB09]. Exclusion [bY93, SGS92]. Executable [Har85, EK11, Sei14]. executed [CXY01]. Execution [Ano06a, DCC85, FAT^{+}23, GS10, GS13, Gre16, JEKC89, Qui90c, RH10, Vol87, dlPZ03, BHR^{+}11, BW93a, BW07a, BW10c, Buz16, GST^{+}97, Gre13, HR03, LS98, RH07, Sri06a]. Execution-Time [Ano06a, GS10, dlPZ03, BW07a, HR03, Sri06a]. Executions [Maz89b, Tai86]. Executive [Ano94f, Ano95e, Ano95f, Ano95g, DZM87, FMS98, Ad93, ABW01, Ear92]. Executors [MMPT16]. Exercise [Huf82, FC91]. Existing [BDD^{+}82, Pys85]. Expedite [Lei99b, Lei00]. Experience [BRW97, Cha00, Dob83, Edg01, FCS83, Gil84, KFS97, KB87, Not80, PDG83, Pys85, RR16, Sch10a, TG09, Buh85, BW07b, CVW03, DR99, Kam98, PW01]. Experiences [Arn86, BTVC99, Bis91, BRF92, DRSK23, Dob93, GS02, Gór20, Hek83, Lea87a, MR87b, Ros04, Ruo05, Sch87a, SSJ85, AW91, BE02]. Experiment [Maz89a]. Experimental [AID05, BKW85, KK03, LW07, LSR^{+}88, WWB99]. Experimenting [Taf11]. Expert [Dob01a, Wal87]. explicit [CAC^{+}13]. Exploitation [Coh82]. exploring [Con97b]. Export [BT88a, BT88b]. Exporting [Ver22]. exposing [Swa07a]. Expressing [Bal95b, Gro86, Yem82]. expressions [Bei92]. Extendable [ML99]. Extended [Ano94f, Ano95g, Bec83, CdN16, CBW^{+}21, Whi85, Gre13, Joh93]. Extending [AH01, Cha82, LYB^{+}10, Low99a, MK91, NS85, RH01, BW03, GLZdlP16, Och09a]. Extensible [KW98, WJS^{+}01, SVK^{+}14]. Extension [SK22, ALB^{+}14, Rui10, Sei91].

Extensions

[Ano00w, RRG15, BD91, TMPM14].

extreme [AC04].

FAA [OS12, San01b, San03b, Sch10a]. FAA-qualifiable [San03b]. Facilitate [And20]. facilities [BHR^{+}11, BN87, BW92, Els91, Wre92]. Facility [CVW03, MC05]. factorial [Mor95b]. Factory [SC87, Hea08c]. Facts [Con90, WFF^{+}87]. fall [Swa10, Off88b]. families [Bur87a]. Fast [Sch87a, KM98]. Faster [WT89, WT88]. Fault [AA88, AA89, DGBMCG97, FD16, GGP^{+}90, Kam99, KU84, Kni87, KR88, LV23, Wol97, BPP06, DB09, GLV97, GdlP02, LYB^{+}10, PV98, PV02, TP98, Wol99]. Fault-Tolerant [KU84, Kni87, PV02]. FC [BD92]. Feasability [HvKPT87]. feather [Dew07a]. Feature [BW97a, TBD22, Taf22a]. Features [AKM^{+}91, BHD98, Bro97, Bro98b, Chr87a, Hou83, Mos22, SW87, Woo87, Chr87b, PMJPA01, TD03, UPRZ07, Wel99, WW01, Gau95]. February [LC86]. Federal [O’L07]. FIFO [Huf82]. FIFO Within Priorities [Ano06d]. Fifth [Ano91c]. figure [Dev17c]. Figures [WFF^{+}87]. Files [RLPD98, Bri09d, Kan12a, Nyb10b]. Filtering [PW97]. final [Ano10a, Gau95]. finalization [Gre99a]. financial [Hai00]. Finding [Lar22, BMT^{+}14]. Fine [BTP22, PMMT15, PMM15]. Fine-Grained [BTP22, PMMT15, PMM15]. First [Bur85a, TPG21, Wol01, Bra85, Sri06c]. First-Class [Wol01]. Fixed [Fro87, AdlPT97]. Fixed-point [Fro87]. Fixing [Bak90c, Taf01b]. Flexibility [LL88, Whi10]. Flexible [Rou85, SB80, BWV03, SLNM04]. Flight [Fri98a, Wai98, BGGS14, Fri98b, ML95a, WBS97]. Floating [Lea87b, Win91]. Floor [ABGH13, BW16b, BW16c, CR18]. flop [Woo99]. Flow [SJ91, ACW04, CH04, TGH13]. fly [BD99]. Fog [DRSK23]. Follies [Ano91b]. Force [Ada88, Gri98, Off88a, Off88b, Off88c]. Forcing [Pap89]. forget [BW10a]. Form [Car90, Ros89, Ano93a]. Formal [AL00, BBH80, Cle82, GSX99, Hum22, KMS82, Lar14, LB80, LNR87, SCD92, Taf20, Wag20, Win13, CHGH19, Dav05, HB96, HM03, Kni09, LA99, SC92, Ven08, Wha13, Pla86]. formalization [CAC^{+}13]. Format [Nyb10b, Bar01, San89]. Formatted [Whi81]. Formatter [Zhu90]. formerly [STF98]. formula [Jac13]. FORTRAN [BH90, PBB^{+}88, Whi81]. FORTRAN-like [Whi81]. Forward [vdL85]. FOSDEM [Cra22]. Foundation [ACM91b, Bro98a, Saï08]. foundational [Sei14]. Fourth [Ano90c]. FrameKit [KM98]. Framework [Gal22, PDN97, Ano88a, Gan03, KM98, MF04, RR14, RC10b, SRC13a, SLNM04, WB07b, KS06]. frameworks [BV13]. Frank [Rog11d]. Free [CM98, Bos13, Car98]. freedom [AC03]. frequently [Col95a, CR97]. freshman [CC98]. Friendly [Deb83, CC98]. Front [BMNS85, Bun85, GW80, Sim82]. Front-End [GW80]. Full [BA82, CG82, TNGC05]. Fully [dB99, dB97a]. fun [MRB06]. Function [Wol84, BA98, Tan91b, Wic86]. functional [Bei92, NBZ^{+}20, Shu93, SSGH^{+}22]. Functionality [BBB^{+}23]. Functions [KS84, Mat87a, Sal92, Dri91c, Dri91a, Dri91b, Dri91d, Dri91e, Duf08a, HR07, Hea08c, ISO91a, ISO91b, Joh93, Squ91a, Squ91b, Squ91c]. fungible [Lev11a]. Further [CC18]. Fusion [WV98]. Future [BDF^{+}85, Bux85a, Bux85b, CMR90, GST^{+}97, Moo96, Boe99, BB02, Dew01, DdlP03, PT99, Trü95, VP03, Wel01, SS94]. Fuzion [Sie21]. FY93 [Ano93i].

gain [LW01]. gains [Lew02]. Game [Hil22, HR07, Lev97a]. Gap [Qui17].

Gateway [DRSK23]. Gem [Ano17c, Ano17a, Ano17b, Bar09b, Bar09c, Bar09d, Bar09e, Bar09f, Bar09g, Bar09h, Bar09i, Bar09j, Bar09k, Bar09l, Bar09m, Bri09d, Bri09a, Bri09b, Bri09c, Bri11a, Bri11b, Bri11c, Bri11d, Bri11e, Bri11f, Bri12b, Bri12c, Bri12d, Bri12e, Bri12a, Cha11, Cha09, Dev17c, DFGZ09, Dew09a, Dew09b, Dew09d, Dew09c, Dew17, Dis09, Duf08b, Duf08c, Duf08a, Duf09d, Duf09c, Duf09a, Duf09b, Duf09e, FM09a, FM09b, Gas08, Hea08b, Hea08d, Hea08c, Hea08a, Kan12a, KW11a, KW11b, KW11c, KW11d, KW11e, KW11f, MC09b, MC09a, Moy11a, Moy11b, Moy11c, Moy11d, Moy17d, Moy17a, Moy17b, Moy17c, Moy17e, Obr09, Obr12a, Obr12b, Och09d, Och09e, Och09c, Och09a, Och09b, Och11, Och12c, Och12a, Och12b, Pan12b, Pan12c, Pan12d, Pan12e, Pan12a, Puc17, Qui11a, Qui11b, Qui11c, Qui12, Qui17, Reb17a, Reb17b, Rog09b, Rog09c]. Gem [Rog09d, Rog11c, Rog11b, Rog12a, Rog12b, dev17a, dev17b]. General [Bry88, SS87, bY93, FC91, MMP13b]. Generalizing [WB10a]. generate [AN05]. Generated [Kim21, HG14]. generating [BV03, Cha09, LZL03, Nyb10a, LRS09]. Generation [CEG23, Hov00, PDV98, Car06a, Lit97, Puk93, PdlPH^{+}07]. Generator [BMNS85, Car00, DS87, HB88, SHLR80, TRT16, WGC17, CS02, FC91]. Generic [HL86, HNS98, Hos90, MS87, PL07, Reh87, SCD92, BH14, Dri91a, Dri91b, Dri91d, Dri91e, Hea08d, ISO91a, ISO91b, NS03, QKP01, Rie98, SC92, Sla95, Squ91a, Squ91b, Squ91c, Tan91b]. Genericity [Gal20, Bak91a]. Generics [Bra83b, YG80, Moo10, Wor97]. genetic [NS03, SN04]. Georegistration [Swa09a]. Georgia [McC06a]. Getting [Rez22]. GKS [HS87]. GKS/Ada [HS87]. GLADE [PW97]. Global [TTRH85, Con97b, SC04b, Trü95]. GNA95GP [KGL98]. GNAT [BOM97, Bri09b, Bri09c, CDG97, Dew07a, GS02, Kir12, MSM^{+}03, MS04, MSK05, MD22, Och09c, Och12c, RTH15, Rog09b, Rog09c, Rog11c, Rui13, RSZ96, dlPRGB99]. GNAT-AJIS [Och09c]. GNATProve [Kan12b]. GNATTest [Kan12b]. GNU [ACW04, LP06]. GNU/Linux [ACW04]. Go [Ano99c, Ano99l, Bri11d, Bri11e, Bri11f, Dew07a, RMT11]. goal [Pio86]. goals [Car94, RSZ96]. Goddard [WBS97]. Going [Dew84, Rui13, Bar14]. gone [Bar14]. Good [Hil22, Har94c]. government [AW91, Hir92, Sma09]. Gprbuild [Kan12a, Bri11a]. GPS [Bri11b, Bri11c, Och12a]. GPU [FAT^{+}23, JARKS22]. Grained [BTP22, PMMT15, PMM15]. Grammar [CF82, Fis84a]. Graphic [Che91b, SGJP89]. Graphical [Gil84, MR87a, Tai86, XWZ^{+}23, Leo85]. Graphics [Car98, Puk88, Bra85, Bro04, Fir91a, MRB06]. GRASP [HCT^{+}98, HCBM98a]. Gripen [Fri98a, Fri98b]. Group [Ano92j, Ano92k, Ano93c, Ano93a, Ano93g, Ano94b, Ano94a, Ano95c, GMO92, Gre16, LWF91, MSW98a, OP85b, Vla93, Vla94, Ano88a, Bak90e, Boy86, Bro96, BP94, Cro90, Dow94, Gar90, Goo90, How86, Joh94, KGW^{+}85, MKP91b, MSW98b, Mun91b, Pen91, Qui90b, Rom88, Sol91b, Sri06a, Taf91b, Van90, Ano92c, Ano92d, Ano92g, Ano92h, Ano92i, Ano94d, BHL^{+}93, Dob01a, Whi95]. Groups [Ano99k, Ano00t, Ano00u, Ano00x, MDPK94, RH07, Ano93j, Ano94g, Ano95h, Ano95i, Ano95j]. GtkAda [MM17]. GUI [CM98, Car99a, Car22b]. Guidance [Wic98, LW07, New99]. Guide [BDV04, Fag00b, Mog91, Plo98]. Guidelines [DF84, FOFY87, NWW82, NW83, NW^{+}84, Off87]. GUIs [MVG99].

HACMS [Fis12]. HAL [Klu87]. HAL/S [Klu87]. Handlers [BA90b, Lev91, RH10].

Handling [Bur87a, BR01, CA89, Gre16, Kru90, Li82, Qui90a, SF82, WV01, Bri09d, GS10, GS13, HM91, KGL98, Moy11c, Och09e, RS01, Rom01, SC01, Var01b, Gau95]. hands [Buh85]. hands-on [Buh85]. happened [HBTW99]. Hard [McC87a, Wei90a, ABW95, BW94, Rog09a, UKDH97]. Hardware [Cas20, Har22, MP98, Riv17, WL98, MMSN09, MMN09, WA02]. Hardware-Based [Riv17]. Hardware/Software [Har22, MP98]. Harmful [Gon91b, Duf09a, Duf09b, Gon91a]. Hartstone [Wei90a]. Hash [Wol84]. HDF [Nyb10b]. headers [Cha09]. Heir [Reb17a]. held [Puk88]. helping [Har94c]. Here [Ano99c, Ano99l]. heterogeneous [GST^{+}97]. Heuristics [SJ91]. hexapod [TT02]. Hi [KSD12, Kan12b]. Hi-Lite [KSD12, Kan12b]. Hibachi [Gro07]. Hidden [BKW82]. Hiding [Cla87b, Pio86]. hierarchical [Bar01, SP07, Nyb10b]. hierarchically [AAAG21]. hierarchically-scheduled [AAAG21]. Hierarchy [BCD83, Rog09b, Rog09c]. High [BM97, DB98, EJ16, GS88, KQT^{+}21, PR98, Tok15, Whi95, ABW01, AW01, Bjo13, BDV04, BWM13, Cha13, Dew06, DB09, Dob01b, Fis12, Gil99b, Jen09, MCS97, PG94, Rog12a, Rog12b, Ros10, Ros11b, UZ07, Wic98, MSW98a]. high-assurance [Jen09]. High-Integrity [DB98, PR98, ABW01, AW01, BWM13, Cha13, Dob01b, Ros11b, UZ07, MSW98a]. High-Performance [EJ16]. high-reliability [Gil99b]. Higher [Ano00w, Ver21]. Highlights [Col95b]. Highly [SS85, Tuc97, BCHR12]. HILT’12 [San12]. History [Ano00d, BDS81]. holes [Dri89a, Dri89b]. HOLWG [Coh81]. Honeywell [Cle86]. HOOD [MVG99]. horizon [Sot06]. Host [Wil83]. Hotel [STF98]. HP [Mat91]. HP/Telegen2 [Mat91]. HRG [MSW98a]. HRT [MVG99]. Hugues [Rog11d]. HW [LKH16]. HW/SW [LKH16]. Hybrid [ALB^{+}14, MDPK94, Moo97]. Hypercube [CM89].

I/O [Deb83, Mat87b, Rog09d]. IBM [Wil87]. icons [Cra95]. idealized [LVTL22]. ideas [Rie98]. Identification [Bac84]. identifiers [Bak93b, Sri06d]. idiom [Hea08b, Rog11b]. Idioms [Hil82]. IDL [NDP00, SV99, ZHP06]. IEC [Plo01, Puk88, Tok15]. IEEE [Moo96, TPG21]. igloos [Oli94]. Ignition [CVW03, MC05]. II [Bla07, Car88b, DH82, FM09b, KR01a]. III [Duf09d]. Illustrating [LHFD13, Lev15b]. Image [FHN83]. imagery [Swa09a]. iMAX [ZW83]. Immediacy [Bak88]. Impact [Rei87, WBS97, Moo93]. Impacts [Car06b, HMZ00, SW87]. Impediments [Fir87a]. imperative [Lau07]. implement [DPP^{+}09]. Implementation [AdlP01, AB15, BCS89, Bei84, Bel80, BBH80, Bra83b, Bro83, BW07b, CSA^{+}87, DZM87, FHN83, Fal82, Fuj87, HB88, Hil82, JEKC89, Jha90, KU84, KVT88a, KVT88b, KGL98, Reh87, RDP97, SGS92, SRC15, San00, SP12, SB99, SGW90a, TBA98, Ves89, Wil85, AdlPT97, BE02, Bur99b, Car99a, CR07, CM90d, GS02, Hos88, Kir12, KM98, KP86b, KP86a, Mah13, MSM^{+}03, MSK05, RSZ96, SRN85, Taf11, Wel03, dlPZR^{+}01]. Implementation-Oriented [BBH80]. Implementations [Ano93f, FRS97, HL86, JA82, BS13, Mic02, SN04, Swa09b, SB11, SB12]. Implemented [GES89, Bos12, GB94]. Implementing [AD82, ABW01, BW94, Car22a, Che91b, GDAG97, HMRF97, KPP97, KR01b, Lav95, PMJPA01, Pow97, RLPD98, SAH01, UPRZ07, WCB16, WT88, WT89, MF04, Pot04]. implementor [How86]. Implications [Bra83b, McE03]. Implicit [LW02, XZ02]. important [GG16]. improve [Mau07]. Improved [CC18, ZHP06]. Improvements [BOM97, Rad94, VW13, dlPP02]. Improving [ACP11a, ACP11b, Bak88, Fra87b]. include [Mic13]. including [Hod91a, Hod91b, Sri06b]. incompatibilities [Dew09d, Moo93]. incomplete [LS98]. incorporated [SC06]. Incorporating [ABGH13, Ber15, RC10b]. incorrect [LS98]. Incremental [HCBM98b]. independence [And05]. independent [BF99, Car99a, Coh94]. index [KP86b, KP86a]. Industrial [AC03, BCMC23, Cha00, DH80, DH82, Win13]. Industry [Har82, Rom05]. inferring [Log13b]. Infinite [Dun98]. influence [AAAG21]. Info [Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u]. Informal [BK85]. Information [Ano01a, Ano06f, CA89, Cla87b, Dav04, Har01, KBT84, Ano10a, BF99, CH04, Faß01, Fus91, LS98, McE03, Pio86]. Infrastructure [JBT^{+}22, Bro09]. Inheritance [Bal95c, Bri94, MD90, Per88, Bal95b, Hir92, Hir94a, Hir94b]. inheritance-based [Hir94a, Hir94b]. Initial [Gau95]. Initialisation [Bur85b]. Initiative [Fis83, Fri83, Eme83]. Input [Bru17, Car89b, KP86b, KP86a, Moy11d]. input-output [KP86b, KP86a]. INRIA [KMS82]. Insertion [Fir87b]. Insertions [Fle86]. Instance [RDP97]. Instances [SCD92]. instantiation [BD91]. Instantiations [Hos90]. instrumentation [HCT^{+}98]. Instruments [LL98]. Insulation [Dru99]. integers [BCS89]. Integrated [HSB^{+}22, MB91, MP98, XRL^{+}88, HBTW99]. Integrating [BBB^{+}23, CH06, Cro95, Wan99, WJS^{+}02, WB07c, TG09]. Integration [BDD^{+}82, Mun91a, Ter87, BP94, Mat91, Mun91b, Sch10a, WRL13, WT03]. Integrations [And20]. Integrity [DB98, KQT^{+}21, NAT20, PR98, Tok15, ABW01, AW01, Bjo13, BDV04, BWM13, Cha13, Dew06, Dob01b, Lan10, Mac96, MCS97, Ros11b, UZ07, Wic98, MSW98a]. Intelligence [Ano94b, Ano94e, Ano95b, Ano95c, BSPK22, JBT^{+}22, SS20, Joh94, Wol85]. intensive [BK22, Mar19]. Inter [GZdlP15]. Inter-partition [GZdlP15]. interaction [ALB^{+}14]. Interactions [Fos20, BW97a]. Interactive [BR94, Che91b, Sta83, Ala13]. interchange [KETT96]. interchangeable [TG09]. Interconnections [Gro86]. Interest [Ano93c]. Interesting [Ano02c]. Interface [ACM89, AKM^{+}91, Ano94a, BST90, Boy89, Col95a, DS87, DeL88a, Fag00a, Gic90, Nyb87, Vla93, Vla94, Ano89c, CM94, CR97, DeL88b, FC91, Puk93, Vok92, Wal94]. Interface-Based [DeL88a, DeL88b]. Interfaces [BDF^{+}85, Cam92, ACM85, Hea08b, Mah13, MSK05, Och09a]. Interfacing [Bot99b, Dor99, Fan84, LMA94, McC87b, Mic07, MC09a, Och09b]. interim [Sch10b]. Interleaving [Moo18]. Intermediate [AD82, RTM82, Lei12b, SV99]. Internal [Taf82, DG97]. International [Ano88b, Ano90c, Ano90d, Ano91c, Ano91a, Ano93h, Ano93k, Ano97, Ano99a, Ano99f, Ano00i, Ano02d, Bar87, Bar88, Bro88, GB87, MR10, Obe94, STF98, ACM87a, Ano93b, BW93b]. interoperability [GST^{+}97]. Interpreter [DFS^{+}80, FRS97, Whe84, Hos88]. Interrupt [Alv87, BA90b, Gre16, Qui90a, GS10, GS13, Lev91, RH10, WD93]. interrupt-driven [WD93]. Interrupts [Hun88, WB15]. Intersection [RLPD98]. Introducing [Bar93, AW91, Bar07a, Bar07b, Kle21, Qui90d]. Introduction [BCMC23, BA07, BW07b, CM90a, Dri91c, Fel09, Fel11, HG07, Lea04, RM07, VR07, Bar09b, Bro09, CHGH19, Fre86a, Obr09, Och09b, Roy90b]. Introductory [CH97, MH98, Pag82, CC98]. intrusion [Lev05a]. intuitive [Gol93]. Invalidation [AP84]. Inversion [CS87, LMP90, Lev88, Lev11a, LSR^{+}88, Nae05]. Investigating [BKWS88, Mah13]. investigation [LSR^{+}88]. Investigative [FHN83]. invitation [Ler03]. invited [Bal99]. Invocation [LW02, XZ02]. IP [Car17, TP98]. IPCP [AB15]. IRTAW [TB02, VP03, dlPU07]. Irvine [OW82]. ISI [KMS82]. ISO [Ano99d, Plo01, Puk88, Tok15]. ISO/IEC [Plo01, Puk88, Tok15]. Isolation [Riv17, MPV10]. Issue [Ano06d, Ano06b, Ano06c, Ano06a, CM90a, Sri06a, Sri06b, Sri06d, Sri06c, Elr89]. Issues [Ano93h, AW01, Bar88, BKWS88, Bur92, BW87, BdlP15, CM90a, CM90c, CG88, GB87, GP18, Jha90, JLM^{+}85, KF98, KW91, Lad89, Mic16, PRQ21, RH16, RR90, VR07, VW18, Whi97, Ad93, Bak90e, Bak91c, Bar87, Bra98, Bro88, Bro07, BW93b, Bur99b, KB97b, LN91, Loc91, Mac86, Plo98, RR13, RdlP13, Van90, VHP10, WA02, Web93, Wel99, WP13, dlPM13, Ano88b, Ano90c, Ano90d, Ano91c, Ano93b, Ano93k]. Iterative [MNG16]. Iterator [Ros89]. iterators [Hea08d]. IVLs [Lei12b].

J [DV01]. Japan [Hag91, Puk88]. Java [Dob01a, Bal97, Bro97, Bro98a, Bro98b, BH02, BF99, CDG97, Dob01a, Dob01b, DV01, Fli98, GSTV97, KPPÉR06, KK03, Mun96, MH97, Nil12a, Nil12b, Och09d, Och09e, Och09c, Och09b, Pot04, RR14, San03a, Sch10a, SC01, TBA98, Wel03, WCB16, Whe97, Woo99]. Javaing [PV99b]. Java^{T M} [BD01, BHR02]. Jérôme [Rog11d]. John [Rog11d, Ano00c]. Journal [Ano99f]. Jovial [Bei84]. JSON [Mos22]. JSON-Like [Mos22]. JTC1 [Puk88]. JTC1/SC24/WG4 [Puk88]. Julia [Ver22].

June [BRC98, Col95b]. Junk [Con90]. just [Ame01]. JVM [GD00].

KAPSE [ILMV83, Tha82, Wil83, Wil85]. Karel [Hos88]. Kernal [Gil84]. Kernel [Leo85, Ros87d, SB99, WL98, MMB^{+}03, UPRZ07, dlPZR^{+}01]. kernels [Wre92, ZdlP02, dlPRGB99, dlPZ03]. Key [Ano99g, Ano00f, Ano00g, Ano00p, Ano00q, Ano06g, Bri11b, Hea08a]. Key-based [Hea08a]. Keynote [Bux85b, Car01, Dew01, Taf01b, Boe99, Bux85a, McC99, Sel99, Lis12]. KEYSTONE [Kie89, Kle89]. Kiasan [BCHR12]. kill [GL89]. kilogram [Puc17]. Kind [LVTL22]. kisses [Bri12b, Bri12c]. Kit [SCD^{+}85, FNS^{+}85]. know [Boo11, Con97d]. Knowing [Hil22]. Knowledge [Ano92b, CG88, MNG16]. Knowledge-Based [Ano92b]. known [JR10].

labels [FBL^{+}10]. laboratory [BTVC99, Wan99]. Lack [Rob92]. Lady [Bri12b, Bri12c]. LALR [CF82, Fis84a]. Landmass [HDHH98]. Language [ACM80, Als83, AB87, Bak86, Bak90a, BYY86, Bon84, Bro82, Bro98a, Bru17, BW10a, Cas20, CMWT21, CG82, Cra82b, Dew84, Gen91, Gor83, Had90, HMZ00, Har85, Har22, HL86, HSB^{+}22, HL85c, Kam83, Ker90b, Ker92b, Ker93a, Ker93b, KBL80, Lin82, Lin83, Mur87, PDG83, Pri82, Puk88, Qui90d, RH16, Rog11a, RTM82, SWR82, TBD22, Taf22a, Tha82, Tok15, VR07, VR16, VW18, WA02, Wau83, WQ83, Whe19, Whi95, ZW83, Abb96, Ame01, Ano89b, Ano10b, Bag98, BT14, BGGS14, Bra85, Bro09, BB02, BV13, Dew01, GBC^{+}14, GST^{+}97, Irw96, Jen09, Ker88a, Ker89, Ker90a, Ker94b, Ker96b, Ker97, MMSN09, Mat96, MK14, Mic13, NKN93, Och09f, PK97, Sei14, Ste12, Taf11, TMPM14, TD03, VHP10, Wal85b, Wel99, WV02, Wic98, Won99, Ker92a, Ker94a, Ker95]. language [Ker96a, Ker98]. Language/CASE [Ker92b, Ker93a, Ker93b, Ker94b, Ker96b, Ker97, Ker92a, Ker94a, Ker95, Ker96a, Ker98]. Languages [Ano00d, Cho19, DoD87a, Mar21, Mic16, SPS88, Sof88, BMT^{+}14, Bro07, DFGZ09, Jac13, Joh93, LMA94, Lei12b, SVK^{+}14, TP09, Ton99, Rog09e]. Large [Bur87a, Gal20, Kru90, MG87, Ros87b, Rou85, Sch87b, Ter87, WV98, ACW04, CVW03, HM91, Ros87c, Sch09]. latching [MRB06]. later [Vau98]. Layered [Taf21, Spi00]. layered-architecture [Spi00]. Lead [Dru82]. Leading [BCHR12, Kan12b]. Leading-edge [BCHR12, Kan12b]. leakproof [Bak93c]. Learn [FGN85]. Learned [SSJ85, BT14, Boo11, Kle21]. Learning [HMZ00, LC22, SBH^{+}98]. legacies [BMW94]. Legacy [BHD98, DeW86, Kle21, Mos06]. legally [Cha82]. Lego [Fag00a]. LEGO(R) [BdlPZ10]. Length [Car89b]. lesson [KW11a, KW11b, KW11c, KW11d, KW11e, KW11f]. Lessons [Buh85, SSJ85, BT14, Kle21]. let [BW10a, Moy11a, Moy11b]. Letter [Bak92, Don90, Har94a, RH96, Bri86, Fir86, PR86, Pla86, Squ86, Tex86]. Letters [MC90]. Level [Ano00w, Bak87b, BOM97, BM97, HSB^{+}22, Pro20, RTM82, Con03b, Dor99, MMSN09, MMN09, Mah11, Mah12a]. Leveraging [HG14]. Lexical [Had90]. LEXICAL ANALYZER G [Had90]. liaison [Bro96]. LibAdalang [Ros21]. liberated [Mor95a]. Libraries [Dun98, MKP91a, Mor87, HG07, MKP91b, RT21]. Library [Ano00c, Dau87, Gre21, MD22, MS87, NS85, PF20, Sol91a, Bal95c, Bos12, CS91, Con03a, CHGH19, LHBK87, Lea04, PS06, Sol91b, Con97b, Con97d, MF04]. Libre [Jen09]. License [Lei99a, GL89]. Life [BK22, BF86, BMGS20, MR83, Mur87, DeW86, San12, Ste12, Lev97a]. Life-Cycle [Mur87, BK22]. Lifecycle [Wag85, Dav04]. Light [MD22]. Lightweight [FMS98, Gal22]. Like [Mos22, Dew07a, Khr95, Lei12b, Whi81]. Limitations [CSL^{+}87]. Limited [Bak91b, Bak93a, Bak93c, Bei92, Duf08b, Duf08b, Duf08c, Duf08a]. Linda [LW97]. Line [Fir88, Gic90, dev17a, dev17b, SAH01]. line-based [SAH01]. Linear [Klu87, Ves90a, Ves90b, EKPPR04]. Linearity [Cam92]. Lines [Win90, BJRW96]. Linkage [FA82]. LINPACK [PG91]. Linux [ACW04, SRC15, SB99]. LISP [DS87, Wal87]. list [Ree85, Ree86, Rom88]. Listing [Wal85a]. Lite [KSD12, Kan12b]. literals [Gas08]. lithography [RLPD98]. Live [MM98, Gre05]. Loader [RDP97]. Loader/Verifier [RDP97]. loading [CR05]. Local [Ano95a, Ano99h, Ano00h, Ano00r, Ano00s, SCD92]. Locally [Lar22]. locating [WW01]. Lock [Bos13, Mal88]. Lock-free [Bos13]. Lockheed [Kle06]. Locking [Ano06d, BW13a, lPB18, Bur01, BW13c]. locks [Rog11b]. Logger [Gal20]. logic [Bal14, EKPPR04, MP91, PL07]. Logical [Saï08, Fir91a]. LOLITA [RTM82]. Long [MM98]. longer [Gre05]. Look [Dew84, Sma09]. Looking [MSW98a, MSW98b, vdL85]. Looks [Hil22]. Lookup [Tro06]. Loop [AW89, Sch87a, AW88, Buz16]. losing [Low99b]. lossless [Bak93b]. Louis [ACM97]. Lovelace [Whe95]. Low [Bak87b, BOM97, RTM82, Dor99]. Low-Level [Bak87b, BOM97, Dor99]. LowerLayer [GBCGDBC97]. Lustre [LVTL22].

M2OS [RT21]. MA [ACM80]. MA1 [McC07]. Machine [Bis80, Fle86, GR80, Lah82, Lis12, CDG97]. Machines [Che91b, San00, VMNM85]. macros [San89]. made [Cro14]. Magnavox [Reh87]. mailboxes [Qui11c]. maintainable [Irw96]. Maintaining [TS20, BMW94]. Maintenance [Ano10b, Bru17, Dew84, HEUV99]. Major [Mun96]. Majors [CH97, CLY98, SS97]. make [RMT11]. Making [EK11, Mah11, Mah12a, Pie90, BF99, Elr89, Plo01]. Management [BK22, Bra82, GS85, Hal83, KBT84, KT87, MKP91a, PVV85, ACW04, Ano89a, Bak90d, Bak93c, Bar09i, Bri92a, Bri92b, Kle06, Med91, MKP91b, Nil12b, PV99a, Van94]. Manager [Car22a, Hum22, Mal88]. Managing [Cel97, HR03, Sch87b, SSGH^{+}22, Bri11c]. Mandate [Har97]. maneuvering [EF01]. Manifestation [Cri01]. manifested [Med91]. Manipulations [DGLM85]. Manual [Fag00b, Ber86b]. Many [VRH21, DFGZ09, MMP13a, PMM13a]. Many-Core [VRH21, MMP13a, PMM13a]. MAP [SC87]. Mapping [NDP00, Taf21, TCRW88, SU91, VE92]. mappings [GG87]. Marching [SS94]. market [Gil99a]. Marketplace [Moo94]. markets [Hai00]. Marsaglia [HB88]. MaRTE [RTH15]. Martin [Kle06]. Massachusetts [ACM87a]. Master [SBH^{+}98, dev17a, dev17b]. Matching [MF91, TBD22, Taf22a]. material [Wic82]. math [CS91]. Mathematical [Moy17e]. Mathematics [Reh87, Mau07]. Matrix [FCS83, Hek83, Ker92b, Ker93a, Ker93b, Hod91a, Hod91b, Ker86, Ker88a, Ker88b, Ker89, Ker90a, Ker92a, Ker94a, Ker94b, Ker95, Ker96a, Ker96b, Ker97]. Matrixr [Ker90b]. mature [Sch09]. maturity [Mog91]. Max [Lar22]. Max-SMT [Lar22]. May [Bar87, BH14]. mbeddr [SVK^{+}14]. McCormick [Rog11d]. meaning [Sei14]. Means [Fri87, LL88]. Measurable [SSFO86]. measure [BC11]. Measurement [GCM90, PDN97, Roy90a, Wei89]. measures [SM92]. Measuring [BW93a, Smi04, XCZ04]. Mechanism [Mun91a, Led95b, VE92]. Mechanisms [Atk90, Coh85, Fer97, ML99, Mun91b]. Mechanization [Hug23]. Medical [LL98, LHFD13, MWRH13]. Meeting [ACM91b, Ano92f, Ano94d, Ano94e, Ano95b, Orb85, Puk88, Bar98, Col95b, How86, MFD85, Obe85, Rob97, Ano92k, Ano95m, BRC98]. Meetings [Ano00j, Ano00i, Ano00k, RH96]. Memory [Lef87, LKSL19, TCRW88, Van94, Bar09i, Bri11d, Bri11e, Bri11f, Nil12b, SLNM05, WMM10]. Mentor [DGLM85]. Mentor-Ada [DGLM85]. MERCURY [MK91]. Mesh [VRH21]. Message [Bro99, Bro00a, Bro00b, Bro00c, Bro00d, Bro01, Col01, Col02, Har94b, Hos89, PDV98]. Meta [PS06]. METAH [Lew02]. metamodel [PdlPH^{+}07]. metamodel-based [PdlPH^{+}07]. metaphysician [Too91]. Method [Car89a, GS88, LP80, SF82, Wei90b, Car88a, Car88b, SU91]. Methodologies [Wag85]. Methodology [Bur85a, Har85, Kie89, Lad89, Lat91, MGB^{+}23, MSW85, Pri82, RG90, Roy90a, SS87, SHR82, de 87, JR10, Ker88b, Kle89, Pul95]. Methods [Boy87, Bry88, Che91a, AW91, Dav05, GSX99, Pla86, Sol91b, Win13]. Metrics [BW91, Pri96, Pri01]. MF1 [Cha07b]. MHP [CXY01]. microcontroller [RC10a]. Microcontrollers [Riv17, ARPT18]. Microprocessor [DH80, DH82]. Microsoft [Bal14, Bot99b, BM97]. Middle [Bro80, Gra83]. Middle-End [Bro80]. middleware [BPP06, QKP01, TG09]. migrate [Mos06]. Migration [MP98]. MIL [RM88, Roa88, Roa89]. MIL-STD-1750A [RM88, Roa88, Roa89]. Military [Ada88, AB98, Off88a, Fis12, Off88b, Off88c]. Mindstorms [BdlPZ10, Fag00a, Fag00b, FME01]. Minicomputer [FHN83].

Minicomputer-Network [FHN83]. Minimal [BCH^{+}19, Wil83, DRF97]. Minimizing [GS88]. Minutes [How86, Pau86, Rob97]. mispredictions [Lat09]. missile [LW07, Spi00]. missing [PMJPA01, Pio86, WB07c]. Mission [Fra87a]. Mission-Critical [Fra87a]. Missions [WCB16]. Mixed [ZDM22]. Mixed-Criticality [ZDM22]. Mixing [Fir88, Ves89]. mixins [Sei92]. MMAIM [Car88a, Car88b, Car89a]. MO [ACM97]. mod [Duf09c]. Mode [Bak93a, BQ90, AdlP01, SRC13a]. Model [ACM89, AB87, BCB^{+}22, BW90d, Cle82, HSB^{+}22, Jam98a, Lap04, LWF91, LKH16, LB80, Mac84, PRQ21, Ros22, SYW85, TRT16, TMPM16, YQZ^{+}23, AP11, Ano89c, BW90b, BW99, Cha13, Dob93, DA13, Fei14, Gan04, Jam98b, LHBK87, LVTL22, LW01, LZL03, LA99, McC99, Moo97, MMP13b, NDP99, New95, Pen91, PQR18, RR14, RH91, RT09, TGH10, TGH13, Ton99, Wha13, CN96]. Model- [HSB^{+}22]. Model-Based [BCB^{+}22, TRT16, Fei14, Wha13]. Model-Driven [YQZ^{+}23]. Modeled [Klu87, LKH16]. Modeling [DRSK23, GDHM02, NDP97, NDP00, Sau05, SSB^{+}20, SS20, SB23, ALB^{+}14, BMT^{+}14, DRH98, GSX99, Glu09, LHFD13, Mah11, Mah12a, NDM98, San12, Sei14, SP07, WV02, Wha13]. Modelling [BM23, Mur90, RD23, XWZ^{+}23]. Models [AL00, FD16, Men87, BW97b, Buz16, CH04, GBC^{+}14, HG14]. modern [HEUV99, Mar19]. modernization [Nil12a]. Modernizing [And20]. modes [RC10b]. Modular [BCD83, Kim21]. Module [Gro86, LV23, SB99, San01b]. Modules [Wat87]. modulo [Bjo13]. Monitor [EHP80, SN94]. Monitoring [BGK^{+}82, BCG^{+}84, BTP22, GHL82, LKSL19, BW93a, DCC85, LYB^{+}10, LS98, MMB^{+}03, NAF05, RH10]. monitors [KPPÉR06]. monotonic [Cro95].

MOPping [MBW01]. Moral [BM85]. Morals [WQ83]. Moretonhampstead [Bar87]. MORPHEMIC [BK22]. Mortem [HS87]. MOSI [Har88]. most [GG16]. Motif [Mat91]. Motion [Tuc97]. Motivation [Lev82b, Ric20]. Motorola [KNB08]. Moving [Ber84, KQT^{+}21, KETT96]. MP1 [Sin07]. MPHF [Tro12]. MS [Puk94]. MS-Windows [Puk94]. Multi [BBH80, Gen91, Had90, JARKS22, Nyb07, Och09f, PV98, ZDM22, FSS87, LYB^{+}10, MKK99, Nae05, Rog12a, Rog12b, Rui10, dB97b]. Multi- [PV98]. Multi-core [JARKS22, Nyb07, LYB^{+}10, Rog12a, Rog12b]. Multi-cores [ZDM22]. Multi-Language [Gen91, Had90, Och09f]. multi-opportunity [Nyb07]. Multi-Processing [BBH80]. multi-processor [FSS87, Rui10]. multi-tasking [Nyb07, dB97b]. multi-threaded [MKK99, Taf13b]. multiagent [Bar09a]. multicast [PVF01, TP98]. Multicore [FAT^{+}23, PM16, BMT^{+}14, PMM13b, Taf12, ZdlP13]. Multilanguage [GD00, HCW04]. Multimicroprocessor [DGCR^{+}84]. Multiple [Rom00, Bri09d, HR03, Hea08b]. multiple-unit [Bri09d]. Multiplication [FCS83, Hek83, Fro87]. multiprocess [VGD^{+}97]. Multiprocessor [Ard87, Bur85b, BW10b, DZM87, RTH15, lPB18, BW10c, BW13a, BW13b, BWM13, Low99a, RR13, SRC13a, WP13]. multiprocessors [GZdlP18, LWB13]. multiprotocol [Gan01]. multitask [San12]. Multitasking [Gon90, KB87, Li82, Yem82, And88]. multithreaded [KR01a, KR01b]. Music [Pie90]. Must [Bak93a]. Mutex [AR95]. Mutual [bY93, Elr89, SGS92, VE92]. my [Bri11d, Bri11e, Bri11f]. Myro [Men09].

Name [Mac87]. Named [WMM10].

Naming [CU89, Ros95]. NASA [Ano89a, WBS97]. National [CVW03, MC05]. Native [Fli98]. Naval [SPS88, Sof88]. NCSA [Bar01]. Need [Dru82]. needed [MWM10]. Needs [Mar21]. Nesting [Bak91b, CWW80]. Net [WGC17, Bot00a, Che92]. Nets [Che97]. Network [Car17, CS94, FHN83, GBCGDBC97, JBT^{+}22, Kie97, SC87, RR14]. Networked [FOSC23, Mar19]. Networks [SCC22, CB07, DRH98, Gan01]. Neumann [Mor95a]. Neural [CS94, SCC22, CB07]. News [Ano92e]. Newsletter [Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u, Ano01a]. next [Bro11, TB02, dlPU07]. nice [FBL^{+}10]. No [WGA90b, Bar14]. node [WGA90b]. Nodes [GA90, Vol90, Gar90]. nodes/distributed [Gar90]. Nomination [Har01]. Nominations [Har99b, Har00]. Non [Bur01, Cam92, CH97, CLY98, Mar86, NBZ^{+}20, SSGH^{+}22, SS97, EK11, HS98, MM21]. Non-Ada [Cam92, Mar86, MM21]. Non-CS [CLY98, SS97]. non-executable [EK11]. Non-functional [NBZ^{+}20, SSGH^{+}22]. Non-Majors [CH97]. Non-preemptive [Bur01]. non-software [HS98]. Notation [Bis86, Che91b, SGJP89, Tai86, Tom97, AdB90, Duf08c]. Note [Tis83, Nyb05]. Notes [Ano02c, Ano02e, Bro83, Cla87b, CG87a, CG87b, PVV85]. November [Ano99l, STF98, ACM97, McC06a]. NRC [Cra97, Taf97]. NT [BBB98, HCBM98a]. NTT [Tan91a]. Nuclear [CCC21]. nuisance [Mor95a]. null [Duf09a, Duf09b]. NUMA [WMAB10]. Number [HB88, AAAG21]. numbers [BMT^{+}14]. numeric [Gas08]. Numerics [Ros22, Squ91c]. NXT [BdlPZ10]. NYU [DFS^{+}80].

O [Deb83, Mat87b, Rog09d]. Object [Ano92j, Atk90, Bak91a, BHD98, Boo82, Boy87, Bro97, Car00, CN96, Col89, Els91, Fir91a, FMG90, GA90, Gre90, Joh93, KF98, Kru90, Lad89, MM98, Moo97, NMT92, NM92, SS87, Sei91, Sei92, Shu91, Tem84, Var01b, WBS97, Wal91, Wel97a, WdlP97, WV01, Yu97, AW91, And05, AdB90, Bar09g, Bar09h, Car94, Fir91b, Gan03, LW01, LZL03, Lit97, MT01, MH09, NDM98, NDP99, Pri96, Pri01, RDS98, Ros10, Ros11b, Sch91, SS91, Shu93, Sot06, WJS^{+}02, dB97b]. Object-Based [Kru90, Wal91]. Object-Oriented [Atk90, BHD98, Boy87, Bro97, Car00, Col89, KF98, Lad89, SS87, Shu91, Tem84, WBS97, Yu97, Bak91a, Fir91a, Moo97, NMT92, NM92, Sei91, Sei92, WdlP97, AW91, AdB90, Car94, Fir91b, Lit97, NDM98, NDP99, Pri96, Pri01, RDS98, Ros11b, SS91, Shu93, WJS^{+}02, dB97b]. ObjectAda [BE02]. Objectives [BSPK22, WG20]. Objects [Cel97, Cla87a, KPP97, LXY98, Ros87b, San00, Wei90b, Wol01, Yeh82, dB99, BD91, CM94, GZdlP18, GSX99, LKN97, Qui11b, Ros87c, WJS^{+}02, dB97a]. OBOSS [VC01]. Observations [Mat87b]. October [ACM82]. officer [EF01]. officers [Whi85]. Offset [Ver21]. Ohio [LC86]. OK [Bar95]. OLE [Bre97]. Omega [LW01]. OMG [Cla97]. Omni [STF98]. OMS [LM94]. On-board [AB98, ML95a]. one [Bar14, WGA90b]. only [Ker96b, Ker97, Ker98, Sel99]. onlywhen [VE92]. onto [MRB06, TCRW88, WD93]. OO [Car06a, LM94]. OO-ERA-RDBMS-OMS [LM94]. OOD [Bro91, Fir90, Hir94c, WD93]. OOP [Car97, WB07c]. OPC [DRSK23]. Open [Gar09, Tok16, KR01a, KR01b, Kle21, MMB^{+}03, RdlP13, dlPZR^{+}01]. Opening [Bak90b]. OpenMP [KQT^{+}21, PQR18, PRQ21, Taf21]. Operating [Fuj87, Mos20, Nyb87, RH07, Whi82, ZW83, Mic07, RC10b]. Operational [AD82, Li82, CVW03]. Operationalized [PF20]. operations [Hea08d, Hod91a, Hod91b]. Operator [SF82]. Opportunity [Mun96, Nyb07]. Optimal [AR95, Tro06]. Optimization [BK22, Bur92, CM90b, KUP^{+}83, LC22, OB97]. Optimizations [Dav82]. optimize [BC11]. Optimized [MF91, Tuc97, LZL03]. Optimizer [TTRH85]. Optimizing [BD99, EH13, RR90, SB05, ZHP06]. Options [AKM^{+}91, DD87]. oracles [HB96]. Oranges [Fir88]. Orbix [Cla97]. Orca [Bal95a]. Orchestrating [MC05]. Order [Whi95, Web93]. Ordering [SGW90b]. organisms [Lav95]. Organization [Kam83]. organized [Bow92]. Organizing [Fuj87, Gan04]. Orientation [WV01, MT01, MH09, Var01b]. Oriented [Ano92j, Atk90, BHD98, BBH80, Boo82, Boy87, Bro97, Car00, Col89, FMG90, GA90, Hai00, KF98, Lad89, Mur87, Sch87b, SS87, Shu91, Tem84, WBS97, Yu97, AW91, AdB90, Bak91a, Bar09g, BS13, Car94, Els91, Fir91a, Fir91b, Joh93, LSP01, Lit97, Moo97, NDM98, NDP99, NMT92, NM92, PC05, Pri96, Pri01, RDS98, Ros10, Ros11b, Sch91, SS91, Sei91, Sei92, Shu93, SK22, Swa07a, Swa07b, Swa09b, SB11, SB12, WdlP97, WJS^{+}02, dB97b, Wel97a]. Origins [Woo87]. orthogonality [WT03]. OSF [Mat91]. OSF/Motif [Mat91]. Other [Cro90, BA07, LLL03, Squ91c, TP09, Ton99, Wel99]. Our [Bod19, BBPT12]. outermost [And05]. outline [Ano10b]. Output [Sla95, Whi81, KP86b, KP86a]. Outstanding [BW90c, PK97, BW90a]. Overhead [BN87, Pau93]. Overload [MF91, Duf09e]. Overloading [PWDD80, SF82]. Overview [Ano90a, Ano90b, Bai20, Bod19, BRKS22, BK85, BKW85, Car22b, CG88, Cou21, Dob01a, Moo98, Rud83, VBF89, Com90, LN91, Lop99, Nil12b, PZ97a, PZ97b, Ryb94, San12].

PACEMAKER [Lar14]. Package [Bak87b, Bar85b, Bru82, Fro15, Gen91, GA90, Had90, Klu87, Mat87a, Pyl84, Reh87, Sal92, SCD92, Dri91a, Dri91b, Dri91d, Dri91e, HD85, ISO91a, ISO91b, Mac96, PG94, Rog09b, Rog09c, SC92, Squ91a, Squ91b, Tan91b]. Packages [Fis84b, HNS98, Lla92, LP80, Mac84, Ros86c, SN88a, vHLKBO85, Hod91a, Hod91b, Sla95, Squ91c, SN88b, XCZ04]. pairs [CXY01]. PAL [Con97d]. Pallada [PGRZ92]. Pamela [Boy87]. Panel [Ano92j, BBPT12, BMT^{+}14, Plo01, HBTW99]. Paper [Als83, Gre18, Mic01, Taf01a, Wek90]. Papers [Ano92b, Ano93h, Ano93o, Ano94c, Ano99f, LC86]. Paradigm [BKS87, BT88a, BT88b, VGD^{+}97]. Paradigms [BN87, MWM10, Mic13]. paradox [Ros09]. Paraffin [Moo11]. Parallel [CBW^{+}21, CM90c, Coh82, GCM90, HR07, Jha90, Moo18, PZ97b, PM16, PV18, PRQ21, SS85, TMPM16, Ver21, Yem82, AP11, CMWT21, KK03, McC07, McC09, McC10, Moo11, PMM13b, Rog11d, RK99, Taf11, Taf13a, Taf13b, TMPM14, WA07, Bur13b]. Parallelism [Moo10, MMP13b, Not80, PMMT15, PMM15]. Parameterization [BYY86, Tra89, Wek90]. parameterized [SS91]. Parameters [Bak93a, SCD92, Led95a, SC92]. ParaSail [Taf11]. Parser [Car00, Car06a]. parsers [Nyb10a]. Parsing [Nyb10b]. Parsl [CBW^{+}21]. Part [Bri09b, Bri09c, Hir94a, Hir94b, Och12a, Och12b, Bri11d, Bri11e, Bri11f, Bri12b, Bri12c, Bri12d, Bri12e, Bri12a, Car88b, Dew09a, Dew09b, Duf09d, Duf09b, FM09a, FM09b, GG16, Kan12a, KR01a, KP86a, Mau07, Moy11a, Moy11b, Obr12a, Obr12b, Pan12c, Pan12d, Pan12e, Pan12a, Qui11c, Qui12, RR13, Rog09b, Rog09c, Rog12a, Rog12b, WP13, KP86b, Moy17a, Moy17b, Moy17c, Whe86, Whe87, dev17a, dev17b]. partial [BD91]. Participation [Ano93l, Ano93m, Ano94h, Ano02e]. partition [GZdlP15, GHVVW93]. Partitioned [JEKC89, Mor87, AAAG21, Dob00, ZdlP13]. Partitioning [Tok03, Bis88]. partitions [Dob93]. parts [HMC88]. Pascal [BD92, AGG^{+}80, MH98]. Pascal-FC [BD92]. Passed [Bak93a]. Passing [Hos89]. Passive [Pie87, Ros89, LMV93]. patents [Wil91]. Path [Dru82, New99]. Pathfinder [RR14]. Pattern [RDP97, TBD22, Taf22a, DB09, GSP^{+}11, KB97a]. Patterns [BHD98, San97, HG07, PdlPH^{+}07, Sel99, Var03]. PC [WD93, Sny91]. PC-based [Sny91]. PDL [Bon84, Gra83, Ker82, Moo96, SWR82, Yav85]. PDL/Ada [Ker82, SWR82]. Peculiarities [Ben84]. Pedestrian [FAT^{+}23]. pennies [Low99b]. Perfect [Wol84]. Performance [BOM97, BFG85, BG90, BH90, CM90a, EJ16, Fra87b, GCM90, Kni90, Pau87, SW87, SM92, Whi97, WHNB91, de 87, AID05, Bur90, GSP^{+}11, KK03, New95, Rog12a, Rog12b, RA91, SC06, Syi95]. Periodic [Qui90c, GB94]. Permissions [Fos20]. persistence [Swa10]. personal [Bar98, Sil98]. Perspective [SYW85, LRS09, Oli94, Sma09, Win13]. perspective-bridged [LRS09]. PFW [KS06]. phased [Mog91]. Philosophers [Age85]. Physical [BCB^{+}22, MGF16, ALB^{+}14, MGB^{+}23, XWZ^{+}23]. pilot [OS12]. Pinching [Low99b]. Pioneering [Fra87a]. PIWG [Ano93e, Gau90a, Gau90b, PC90, RG90, Roy90a, Squ86]. Place [Coh86, Wal85b]. Plan [Har97, Con03a]. Planning [MFD85, LS98]. Plans [RSC16, TB02, dlPU07]. platform [Bro03, BF99, RTH15]. Platforms [JARKS22, BW10c, BW13b, KETT96, PMM13a]. Platinum [Rog21]. Plato [GG16]. plenary [Gil99b]. plug [CR05]. plug-in [CR05]. Plugging [Dri89a, Dri89b]. PM [Ano99l]. Point [Har88, Lea87b, Fro87, Win91]. pointers [Bar09e, Gre99b]. Pointing [Gre90]. Policies [Ano06d, Ano06b, Asp01, Bur01, BW13a, KPPÉR06, TG09, WT03]. policing [NAF05]. Policy [Ano99e, Ano00e, Ano00n, Ano00o, Car02, DoD87a, Sri06e, AR95]. polymorphism [Hir92]. pool [WMM10]. Portability [BOM97, Mat87b, NWW82, Lew02]. Portable [AD82, BM97, CM98, FG82, KT87, TBA98, KP86b, KP86a, LHBK87, Tan91b, Vok92, WGA90b]. porting [ACW04]. Ports [VRH21]. Position [Als83, Gre18, Mic01, RH10, Taf01a]. positioning [Trü95]. POSIX [AH01, GDAG97, HMRF97, Pow97, RH01, dlPRGB99]. possibly [Moy17d]. Post [HS87, BH14, MWM10]. postconditions [Dew09c]. Power [FAT^{+}23]. PQCC [Bro80]. Practical [Col87, Log13a, LP80, Mic02, Buh85, Led95a, LG88, Pot04, Ven08]. Practice [MM17]. pragma [Dis09, Tok03]. PragmAda [Car04]. Pragmatic [Fir87b, Pul95]. Pre [Cha82, BH14]. Pre-Processors [Cha82]. pre/post [BH14]. Precise [ZdlP02]. Precision [Lea87b, Ver21]. precluded [PJPD11]. preconditions [Dew09c]. preconditions/postconditions [Dew09c]. Predictable [LVM90]. Predicting [Boe99]. Predictive [LWF91]. preemptive [Bur01]. Preface [Ano91d]. Preliminary [Ano92f, Ano02a, Ano02e, PWDD80, Cro95]. premature [WBCS13]. Preprocessor [Bak90a]. presentation [Bal99, Lis12]. price [Fav91]. primitive [Dri91b, Dri91e, ISO91b]. principles [HEUV99]. Priorities [Ano06c, MD90, BW97a, MSM^{+}03, RW99, RLC01]. Prioritized [Els90a]. Prioritizing [GH99, GG99]. Priority [Alv87, Bri94, Bur87a, CS87, GS88, LMP90, Lev88, Lev11a, LSR^{+}88, MD90, Nae05, RSC16, AdlPT97, Sri06b, CR07]. PRISM [Wel97b]. Privacy [Car96]. Private [Bak91b, Bak93a, Gar84, Bei92, Gon91a]. Problem [Age85, Ano92j, Bel82, BW90c, CM90e, CM90g, Fuj87, SS89, SS97, WKT84, WQ83, bY93, BW90a, WGA90b]. Problems [Als83, Bak90c, LV87, Paz90, VMNM85, de 88, Bar09a, JR10, LS98, RK99, RSZ96]. procedure [GH99, GG99]. Procedures [Off87]. Proceedings [ACM82, ACM91a, ACM91b, ACM97, Ano93a, Ano02d, STF98, BHL^{+}93, ACM80, Bar87, Obe94]. Process [Dow94, Mog91, MNG16, SYW85, Con97b, Cro95, WRL13, Dob01a, Sil98]. Processes [Ves89, Fer97]. Processing [BBH80, Cra98, GPZdlP21, Jam98b, McC07, McC09, PL07]. Processor [SK22, FSS87, Nae05, Rui10, SC06]. Processors [Cha82, VRH21, MMP13a, WB07a]. producing [Con03a]. product [BB85, SAH01, WW01]. Productive [CBW^{+}21]. products [Ker98, Rom88]. products-updates [Ker98]. Profession [Ber86a]. Profile [Car17, DB98, GZdlP15, Hug23, RRG15, ARPT18, AdlP01, BB02, Bur13a, BV13, BWM13, Dob00, Dob01b, DdlP03, GLZdlP16, Gre13, LA99, MPV10, Mic01, Ros11b, TGH13, Tok03, VC01, Var03, Wel01, BE02, Bur99a, Bur99b, BDV04, DR99, Mic02, RdlPZFM01]. Profiles [ARG18, VR16, BBV97]. Program [Als83, Ano02a, BYY86, Bon84, DGLM85, Fri87, Gor83, Hil22, KF98, Lei12b, Lin82, Lin83, NS85, RS91, Ric20, Ala13, Edg01, Gar09, HS98, KSD12, Kle21, KK03, LSP01, LT99, Plo92, Sch10a, SC04a, SB05, WBCS13, Gri95]. Programmable [Cas20]. Programmed [Bur85b, Faß01]. programmer [Ker99]. programmers [MK91]. Programming [ACM80, Alv87, Ano00d, Bak91b, Bru17, BW89, BQ90, BW07a, CBW^{+}21, Cho19, Coh82, Col89, DF84, DeL88a, DGBMCG97, DoD87a, Dru82, FG82, GD00, GBCGDBC97, Hai00, HMZ00, HG07, Har22, HL86, Hou83, HSW87, Jha90, KFS97, Leb82, Lis12, MB91, Mic13, Mic16, NMT92, PDG83, PVF01, PV18, Rog09e, Rou85, Sac89, Sch87a, SHR82, SCD^{+}85, Ste12, Tok15, Wau83, WBCS13, Whi97, XRL^{+}88, AP11, AC04, Ano10b, Bag98, Bak91a, Bar09g, BMT^{+}14, BGGS14, Buh85, BWK^{+}01, CC98, Car94, DeL88b, Els91, FNS^{+}85, Gol93, HCW04, Joh93, MMP13a, NKN93, NM92, Och09f, Pan12c, Pan12d, Pan12e, Pan12a, PC05, Rog12a, Rog12b, San03a, Sei91, Sei92, SV99, Taf12, Taf13a, TMPM14, TP09, TT02, Ton99, WdlP97, WJS^{+}02, Wic98, dlPRGB99]. Programs [AG88, BHN20, Bur87b, CAU88, Col87, Cor83, CDM87, DB98, Fan84, GS85, HvKPT87, JEKC89, Kam83, KR88, KBL80, LSH98, LBO84, LP80, Men87, Mic16, Moy17e, MP89, NWW82, Pau87, Pyl84, SGJP89, Tai86, Tic82, VMNM85, WGC17, AID05, AD03, BW99, CM90d, Dob01b, Ehr94, EGC13, EKPPR04, GB94, GG87, HM03, Lau07, Lei12a, Mar99, RR14, San89, Taf13b, TNGC05]. Project [BGK^{+}82, FMG90, KMS82, OP85a, OP85b, Pie85, Plo84, Spu86, Ter87, BF86, Bow92, BTB^{+}10, Fre86a, Mat91, BCB^{+}22, BSPK22, Con97a, Con98, Fal91, JBT^{+}22, Kan12b]. project-wide [Bow92]. Projects [Bra82, AW91, Gri98, Moo93]. Promote [BBB97]. pronounce [LM94]. Proof [PD82, Mah13, Mau07]. Propagation [BS01, NDP97, NDP00, NDM98, NDP99, San01a]. proper [Fir87a]. Properties [Moy17e, EKPPR04]. Proposal [ARPT18, Cla87c, KS84, DV01, WJS^{+}01]. proposals [Mic13]. Proposed [Cra95, Dri91a, Dri91b, FG82, Hod91a, ISO91a, ISO91b, Sal92, Squ91a, Dri91c, Dri91d, Dri91e, Hod91b, Squ91b]. Protected [Bak90d, Jam98a, KPP97, Kam91, KW98, Led95a, LXY98, MM98, RCWB02, San00, Wre92, Bos13, BD92, GZdlP18, Led95b, LMV93, Nae05, WJS^{+}01, WJS^{+}02].

Protecting [DG97]. Protection [Riv17]. Protection/Isolation [Riv17]. Protocol [BW16c, CR18, GS88, LSRM12, LG88, ZBW07, ABGH13, BW16b, CR07]. protocols [BW13c, GZdlP18, WP13]. Prototype [CSA^{+}87, LRS09, LZL03]. Prototypes [KBT84]. Prototyping [MK83, Vas91]. proud [Woo99]. Provably [MGB^{+}23]. Provide [LL88]. Provided [KPP97]. Providing [Whi10]. Proving [MD22, Lei12b, Taf13b]. PSP [Sil98]. Pthreads [Paz90]. Public [Con97b, Con97d]. publications [Rom86, Rom88]. Publisher [KS06]. purpose [FC91]. Purposes [Pag82]. putting [Cha07a]. pyramids [Oli94]. Python [Bri12b, Bri12c, Mar21, Ver22].

qualifiable [San03b]. Quality

[Ano93f, BD91, Mar19, Mol83, SCC22, ACP11a, ACP11b, Med91, Rad94]. Quantitative [Rei87]. Quasar3 [EKPPR04]. queries [LSP01]. questions [Col95a, CR97, Mat96]. Queues [Huf82, BW02]. Queuing [VRH21, KPPÉR06]. Quick [Smi84]. Quicksort [Coh82]. Quiz [Reb17a, Reb17b, Och11].

R [Roa88]. R1000 [Wil87]. Radar

[HDHH98]. radio [LSRM12]. railroading [McC99]. Railway [BCMC23]. Raleigh [Fis83]. Ramifications [Qui90d]. Random [HB88]. range [ACP11a, ACP11b]. Rapid [KBT84, Vas91, CM98]. Rapporteur [MSW98a, MSW98b]. rate [Cro95, Ear92]. Rational [Ano92k, Wil87]. Rationale [Dri91d, Dri91e, GES89, Hod91b, Squ91b, Wei89, CM90d, Taf97]. RAVEN [BE02]. Ravenscar [BDV04, AdlP01, AD03, ABW01, AW01, BE02, Bur99a, Bur99b, BB02, Bur13a, BWM13, CC18, Car17, DB98, DR99, Dob00, Dob01b, DdlP03, GZdlP15, GLZdlP16, Gre13, Hug23, LA99, MMB^{+}03, MPV10, Mic01, Mic02, MMP13b, PV13, PV02, RSC18, RM18, RRG15, RdlPZFM01, Rui10, Sri06d, TGH13, UZ07, VC01, Var03, Wel01, ZdlP02, dlPZR^{+}01, dlPZ03]. Ravenscar-EDF [CC18]. RCLAda [Mos20]. RDBMS [LM94, Vok92]. Re [BT88a, BT88b, Qui90d, Rob92, SC04b, LRS09]. re-ADA [LRS09]. Re-engineering [SC04b]. Re-Export [BT88a, BT88b]. Re-introducing [Qui90d]. Re-usable [Rob92]. Reaction [Cra97]. Reactive [Che91b, WBCS13]. Reactor [CCC21]. readability [Car97]. reader [Plo98]. Readers [Lev01a, SS89]. Readers-Writers [SS89]. Real [All87, Alv87, Ano88b, Ano90c, Ano90d, Ano91c, Ano93h, Ano93k, Ano97, Ano00i, Ano02d, Ard87, Bak87a, BM85, Bar87, BA90a, BTP22, BdlPZ10, Bri94, BD01, BW90a, BW15, Chr87a, CSL^{+}87, DB98, Fan84, Fri87, Gal20, Goo90, HSW87, Mac80, McC87a, MMP13a, MMPT16, Nil12a, Pau87, PS84, PMMT15, PRQ21, PR90, San03a, SW87, SCC22, Taf91a, Wei90a, Wel90, Wic82, de 87, dlPRGB99, AH01, ABW95, Ad93, AdlPT97, AAAG21, BTVC99, BCF94, Bos13, Bri92a, Bri92b, Bro88, BHR02, BH02, Buh85, BKW^{+}94, BW92, BW93b, BW94, CS91, Chr87b, Col99b, DV01, Ear92, Fer97, GH01, GB94, GHV03, GDAG97, GdlP02, GDHM02, HMRF97, Har99a, HP01, HMC88, Hod91a, Hod91b, HM03, LN91, LSRM12, LG88, LVM90, LT99, Mac86, MMB^{+}03, McC99, McC07]. real [McC09, McC10, MS11, Moo97, MKK99, MP91, New95, New99, Pan12c, Pan12d, Pan12e, Pan12a, Pet10, PV98, PV99b, PV99a, PV02, Pot04, RH01, Rog09a, Rog11d, Rui13, Sel99, SLNM04, Sin07, Taf91b, TGH10, TPG21, UKDH97, UPRZ07, VGD^{+}97, WWB99, WD93, WdlP97, Wel03, WB07b, Whi10, Wre92, ZEdlP13, ZdlP13, Ano93b, ACWB89, Bar88, BKWS88, Bur87b, BW87, BW90c, Col87, Dob01a, Dom87, GB87, LD87, Mea87, Rog09e, VMNM85, de 87]. Real-Time [All87, Alv87, Ano88b, Ano90c, Ano90d, Ano91c, Ano93h, Ano93k, Ano97, Ano00i, Ard87, Bak87a, Bar87, BA90a, BTP22, Bri94, BW15, Chr87a, CSL^{+}87, DB98, HSW87, Mac80, McC87a, MR10, Pau87, PMMT15, PR90, SW87, SCC22, Taf91a, Wei90a, de 87, BdlPZ10, BD01, BW90a, Gal20, Goo90, MMP13a, MMPT16, Nil12a, PRQ21, San03a, Wel90, dlPRGB99, AH01, ABW95, Ad93, AdlPT97, AAAG21, BTVC99, Bos13, Bri92a, Bri92b, Bro88, BHR02, BH02, Buh85, BKW^{+}94, BW92, BW93b, BW94, CS91, Chr87b, Col99b, DV01, Ear92, Fer97, GH01, GB94, GHV03, GDAG97, GdlP02, GDHM02, HMRF97, Har99a, HP01, HMC88, LN91, LSRM12, LG88, LVM90, LT99, McC99, McC07, McC09, McC10, MS11, Moo97, MKK99, MP91, New95, New99, Pan12c, Pan12d, Pan12e]. real-time [Pan12a, Pet10, PV98, PV99b, PV99a, PV02, Pot04, RC10b, RH01, Rog09a, Rog11d, Rui13, SRC13a, Sel99, Taf91b, TGH10, TPG21, UKDH97, UPRZ07, VGD^{+}97, WD93, WdlP97, Wel03, WB07b, Whi10, Wre92, ZEdlP13, ZdlP13, Ano93b, ACWB89, Bar88, BKWS88, Bur87b, BW87, BW90c, Col87, Dob01a, Dom87, GB87, LD87, Mea87, Rog09e, VMNM85, de 87]. Reality [Cra82a]. realized [Lew02]. really [Mor95a]. Realtime [MWM10, DRF97]. reasoning [Lau07]. Reasons [Men88]. reckoning [EF01]. Reclamation [Lef87, Men87]. Recognition [SN94, GSP^{+}11]. Recommendation [Har88, Vau98]. Recommendations [CMR90, Ano89a, Cra97, Taf97]. recommended [ML91]. Reconsidered [Lev91, Pau93]. record [And05, Coh94, Mar99]. records [Bak90d, Kam91, LMV93]. recovery [Nyb05]. Recursion [Mor95b, Moo11]. Reddo [DA13]. Redefinition [Rob92].

Redistribution [Jam99]. Reducing [HEUV99, Maz89b]. Reduction [TMPM16]. redundancy [Due97]. redundant [Gar09, Sri06d]. Reengineering [BHD98, Faß01]. Refactoring [PS06, And04]. Reference [Bak93a, Fag00b, Smi84, Ber86b, Bri12d, Bri12e, Bri12a, Pen91]. references [Bri12a]. Refinement [HCBM98b, dNW23, KPPÉR06]. Reflection [Gal20]. Reflections [BDS81, Var03]. register [Mah11, Mah12a]. rehabilitated [Bak91a]. Rehost [WD93]. rehosting [Cle86]. Reimplementing [VGD^{+}97]. Rejuvenation [LV23]. Related [Bak90c, Bak91c, Bar09a, FG86]. Relating [Bur92]. Relational [McC87b, PVV85, DCC85]. relationship [Lei02]. Relationships [MSW85, Bal95b]. relaxed [Yav85]. Relaxing [Bei92]. Reliability [KPP97, LBO84, Sac89, Gil99b, Ros10]. Reliable [Ano99i, LC22, BC11, BWK^{+}01, BWM13, Sch09]. religion [Syi95]. remote [GH99, GG99, WGA90b]. Rendezvous [EHP80, Gil92a, Gil92b, Gil92c, Gil93a, Gil93b, Gil93c, Gil93d, Gil94a, Gil94b, JA82, MM98, PD82, RB85, LVM90, LW97, SM92]. Replacement [Tin90]. Replacing [LMV93]. Replay [NPT97]. Replica [PV99a]. replicAda [DGBMCG97]. Replication [Tro20, Wol99]. Report [Ano92g, Ano92h, Ano92j, Ano92i, Ano93a, Ano93e, Ano93g, Ano93i, Ano99l, Bar85a, Bel80, BWV03, BV03, DV01, Fis83, GHV03, GMO92, HvKPT87, McC06b, Moo85, Mun91b, Off88c, Puk88, RC01, Tas88, WV02, Bar98, Boy86, Bro88, Bro96, Edg01, GS02, KGW^{+}85, Kam98, MSM^{+}03, Off88b, PW01, Sch10a, Sch10b, Sol91b, BRC98, Off88a]. Reporting [Gau90b, GR90, DR99]. Reports [Tok15]. Repositories [Ano92l]. repository [Gic91]. Representation [HLRS80, Nyb87, Sol91a, Taf82, Coh94, Dew09a, Dew09b, Mar99, Sol91b]. Reproducing [Lom83, Lav95]. request [Mah12b]. Requests [Bur87a, Gau95]. requeue [VE92, WB07c]. requirement [Bur13b]. Requirements [BA90a, BYY86, FMG90, GG16, MNG16, SSGH^{+}22, Wag20, Wei90a, Wel90, Bai10, Car99b, Fir91a, NBZ^{+}20, Shu93, SLNM05]. Research [Ano00d, Sch87a, WV98, Bal14]. Reselect [LCN91]. Reserved [Tro06, Wol84]. Resilience [Tro20, Whe19, ZDM22]. Resilience-Aware [ZDM22]. Resiliency [Bod19]. Resolute [GBC^{+}14]. Resolution [Bel80, FG86, Lev01a, MF91, PC90, Duf09e, PG94]. Resource [KPP97, San97, WKT84, Bak93c, LWB13, LCB09, WP13]. resources [Lev11a]. Response [Ada88, Bak92, Che91b, Mah12b, Off88a, ZdlP02]. Responses [Ree88]. Responsibility [KQT^{+}21]. restated [LRS09]. Restricted [BW97b, SB99]. restriction [Sri06d]. restrictions [UZ07]. restructuring [BR94]. result [BA98]. Results [CC18, Gau90a, Gau90b, GR90, PG91, Roy90b, LW07]. Retargeting [Cle86]. Rethinking [Rym98]. retrospective [Sch09]. Return [Ano17b]. Reusability [JLM^{+}85, PDN97, Fav91, KB97b]. Reusable [Ad93, Car90, Car91, Dau87, Dun98, Fai94, FMS98, GES89, Lev90, Lev92a, Lev92b, Lev93a, Lev93b, Lev93c, Lev93d, Lev93e, Lev94a, Lev94b, Lev94c, Lev95a, Lev95b, Lev95c, Lev95d, Lev96a, Lev96b, Lev97b, Lev97c, Lev98b, Lev98c, Lev99a, Lev99b, Lev00, Lev01b, Lev02a, Lev02b, Lev04, Lev05d, Lev05b, Lev05c, Lev06, Lev08, Lev09b, Lev10, Lev11b, Lev11c, Lev13, Lev15a, LM83a, LM83b, MK87, SSFO86, Yu97, dB99, Car92, Car04, HMC88, Mac96, SU91, Vok92, dB97a, dB97b]. Reuse [BBB97, Lat91, MDPK94, Moo94, SS94, AdB90, BBB98, Bow92, Con97b, FC91, Hir94a, Hir94b, PB98, RH91, Sol91b, Wad92, Yu98, BBB97, PB98, Ano92a, Con98]. ReUSE/Ada [BBB97]. ReUse/Web [PB98]. Reuse System [Gic91]. reversal [And05]. Reverse [YQZ^{+}23, Wel97b]. Review [Led92, Orb85, Rog97, Rog09e, Rog11d, DeW86, Obe85]. Reviews [Har97]. Revising [Gre16]. Revision [Bru17, Ano10b, FG86]. revisited [Hek89]. Revisiting [BP13]. Right [McC00, WB10b]. Rigorous [Taf22a]. rise [Swa10]. Risk [DM91]. road [MS04]. Roberts [KM81]. robin [Sri06b]. Robot [BM23, GDAG97, HMRF97, Mos20]. Robotics [Gre21, FME01, Men09]. Robots [Cra98, Men09, ML95b]. robust [Kir12]. Role [Boy89, PS84, LT99]. ROLM [Ell83]. Room [Cra22]. ROS [SB23]. rotate [Cha11]. round [Sri06b]. route [OWSB08]. Routines [Bur85b]. routing [Gan03]. RT [Dob01a]. RT-Java [Dob01a]. RTEMS [CSSW10]. RTSJ [Wel03, WT03]. Rules [Bac84, Wei89, Bar95]. Run [All87, Ano93c, Ano93a, Ano94d, CU89, DM91, FG82, Bur13a, CAC^{+}13, EK12, KGW^{+}85, LHBK87, ML95b, RC10a, BHL^{+}93]. Run-Time [All87, Ano93a, CU89, FG82, DM91, Bur13a, CAC^{+}13, EK12, KGW^{+}85, LHBK87, RC10a, BHL^{+}93]. Runtime [ACM87b, ACM89, Ano92c, Ano92d, Bak87a, BTP22, Fal82, HL85a, HL85b, HLRS80, Kam83, LV87, MD22, RB85, Riv17, Ros87d, AKM^{+}91, Ano87, Ano88a, Ano89c]. Russia [Ryb94]. Rust [Har22, MK14].

S [Klu87]. S3 [VGGS20]. SA [Bro91, Hir94c]. SA/OOD [Hir94c]. SA/SD [Bro91]. SA1 [Bar07a]. SA2 [Bro07]. Safe [Bak93c, Gre99b, TMPM14, Bar09b, Bar09c, Bar09d, Bar09e, Bar09f, Bar09g, Bar09h, Bar09i, Bar09j, Bar09k, Bar09l, Bar09m, BMT^{+}14, Cro14, DRF97, Mar99, Men09, Moo11, Taf13a, Wic93]. SAFECode [Bai20]. SafeProver [EJ16].

Safety [Ano93a, AL00, BMGS20, LFT12, LV23, MGF16, MSW98a, Pro20, Sie21, WCB16, YQZ^{+}23, BMT^{+}14, Bri12e, Bro07, Bro11, BHL^{+}93, Car99b, CH04, Col99b, Gar09, LHFD13, Mar19, MSW98b, Nil12b, Rog11a, San03a, SG06, Taf13b, dlPP02]. Safety-Critical [LV23, WCB16, YQZ^{+}23, MGF16, Bro07, Car99b, Col99b, LHFD13]. SafetyChip [NAF05]. Saga [BM85]. Sample [Ano92j]. Satisfiability [Bjo13]. SAVI [WRL13]. Saving [LP85]. SAW [CFH^{+}13]. SC24 [Puk88]. Scale [Gal20, SC87]. scaling [Wha13]. Scanning [Tis83, Gau96]. Scenario [VGGS20]. Scenario-based [VGGS20]. schedulability [AAAG21, GDHM02, LSRM12]. Scheduled [RSC16, AAAG21]. Scheduler [Taf22b, Ear92, LP06]. schedulers [SP07]. Scheduling [CHHB90a, CHHB90b, Coh88, CSL^{+}87, Elr88, LL88, LV87, Loc91, MD90, McC87a, RSC16, RSC18, RM18, RK99, SLNM05, ZDM22, de 88, AH01, Asp01, BWV03, BW03, GB94, HHBC90, RH01, RH02, RH03, SRC13b, SC13, SLNM04, Sin07, Sri06c, TG09, TPG21, WV02, WT03, WB10a]. scheduling/dispatching [Asp01]. Schemata [Bak86]. Scheme [The90]. Schemes [Ano17c, GS85]. Schizophrenic [BPP06]. Science [Ada88, Ano99f, MH98, Off88a, Off88b, Off88c, CC98, FME01, LC86, SBH^{+}98, Toa96]. Sciences [OW82]. Scientific [LL98, Whi97, Mac96]. SCOPE [Gar09, NS85, Rog11b]. script [Abb96]. scripting [Bri09b, Bri09c]. SD [Bro91]. SDSAWG [GMO92, Ano92i, Ano93g, Fir86]. Search [BM85, WT89, Bri09a, WT88]. searching [Hea08a]. SEATECS [Mye85]. Second [Bar88, Obe85, Obe94, Orb85, Ano88b]. section [Bra98]. sector [Gil99b]. Secure [MGB^{+}23, Bar09b, Bar09c, Bar09d, Bar09e, Bar09f, Bar09g, Bar09h, Bar09i, Bar09j, Bar09k, Bar09l, Bar09m]. Security [BCH^{+}19, BMGS20, Cas20, Cho19, FOSC23, ICS22, Pro20, Taf20, CH04, Cha07b, Dav04, HSWP12, KNB08, Mar19, MSW98b, Moy11c, Moy11d, RDS98, Saï08]. see [Dew07a, BMW94, Pen91]. SEI [Fel86, Rob86]. Select [The90]. Select-And [The90]. Selected [Taf97]. Selection [NW83, NW^{+}84, TR87]. Selective [LMP90, LCN91]. Self [Fuj87, Lom83, RLPD98, Gan04, Lav95]. Self-Intersection [RLPD98]. Self-Organizing [Fuj87, Gan04]. Self-Reproducing [Lom83, Lav95]. SEMANOL [BBH80]. Semantic [Ano94a, Col95a, SB80, Vla93, Vla94, vHLKBO85, CR97, RT09]. Semantics [KMS82, Li82, CAC^{+}13, Goo90, Lar14, RLC01]. Semaphores [bY94, Rog11c]. sensor [BC95]. separate [Khr95]. September [Off88c]. Sequence [FHN83]. Sequencing [HL85c]. Sequential [Moo18, KP86b, KP86a]. Server [Ano95k, CS87, Obr09, Obr12a, Qui11a, Ano95l]. servers [BW07a]. Service [BS13, KPP97, Swa09b, SB11, SB12, Lev09a, Swa07a, Swa07b]. Service-oriented [BS13, SB11, SB12, Swa07a, Swa07b]. services [AH01, PQT99, RH01, Swa07a, ZEdlP13]. Serving [LXY98]. Session [ARG18, Asp01, BH02, BB02, BV13, BW13c, BdlP15, BW16c, CR18, DdlP03, GdlP02, GP18, HP01, MdlP16, PMM13b, PMM15, PM16, PV18, RR13, RdlP13, RR16, RM18, RH16, TB02, TD03, VP03, VHP10, VW13, VR16, VW18, WT03, WP13, WR15, dlPP02, dlPM13, lPB18, BBV97, Bur99b, BWV03, BV03, BW10b, DV01, GLV97, Gil99b, GHV03, Har99a, HBTW99, Kam99, PK97, WdlP97, Wel99, Wel01, WV02, Dob01a]. Set [MP89, Hea08a, MP91, San89]. SETA1 [LWF91, MKP91b, Taf91b]. SETA2 [Obe94, BP94, Dow94, MDPK94]. Sets [Lar22, RSC16, SGW90a]. setting [SRC13b, SC13]. seventeenth [LC86]. Seventh [Ano93h]. Shared [Els90b]. Sharing [San97, LWB13, Mar05]. Sheet [Smi84]. SHell [Wes97a, Wes97b]. shift [Cha11]. Ship [KS01]. Shoreham [STF98]. shortcuts [Bri11b]. shots [MC05]. Should [CS87, Ker82, BBPT12, Con97d, Taf06]. sic [JF98b, ML99]. side [SC01]. side-by-side [SC01]. sides [Sma09]. Sieve [And88, Col98, Dri89a, Dri89b, Hek89]. SIG [Whi85]. SIGAda [Ano93c, Ano93a, Ano95m, MH20, STF98, ACM87a, ACM91b, Ano92f, Ano92i, Ano93g, Ano93i, Ano93j, Ano94e, Ano94f, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano95l, Ano99h, Ano99j, Ano99k, Ano00h, Ano00k, Ano00r, Ano00s, Ano00t, Ano00u, Ano00v, Ano00w, Ano00x, Ano01b, Ano02b, Ano02e, Ano06f, Bar85a, GMO92, Gri95, Har94c, Har99b, Har00, Har01, Lei99b, Lei00, Lei02, McC06a, McC06b, RH96, RC01, Ano02c, Col90, Ano94g]. SIGAda’98 [Ano99l]. SIGCSE [LC86]. Signal [GPZdlP21, Gar09, PL07]. Signaling [BA90b, Lev91]. Signals [Moo18]. SIGPLAN [ACM80]. SIMD [SK22]. Simple [AP84, FGN85, Gic90, SJ91, Hof86, LHFD13, Qui11a, SP12, WBCS13, Yav85]. SimpleGraphics [MKK99]. Simplest [Age85]. Simplicity [Sie21]. Simplified [Hir94c, SGJP89]. simulate [DPP^{+}09]. Simulating [Per88]. Simulation [AS87, Bru82, Buz16, CCC23, MG87, SC87, Sho87, Abb96, Gan01, MMN09, Mah13, WD93, HDHH98]. simulations [PL07]. simulator [Bro03, ML95b, SC06]. Simulink [XWZ^{+}23]. Simulink/Stateflow [XWZ^{+}23]. Singhoff [Rog11d]. single [HR03]. situated [LS98]. situational [SG06]. Sixth [Ano92k]. size [AAAG21]. skeletons [NLA05]. slicer [SC04a, SB05]. SlowSort [Con90]. Small [BA90a, Bun85, ARPT18]. Smallest [Lar22]. Smalltalk [BMW94]. smart [Och12a, Och12b, DRF97]. SMP [KK03, WB07a]. SMT [Lar22]. SOA [BS13, Swa07a, Swa09b, SB12, SB11]. SOAP [Obr12a, Obr12b]. SOAP/WSDL [Obr12a, Obr12b]. Soaring [Bak91b]. societies [Sot06]. Socket [Cri01]. Socket-Based [Cri01]. Software [ACM91b, Ada88, Ano92a, Ano92b, Ano92i, Ano92l, Ano93a, Ano93g, Ano99a, Ano99i, Ano00d, AC85, BM85, BT88a, BT88b, BGK^{+}82, BCG^{+}84, Ben94, Ber86a, BRW97, Car89a, Cra82a, Eme83, Fal91, FAT^{+}23, FOSC23, FMN80, Fra87a, Fri83, Gar83, Gib00, Gon90, GMO92, Har22, Har82, Har97, JLM^{+}85, KB97b, Lev92b, Lev93b, Lev93c, Lev93e, Lev94b, Lev99a, Lev00, Lev01b, Lev02a, Lev10, Lev15a, Lew02, LV23, LNR87, MK83, McC00, McD88b, MP98, Moo94, NAT20, PJPD11, RH91, RDP97, Rob92, Sch87b, SSJ85, SS87, Sil98, SSFO86, Tem84, Ter87, Ver21, Ver22, Wil91, WL98, YQZ^{+}23, vdL84, ACP11a, ACP11b, Ame01, Ano89a, AdB90, Bar09b, Bar09c, Bar09d, Bar09e, Bar09f, Bar09g, Bar09h, Bar09i, Bar09j, Bar09k, Bar09l, Bar09m, Bar08, BGGS14, Boe99, Bro07, BC11]. software [BHL^{+}93, BTB^{+}10, Buz16, Car99b, Car88a, Car88b, CFH^{+}13, Cha13, Cha07a, Che92, Col99b, Con97b, Dav05, DA13, Edg01, Fai94, FBL^{+}10, FC91, Fre86b, Gic91, Gil99b, HB96, HS98, HCBM98a, HEUV99, Irw96, Jar07, Jen09, Lan10, LW07, LFT12, Lev90, Lev92a, Lev93a, Lev93d, Lev94a, Lev94c, Lev95a, Lev95b, Lev95c, Lev95d, Lev96a, Lev96b, Lev97b, Lev97c, Lev98b, Lev98c, Lev99b, Lev02b, Lev04, Lev05d, Lev05b, Lev05c, Lev06, Lev08, Lev09b, Lev11b, Lev11c, Lev13, LSRM12, Mar19, McC99, Mic02, MY98, MP91, OS12, Off88b, Off88c, Pet10, Pul95, Rad94, San12, San01b, SS91, SBH^{+}98, Sny91, SG06, SVK^{+}14, Taf01b, Ven08, Wan99, Yu98, Fis83, Mye85, Off88a, SS94, Tas88]. Software-Based [FOSC23]. software-in-the-loop [Buz16]. software-intensive [Mar19]. Solution [Age85, Dob90, Hir94c, bY93, And88, Shu93, WGA90b]. solutions [BCF94, Col98]. solve [Bar09a]. Solving [LS98, SS97]. SOM [CN96]. Some [Bak90c, Hek83, VMNM85, Led95a]. Songbook [Ano91b]. Soundness [LKSL19]. Source [AGG^{+}80, Wal85a, WB89, Bar08, Bri09d, Gar09, Kle21, Con97a]. Source-to-Source [AGG^{+}80]. Sources [Ano17b]. SP1 [Bar07b]. SP2 [Swa07a]. Space [CM90e, JARKS22, SK22, Tok03, VC01]. Spacecraft [BC16, Trü95]. spaceport [Bar14]. SPAIDS [RDP97]. Spare [Reb17a]. SPARK [Ano10a, Bar00, Bar09m, BHR^{+}11, BC16, Cha00, Cha11, CAC^{+}13, CHGH19, Cro14, Dro22, EH13, Gre21, HG14, Hum22, Jen09, Lau07, LW07, LCB09, Moy11a, Moy11b, NAT20, PJPD11, Rog21, Ruo05, Sau05, SB05, Taf13a, Taf20]. SPARK.Specific [Ano10a]. speaks [DFGZ09]. SpeAR [WG20, Wag20]. Special [Ano93a, CM90a, McC06b, Bra98, WGA90a]. specialised [dlPRGB99]. specific [Jac13, Nyb10a, Sri06b]. Specification [Ano94a, BH14, BG90, Col95a, Dro22, Fle86, LNR87, NW83, NW^{+}84, PDV98, Vla93, Vla94, Wag20, vHLKBO85, BHR02, BH02, CR97, Dob01a, Lar14, Log13a, Sol91b, Taf11]. Specifications [BCH^{+}19, HB96, Puk93]. Specifying [BKC91, Che91b, Moy17e, Pyl84]. Spectroscopy [CA89]. speed [DB09]. speeding [MRB06]. speedy [Cha11]. SPERBER [Plo84]. sponsored [Hir92]. Sporadic [ABW95, BW94]. Spot [BGGS14]. SQL [BST90, Bry88, DD87, Lop99, Moo91]. SQL ArmAda [BST90]. St. [ACM97]. stable [KS01]. Stack [Car17, Moo11, Och12c]. Stacks [LV23].

Stand [Pow90]. Stand-alone [Pow90]. Standard [Ano99d, KS84, MF04, Rob92, Ros86b, Sal92, Smi84, Bro11, Bur90, Dri91c, Dri91a, Dri91b, Dri91d, Dri91e, Hod91a, Hod91b, ISO91a, ISO91b, Moo96, Ros86a, Spi00, Squ91a, Squ91b, Squ91c, CHGH19, The90]. standard-missile [Spi00]. standardization [Moo98]. Standardized [Gic90, Mat96]. Standards [Ano92i, Ano93g, DF84, Van86, BA07, Ros11a, GMO92]. STAR [Zhu90]. Started [Rez22]. startup [Bar09j]. State [HPT81, San00, Bal99, DG97]. Stateflow [XWZ^{+}23]. Statement [LCN91, The90, GL89, Mor95a, RH10]. Statements [Bak86, Reb17b, CXY01]. States [Gri98]. Static [AD03, AC04, And20, Bla07, CBW94, Ehr94, KNB08, Mar21, PR98, Bar08, Dew07b, GG87, JR10, Saï08, Ven08]. Statistics [ZW83]. Status [Ano93e, Wel01, DdlP03, MB08, WJS^{+}01]. STD [Buc87, FG86, GG87, RM88, Roa88, Ros86b, Ros86a, Roa89]. Steal [Bak93a]. Stealing [Taf22b, Taf12]. Steelman [Whe97]. Stein [DeW86]. Stephe [Lea04]. steps [Bis88, TPG21]. Stereo [RLPD98]. Stereo-lithography [RLPD98]. Stimulus [Che91b]. Stimulus-Response [Che91b]. STL [Hea04]. Storage [GS85, KT87, Men87]. Store [Kim21]. Strategies [Bak93b, Hil82, Wil85]. strategy [OWSB08, RSZ96]. stream [Rog09d, WA07]. Streams [Cri01, PW97]. strength [AC03]. strengths [SHT^{+}23]. String [Car89b, WT89, OWSB08, WT88]. Strings [SGW90b, Bak93b]. Strong [BYY86]. Strongly [Sal92]. Structure [Bec83, Cam92, DCBM97, JF98b, Moo94, Mos22, Win84, BL86, GG87, JF98a]. Structured [Bak86, Bak91b, Fir91b, KBT84, Pri82, Shu91, Wel85]. Structures [Cel97, Dau87, Dun98]. Studies [HF84, HHR^{+}86]. studio [CH06]. Study [BCMC23, Dob83, HvKPT87, JF98b, KPP97, MP84, NAT20, Rog21, Shu87, Tra89, Cle86, DPB^{+}97, Fav91, Fre86b, JF98a, KPPÉR06, KB97a, LVM90, Sch91, Sum87, Wad92, Wek90]. Style [SJ91, ER86, HHR^{+}86, Khr95]. subclasses [DG97]. Subgroup [Mun91a, Sol91a, Sol91b]. subject [Hof86]. Sublanguages [BCD83]. subset [Hir94a, Hir94b, San03b, Taf13a]. Subunits [Bur92]. successful [Spi00]. such [BB02]. Suggested [Dob90]. Suggestions [WA07]. Suitability [Yem82]. Suite [PC90, RS91, Pri01, Tan91b]. Summary [ARG18, Ano93k, Bro82, BW93b, BdlP15, BW16c, CR18, Eme83, Gil92a, Gil92b, Gil92c, Gil93a, Gil93b, Gil93c, Gil93d, Gil94a, Gil94b, GP18, Kam95, LWF91, MdlP16, PMM15, PM16, PV18, RR16, RM18, RH16, SPS88, VR16, VW18, WR15, dlPU07, lPB18, Ben94, BMT^{+}14, Bro88, BH02, BP94, BBV97, Bur99b, BB02, BW10b, BV13, BW13c, Dow94, GLV97, Har99a, HP01, Kam99, MDPK94, PK97, Pen91, PMM13b, RR13, RdlP13, Rob86, Sof88, SHT^{+}23, TB02, TD03, VP03, VHP10, VW13, Wal94, WdlP97, Wel99, Wel01, WT03, WP13, dlPP02, dlPM13, Dob01a]. Summer [ACM91b, Ano92f, Ano95m]. summit [Bla07]. Sun [Dob01a]. Sunday [Ano99l]. Supervisor [Fal82, RB85]. Supervisors [Ros87d]. Support [Bak87a, BOM97, Bra82, BKC91, BW13b, DGCR^{+}84, DeL88a, Dru82, Fai80, Gór20, Gre16, HCBM98b, Hou83, MB91, MR83, MK91, NDP00, Pie85, PR90, RSC18, RB85, RdlPZFM01, RSK^{+}19, SK22, TGH10, Wag85, Wel91, BPP06, BBB98, BW92, BW03, BWM13, CMWT21, CBB^{+}97, Cro90, DeL88b, GLZdlP16, Gre18, LYB^{+}10, PV98, PV02, RH07, SRC13a, Sri06c, Taf01a, WB10a]. Supporting [BW10c, Dun98, HSB^{+}22, HW88a, HW88b, JEKC89, AdB90, ER86, Gan03]. suppress [Dis09]. suppressed [EK12]. Surveillance [LT99]. Survey [Ano92l, AC85, Che91a, Lad89, Lin82, Lin83, Seb87, Gil99a]. Survivable [Cor83]. suspending [WGA90b]. SW [LKH16]. Swarm [SS20]. Swarms [SSB^{+}20, SS20]. Sweden [BRC98]. SweetAda [Gal22]. SWIM [Sch10a]. switches [SC06]. symbiotic [Lei02]. Symbol [Cra98]. Symbolic [dNW23, BHR^{+}11]. Symposium [ACM80, ACM91b, Ano91a, Obe94, BHL^{+}93, LC86, Ano93a, Moo85]. Symposium/Summer [ACM91b]. Synchronization [Bos12, dB99, Bal95a, Elr89, GSX99, dB97a]. synchronized [MSK05]. Synchronous [BW16a, Moo18]. Syntax [Gen91, Gra83, Leb82, Bar09c, Yav85]. SYNTAX ANALYSER G [Gen91]. Synthetic [HF84, Wei90a]. SysML [RD23]. System [ACM89, AB98, BHD98, BCMC23, CA89, Cor83, Deb83, FG82, Fri98a, Fuj87, Gil84, Jam98a, Kam83, Kie89, Lev82a, Lev82b, MMN09, MG87, MK91, NAT20, Nyb87, PGRZ92, PVV85, PF20, Pro20, Rud83, Sch87a, Sch87b, Tha82, Tok16, Whe86, Whe87, Whe19, Whi82, Wil87, WV98, WB89, ZW83, AID05, Ano89c, BBB98, BdlPZ10, BF99, Buh85, BKW^{+}94, CVW03, CM94, Cle86, Faß01, Fri98b, Goo13, HB96, KS01, Kle89, Lar14, LW07, LG88, LCB09, MMSN09, MWRH13, NKN93, OWSB08, OS12, Pot04, RH07, Ros10, SP12, Trü95, Aus22, Bra94, CN96, Leo85, Mos20, Nil12a]. system-critical [HB96]. system-level [MMSN09]. System-Oriented [Sch87b]. SystemAda [MMSN09, MMN09, Mah12b, Mah13]. systematically [Mar19]. SystemC [LKH16, Mah13]. Systems [Alv87, Ano99f, AL00, BKS87, BCB^{+}22, Bak87a, BSPK22, Bal97, BA90a, BDD^{+}82, Bod19, BTP22, BMGS20, Bri94, Bur85b, Che97, Che91b, CG88, Col87, DGBMCG97, DoD87b, FMS98, Gal20, GG16, Jan88, Kim21, KBT84, KQT^{+}21, KU84, Kni87, Kru90, Lan10, Mac80, MGB^{+}23, MGF16, Mea87, MMPT16, Mic16, Mye85, PM16, PR90, PR98, Rog09e, Ros87b, Rou85, Sac89, Sch87b, Taf91a, TCRW88, Tok15, TBA98, Wag85, Wal87, Wel97a, XWZ^{+}23, ZDM22, de 87, AH01, ABW95, AdlPT97, Ame01, AAAG21, AW01, Ber05, Boe99, Bri92a, Bri92b, BDV04, BW10b, CSSW09, CSSW10, CBB^{+}97, Dav04, DPP^{+}09, Dew06, DPB^{+}97, Fis12, Fus91, Gan04, GH99, GH01, Gar90, GLV97, Gid96, Glu09, GDHM02, GG99, HM91, IMM85, Kam95, KK03, LRS09, MM21, MVG99, Mar19, MDPK94, MCS97]. systems [Mic07, Moo97, Nae05, New95, PZ97a, PT99, Pet10, PV98, PV99b, PMM13b, Qui11a, Qui11b, Qui11c, Qui12, RH01, Rog09a, Ros87c, Ros11b, Rui10, RK99, Sau05, Sch09, Sel99, Swa09a, Taf91b, TP98, TPG21, UKDH97, UZ07, VGD^{+}97, WA07, WRL13, Wea10, Wel91, Wel03, WB07a, WBCS13, Wic98, ZdlP13].

T [DRF97]. T-SMART [DRF97]. Table [Tro06]. Tactical [Mye85]. Taft [The90]. Tailored [All87]. Tailoring [Wai98]. tainted [Moy11c]. tall [Puc17]. Taming [Pag82]. Tapestry [Con98]. Target [Ber84]. Targeting [CDG97, EJK89, Gan01]. Targets [AC85, DGCR^{+}84, Mid87, TR87]. TASH [Wes97a, Wes97b]. Task [Ada88, Ber15, BJRW96, BN87, BW03, BW16a, Che97, Cla87c, Coh88, CS87, Fal82, HPT81, HL85c, KVT88a, Lla92, LV87, Nie86, Off88a, Off88b, Off88c, RSC16, Sac89, Tas88, WBP97, Bri12e, DRF97, HR03, KVT88b, ML99, Che92]. task-safe [DRF97]. Tasking [Bak87b, Bak90b, BOM97, BN87, BW90d, BBV97, CAU88, Che90, Che91a, Cle82, Col98, DB98, DR99, Elr88, Fra87b, GHL82, Gon88, HL85a, Hil82, Lef87, LB80, MT01, Mur90, OB97, RB85, Ros87d, SB99, Shu87, Ste80, TNGC05, Ves89, Wel85, BW90b, BW97b, EGC13, Goo90, HL85b, Kie99, KR01a, LA99, Nyb07, Sum87, Tom97, WB07c, dB97b]. tasking-model [BW90b]. tasklet [PQR18]. Tasks [Ber15, CU89, Coh85, FCS83, GS88, Hek83, KPP97, LXY98, Lom83, Mal88, Pap89, Pie87, Qui90c, Rom00, San00, SN94, ABW95, BW94, FSS87, GB94, Lev97a, LVM90, LMV93, RT21, WB07a]. TASTE [BBB^{+}23]. Taxonomy [CM90f, PF20, SN88a, Fer97, Hou83, SN88b]. Tcl [MVG99, MKK99, Wes97a, Wes97b]. Tcl-Tk [MVG99]. Tcl/Tk [MKK99]. TCOL [Bro80]. TCOL-Ada [Bro80]. Teach [SS97, Bag98]. Teaching [Bro98a, Bro04, DRH98, FME01, Gib00, GBCGDBC97, Lea87a, Pag82, Bra85, Buh85, Won99]. Team [McD89, McD88a, McD88b]. Teams [MK91]. Technical [Bak92, Tok15, LC86]. Techniques [Col89, Sch87a, Yu97, dB97b]. Technologies [Ano99i, BCHR12, Bot99b, Kan12b, Ros10]. Technology [AW91, Boy89, DDJ98, Fis83, Log13b, OW82, Weg82, KSD12, Kle21, PW01, Wel03]. Telegen2 [Mat91]. Telesoft [Mat91]. Temporal [BKC91, KB87, MPV10, NLA05, EKPPR04]. termination [FSS87, WBP97, WBCS13]. terms [Whi85]. Test [AP84, CEG23, Gau90a, Gau90b, GR90, HB96, ML91, Tan91b]. Testbed [BKWS88, LT99, PW01, WWB99]. Testing [BW15, Fai80, FRS97, HNS98, KPR93, KMS82, Taf91a, Kan12b, Rym98, San01b, Taf91b]. tests [EK11, OWSB08]. Text [Zhu90, Bri09a]. theater [Con97b]. Theme [FA82]. Theoretical [PD82]. theories [Bjo13]. theory [Sin07]. There [EHP80]. Third [Ano90d]. thread [RH07]. threaded [MKK99, Taf13b]. threads [dlPRGB99]. Three [Bis88, Men88]. Tidbits [Bal94]. Time [All87, Alv87, Ano88b, Ano90c, Ano90d, Ano91c, Ano93c, Ano93a, Ano93h, Ano93k, Ano94d, Ano97, Ano00i, Ano02d, Ano06a, Ard87, Bak87a, Bak90c, Bak90e, Bak91c, Bar87, BA90a, BTP22, Bri92a, Bri92b, Bri94, BW15, CU89, Chr87a, CM90g, CSL^{+}87, DB98, FG82, FAT^{+}23, Gre16, HSW87, Mac80, McC87a, MR10, MdlP16, Mic16, Pau87, PS84, PMMT15, PR90, RSC16, RSC18, RM18, SW87, Sot06, SCC22, Taf91a, Tok03, Wei90a, de 87, AH01, ABW95, Ad93, AdlPT97, AAAG21, Bak90d, BTVC99, BCF94, Bos13, BdlPZ10, BJRW96, Bro88, BD01, BHR02, BH02, Buh85, BKW^{+}94, BW90a, BW92, BW93a, BW93b, BW94, BW07a, Bur13a, CS91, Chr87b, Col99b, CAC^{+}13, DM91, DV01, Ear92, EK12, EKPPR04, Fer97, Gal20, GH01, GB94, GHV03, GDAG97, GdlP02]. time [Goo90, GS10, Gre13, GS13, GDHM02, HMRF97, Har99a, HP01, HR03, HMC88, HM03, KGW^{+}85, LHBK87, LN91, LSRM12, LG88, LVM90, LT99, Mah13, MMB^{+}03, McC99, McC07, McC09, McC10, MS11, MMP13a, MMPT16, Moo97, MKK99, MP91, NAF05, NLA05, New95, New99, Nil12a, Pan12c, Pan12d, Pan12e, Pan12a, Pet10, PV98, PV99b, PV99a, PV02, PRQ21, Pot04, RC10a, RC10b, RH01, RH07, RH10, Rog09a, Rog11d, Rui13, SRC13a, San03a, Sel99, SLNM04, Sin07, Sri06a, Taf91b, TGH10, TPG21, UKDH97, UPRZ07, VGD^{+}97, WWB99, WD93, Wel90, WdlP97, Wel03, WB07b, WB10b, Whi10, Wre92, ZdlP02, ZEdlP13, ZdlP13, dlPRGB99, dlPZ03, Ano93b, ACWB89, Bar88, BKWS88, BHL^{+}93, Bur87b, BW87, BW90c, Col87, Dob01a, Dom87, GB87, LD87, Mea87, Rog09e, VMNM85, de 87]. time-partitioned [AAAG21]. Time-Related [Bak90c, Bak91c]. Time-Triggered [RSC16, RSC18]. TimeBench [BKW^{+}94]. timer [PG94]. Timers [Gre16, GS13, HR03]. Timing [AW88, AW89, CB07, CdN16, HF84, Lev15b, SRC15, WB15, CBW94]. Timing-Event [SRC15]. Tips [Bal94]. title [WGA90b]. Tk [MVG99, MKK99]. TLM [Mah12b]. TLM2.0 [Mah13]. TLM FIFO [Mah13]. TM [Bro97]. together [RT21]. tokeneer [KW11a, KW11b, KW11c, KW11d, KW11e, KW11f]. Tokyo [Puk88]. Tolerance [GGP^{+}90, KR88, LV23, BPP06, DB09, GdlP02, Kam99, LYB^{+}10, PV98, Wol97, Wol99]. Tolerant [AA88, AA89, DGBMCG97, KU84, Kni87, GLV97, PV02, TP98]. too [Har94c]. Tool [Ano93f, BBB97, CM98, Con97a, DGLM85, EJ16, FMN80, Hou83, MR87a, MNG16, Mur90, PDV98, PDN97, PR98, RS91, RSK^{+}19, Sch87b, SCD^{+}85, SS97, WHNB91, And04, BJRW96, BKW^{+}94, Car99a, CH04, CBB^{+}97, Dew07b, DCC85, Fre86b, GSP^{+}11, Gic91, GB94, LSP01, MP91, PS06, SG06]. tool-oriented [LSP01]. Tools [And20, Ano91a, FGN85, Hov00, Obe94, PBB^{+}88, Con97b, DPB^{+}97, ER86, KNB08, Sol91b]. toolset [DRF97, DA13, Jen09, Wel97b, Gro07]. toolsets [GST^{+}97]. topic [WGA90a]. Total [Med91]. Tour [Con97c]. tracer [EF01]. Traces [LP85]. Track [McC00]. Tracz [Wek90]. Traditional [EJK89]. Traffic [Aus22, ACW04, Kle06, OWSB08]. Training [AB87, Bra83a, Seb87, BB85, HS98, Mac86, McD88b]. transaction [Kie99, Mah11, Mah12a]. transactional [TGH10]. transactions [BP13, KR01a, KR01b, PMJPA01]. Transfer [Qui90a, Tv88, Weg82, de 88, AW91, AV93, BHR02, BWD90, Mah11, Mah12a, Qui90b]. Transformation [Bak86]. Transformational [KB83]. Transforming [LXY98, SJ91]. Transition [Coh81, FMN80, Woo88a, Woo88b, Wal85b]. Transitioning [CH97, Har82, Wis99, LRS09]. Transitions [HPT81]. Translating [GHVVW93, HvKPT87, Ste80, Men09]. Translation [AGG^{+}80, AB87, Led95b, PBB^{+}88, PDV98, The90, Hir94a, Hir94b]. Translator [DFS^{+}80]. Transparent [PW97, Wol99]. Transporting [Fre86b]. Traps [SS89]. Tree [FD16, BD91]. Trends [CMR90]. TRI [ACM91a, ACM97, Ano92m, Ano92j, Ano93l, Ano93m, Ano94h, Rob97]. TRI-Ada [ACM91a, Ano92m, Ano92j, Ano93l, Ano93m, Ano94h]. Tri-Ada’96 [Rob97]. TRI-Ada’97 [ACM97]. TriAda [STF98]. Trig [Sal92]. Triggered [RSC16, RSC18, RM18]. truly [Car99a]. Trust [Har22, TRT16, TS20, BBPT12]. truth [Moy17d]. TSL [HL85c]. TTF [BWM13]. TTF-Ravenscar [BWM13]. Tucker [The90]. Tunnel [Ben94]. Turing [Lis12]. Turtle [Bra85, MRB06]. Tutorial [Nil12b, Taf12, Taf13b, Wic82, San12, Whe95]. Two [BM85, Boy87, ER86, Fir87a, Gib00, WQ83]. Type [Bac82, Bel80, MF91, WQ83, Hod91a, Hod91b, KETT96, Led95b, Men09, Moy11c, Moy11d, Sei91]. type-based [Moy11c, Moy11d]. type-safe [Men09]. Typed [Sal92]. Types [Bak91b, Bak93a, Car91, Cla87c, Gar84, GES89, GA90, HLRS80, Hof86, Jam98a, KW98, KVT88a, Ler01, Lla92, SHR82, Wic82, Yeh82, And05, Bak93c, Bei92, Bos13, BD92, Duf08b, Duf08c, Duf08a, EGC13, Gon91a, Hod91a, Hod91b, Kir12, KVT88b, Led95a, LBO84, Och11, Rog09d, WJS^{+}01]. typical [Ros04]. Typing [BYY86, Bar09d].

UA [DRSK23]. UDP [RR14]. UK

[Bar87, Gil99b]. Ultracomputer [SS85]. UML [Faß01, Pet10, Sau05, Sei14]. Undergraduate [BRW97, Ruo05]. Underneath [Bar98]. Understanding [Wor97, Nil12b]. Unified [XWZ^{+}23]. uniform [LW01]. Uniformity [KW91]. Unify [WL98]. Uninitialized [Dew17]. unit [Bri09d]. United [Gri98]. Units [Mud87, Vol90, Bal95c]. unity [HD85]. Universal [Fis84b, Fro15, HB88]. UNIVERSAL FILE NAMES [Wan90]. UNIX [ER86, SHLR80]. Unlimited [LBO84]. Unmanned [CSSW09, CSSW10, Hum22, Wea10, SG06, Swa09a]. Unorthogonalities [Bac84]. Unpredictability [Maz89b]. unsigned [BCS89]. until [BRF92, LA99]. Update [Lin83, MC22, Tok15, BH02, Ker86, MB08, Ree86]. Updated [Tro12]. updates [Ker96b, Ker97, Ker98]. Updating [Coh86]. Uppsala [BRC98]. USA [ACM80, STF98]. Usability [BW90b, BW90d]. usable [Rob92]. USAF [SCFG04]. Usage [BCMC23, BG90, Cel97, Fri98b, Seb87, BW93a]. Usage/Performance [BG90]. USC [KMS82]. USC-ISI [KMS82]. Use [BYY86, BC16, Bur85a, BQ90, Car90, DoD87b, FAT^{+}23, Fos20, FOFY87, Gar84, HDHH98, KBT84, Kle06, KU84, Lei99b, LCB09, Men88, MMPT16, Mos22, Pie87, Rac89, Rom00, Ros10, SSB^{+}20, Tok15, WGC17, Wil87, BDV04, EK12, Fir87a, IMM85, Lei00, Rac88, Ros87a, Sin07, Var03, Wic98]. used [BC95, Fer97, ML95a, ML95b, Trü95]. User [ACM85, Ano92k, BE02, BDF^{+}85, CM94, Deb83, Fag00b, Fri83, Mac84, Rob92, WB10b, Wal94, Fos20]. User-defined [WB10b]. User-Friendly [Deb83]. Users [Ano92g, Ano92h, KQT^{+}21, Con97d, Bar85a, Gau95]. Using [ACM87a, AN05, Bag98, BT88b, BHD98, Bur87a, BH90, CLY98, DGCR^{+}84, DDJ98, Dru99, DH80, DH82, FCS83, Fli98, Gal20, Gar83, Gib00, Gór20, HB96, HF84, Hek83, Hir92, Jam98a, Lau07, MK87, Mac87, Mal88, MM17, MK83, Mau07, MR87b, MG87, MCS97, NAT20, Nyb87, PV02, Sal92, Sny91, SS97, Swa07b, Taf01c, Tan91a, Toa96, Tom97, VC01, Vas91, Win84, WV98, Yu97, ABW01, AW01, Bak93c, BTVC99, Bar09a, BHR^{+}11, BCHR12, BdlPZ10, Bro04, Car06a, CXY01, Col99b, CAC^{+}13, DPP^{+}09, DCC85, FME01, Faß01, Fuj87, Gid96, Gri98, Hov00, Jam98b, JR10, Lar22, LHFD13, Lei12b, Lit97, LVM90, LS98, Mic02, MY98, Moo97, NDM98, NDP99, Och09c, PMJPA01, Pet10, Plo92, Pow97, PL07, Ros11b, Ruo05, SS89, Swa07a, Swa09a, Taf06, Taf12, TP98, TS20]. using [Wag20, WD93, Wha13, dB97b]. utilities [WB07b]. utilization [HCT^{+}98].

v.2 [LHFD13]. v2 [RD23]. VADS [MB91]. Validate [DPP^{+}09]. validating [MMB^{+}03, Moy11d]. Validation [Goo80, Off87, PDV98, RS91, VGGS20, Bra99, HMC88, Squ91c]. Value [CEG23]. Values [Gre90]. VANETs [TS20]. Variabilities [Sal89]. Variable [Car89b, Sal89]. Variable-Length [Car89b]. Variables [Els90b, HLRS80, DG97, SC04b]. Variant [Mor87]. variation [AW88]. Variations [AW89, FA82]. VAX [Mal88, SHLR80]. VAX/VMS [Mal88]. VAX^{T M} [Fri87]. vector [Hod91a, Hod91b]. Vectorization [GPZdlP21]. Vehicle [LV23, SG06]. Vehicles [Hum22, ICS22, LC22]. Vendors [KQT^{+}21]. Venue [Ano02c, Ano02e]. verifiable [Taf13a]. Verification [BHN20, Car99b, CdN16, EJ16, HSB^{+}22, Hum22, Taf20, VGGS20, XWZ^{+}23, YQZ^{+}23, YG80, Ala13, AC04, Bal14, BCHR12, CHGH19, EH13, HM03, KSD12, Kan12b, Kni09, LMA94, Lei12b, Log13a, MWRH13, Ven08]. Verified [LW07, BGGS14, Lei12a]. Verifier [RDP97]. verify [BW99, Tom97]. Verifying [EKPPR04, LP80, MMB^{+}03, BWK^{+}01, NLA05]. Version [ACM89, Lei99a, MKP91a, Off87, Wei89, MKP91b, Wis99, Ano89c]. Versus [BH90, Ala13, WT03, dlPRGB99]. Vetronics [PW01]. VHDL [MP98]. Via [Bar00, HL86, Bal14, Cha82, LZL03, SBH^{+}98]. Vice [RH96]. Vice-Chair [RH96]. Video [Ano93p]. View [Har88, PD82, Ker99, VBF90]. Viewing [SYW85]. views [Hea08b]. viral [RMT11]. Virginia [ACM82]. Virtual [CDG97, Gar90, GA90, GR80, Vol90, Whi82, Joh93, WRL13]. virtualization [ZEdlP13]. visitor [CS02]. visitors [Car06a]. Visual [HCBM98b, BC95, CH06, Dul03]. Visualization [DCBM97, MKK99]. VMS [Mal88]. Void [Vol87]. vs [Bro91, Car97, Hea08b, Ker99, PV99b, Ros21, Syi95, Whe97, Yeh82]. Vulnerabilities [BCH^{+}19, Mar19, MdlP16, Mic16, Ano10a, BTB^{+}10, BW10a, Mic13, PJPD11].

WADAS [ACM91b, Ano92n, Ano92o, Ano93p, Ano93n, Ano93o]. Wait [LCN91]. Waits [LMP90]. walking [TT02]. Walnut [Con97c]. want [Mor95a]. Wanted [Jar07]. Washington [ACM91b, Ano99l, STF98, Moo85]. Way [Bar00, Gra83]. weak [Bri12a]. Weakness [Mar19, MB08]. weaknesses [SHT^{+}23]. Weapon [DoD87b, Nil12a]. Weaving [CSH03]. Web [Obr09, DDJ98, JF98a, JF98b, PB98, Ros04, Swa07a]. Web-based [JF98a, PB98, JF98b]. Web/database [Ros04]. WebAda [Smi97]. weights [Tro12]. Wellings [Rog97, Rog09e]. We’re [Mac87]. WG [Ano94e, Ano95b]. WG4 [Puk88]. WG9 [BRC98]. Where [Ano99c, Ano99l, Dru82, Bar14, Bri11d, Bri11e, Bri11f, Dew07a]. Whetstone [HF84]. which [PMJPA01]. while [Low99b]. Who [Fos20]. whole [Moy17d]. Wholesale [And05]. Why3 [Lei12b]. Wide [DDJ98, Bow92]. Will [Wek90]. windows [AAAG21, Ano00c, BBB98, BM97, HCBM98a, Nyb05, Puk94]. Winners [Har99b, Har00, MH20]. within [BA90b, Har94c, Lev91]. Without [Hil22]. Words [Tro06, Wol84]. Work [Ell83, Taf22b, Wai98, CN96, GG16, Taf12]. Work-bench [Wai98]. Work-Stealing [Taf22b]. workbench [CFH^{+}13]. Working [Ano92c, Ano92d, Ano92g, Ano92h, Ano92j, Ano92i, Ano93a, Ano93g, Ano93j, Ano94b, Ano94a, Ano94d, Ano94g, Ano95c, Ano95h, Ano95i, Ano95j, Ano99k, Ano00t, Ano00u, Ano00x, BHL^{+}93, Che09, GMO92, LWF91, OP85b, Sol91b, Vla93, Vla94, Whi95, Ano88a, Bak90e, Boy86, Bro96, BP94, Cro90, Dow94, Gar90, Goo90, Joh94, KGW^{+}85, MDPK94, MKP91b, Mun91b, Pen91, Qui90b, RT21, Rom88, Taf91b, Van90]. works [MH09]. Workshop [Ano88b, Ano90c, Ano90d, Ano91c, Ano92a, Ano93k, Ano99l, Ano00w, Bar87, Bar88, BDF^{+}85, Bux85b, GB87, Lei99b, Lei06, Wal94, Bro88, Bux85a, Kam95, Lei00, Lei02, Rob86, SHT^{+}23, Taf01a, Ano93b, Ano93h, Ano97, Ano00i, Ano02d, BW93b, Fis83, MR10, RC01, SPS88, Sof88]. workspace [Bri11c]. World [Ano99b, Ano00a, Ano00l, Ano00m, Har94a, DDJ98]. Worse [Har97]. Worst [FAT^{+}23, CBW94]. worst-case [CBW94]. would [Dew07a]. Wouldn’t [FBL^{+}10]. WOW [Ano02b]. Writers [Lev01a, SS89]. Writing [Bre97, vdL84]. Writtein [Cor83]. Written [KBT84, Whe86, Whe87]. Wrong [Mac87]. WSDL [Obr12a, Obr12b]. WWW [Ano95l, Ano95k, MH97].

XAda [Bur85a, Har85]. XERIS [Wai21]. XERIS/APEX [Wai21]. XML [Lei02, LLL03, Nyb10a].

year [Vau98]. yearbook [Lof93]. years [BT14]. York [WFF^{+}87].

zealot [Car01]. Zero [Har22].

References

Arevalo:1988:FTD

[AA88] Sergio Arevalo and Angel Alvarez. Fault tolerant distributed Ada. ACM SIG-ADA Ada Letters, 8(7): 118–122, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Arevalo:1989:FTD

[AA89] Sergio Arevalo and Angel Alvarez. Fault tolerant distributed Ada. ACM SIG-ADA Ada Letters, 9(5):54– 59, July/August 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Amurrio:2021:HWS

[AAAG21] A. Amurrio, E. Azketa, M. Aldea, and J. J. Gutiérrez. How windows size and number can influence the schedulability of hierarchicallyscheduled time-partitioned distributed real-time systems. ACM SIGADA Ada Letters, 41(1):64–68, June 2021. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL

https://dl.acm.org/doi/
10.1145/3570315.3570319.

Arnett:1987:ALT

[AB87] Kirk P. Arnett and Charles M.

Butler. Ada language training with a COBOL translation model. ACM SIGADA Ada Letters, 7(1):82–88, January/February 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Albertini:1998:ABM

[AB98] Victor D. Albertini and Craig J. Berrett. Ada in an on-board military communication system. ACM SIG-ADA Ada Letters, 18(6):132– 136, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Audsley:2015:EII

[AB15] N. C. Audsley and A. Burns. Efficient implementation of IPCP and DFP. ACM SIGADA Ada Letters, 35(1): 9–16, April 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Abbink:1996:ABS

[Abb96] H. J. Abbink. An Ada-based script language for simulation applications. ACM SIG-ADA Ada Letters, 16(5):35– 47, September/October 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Aldea:2013:IDF

[ABGH13] Mario Aldea, Alan Burns, Marina Gutiérrez, and Michael González Harbour. Incorporating the Deadline Floor Protocol in Ada. ACM SIG-ADA Ada Letters, 33(2): 49–58, August 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Allen:1995:STH

[ABW95] R. K. Allen, A. Burns, and A. J. Wellings. Sporadic tasks in hard real-time systems. ACM SIGADA Ada Letters, 15(5):46–51, September/October 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Audsley:2001:IHI

[ABW01] Neil Audsley, Alan Burns, and Andy Wellings. Implementing a high-integrity executive using Ravenscar. ACM SIGADA Ada Letters, 21(1):40–45, March 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Armitage:1985:ASD

[AC85] James W. Armitage and James V. Chelini. Ada software on distributed targets: a survey of approaches. ACM SIGADA Ada Letters, 4(4):32–37, January/February 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Amey:2003:ISE

[AC03] Peter Amey and Roderick Chapman. Industrial strength exception freedom. ACM SIGADA Ada Letters, 23(1):1–9, March 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Amey:2004:SVE

[AC04] Peter Amey and Roderick Chapman. Static verification and extreme programming. ACM SIGADA Ada Letters, 24(1):4–9, March 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

ACM:1980:PAS

[ACM80] ACM, editor. Proceedings of the ACM-SIGPLAN Symposium on the Ada Programming Language. Boston, MA, USA, 9–11 December, 1980, volume 15(11) of ACM SIG-PLAN Notices. ACM Press, New York, NY, USA, November 1980. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

ACM:1982:PAC

[ACM82] ACM, editor. Proceedings of the AdaTEC Conference on Ada, Arlington, Virginia, October 6–8, 1982. ACM Press, New York, NY, USA, October 1982. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Adatec:1985:UI

[ACM85] ACM Adatec Future Ada and Environment Workshop and Working Group 6. User interfaces. ACM SIG-ADA Ada Letters, 4(5):90– 96, March/April 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

ACM:1987:UAA

[ACM87a] ACM, editor. Using Ada:

ACM SIGAda international conference, Boston, Massachusetts, December 8–11, 1987. ACM Press, New York, NY, USA, 1987. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

ASA:1987:CAR

[ACM87b] ACM SIGAda ARTEWG.

The challenge of Ada runtime environments. ACM SIGADA Ada Letters, 7(5): 113–127, September/October 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

ASA:1989:MRS

[ACM89] ACM SIGAda ARTEWG.

A model runtime system interface for Ada, version 2.3. ACM SIGADA Ada Letters, 9(1):84–132, January/February 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

ACM:1991:TAP

[ACM91a]  ACM, editor.  TRI-Ada

’91 Proceedings. ACM Press, New York, NY, USA, 1991. ISBN 0-89791-445-7. LCCN ????

ACM:1991:WSS

[ACM91b]  ACM, editor.  WADAS

’91/Summer SIGAda Meeting. Eighth Annual Washington Ada Symposium/Summer SIGAda Meeting Software: Foundation for Competitiveness. Proceedings. ACM Press, New York, NY, USA, 1991. ISBN 0-89791-393-0. LCCN ????

ACM:1997:PTA

[ACM97] ACM, editor. Proceedings of the TRI-Ada’97 Conference, November 9–13, 1997, St. Louis, MO. ACM Press, New York, NY, USA, 1997. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Abraham:2011:IQAa

[ACP11a] Jay Abraham, Jeff Chapple, and Cyril Preve. Improving quality of Ada software with range analysis. ACM SIGADA Ada Letters, 31(3): 7–8, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Abraham:2011:IQAb

[ACP11b] Jay Abraham, Jeff Chapple, and Cyril Preve. Improving quality of Ada software with range analysis. ACM SIGADA Ada Letters, 31(3): 69–74, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Allaert:2004:EAT

[ACW04] Gaetan Allaert, Dirk Craeynest, and Philippe Waroquiers. European air traffic flow management: porting a large application to GNU/Linux. ACM SIGADA Ada Letters, 24(1):29–37, March 2004. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Asplund:1989:RTA

[ACWB89] L. Asplund, M. Carlsson, D. Wengelin, and G. Bray. Real-Time Ada compilers for the 68020. ACM SIGADA Ada Letters, 9(7):102–113, November/December 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Appelbe:1982:ODI

[AD82] B. Appelbe and G. Dismukes. An operational definition of intermediate code for implementing a portable Ada compiler. In ACM [ACM82], pages 266– 274. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Alonso:1993:RRT

[Ad93] Alejandro Alonso and Juan A. de la Puente. Reusable real-time executive in Ada. Design issues. ACM SIG-ADA Ada Letters, 13(2):44– 53, March/April 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Amey:2003:SAR

[AD03] P. N. Amey and B. J. Dobbing. Static analysis of Ravenscar programs. ACM SIGADA Ada Letters, 23(4): 58–64, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ada:1988:RDS

[Ada88] Ada Board. Response to the defense science board task force on military software. ACM SIGADA Ada Letters, 8(4):47–68, July/ August 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Atkinson:1990:DOO

[AdB90] C. Atkinson, Andrea di Maio, and R. Bayan. Dragoon: an object-oriented notation supporting the reuse and distribution of Ada software. ACM SIGADA Ada Letters, 10(9):50–59, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Alonso:2001:IMC

[AdlP01] Alejandro Alonso and Juan Antonio de la Puente. Implementation of mode changes with the Ravenscar profile. ACM SIGADA Ada Letters, 21(1):27–32, March 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Alonso:1997:CIF

[AdlPT97] Alejandro Alonso, Juan Antonio de la Puente, and Ken Tindell. Components for the implementation of fixed priority real-time systems in Ada. ACM SIGADA Ada Letters, 17(5):18–23, September/October 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Atkinson:1988:CBA

[AG88] C. Atkinson and S. J. Goldsack. Communication between Ada programs in DI-ADEM. ACM SIGADA Ada Letters, 8(7):86–96, Fall 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Agerberg:1985:SAS

[Age85] Jonas Agerberg. The simplest? Ada solution to the dining philosophers problem. ACM SIGADA Ada Letters, 5(1):42–48, July/ August 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Albrecht:1980:STA

[AGG^{+}80] Paul F. Albrecht, Phillip E. Garrison, Susan L. Graham, Robert H. Hyerle, Patricia Ip, and Bernd Krieg-Bruekner. Source-to-source translation: Ada to Pascal and Pascal to Ada. In ACM [ACM80], pages 183– 193. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

AldeaRivas:2001:EAR

[AH01] Mario Aldea Rivas and Michael González Harbour. Extending Ada’s real-time systems annex with the POSIX scheduling services. ACM SIGADA Ada Letters, 21(1):20–26, March 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Alexandr:2005:EPA

[AID05] Korochkin Alexandr, Salah Imad, and Korochkin Dmitry. Experimental performance analysis of Ada programs in cluster system. ACM SIG-ADA Ada Letters, 25(4):31– 36, December 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Allen:1991:CIF

[AKM^{+}91] D. Allen, M. Kamrad, C. McKay, R. Powers, and P. Rogers. Catalogue of interface features and options for the Ada runtime environment. ACM SIG-ADA Ada Letters, 11(8): 177–??, Fall 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Asplund:2000:SCS

[AL00] Lars Asplund and Kristina Lundqvist. Safety critical systems based on formal models. ACM SIGADA Ada Letters, 20(4):32–39, December 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/dec2000/asplund-

paper.pdf. Special Issue: Presentations from SIGAda 2000.

Alagic:2013:AVI

[Ala13]  Suad Alagic.  Automatic

versus interactive program verification. ACM SIG-ADA Ada Letters, 33(3):87– 88, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ahmad:2014:HAA

[ALB^{+}14] Ehsan Ahmad, Brian R. Larson, Stephen C. Barrett, Naijun Zhan, and Yunwei Dong. Hybrid annex: an AADL extension for continuous behavior and cyber-physical interaction modeling. ACM SIG-ADA Ada Letters, 34(3):29– 38, December 2014. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Allen:1987:TRT

[All87] Dock Allen. Tailored runtime environments for realtime applications. ACM SIGADA Ada Letters, 7(6): 13–14, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Alstad:1983:PAP

[Als83] James P. Alstad. Problems with Ada as a program design language: a position paper. ACM SIGADA Ada Letters, 2(6):51–52, May/June 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Alvarez:1987:RTP

[Alv87] Angel Alvarez. Real-time programming and priority interrupt systems. ACM SIGADA Ada Letters, 7(6): 97–100, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Amey:2001:LSJ

[Ame01] Peter Amey. A language for systems not just software. ACM SIGADA Ada Letters, 21(4):3–11, December 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ausden:2005:UAG

[AN05] Howard Ausden and Karl Nyberg. Using ASIS to generate C++ bindings. ACM SIGADA Ada Letters, 25(4): 23–30, December 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anderson:1988:AMS

[And88] G. E. Anderson. An Ada multitasking solution for the Sieve of Eratosthenes. ACM SIGADA Ada Letters, 8(5):71–74, September/October 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anderson:2004:RTA

[And04] Paul Anderson. A refactoring tool for Ada 95. ACM SIGADA Ada Letters, 24(4): 23–28, December 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Andress:2005:WBR

[And05] Randal P. Andress. Wholesale byte reversal of the outermost Ada record object to achieve endian independence for communicated data types. ACM SIGADA Ada Letters, 25(3):19–27, September 2005. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anderson:2020:MSA

[And20] Paul Anderson. Modernizing static analysis tools to facilitate integrations. ACM SIGADA Ada Letters, 39(1): 101–108, January 2020. ISSN 0736-721X. URL https:

//dl.acm.org/doi/abs/10.
1145/3379106.3379119.

Anonymous:1987:CAR

[Ano87] Anonymous. The challenge of Ada runtime environments (ARTEWG). ACM SIGADA Ada Letters, 7(5):113–127, September/October 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1988:ARE

[Ano88a] Anonymous. Ada runtime environment working group — a framework for describing Ada runtime environment. ACM SIGADA Ada Letters, 8(3):51–68, May/June 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1988:SIW

[Ano88b] Anonymous. Second International Workshop on Real-Time ADA Issues. ACM SIGADA Ada Letters, 8 (7):??, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1989:ASM

[Ano89a] Anonymous. Ada and software management in NASA: assessment and recommendations. ACM SIGADA Ada Letters, 9(6):53–66, September/October 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1989:AAL

[Ano89b]  Anonymous.  Approved

Ada language commentaries. ACM SIGADA Ada Letters, 9(3):1–341, Spring 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1989:MRS

[Ano89c] Anonymous. A model runtime system interface for Ada Version 2.3. ACM SIG-ADA Ada Letters, 9(1):84– 132, January/February 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1990:ACEa

[Ano90a]  Anonymous.  Ada Com-

piler Evaluation Capability (ACEC): An overview. ACM SIGADA Ada Letters, 10(3): 101–110, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1990:ACEb

[Ano90b]  Anonymous.  Ada Com-

piler Evaluation Capability (ACEC) data analysis: An overview. ACM SIG-ADA Ada Letters, 10(3):111– 125, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1990:FIW

[Ano90c] Anonymous. Fourth International Workshop on Real-Time Ada Issues. ACM SIGADA Ada Letters, 10 (9):??, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1990:TIW

[Ano90d] Anonymous. Third International Workshop on Real-Time Ada Issues. ACM SIGADA Ada Letters, 10 (4):??, Spring 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1991:ISE

[Ano91a] Anonymous. 1st International Symposium on Environments and Tools for Ada. ACM SIGADA Ada Letters, 11(3):??, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1991:AFS

[Ano91b]  Anonymous.  Ada follies

songbook. ACM SIG-ADA Ada Letters, 11(4):99– ??, May/June 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1991:FIW

[Ano91c] Anonymous. Fifth International Workshop on Real-Time Ada Issues. ACM SIGADA Ada Letters, 11(6): ??, September/October 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1991:PPI

[Ano91d] Anonymous. Preface. ACM SIGADA Ada Letters, 11 (3):iii, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:AWS

[Ano92a] Anonymous. 5th Annual Workshop on Software Reuse. ACM SIGADA Ada Letters, 12(3):43–??, May/June 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:KBS

[Ano92b] Anonymous. 7th Knowledge-Based Software Engineering Conference: Call for papers. ACM SIGADA Ada Letters, 12(2):28–??, March/April 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:AARa

[Ano92c] Anonymous. Activities of the Ada Runtime Environment Working Group (ARTEWG). ACM SIGADA Ada Letters, 12(3):50–??, May/June 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:AARb

[Ano92d] Anonymous. Activities of the Ada Runtime Environment Working Group (ARTEWG). ACM SIGADA Ada Letters, 12(5):30–??, September/October 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:ECN

[Ano92e] Anonymous. Education committee news. ACM SIG-ADA Ada Letters, 12(3):65– ??, May/June 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:PSS

[Ano92f]  Anonymous.  Preliminary

Summer ’92 SIGAda meeting. ACM SIGADA Ada Letters, 12(2):33–??, March/ April 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:RCAa

[Ano92g] Anonymous. Report from the Commercial Ada Users Working Group (CAUWG). ACM SIGADA Ada Letters, 12(2):29–??, March/April 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:RCAb

[Ano92h] Anonymous. Report from the Commercial Ada Users Working Group (CAUWG). ACM SIGADA Ada Letters, 12(3):64–??, May/June 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:RSS

[Ano92i] Anonymous. Report from the SIGAda Software Development Standards and Ada Working Group (SD-SAWG). ACM SIGADA Ada Letters, 12(2):31–??, March/April 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:ROO

[Ano92j]  Anonymous.  Report of

the object oriented working group and sample problem for Tri-Ada 92 panel. ACM SIGADA Ada Letters, 12(5):37–??, September/October 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:SAR

[Ano92k]  Anonymous.  Sixth An-

nual Rational Users’ Group Meeting. ACM SIGADA Ada Letters, 12(3):42–??, May/June 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:SRS

[Ano92l] Anonymous. Software repositories — survey. ACM SIG-ADA Ada Letters, 12(5):14– ??, September/October 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:TA

[Ano92m] Anonymous. TRI-Ada ’92. ACM SIGADA Ada Letters, 12(4):16–??, July/August 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:Wa

[Ano92n] Anonymous. WADAS ’92. ACM SIGADA Ada Letters, 12(2):25–??, March/April 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1992:Wb

[Ano92o] Anonymous. WADAS ’92. ACM SIGADA Ada Letters, 12(3):40–??, May/June 1992.

CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:ARA

[Ano93a] Anonymous. 1991 annual report for the ACM Special Group for Ada (SIG-Ada): Ada Run-Time Environment Working Group Proceedings form the Software Safety Symposium. ACM SIGADA Ada Letters, 13(1): 35–??, January 1, 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:IWR

[Ano93b] Anonymous. 6th International Workshop on Real-Time Ada Issues. ACM SIGADA Ada Letters, 13(2): ??, March/April 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:AAR

[Ano93c] Anonymous. Activities of the Ada Run Time Environment Interest Group for Ada (SIGAda). ACM SIG-ADA Ada Letters, 13(1):30– ??, January 1, 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:EA

[Ano93d] Anonymous. Evolution of Ada 9X. ACM SIGADA Ada Letters, 13(6):66–158, November/December 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:PSR

[Ano93e] Anonymous. PIWG: a status report. ACM SIG-ADA Ada Letters, 13(3):42– ??, May/June 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:QAT

[Ano93f] Anonymous. Quality assessment tool for implementations of Ada. ACM SIGADA Ada Letters, 13(6):26–??, November/December 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:RSS

[Ano93g] Anonymous. Report from the SIGAda software development standards and Ada working group (SD-SAWG). ACM SIGADA Ada Letters, 13(4):22–??, July/August 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:SIR

[Ano93h] Anonymous. Seventh International Real-Time Ada Issues Workshop: Call for papers. ACM SIGADA Ada Letters, 13(6):32–??, November/December 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:SAR

[Ano93i] Anonymous. SIGAda annual report for FY93. ACM SIGADA Ada Letters, 13(6): 13–??, November/December 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:SWG

[Ano93j] Anonymous. SIGAda Working Groups. ACM SIG-ADA Ada Letters, 13(1):4– ??, January 1, 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:SIW

[Ano93k] Anonymous. Summary of the 6th International Workshop on Real-Time Ada Issues. ACM SIGADA Ada Letters, 13(2):20–??, March/April 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:TACa

[Ano93l] Anonymous. Tri-Ada ’93:

Call for participation. ACM SIGADA Ada Letters, 13 (2):17–??, March/April 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:TACb

[Ano93m] Anonymous. Tri-Ada ’94: Call for participation. ACM SIGADA Ada Letters, 13(6): 33–??, November/December 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:W

[Ano93n] Anonymous. WadaS ’93. ACM SIGADA Ada Letters, 13(3):18–??, May/June 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:WCP

[Ano93o] Anonymous. WadaS ’93: Call for papers. ACM SIG-ADA Ada Letters, 13(2):15– ??, March/April 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:WDV

[Ano93p] Anonymous. WAdaS ’93 debate video. ACM SIGADA Ada Letters, 13(6):27–??, November/December 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1994:AAS

[Ano94a] Anonymous. Activities of the Ada semantic interface specification working group (ASISWG). ACM SIG-ADA Ada Letters, 14(2):54– ??, March/April 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1994:AAI

[Ano94b] Anonymous. Activities of the artificial intelligence working group. ACM SIG-ADA Ada Letters, 14(2):50– ??, March/April 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1994:AEC

[Ano94c] Anonymous. Ada in Europe:

Call for papers. ACM SIG-ADA Ada Letters, 14(2):18– ??, March/April 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1994:ART

[Ano94d] Anonymous. Ada Run Time Environment Working Group (ARTEWG) meeting. ACM SIGADA Ada Letters, 14(3): 18–??, May/June 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1994:SAI

[Ano94e] Anonymous. SIGAda artificial intelligence WG meeting. ACM SIGADA Ada Letters, 14(3):16–??, May/June 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1994:SEE

[Ano94f] Anonymous. SIGAda Extended Executive Committee. ACM SIGADA Ada Letters, 14(6):3–??, November 1, 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1994:SWG

[Ano94g] Anonymous. SIGAda Working Groups. ACM SIG-ADA Ada Letters, 14(6):4– ??, November 1, 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1994:TAC

[Ano94h] Anonymous. Tri-Ada ’94: Call for participation. ACM SIGADA Ada Letters, 14 (2):20–??, March/April 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1995:LSC

[Ano95a]  Anonymous.  Local SIG-

Ada chapters. ACM SIG-ADA Ada Letters, 15(6):7– ??, November 1, 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1995:SAIa

[Ano95b] Anonymous. SIGAda Artificial Intelligence WG meeting. ACM SIGADA Ada Letters, 15(3):39–??, May/ June 1, 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1995:SAIb

[Ano95c] Anonymous. SIGAda Artificial Intelligence Working Group Charter. ACM SIG-ADA Ada Letters, 15(3):40– ??, May/June 1, 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1995:SC

[Ano95d] Anonymous. SIGAda at a crossroads? ACM SIG-ADA Ada Letters, 15(4):12– ??, July/August 1, 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1995:SECa

[Ano95e] Anonymous. SIGAda Executive Committee. ACM SIG-ADA Ada Letters, 15(3):3– ??, May/June 1, 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1995:SECb

[Ano95f] Anonymous. SIGAda Executive Committee. ACM SIG-ADA Ada Letters, 15(6):4– ??, November 1, 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1995:SEE

[Ano95g] Anonymous. SIGAda Extended Executive Committee. ACM SIGADA Ada Letters, 15(1):3–??, January 1, 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1995:SWGa

[Ano95h] Anonymous. SIGAda Working Groups. ACM SIG-ADA Ada Letters, 15(1):4– ??, January 1, 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1995:SWGb

[Ano95i] Anonymous. SIGAda Working Groups. ACM SIG-ADA Ada Letters, 15(3):4– ??, May/June 1, 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1995:SWGc

[Ano95j] Anonymous. SIGAda Working Groups. ACM SIG-ADA Ada Letters, 15(6):5– ??, November 1, 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1995:SWSa

[Ano95k] Anonymous. SIGAda WWW server. ACM SIGADA Ada Letters, 15(3):19–??, May/June 1, 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1995:SWSb

[Ano95l] Anonymous. SIGAda WWW Server. ACM SIGADA Ada Letters, 15(5):18–??, September 1, 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1995:SSM

[Ano95m] Anonymous. Summer ’95 SIGAda Meeting. ACM SIG-ADA Ada Letters, 15(3):35– ??, May/June 1, 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1997:EIR

[Ano97] Anonymous. Eighth International Real-Time Ada Workshop. ACM SIGADA Ada Letters, 17(5):??, September/October 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1999:ICS

[Ano99a] Anonymous. The 21^{st} international conference on software engineering. ACM SIGADA Ada Letters, 19(1): 18–??, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1999:AAW

[Ano99b] Anonymous. Ada around the world. ACM SIG-ADA Ada Letters, 19(1): 11–??, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1999:AWD

[Ano99c] Anonymous. ASIS — where do we go from here? ACM SIGADA Ada Letters, 19(1): 42–47, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1999:ABA

[Ano99d] Anonymous. ASIS has been approved as ISO standard. ACM SIGADA Ada Letters, 19(1):40–41, March 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1999:EP

[Ano99e] Anonymous. Editorial policy. ACM SIGADA Ada Letters, 19(1):5–??, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1999:IJC

[Ano99f] Anonymous. International journal of computer systems: Science and engineering call for papers. ACM SIG-ADA Ada Letters, 19(1):

16–??, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1999:KC

[Ano99g] Anonymous. Key contacts.

ACM SIGADA Ada Letters, 19(1):6–??, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1999:LSC

[Ano99h]  Anonymous.  Local SIG-

Ada chapter. ACM SIG-ADA Ada Letters, 19(1): 9–??, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1999:RST

[Ano99i] Anonymous. Reliable software technologies: Ada-Europe ’99. ACM SIG-ADA Ada Letters, 19(1): 15–??, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1999:S

[Ano99j] Anonymous. SIGAda ’99.

ACM SIGADA Ada Letters, 19(1):13–??, March 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1999:SWG

[Ano99k] Anonymous. SIGAda working groups. ACM SIG-ADA Ada Letters, 19(1): 7–??, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1999:WRA

[Ano99l] Anonymous. Workshop report: ASIS — where do we go from here? 6–10 PM, Sunday, 8 November 1998 SIGAda’98, Washington DC. ACM SIG-ADA Ada Letters, 19(1): 42–47, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2000:AAW

[Ano00a] Anonymous. Ada around the world. ACM SIG-ADA Ada Letters, 20(1): 10–11, March 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2000:AE

[Ano00b] Anonymous. Ada Europe. ACM SIGADA Ada Letters, 20(1):16–17, March 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2000:AJE

[Ano00c]  Anonymous.  Announce-

ments: John English Windows library. ACM SIG-ADA Ada Letters, 20(2): 18, June 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.acm.org/sigada/ada_
letters/june2000/jewl.pdf.

Anonymous:2000:ARH

[Ano00d]  Anonymous.  Announce-

ments: Research in the history of programming languages and software engineering. ACM SIGADA Ada Letters, 20(2):17, June 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/june2000/plresearch.
pdf.

Anonymous:2000:EP

[Ano00e] Anonymous. Editorial policy.

ACM SIGADA Ada Letters, 20(1):3–4, March 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2000:KCa

[Ano00f] Anonymous. Key contacts.

ACM SIGADA Ada Letters, 20(1):5, March 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2000:KCb

[Ano00g] Anonymous. Key contacts.

ACM SIGADA Ada Letters, 20(4):80–??, December 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Special Issue: Presentations from SIGAda

Anonymous:2000:LSC

[Ano00h]  Anonymous.  Local SIG-

Ada chapters. ACM SIG-ADA Ada Letters, 20(1): 8–9, March 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2000:MIR

[Ano00i] Anonymous. Meetings: 10^{th} International Real-Time Ada Workshop. ACM SIG-ADA Ada Letters, 20(2): 14, June 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

                                        www.acm.org/sigada/ada_
                                        letters/june2000/irtaw.
                                        pdf.

Anonymous:2000:MAE

[Ano00j] Anonymous. Meetings: Ada Europe 2001. ACM SIG-ADA Ada Letters, 20(2): 15–16, June 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

                                        www.acm.org/sigada/ada_
                                        letters/june2000/ada_europe_
                                        2001.pdf.

Anonymous:2000:MS

[Ano00k] Anonymous. Meetings: SIG-Ada 2000. ACM SIG-ADA Ada Letters, 20(2): 11–13, June 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.acm.org/sigada/ada_
letters/june2000/sigada_
2000.pdf.

Anonymous:2000:NIAa

[Ano00l]  Anonymous.  Newsletter

info: Ada around the world. ACM SIGADA Ada Letters, 20(2):10, June 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/june2000/newsletter_
info.pdf.

Anonymous:2000:NIAb

[Ano00m] Anonymous. Newsletter info:

Ada around the world. ACM SIGADA Ada Letters, 20 (3):10–11, September 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/sept2000/newsletter_
info.pdf.

Anonymous:2000:NIEa

[Ano00n]  Anonymous.  Newsletter

info: Editorial policy. ACM SIGADA Ada Letters, 20 (2):3–4, June 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.acm.org/sigada/ada_
letters/june2000/newsletter_
info.pdf.

Anonymous:2000:NIEb

[Ano00o] Anonymous. Newsletter info:

Editorial policy. ACM SIG-ADA Ada Letters, 20(3):3– 4, September 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.acm.org/sigada/ada_
letters/sept2000/newsletter_
info.pdf.

Anonymous:2000:NIKa

[Ano00p]  Anonymous.  Newsletter

info: Key contacts. ACM SIGADA Ada Letters, 20 (2):5, June 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.acm.org/sigada/ada_
letters/june2000/newsletter_
info.pdf.

Anonymous:2000:NIKb

[Ano00q] Anonymous. Newsletter info: Key contacts. ACM SIG-ADA Ada Letters, 20(3): 5, September 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.acm.org/sigada/ada_
letters/sept2000/newsletter_
info.pdf.

Anonymous:2000:NILa

[Ano00r]  Anonymous.  Newsletter

info: Local SIGAda chapters. ACM SIGADA Ada Letters, 20(2):8–9, June 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_

letters/june2000/newsletter_
info.pdf.

Anonymous:2000:NILb

[Ano00s]  Anonymous.  Newsletter

info: Local SIGAda chapters. ACM SIGADA Ada Letters, 20(3):8–9, September 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/sept2000/newsletter_
info.pdf.

Anonymous:2000:NISa

[Ano00t] Anonymous. Newsletter info:

SIGAda working groups. ACM SIGADA Ada Letters, 20(2):6–7, June 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/june2000/newsletter_
info.pdf.

Anonymous:2000:NISb

[Ano00u] Anonymous. Newsletter info:

SIGAda working groups. ACM SIGADA Ada Letters, 20(3):6–7, September 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/sept2000/newsletter_
info.pdf.

Anonymous:2000:S

[Ano00v] Anonymous. SIGAda 2000.

ACM SIGADA Ada Letters, 20(1):18, March 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2000:SWA

[Ano00w] Anonymous. SIGAda ’99 workshop: ASIS — extensions for higher level abstractions. ACM SIG-ADA Ada Letters, 20(1): 19–24, March 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2000:SWG

[Ano00x] Anonymous. SIGAda working groups. ACM SIG-ADA Ada Letters, 20(1): 6–7, March 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2001:NI

[Ano01a]  Anonymous.  Newsletter

information. ACM SIG-ADA Ada Letters, 21(2): 3–4, June 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2001:SA

[Ano01b] Anonymous. SIGAda 2001 announcement. ACM SIG-ADA Ada Letters, 21(2): 11, June 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2002:AEP

[Ano02a] Anonymous. Ada Europe 2002 preliminary program. ACM SIGADA Ada Letters, 22(1):39–42, March 2002. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2002:AWS

[Ano02b] Anonymous. Ada WOW from SIGAda 2001. ACM SIGADA Ada Letters, 22(1): 43–60, March 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2002:INV

[Ano02c]  Anonymous.  Interesting

notes on the venue for SIGada 2002. ACM SIG-ADA Ada Letters, 22(1): 62–63, March 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2002:PIR

[Ano02d] Anonymous. Proceedings of the 11^{th} International Real Time Ada Workshop. ACM SIGADA Ada Letters, 22(4): ??, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2002:SPC

[Ano02e] Anonymous. SIGAda 2002 preliminary call for participation and notes on venue.

ACM SIGADA Ada Letters, 22(1):61, March 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2006:AIE

[Ano06a] Anonymous. Ada issue 307 — execution-time clocks. ACM SIGADA Ada Letters, 26(1): 31–44, April 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2006:AIDa

[Ano06b] Anonymous. Ada issue 321 — definition of dispatching policies. ACM SIGADA Ada Letters, 26(1):45–55, April 2006. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2006:AIDb

[Ano06c] Anonymous. Ada issue 327 — dynamic ceiling priorities. ACM SIGADA Ada Letters, 26(1):56–63, April 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2006:AIA

[Ano06d] Anonymous. Ada issue 333 — additional locking policies with FIFO Within Priorities. ACM SIGADA Ada Letters, 26(1):64–65, April 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2006:CAA

[Ano06e]  Anonymous.  Conference

announcements: Ada Europe 2006 CFP. ACM SIGADA Ada Letters, 26 (1):66, April 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2006:CAS

[Ano06f]  Anonymous.  Conference

announcements: SIGAda 2006 information. ACM SIGADA Ada Letters, 26 (1):67, April 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2006:KC

[Ano06g] Anonymous. Key contacts.

ACM SIGADA Ada Letters, 26(1):4–6, April 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2010:ASF

[Ano10a] Anonymous. Annex SPARK — final draft: SPARK.Specific information for vulnerabilities. ACM SIGADA Ada Letters, 30(2):53–66, August 2010. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2010:MRA

[Ano10b] Anonymous. Maintenance and revision of the Ada programming language: outline announcement. ACM SIGADA Ada Letters, 30(2): 25–26, August 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:2017:GEA

[Ano17a] Anonymous. Gem #142: Exception-ally. ACM SIG-ADA Ada Letters, 37(2):9– 12, December 2017. CODEN AALEE5. ISSN 0736-721X. Anonymous:2017:GRS

[Ano17b] Anonymous. Gem #143: Return to the sources. ACM SIGADA Ada Letters, 37(2): 13–17, December 2017. CO-DEN AALEE5. ISSN 0736-721X.

Anonymous:2017:GBB

[Ano17c] Anonymous. Gem #144: a bit of bytes: Characters and encoding schemes. ACM SIGADA Ada Letters, 37(2): 18–22, December 2017. CO-DEN AALEE5. ISSN 0736-721X.

Ardo:1984:SAC

[AP84] Anders Ardo and Lars Philipson. A simple Ada compiler invalidation test. ACM SIG-ADA Ada Letters, 3(5):69– 74, March/April 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ali:2011:PPM

[AP11] Hazem Ismail Ali and Luís Miguel Pinho. A parallel programming model for Ada. ACM SIGADA Ada Letters, 31(3): 19–26, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Abu-Ras:1995:OMP

[AR95]  Jim Abu-Ras.  Optimal

Mutex policy in Ada 95. ACM SIGADA Ada Letters, 15(6):46–56, November/December 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ardo:1987:RTE

[Ard87] Anders Ardo. Real-time efficiency of Ada in a multiprocessor environment. ACM SIGADA Ada Letters, 7(6): 40–42, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Aldea-Rivas:2018:SSP

[ARG18] Mario Aldea-Rivas and Kristoffer Nyborg Gregertsen. Session summary: Profiles. ACM SIGADA Ada Letters, 38(1):62–65, June 2018. CO-DEN AALEE5. ISSN 0736-721X.

Arndt:1986:CBE

[Arn86]  Douglas Arndt.  Char-

acter building experiences. ACM SIGADA Ada Letters, 6(1):63–71, January/ February 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Aldea-Rivas:2018:PNA

[ARPT18] Mario Aldea-Rivas and Héctor Pérez-Tijero. Proposal for a new Ada profile for small microcontrollers. ACM SIG-ADA Ada Letters, 38(1):34– 39, June 2018. CODEN AALEE5. ISSN 0736-721X.

Amiguet:1987:DSA

[AS87] C. Amiguet and A. Schiper. Discrete-event simulation in Ada. In ACM [ACM87a], pages 133–140. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Asplund:2001:SNS

[Asp01] Lars Asplund. Session: new scheduling/dispatching policies. ACM SIGADA Ada Letters, 21(1):11–13, March 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Atkinson:1990:OOM

[Atk90]  Colin Atkinson.  Object-

oriented mechanisms. ACM SIGADA Ada Letters, 10 (9):35–38, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ausden:2022:AAE

[Aus22] Howard Ausden. Achieving 100% availability: In the ERAM Air Traffic Control System. ACM SIG-ADA Ada Letters, 42(2):89– 91, December 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3591335.3591345.

Antonelli:1993:AAT

[AV93] Charles J. Antonelli and Richard A. Volz. An alternative to asynchronous transfer of control in Ada 9X. ACM SIGADA Ada Letters, 13 (2):37–43, March/April 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Altman:1988:TVD

[AW88] N. Altman and N. Weiderman. Timing variation in dual loop benchmarks. ACM SIGADA Ada Letters, 8(3): 98–106, May/June 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Altman:1989:TVD

[AW89] N. Altman and Nelson Weiderman. Timing variations in dual loop benchmarks. ACM SIGADA Ada Letters, 8(3): 98–106, May/June 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anderson:1991:TTE

[AW91] J. A. Anderson and E. S.

Ward. Technology transfer: experiences in introducing object-oriented methods to government projects. In ACM [ACM91b], pages 10– 15. ISBN 0-89791-393-0. LCCN ????

Audsley:2001:IUR

[AW01] Neil Audsley and Andy Wellings. Issues with using Ravenscar and the Ada distributed systems annex for high-integrity systems. ACM SIGADA Ada Letters, 21(1): 33–39, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ben-Ari:1982:CFA

[BA82] Mordechai Ben-Ari. The case for full Ada. ACM SIGADA Ada Letters, 2(3):34–37, November/December 1982. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ben-Ari:1990:ARS

[BA90a]  M. Ben-Ari.  Ada re-

quirements for small realtime systems. ACM SIG-ADA Ada Letters, 10(4): 159–165, Spring 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ben-Ari:1990:SWI

[BA90b] M. Ben-Ari. Signaling from within interrupt handlers. ACM SIGADA Ada Letters, 10(1):100–103, January/February 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ben-Ari:1998:DFR

[BA98] Mordechai Ben-Ari. Dispatching on the function result. ACM SIGADA Ada Letters, 18(4):101–106, July 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:2007:AOS

[BA07] Ben Brosgol and Mario Aldea. Ada and other standards: Introduction. ACM SIGADA Ada Letters, 27(2): 88–89, August 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bach:1982:TCA

[Bac82] Ivan Bach. On the type concept of Ada. ACM SIGADA Ada Letters, 2(3):38–50, November/December 1982. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bach:1984:UIR

[Bac84] Ivan Bach. Unorthogonalities in the identification rules in Ada. ACM SIGADA Ada Letters, 4(3):37–43, November/December 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bagert:1998:UAT

[Bag98]  Donald J. Bagert.  Us-

ing Ada to teach programming language design concepts. ACM SIGADA Ada Letters, 18(1):54–64, January/February 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bail:2010:ERE

[Bai10]  William Bail.  Effective

requirements engineering. ACM SIGADA Ada Letters, 30(3):1–2, December 2010. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baize:2020:SO

[Bai20]  Eric Baize.  SAFECode

overview. ACM SIGADA Ada Letters, 39(1):17–19, January 2020. ISSN 0736-

                                        721X. URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3379106.3379109.

Baker:1986:TSD

[Bak86] Paul L. Baker. Transformation of structured data schemata into Ada language statements. ACM SIG-ADA Ada Letters, 6(4):66– 74, July/August 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1987:ARS

[Bak87a] Ted Baker. Ada runtime support environments to better support real-time systems. ACM SIGADA Ada Letters, 7(6):85–87, Fall 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1987:LTP

[Bak87b] Ted P. Baker. A low-level tasking package for Ada. In ACM [ACM87a], pages 141– 146. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Baker:1988:IIA

[Bak88] Ted Baker. Improving immediacy in Ada. ACM SIGADA Ada Letters, 8(7): 50–56, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1990:APL

[Bak90a] Paul L. Baker. Ada as a preprocessor language. ACM SIGADA Ada Letters, 10 (1):83–91, January/February 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1990:OAT

[Bak90b]  T. Baker.  Opening up

Ada tasking. ACM SIG-ADA Ada Letters, 10(9): 60–64, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1990:FST

[Bak90c] Ted Baker. Fixing some timerelated problems in Ada. ACM SIGADA Ada Letters, 10(4):136–143, Spring 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1990:PRT

[Bak90d]  Ted Baker.  Protected

records, time management, and distribution. ACM SIGADA Ada Letters, 10 (9):17–28, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1990:TIW

[Bak90e] Ted Baker. Time issues working group. ACM SIG-ADA Ada Letters, 10(4):119– 135, Spring 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1991:OOP

[Bak91a] Henry G. Baker. Objectoriented programming in Ada83—genericity rehabilitated. ACM SIGADA Ada Letters, 11(9):116–127, November/December 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1991:SPL

[Bak91b] Henry G. Baker. Structured programming with limited private types in Ada: Nesting is for the soaring eagles. ACM SIGADA Ada Letters, 11 (5):79–90, July/August 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1991:TRI

[Bak91c] Ted Baker. Time-related issues in Ada 9X. ACM SIG-ADA Ada Letters, 11(6):54– 60, September/October 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1992:RLT

[Bak92] P. Baker. Response letter from the technical editor. ACM SIGADA Ada Letters, 12(6):46–??, November/December 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1993:HSL

[Bak93a] Henry G. Baker. How to steal from a limited private account — why mode IN OUT parameters for limited types must be passed by reference. ACM SIGADA Ada Letters, 13(3):91–95, May/June 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1993:SLE

[Bak93b] Henry G. Baker. Strategies for the lossless encoding of strings as Ada identifiers. ACM SIGADA Ada Letters, 13(5):43–47, September/October 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1993:SLR

[Bak93c] Henry G. Baker, Jr. Safe and leakproof resource management using Ada83 limited types. ACM SIGADA Ada Letters, 13(5):32–42, September/October 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Balfour:1994:ATT

[Bal94] Brad Balfour. Ada 9X: Tips and tidbits. ACM SIG-ADA Ada Letters, 14(5):65– 70, September/October 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bal:1995:CDS

[Bal95a] Henri E. Bal. Comparing data synchronization in Ada 9X and Orca. ACM SIG-ADA Ada Letters, 15(1):50– 63, January/February 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Balfour:1995:EDI

[Bal95b] Brad Balfour. Expressing design inheritance relationships in Ada 95. ACM SIG-ADA Ada Letters, 15(3):71– 75, May/June 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Balfour:1995:ICL

[Bal95c] Brad Balfour. Inheritance and child library units. ACM SIGADA Ada Letters, 15 (4):29–35, July/August 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Balfour:1997:AJB

[Bal97] B. Balfour. Ada 95, Java byte code, and the distributed systems annex. In ACM [ACM97], pages 247– 262. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Balfour:1999:CSC

[Bal99] Brad Balfour. The current state of CORBA (invited presentation). ACM SIGADA Ada Letters, 19(3): 223, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ball:2014:CCL

[Bal14] Thomas Ball. Correctness via compilation to logic: a decade of verification at Microsoft Research. ACM SIG-ADA Ada Letters, 34(3):69– 70, December 2014. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bardin:1985:RSU

[Bar85a]  Bryce M. Bardin.  Re-

port from the SIGAda Users’ Committee chairperson. ACM SIGADA Ada Letters, 5(3–6):61–62, November/December 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bardin:1985:DPA

[Bar85b] Bryce M. Bardin. A “To Be Determined” package for Ada development. ACM SIGADA Ada Letters, 5(3–6):45–56, November/December 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:1987:PIW

[Bar87] John Barnes, editor. Proceedings of the International Workshop on Real-Time Ada issues, Moretonhampstead, Devon, UK, 13–15 May 1987, For parts, see ACM SIGADA Ada Letters vol. 7, no. 6. ACM Press, New York, NY, USA, 1987. ISBN 0-89791-240-3. LCCN QA76.73.A35 A3 v.7:6. US$14.

Barnes:1988:SIW

[Bar88] John Barnes. Second international workshop on Real-Time Ada issues. ACM SIGADA Ada Letters, 8 (7):??, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:1993:IA

[Bar93] John Barnes. Introducing Ada 9X. ACM SIGADA Ada Letters, 13(6):61–132, November/December 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:1995:ARO

[Bar95] John Barnes. Accessibility rules OK! (Ada 9X). ACM SIGADA Ada Letters, 15(1):39–49, January/ February 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:1998:UAP

[Bar98] John Barnes. Underneath the arch: a personal report of ARG meeting. ACM SIGADA Ada Letters, 18(2): 36–41, March 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2000:SWC

[Bar00] John Barnes. The SPARK way to correctness is via abstraction. ACM SIG-ADA Ada Letters, 20(4):69– 79, December 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.acm.org/sigada/ada_
letters/dec2000/barnes-

paper.pdf. Special Issue: Presentations from SIGAda 2000.

Barkstrom:2001:ABN

[Bar01] Bruce R. Barkstrom. Ada 95 bindings for the NCSA hierarchical data format. ACM SIGADA Ada Letters, 21(4): 27–30, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2007:SIBa

[Bar07a] John G. P. Barnes. SA1: introducing the best of Ada. ACM SIGADA Ada Letters, 27(3):1, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2007:SIBb

[Bar07b] John G. P. Barnes. SP1: introducing the best of Ada 2005. ACM SIG-ADA Ada Letters, 27(3): 3, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bartholomew:2008:ESS

[Bar08] Redge Bartholomew. Evaluation of static source code analyzers for avionics software development. ACM SIGADA Ada Letters, 28(1): 83–87, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barkstrom:2009:UAS

[Bar09a] Bruce R. Barkstrom. On using Ada to solve problems in computational economics and related disciplines with concurrent, multiagent algorithms. ACM SIG-ADA Ada Letters, 29(3):61– 72, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2009:GSSa

[Bar09b] John Barnes. Gem #30:

safe and secure software: introduction. ACM SIG-ADA Ada Letters, 29(1): 45–47, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2009:GSSb

[Bar09c] John Barnes. Gem #32: safe and secure software: chapter 1, safe syntax. ACM SIGADA Ada Letters, 29 (1):50, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2009:GSSc

[Bar09d] John Barnes. Gem #34: safe and secure software: chapter 2, safe typing. ACM SIGADA Ada Letters, 29 (1):53, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2009:GSSd

[Bar09e] John Barnes. Gem #36: safe and secure software: chapter 3, safe pointers. ACM SIGADA Ada Letters, 29 (1):57, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2009:GSSe

[Bar09f] John Barnes. Gem #38: safe and secure software: chapter 4, safe architecture. ACM SIGADA Ada Letters, 29 (1):61, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2009:GSSf

[Bar09g] John Barnes. Gem #40: safe and secure software: chapter 5, safe object oriented programming. ACM SIGADA Ada Letters, 29 (1):65, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2009:GSSg

[Bar09h] John Barnes. Gem #42: safe and secure software: chapter 6, safe object construction. ACM SIGADA Ada Letters, 29(1):69, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2009:GSSh

[Bar09i] John Barnes. Gem #43: safe and secure software: chapter 7, safe memory management. ACM SIGADA Ada Letters, 29(1):70, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2009:GSSi

[Bar09j] John Barnes. Gem #45: safe and secure software: chapter 8, safe startup. ACM SIGADA Ada Letters, 29 (1):74, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2009:GSSj

[Bar09k] John Barnes. Gem #47: safe and secure software: chapter 9, safe communication. ACM SIGADA Ada Letters, 29(1):77, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2009:GSSk

[Bar09l] John Barnes. Gem #49: safe and secure software: chapter 10, safe concurrency. ACM SIGADA Ada Letters, 29 (1):80, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2009:GSSl

[Bar09m] John Barnes. Gem #51: safe and secure software: chapter 11, certified safe with SPARK. ACM SIG-ADA Ada Letters, 29(2): 36, August 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2014:ASA

[Bar14] John Barnes. From Ada 9x to spaceport America: going where no one has gone before. ACM SIGADA Ada Letters, 34(3):1–2, December 2014. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Beretz:1985:DAA

[BB85] Rene Beretz and Benjamin M. Brosgol. Developing an automated Ada training product. ACM SIGADA Ada Letters, 5(2):229–240, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Burns:2002:SSF

[BB02] Alan Burns and Ben Brosgol. Session summary: future of the Ada language and language changes such as the Ravenscar profile. ACM SIG-ADA Ada Letters, 22(4):113– 119, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Battaglia:1997:RAT

[BBB97] D. Battaglia, A. Burke, and J. Beidler. ReUSE/Ada: a tool to promote code reuse. In ACM [ACM97], pages 113– 116. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Battaglia:1998:ARS

[BBB98] David Battaglia, Austin Burke, and John Beidler. An ADA reuse support system for Windows 95/ NT. ACM SIGADA Ada Letters, 18(1):78–85, January/February 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bombardelli:2023:CIC

[BBB^{+}23] A. Bombardelli, A. Bonizzi, M. Bozzano, R. Cavada, A. Cimatti, A. Griggio, M. Nazaria, E. Nicolodi, S. Tonetta, and G. Zampedri. COMPASTA: Integrating COMPASS functionality into TASTE. ACM SIGADA Ada Letters, 43(1):54–57, June 2023. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL

https://dl.acm.org/doi/
10.1145/3631483.3631489.

Belz:1980:MIF

[BBH80] F. C. Belz, E. K. Blum, and D. Heimbigner. A multiprocessing implementationoriented formal definition of Ada in SEMANOL. In ACM [ACM80], pages 202– 212. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Beringer:2012:PCC

[BBPT12] Lennart Beringer, Randall Brukardt, Thomas Plum, and S. Tucker Taft. Panel on compiler certification: should we trust our compiler? ACM SIGADA Ada Letters, 32 (3):103–104, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Burns:1997:TPS

[BBV97] Alan Burns, Ted Baker, and Tullio Vardenaga. Tasking profiles (session summary). ACM SIGADA Ada Letters, 17(5):5–7, September/October 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Botting:1995:AUD

[BC95] Paul Botting and Eugene Clayton. Ada used to develop visual and sensor displays. ACM SIGADA Ada Letters, 15(4):19–21, July/August 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Broster:2011:HMO

[BC11] Ian Broster and Andrew Coombes. How to measure and optimize reliable embedded software. ACM SIG-ADA Ada Letters, 31(3):1– 2, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brandon:2016:USC

[BC16] Carl Brandon and Peter Chapin. The use of SPARK in a complex spacecraft. ACM SIGADA Ada Letters, 36(2):18–21, December 2016. CODEN AALEE5. ISSN 0736-721X.

Bagnato:2022:AAM

[BCB^{+}22] Alessandra Bagnato, Antonio Cicchetti, Luca Berardinelli, Hugo Bruneliere, and Romina Eramo. AIaugmented model-based capabilities in the AIDOaRt Project: Continuous development of cyber-physical systems. ACM SIGADA Ada Letters, 42(2):99–103, December 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3591335.3591349.

Bossi:1983:MDA

[BCD83] A. Bossi, N. Cocco, and S. Dulli. Modular decomposition of Ada into a hierarchy of sublanguages. ACM SIGADA Ada Letters, 2(6): 53–58, May/June 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Blazquez:1994:AAS

[BCF94] V. Blázquez, A. Correa, and J. L. Freniche. Advancing Ada 9X solutions in real time avionics computers. ACM SIGADA Ada Letters, 14(5):80–87, September/October 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Basili:1984:MAS

[BCG^{+}84] Victor R. Basili, Shih Chang, John Gannon, Elizabeth Katz, N. Monina Panlilo-Yap, Connie Loggia Ramsey, Marvin Zelkowitz, John Bailey, Elizabeth Kruesi, and Sylvia Sheppard. Monitoring an Ada software development. ACM SIG-ADA Ada Letters, 4(1):32– 39, July/August 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Berns:2019:MSD

[BCH^{+}19] Andrew Berns, James Curbow, Joshua Hilliard, Sheriff Jorkeh, and Miho Sanders. Minimal specifications for detecting security vulnerabilities. ACM SIGADA Ada Letters, 38(2):109–114, December 2019. ISSN 0736-721X.

URL https://dl.acm.org/
doi/abs/10.1145/3375408.
3375417.

Belt:2012:LEA

[BCHR12] Jason Belt, Patrice Chalin, John Hatcliff, and Robby. Leading-edge Ada verification technologies: highly automated Ada contract checking using Bakar Kiasan. ACM SIGADA Ada Letters, 32(3): 3–4, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Blouin:2023:IAU

[BCMC23] Dominique Blouin, Paolo Crisafulli, Cristian Maxim, and Francoise Caron. An introduction to ALISA and its usage for an industrial railway system case study. ACM SIGADA Ada Letters, 43(1):69–72, June 2023. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3631483.3631493.

Bardin:1989:IUI

[BCS89] B. Bardin, C. Colket, and D. Smith. Implementation of unsigned integers in Ada. ACM SIGADA Ada Letters, 9(1):47–70, January/February 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Basson:1991:QTE

[BD91] H. Basson and J. C. Derniame. Quality tree extensions and partial instantiation for Ada objects. In ACM [ACM91b], pages 156– 171. ISBN 0-89791-393-0. LCCN ????

Burns:1992:APT

[BD92] A. Burns and G. L. Davies. Ada 9X protected types in pascal-FC. ACM SIGADA Ada Letters, 12(6):59–74, November/December 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bernstein:1999:OAF

[BD99] Sheri J. Bernstein and Robert S. Duff. Optimizing Ada on the fly. ACM SIG-ADA Ada Letters, 19(3):169– 179, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:2001:RTC

[BD01] Ben Brosgol and Brian Dobbing. Real-time convergence of Ada and Java^{T M}. ACM SIGADA Ada Letters, 21(4): 11–26, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bever:1982:IED

[BDD^{+}82] M. Bever, M. Dausmann, S. Drossopoulou, W. Kirchgassner, P. C. Lockemann, G. Persch, and G. Winterstein. The integration of existing database systems in an Ada environment. In ACM [ACM82], page ?? ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Braesicke:1985:FAE

[BDF^{+}85] Carl Braesicke, Jeff Dean, Dave Fisher, Jim Holder, Rand McKinney, Panna Nagarsenker, Dewayne Perry, Phil Rossomando, Tim Standish, and Dick Wisehart. Future Ada environments workshop: User interfaces. ACM SIGADA Ada Letters, 4(5): 90–96, March/April 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2015:SSC

[BdlP15] Alan Burns and Juan Antonio de la Puente. Session summary: Conformance issues. ACM SIG-ADA Ada Letters, 35(1): 95–96, April 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bradley:2010:RTS

[BdlPZ10] Peter J. Bradley, Juan A. de la Puente, and Juan Zamorano. Real-time system development in Ada using LEGO(R) Mindstorms(R) NXT. ACM SIGADA Ada Letters, 30(3):37–40, December 2010. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Buxton:1981:RHA

[BDS81] John N. Buxton, Larry E. Druffel, and Thomas A. Standish. Reflections on the history of Ada environments. ACM SIG-ADA Ada Letters, 1(1):16– 21, July/August 1981. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brukardt:1999:ACA

[BDT99] Randall Brukardt, Steven Deller, and Joyce L. Tokar. Ada 95 conformity assessment. ACM SIGADA Ada Letters, 19(1):52–57, March 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2004:GUA

[BDV04] Alan Burns, Brian Dobbing, and Tullio Vardanega. Guide for the use of the Ada Ravenscar Profile in high integrity systems. ACM SIGADA Ada Letters, 24 (2):1–74, June 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1991:AA

[BE91] Alan Burns and William Eventoff. Asynchronism in Ada 9X. ACM SIGADA Ada Letters, 11(6):66–68, September/October 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brach:2002:UEA

[BE02] David Brach and P. Eng.

User experiences with the Aonix ObjectAda RAVEN: Ravenscar Profile implementation. ACM SIG-ADA Ada Letters, 22(4):10– 21, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Becker:1983:AES

[Bec83] Lee A. Becker. Ada — extended structure charts. ACM SIGADA Ada Letters, 3(2):93–97, September/October 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bein:1984:ADJ

[Bei84]  Edward Bein.  Ada de-

sign, Jovial implementation. ACM SIGADA Ada Letters, 3(4):62–69, January/February 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Beidler:1992:RCA

[Bei92] John Beidler. Relaxing the constraints on Ada’s limited private types through functional expressions. ACM SIGADA Ada Letters, 12 (2):57–61, March/April 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Beidler:1997:AC

[Bei97] Jack Beidler. Ada in concert. ACM SIGADA Ada Letters, 17(3):57–66, May/June 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Belmont:1980:TRA

[Bel80] Peter A. Belmont. Type resolution in Ada: An implementation report. In ACM [ACM80], pages 57– 61. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Belmont:1982:APA

[Bel82] P. A. Belmont. On the accessbefore-elaboration problem in Ada. In ACM [ACM82], pages 112–119. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Bengel:1984:PA

[Ben84] G. G. Bengel. Peculiarities of Ada. ACM SIG-ADA Ada Letters, 3(5):75– 81, March/April 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bennett:1994:SDC

[Ben94] P. A. Bennett. Software development for the Channel Tunnel: a summary. ACM SIGADA Ada Letters, 14(6):73–76, November/December 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Berard:1983:EA

[Ber83] Edward V. Berard. Engineering Ada. ACM SIGADA Ada Letters, 3(3):33–44, November/December 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Berard:1984:AEM

[Ber84] Edward V. Berard. Ada education is a moving target. ACM SIGADA Ada Letters, 4(1):45–49, July/ August 1984. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Berard:1986:TSP

[Ber86a] Edward V. Berard. Towards a software profession. ACM SIGADA Ada Letters, 6(1):29–40, January/February 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bernard:1986:DRM

[Ber86b] L. Bernard. Dereference the reference manual. ACM SIG-ADA Ada Letters, 6(3):56– 60, May/June 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Berns:2005:CCA

[Ber05] Andrew Berns. A comparison of CORBA and Ada’s distributed systems annex. ACM SIGADA Ada Letters, 25(4):103–108, December 2005. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bernardi:2015:ICT

[Ber15] Patrick Bernardi. Incorporating cyclic task behaviour into Ada tasks. ACM SIG-ADA Ada Letters, 35(1): 59–73, April 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baskette:1986:LCA

[BF86] Jerry Baskette and John Foreman. Life cycle analysis of the AIM project. ACM SIGADA Ada Letters, 6(2): 86–90, March/April 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Buhler:1999:AAJ

[BF99] Gerhard Bühler and Heinz Faßbender. Applying Ada, Java and CORBA for making a command and control information system platform independent. ACM SIG-ADA Ada Letters, 19(3):83– 88, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bassman:1985:AEP

[BFG85] Mitchell J. Bassman, Gerald A. Fisher, Jr., and Anthony Gargaro. An approach for evaluating the performance efficiency of Ada compilers. ACM SIGADA Ada Letters, 5(2):151–163, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Borger:1990:AUP

[BG90] M. W. Borger and J. B. Goodenough. Ada usage/ performance specification. ACM SIGADA Ada Letters, 10(9):65–69, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bocchino:2014:SPL

[BGGS14] Robert L. Bocchino, Edward Gamble, Kim P. Gostelow, and Raphael R. Some. Spot: a programming language for verified flight software. ACM SIGADA Ada Letters, 34(3): 97–102, December 2014. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Basili:1982:MAS

[BGK^{+}82] Victor Basili, John Gannon, Elizabeth Katz, Marvin Zelkowitz, John Bailey, Elizabeth Kruesi, and Sylvia Sheppard. Monitoring an Ada software development project. ACM SIG-ADA Ada Letters, 2(1):58– 61, July/August 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Byrne:1990:AVF

[BH90] Dan J. Byrne and Richard C. Ham. Ada versus FOR-TRAN: Performance analysis using the ACPS. ACM SIGADA Ada Letters, 10(3):

139–145, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:2002:SSU

[BH02] Ben Brosgol and Michael González Harbour. Session summary: update on the real-time specification for Java. ACM SIG-ADA Ada Letters, 22(4):128– 130, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bagge:2014:SGA

[BH14] Anya Helene Bagge and Magne Haveraaen. Specification of generic APIs, or: why algebraic may be better than pre/post. ACM SIG-ADA Ada Letters, 34(3):71– 80, December 2014. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barkataki:1998:RLS

[BHD98] Shan Barkataki, Stu Harte, and Tong Dinh. Reengineering a legacy system using design patterns and Ada 95 object-oriented features. ACM SIGADA Ada Letters, 18(6):148–151, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brown:1993:ART

[BHL^{+}93] Mike Brown, Walter Heimerdinger, Nancy Leveson, John McHugh, Arch McKinlay, and George Romanski. Ada Run-Time Environment Working Group: proceedings from the software safety symposium. ACM SIGADA Ada Letters, 13(1):35–59, January/February 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Beyene:2020:VAP

[BHN20] Tewodros A. Beyene, Christian Herrera, and Vivek Nigam. Verification of Ada programs with AdaHorn. ACM SIGADA Ada Letters, 39(2):29–34, April 2020. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3394514.3394517.

Brosgol:2002:ATC

[BHR02] Benjamin M. Brosgol, Ricardo J. Hassan, II, and Scott Robbins. Asynchronous transfer of control in the real-time specification for Java^{T M}. ACM SIG-ADA Ada Letters, 22(4):95– 112, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Belt:2011:ESC

[BHR^{+}11] Jason Belt, John Hatcliff, Robby, Patrice Chalin, David Hardin, and Xianghua Deng. Enhancing SPARK’s contract checking facilities using symbolic execution. ACM SIG-ADA Ada Letters, 31(3):47– 60, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bishop:1980:EMD

[Bis80] Judy M. Bishop. Effective machine descriptors for Ada. In ACM [ACM80], pages 235–242. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Bishop:1986:CNA

[Bis86] Judy M. Bishop. A complete notation for Ada charts. ACM SIGADA Ada Letters, 6(6):49–53, November/December 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bishop:1988:TSD

[Bis88] Judy M. Bishop. Three steps to distribution: partitioning, configuring, and adapting. ACM SIGADA Ada Letters, 8(7):97–100, Fall 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bishop:1991:DAD

[Bis91] J. Bishop. Distributed Ada: Developments and experiences. ACM SIGADA Ada Letters, 11(1):121–??, January/February 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bjorner:2013:SMT

[Bjo13] Nikolaj Bjorner. Satisfiability modulo theories for high integrity development. ACM SIGADA Ada Letters, 33(3): 5–6, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briggs:1996:TTL

[BJRW96] J. S. Briggs, S. D. Jamieson, G. W. Randall, and I. C. Wand. Task time lines as a debugging tool. ACM SIG-ADA Ada Letters, 16(2):50– 69, March/April 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Buhr:1985:IOC

[BK85] R. J. A. Buhr and G. M. Karam. An informal overview of CADA: a design environment for Ada. ACM SIG-ADA Ada Letters, 4(5):49– 58, March/April 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bagnato:2022:MOD

[BK22] Alessandra Bagnato and Józefina Krasnodebska. MOR-PHEMIC — optimization of the deployment and lifecycle management of dataintensive applications in the cloud computing continuum. ACM SIGADA Ada Letters, 42(2):104–108, December 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL

https://dl.acm.org/doi/
10.1145/3591335.3591350.

Buhr:1991:SST

[BKC91] R. J. A. Buhr, G. M. Karam, and R. Casselman. Support for specifying temporal behavior in Ada designs. ACM SIGADA Ada Letters, 11(3): 91–101, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Berecz:1985:DE

[BKL85] Vic Berecz, Jack Kramer, and Carol LeDoux. Distributed environments. ACM SIG-ADA Ada Letters, 4(5):84– 89, March/April 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Future Ada Environment Workshop.

Back:1987:NPD

[BKS87] Ralph Back and Reino Kurki-Suonio. A new paradigm for the design of concurrent systems. ACM SIG-ADA Ada Letters, 7(6): 110–112, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bennett:1982:HCA

[BKW82] David A. Bennett, Brent D. Kornman, and James R. Wilson. Hidden costs in Ada. ACM SIGADA Ada Letters, 1(4):9–20, May/June 1982. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Buhr:1985:OEA

[BKW85] R. J. A. Buhr, G. M. Karam, and C. M. Woodside. An overview and example of application of CAEDE: a new, experimental design environment for Ada. ACM SIGADA Ada Letters, 5(2):173–184, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Buhr:1994:TCT

[BKW^{+}94] R. J. A. Buhr, G. M. Karam, C. M. Woodside, R. Casselman, G. Franks, H. Scott, and D. Bailey. TimeBench: a CAD tool for real-time system design. ACM SIGADA Ada Letters, 14(Special Issue):3–15, Fall 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Borger:1988:TIR

[BKWS88] Mark Borger, Mark Klein, Nelson Weiderman, and Lui Sha. A testbed for investigating Real-Time Ada issues. ACM SIGADA Ada Letters, 8(7):7–11, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burkhard:1986:DAS

[BL86] B. Burkhard and M. Lee.

Drawing Ada structure charts. ACM SIGADA Ada Letters, 6(3):71–80, May/June 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Black:2007:SAS

[Bla07] Paul E. Black. Static analysis summit II. ACM SIG-ADA Ada Letters, 27(3):101– 107, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bardin:1985:SRA

[BM85] Bryce M. Bardin and Marion F. Moon. In search of “real” Ada: a software saga with a moral or two. ACM SIGADA Ada Letters, 5(2):217–228, September/October 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Brukhardt:1997:CHL

[BM97] R. Brukhardt and T. Moran. CLAW, a high level, portable, Ada 95 binding for Microsoft Windows. In ACM [ACM97], pages 91–104. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Bardaro:2023:MRA

[BM23] Gianluca Bardaro and Matteo Matteucci. Modelling robot architectures with AADL. ACM SIG-ADA Ada Letters, 43(1): 59–63, June 2023. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3631483.3631491.

Bramberger:2020:CES

[BMGS20] Robert Bramberger, Helmut Martin, Barbara Gallina, and Christoph Schmittner. Coengineering of safety and security life cycles for engineering of automotive systems. ACM SIGADA Ada Letters, 39(2):41–48, April 2020. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3394514.3394519.

Barbacci:1985:AFE

[BMNS85] M. R. Barbacci, W. H. Maddox, T. D. Newton, and R. G. Stockton. The Ada+ front end and code generator. ACM SIGADA Ada Letters, 5(2):343–354, September/October 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Bocchino:2014:PSF

[BMT^{+}14] Robert Bocchino, Nicholas Matsakis, S. Tucker Taft, Brian Larson, and Ed Seidewitz. Panel summary: finding safety in numbers: new languages for safe multicore programming and modeling. ACM SIGADA Ada Letters, 34(3):105–106, December 2014. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barry:1994:DSS

[BMW94] Brian M. Barry, James McGugan, and Mike Wilson. DIR/SEE: a Smalltalk environment for developing Ada applications and maintaining legacies. ACM SIGADA Ada Letters, 14(Special Issue):26–35, Fall 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burger:1987:AOA

[BN87] Thomas. M. Burger and Kjell W. Nielsen. An assessment of the overhead associated with tasking facilities and task paradigms in Ada. ACM SIGADA Ada Letters, 7(1):49–58, January/February 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bodeau:2019:CRO

[Bod19] Deborah Bodeau. Cyber resiliency overview: What is it, and how do we build it into our systems? ACM SIGADA Ada Letters, 38(2):58–63, December 2019. ISSN 0736-

                                        721X. URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3375408.3375412.

Boeing:1990:ACE

[Boe90]  Boeing.  The Ada com-

piler evaluation capability (ACEC). ACM SIG-ADA Ada Letters, 10(3): 101–??, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Boehm:1999:PFC

[Boe99] Barry Boehm. Predicting the future of computer systems and software engineering (keynote address). ACM SIGADA Ada Letters, 19(3): 227, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Baker:1997:LLA

[BOM97] T. P. Baker, Dong-Ik Oh, and Seung-Jin Moon. Lowlevel Ada tasking support for GNAT — performance and portability improvements. ACM SIGADA Ada Letters, 17(3):36–44, May/June 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bond:1984:APD

[Bon84] Rodney M. Bond. Ada as a program description language (PDL). ACM SIG-ADA Ada Letters, 4(1):67– 73, July/August 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Booch:1982:OOD

[Boo82] Grady Booch. Object oriented design. ACM SIG-ADA Ada Letters, 1(3):64– 76, March/April 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Booch:2011:EKL

[Boo11] Grady Booch. Everything I know I learned from Ada. ACM SIGADA Ada Letters, 31(3):17–18, December 2011. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bosch:2012:SCI

[Bos12] Geert Bosch. Synchronization cannot be implemented as a library. ACM SIG-ADA Ada Letters, 32(3):73– 80, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Bosch:2013:LFP

[Bos13]  Geert Bosch.  Lock-free

protected types for realtime Ada. ACM SIG-ADA Ada Letters, 33(2): 66–74, August 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Botton:1999:DA

[Bot99a] David Botton. Dear Ada. ACM SIGADA Ada Letters, 19(1):108–??, March 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Botton:1999:IAM

[Bot99b] David Botton. Interfacing Ada 95 to Microsoft COM and DCOM technologies. ACM SIGADA Ada Letters, 19(3):9–14, September 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Botton:2000:AN

[Bot00a]  David Botton.  Ada on

the NET! ACM SIG-ADA Ada Letters, 20(3):50– 52, September 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.acm.org/sigada/ada_
letters/sept2000/ada_on_
the_net.pdf.

Botton:2000:DA

[Bot00b] David Botton. Dear Ada.

ACM SIGADA Ada Letters, 20(3):53–56, September 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/sept2000/dear_ada_
sep2000.pdf.

Bowen:1992:ODP

[Bow92] Gregory M. Bowen. An organized, devoted, project-wide reuse effort. ACM SIG-ADA Ada Letters, 12(1):43– 52, January/February 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Boyd:1986:ABW

[Boy86] Stowe Boyd. APSE builders’ working group report. ACM SIGADA Ada Letters, 6(2): 79–82, March/April 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Boyd:1987:OOD

[Boy87] Stowe Boyd. Object-oriented design and Pamela: a comparison of two design methods for Ada. ACM SIG-ADA Ada Letters, 7(4):68– 78, July/August 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Boyd:1989:RAC

[Boy89] Stowe Boyd. The role of Ada in contemporary interface technology. ACM SIG-ADA Ada Letters, 9(5):115– 122, July/August 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brown:1994:EIW

[BP94] Alan W. Brown and Maria H. Penedo. “environment integration” working group summary SETA2. ACM SIG-ADA Ada Letters, 14(Special Issue):85–92, Fall 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barros:2013:RTA

[BP13] António Barros and Luís Miguel Pinho. Revisiting transactions in Ada. ACM SIG-ADA Ada Letters, 33(1): 84–92, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barbaria:2006:SMS

[BPP06] Khaled Barbaria, Laurent Pautet, and Isabelle Perseil. Schizophrenic middleware support for fault tolerance. ACM SIGADA Ada Letters, 26(3):51–60, December 2006. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1990:EUA

[BQ90] A. Burns and T. J. Quiggle. Effective use of abort in programming mode changes. ACM SIGADA Ada Letters, 10(6):61–67, July/August 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bruno:1994:ICR

[BR94] Jeanette M. Bruno and Daniel J. Rosenkrantz. Interactive control restructuring. ACM SIGADA Ada Letters, 14(Special Issue): 36–53, Fall 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2001:HEE

[BR01] Alan Burns and Alexander Romanovsky. How to evolve exception handling in Ada. ACM SIGADA Ada Letters, 21(3):16–18, September 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bray:1982:ASM

[Bra82] Gary Bray. AIE support for management of embedded computer projects. ACM SIGADA Ada Letters, 2(1): 33–49, July/August 1982. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Braun:1983:ATC

[Bra83a] Christine L. Braun. Ada training considerations. ACM SIGADA Ada Letters, 2(5): 42–55, March/April 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bray:1983:IIA

[Bra83b] Gary Bray. Implementation implications of Ada generics. ACM SIGADA Ada Letters, 3(2):62–71, September/October 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brandon:1985:TGT

[Bra85] C. Brandon. Turtle graphics for teaching Ada as a first language. ACM SIGADA Ada Letters, 5(3–6):100, November/December 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brashear:1994:ACE

[Bra94] Phil Brashear. The Ada Compiler Evaluation System.

ACM SIGADA Ada Letters, 14(2):68–79, March/April 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brashear:1998:AIS

[Bra98] Phil Brashear. The Ada issues: a special section. ACM SIGADA Ada Letters, 18(3):17, May 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brashear:1999:AVA

[Bra99] Phil Brashear. Ada validation := Ada conformity assessment. ACM SIG-ADA Ada Letters, 19(1): 48–51, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Blake:1998:ARW

[BRC98] Stephen Blake, Clyde G.

Roby, Jr., and William Currie Colket. ASIS Report for WG9 Meeting on 12 June 1998, Uppsala, Sweden. ACM SIGADA Ada Letters, 18(4): 111–113, July 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bremmon:1997:WOA

[Bre97] Chad Bremmon. Writing an OLE automation controller in Ada95. ACM SIG-ADA Ada Letters, 17(3):45– 56, May/June 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Blazquez:1992:EDU

[BRF92] V. Blázquez, L. Redondo, and J. L. Freniche. Experiences with “delay until” for Avionics computers. ACM SIGADA Ada Letters, 12(1):65–72, January/February 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brintzenhoff:1986:CL

[Bri86] Alton L. Brintzenhoff. Chairperson’s letter. ACM SIG-ADA Ada Letters, 6(2):53– 56, March/April 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briand:1992:TMA

[Bri92a] L. Briand. Time management for Ada real-time systems. ACM SIGADA Ada Letters, 12(5):84–95, September/October 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briand:1992:TMR

[Bri92b] Loïc Briand. Time management for real-time systems. ACM SIGADA Ada Letters, 12(5):84–95, September/October 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briand:1994:ART

[Bri94]  Loïc Briand.  Ada real-

time systems and basic priority inheritance. ACM SIG-ADA Ada Letters, 14(3):105– 112, May/June 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2009:GHS

[Bri09a] Emmanuel Briot. Gem #25:

how to search text. ACM SIGADA Ada Letters, 29(1): 29–32, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2009:GSCa

[Bri09b]  Emmanuel Briot.  Gem

#52: scripting capabilities in GNAT (part 1). ACM SIGADA Ada Letters, 29(2): 37–39, August 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2009:GSCb

[Bri09c]  Emmanuel Briot.  Gem

#54: scripting capabilities in GNAT (part 2). ACM SIGADA Ada Letters, 29(2): 40–42, August 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2009:GHM

[Bri09d]  Emmanuel Briot.  Gem

#64: handling multiple-unit source files. ACM SIG-ADA Ada Letters, 29(2): 68–70, August 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2011:GG

[Bri11a]  Emmanuel Briot.  Gem

#65: gprbuild. ACM SIG-ADA Ada Letters, 31(1): 11–13, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2011:GGK

[Bri11b] Emmanuel Briot. Gem #66: GPS’s key shortcuts editor. ACM SIGADA Ada Letters, 31(1):14–15, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2011:GMG

[Bri11c]  Emmanuel Briot.  Gem

#67: managing the GPS workspace. ACM SIG-ADA Ada Letters, 31(1): 16–18, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2011:GWDa

[Bri11d]  Emmanuel Briot.  Gem

#77: where did my memory go? (part 1). ACM SIGADA Ada Letters, 31(2): 23–24, August 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2011:GWDb

[Bri11e]  Emmanuel Briot.  Gem

#78: where did my memory go? (part 2). ACM SIGADA Ada Letters, 31(2): 25–27, August 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2011:GWDc

[Bri11f]  Emmanuel Briot.  Gem

#79:: where did my memory go? (part 3). ACM SIGADA Ada Letters, 31(2): 28–29, August 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2012:GRCc

[Bri12a] Emmanuel Briot. Gem #100:

reference counting in Ada — part 3: weak references. ACM SIGADA Ada Letters, 32(2): 33–34, August 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2012:GLAa

[Bri12b] Emmanuel Briot. Gem #105:

Lady Ada kisses Python — part 1. ACM SIG-ADA Ada Letters, 32(2): 45–46, August 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2012:GLAb

[Bri12c] Emmanuel Briot. Gem #106: Lady Ada kisses Python — part 2. ACM SIG-ADA Ada Letters, 32(2): 47–49, August 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2012:GRCa

[Bri12d] Emmanuel Briot. Gem #97: reference counting in Ada — part 1. ACM SIG-ADA Ada Letters, 32(2):24– 27, August 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2012:GRCb

[Bri12e] Emmanuel Briot. Gem #99: reference counting in Ada — part 2: task safety. ACM SIGADA Ada Letters, 32(2): 31–32, August 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bordis:2022:CCO

[BRKS22] Tabea Bordis, Tobias Runge, Alexander Kittelmann, and Ina Schaefer. Correctness-by-construction: an overview of the CorC ecosystem. ACM SIGADA Ada Letters, 42 (2):75–78, December 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3591335.3591343.

Brosgol:1980:TMP

[Bro80] Benjamin M. Brosgol. TCOL-Ada and the “middle-end” of the PQCC Ada compiler. In ACM [ACM80], pages 101–112. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Brosgol:1982:SAL

[Bro82] Benjamin Brosgol. Summary of Ada language changes. ACM SIGADA Ada Letters, 1(3):34–43, March/April 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:1983:AIN

[Bro83] Ben Brosgol. Ada implementation notes: Constraint check elimination. ACM SIG-ADA Ada Letters, 2(4):54– 57, January/February 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:1988:IWR

[Bro88] Benjamin Brosgol. International workshop on realtime Ada issues: summary report. ACM SIGADA Ada Letters, 8(1):91–107, January/February 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brookman:1991:SSV

[Bro91] David Brookman. SA/SD vs. OOD. ACM SIGADA Ada Letters, 11(9):96–99, November/December 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:1996:ACW

[Bro96] Benjamin M. Brosgol. Ada-COBOL working group liaison report. ACM SIG-ADA Ada Letters, 16(1):36– 43, January/February 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:1997:COF

[Bro97] B. M. Brosgol. A comparison of the object-oriented features of Ada 95 and Java[TM]. In ACM [ACM97], pages 213–230. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Brosgol:1998:CAJ

[Bro98a] Benjamin M. Brosgol. A comparison of Ada and Java as a foundation teaching language. ACM SIGADA Ada Letters, 18(5):12–38, September/October 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:1998:CCF

[Bro98b] Benjamin M. Brosgol. A comparison of the concurrency features of Ada 95 and Java. ACM SIGADA Ada Letters, 18(6):175–192, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:1999:MC

[Bro99] Ben Brosgol. Message from the Chair. ACM SIG-ADA Ada Letters, 19(1): 1–??, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:2000:MCa

[Bro00a] Ben Brosgol. Message from the Chair. ACM SIG-ADA Ada Letters, 20(1): 1–2, March 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:2000:MCb

[Bro00b] Ben Brosgol. Message from the Chair. ACM SIG-ADA Ada Letters, 20(2): 1–2, June 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.acm.org/sigada/ada_
letters/june2000/ChairLetterJune2000AdaLetters.
pdf.

Brosgol:2000:MCc

[Bro00c] Ben Brosgol. Message from the Chair. ACM SIG-ADA Ada Letters, 20(3):1– 2, September 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:2000:MCd

[Bro00d] Ben Brosgol. Message from the Chair. ACM SIG-ADA Ada Letters, 20(4):1– 2, December 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:2001:MC

[Bro01] Ben Brosgol. Message from the Chair. ACM SIG-ADA Ada Letters, 21(2): 1–2, June 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brooke:2003:DDC

[Bro03] Thomas C. Brooke. Development of a distributed, crossplatform simulator. ACM SIGADA Ada Letters, 23(1): 12–21, March 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brown:2004:TGU

[Bro04] C. Wayne Brown. Teaching graphics using Ada. ACM SIGADA Ada Letters, 24(4):

47–50, December 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:2007:SLS

[Bro07] Ben Brosgol. SA2: languages for safety-critical software: issues and assessment. ACM SIGADA Ada Letters, 27(3):2, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:2009:ICL

[Bro09] Ben Brosgol. An introduction to the C# language and .NET infrastructure. ACM SIGADA Ada Letters, 29(3): 3–4, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:2011:DNA

[Bro11]  Benjamin Brosgol.  DO-

178C: the next avionics safety standard. ACM SIG-ADA Ada Letters, 31(3):5– 6, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bruno:1982:APD

[Bru82] G. Bruno. An Ada package for discrete event simulation. In ACM [ACM82], pages 172–180. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Brukardt:2017:CIM

[Bru17] Randy Brukardt. Community input for the maintenance and revision of the Ada programming language. ACM SIGADA Ada Letters, 37(1):54, June 2017. CODEN AALEE5. ISSN 0736-721X.

Blair:1997:UCS

[BRW97] J. R. S. Blair, E. K. Ressler, and T. D. Wagner. The undergraduate Capstone software design experience. In ACM [ACM97], pages 41–50. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Brykczynski:1988:MBA

[Bry88] Bill Brykczynski. Methods of binding Ada to SQL: a general discussion. ACM SIG-ADA Ada Letters, 8(1):38– 51, January/February 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bryan:1990:DAa

[Bry90a] Doug Bryan. Dear Ada. ACM SIGADA Ada Letters, 10(5): 41–47, May/June 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bryan:1990:DAb

[Bry90b] Doug Bryan. Dear Ada. ACM SIGADA Ada Letters, 10(8):24–33, November/December 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bail:2001:EP

[BS01] William Bail and Bo I.

Sandén. Exception propagation. ACM SIGADA Ada Letters, 21(3):8–10, September 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Boleng:2013:SOA

[BS13] Jeff Boleng and Ricky Sward.

Service-oriented architecture (SOA) concepts and implementations. ACM SIG-ADA Ada Letters, 33(3):11– 12, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Balador:2022:DPD

[BSPK22] Ali Balador, Sima Sinaei, Mats Pettersson, and Ilhan Kaya. DAIS Project — distributed artificial intelligence systems: Objectives and challenges. ACM SIGADA Ada Letters, 42 (2):96–98, December 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3591335.3591348.

Bar:1990:SAA

[BST90] Dieter Bär, Klaus Sum, and Léon Treff. SQL ArmAda:

An Ada-appropriate interface to SQL. ACM SIG-ADA Ada Letters, 10(2):64– 83, March/April 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bardin:1988:CAS

[BT88a] Bryce Bardin and Christopher Thompson. Composable Ada software components and the re-export paradigm. ACM SIGADA Ada Letters, 8(1):58–79, January/February 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bardin:1988:URE

[BT88b] Bryce M. Bardin and Christopher J. Thompson. Using the re-export paradigm to build composable Ada software components. ACM SIG-ADA Ada Letters, 8(2):39– 54, March/April 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Barnes:2014:AAL

[BT14] John Barnes and S. Tucker Taft. Ada 83 to Ada 2012: lessons learned over 30 years of language design. ACM SIGADA Ada Letters, 34(3): 3–4, December 2014. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2010:ASV

[BTB^{+}10] Alan Burns, Joyce L. Tokar, Stephen Baird, John Barnes, Rod Chapman, Gary Dismukes, Michael Gonzales-Harbour, Stephen Michell, Brad Moore, Miguel Pinho, Erhard Ploedereder, Jorge Real, J. P. Rosen, Ed Schonberg, S. Tucker Taft, and T. Vardanega. Ada and the software vulnerabilities project. ACM SIG-ADA Ada Letters, 30(2): 27–52, August 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Boukili:2022:FGR

[BTP22] Zineb Boukili, Hai Nam Tran, and Alain Plantec. Finegrained runtime monitoring of real-time embedded systems. ACM SIGADA Ada Letters, 42(1):105, June 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3577949.3577970.

Ballbastre:1999:EUA

[BTVC99] P. Ballbastre, S. Terrasa, J. Vila, and A. Crespo. Experiences using Ada in a real-time and distributed laboratory. ACM SIG-ADA Ada Letters, 19(3):145– 155, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Buchman:1987:DAA

[Buc87]  Brett Buchman.  Design

automation for Ada development under DOD-STD-2167 (and beyond). In ACM [ACM87a], pages 75– 80. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Buhr:1985:LPE

[Buh85] R. J. A. Buhr. Lessons from practical experience teaching hands-on, real-time, embedded system programming with Ada. ACM SIGADA Ada Letters, 5(2):210–216, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Bundgaard:1985:DAF

[Bun85]  J. Bundgaard.  The de-

velopment of an Ada front end for small computers. ACM SIGADA Ada Letters, 5(2):321–328, September/October 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Burkhardt:1985:FUX

[Bur85a] Bonnie Burkhardt. First use of XAda methodology. ACM SIGADA Ada Letters, 5(1): 79–88, July/August 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1985:EIR

[Bur85b] A. Burns. Efficient initialisation routines for multiprocessor systems programmed in Ada. ACM SIGADA Ada Letters, 5(1):55–60, July/August 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1987:ULF

[Bur87a] A. Burns. Using large families for handling priority requests. ACM SIGADA Ada Letters, 7(1):97–104, January/February 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1987:CDR

[Bur87b] Greg Burns. Cross-debugging Real-Time Ada programs. ACM SIGADA Ada Letters, 7(6):21–23, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1990:PSA

[Bur90] A. Burns. A performance standard for Ada 9X. ACM SIGADA Ada Letters, 10 (9):70–74, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burger:1992:OIR

[Bur92] Tom Burger. Optimization issues relating to subunits. ACM SIGADA Ada Letters, 12(3):99–109, May/June 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1999:RP

[Bur99a] Alan Burns. The Ravenscar Profile. ACM SIG-ADA Ada Letters, 19(4):49– 52, December 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1999:RPI

[Bur99b] Alan Burns. The Ravenscar Profile and implementation issues (session summary). ACM SIGADA Ada Letters, 19(2):12–14, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2001:NPD

[Bur01] Alan Burns. Non-preemptive dispatching and locking policies. ACM SIGADA Ada Letters, 21(1):46–47, March 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2013:ERT

[Bur13a] A. Burns. An EDF run-time profile based on Ravenscar. ACM SIGADA Ada Letters, 33(1):24–31, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2013:PAR

[Bur13b] A. Burns. Parallel Ada: a requirement for Ada 2020. ACM SIGADA Ada Letters, 33(2):9–13, August 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Buxton:1985:FAE

[Bux85a]  J. N. Buxton.  Future

Ada environment workshop: keynote address. ACM SIG-ADA Ada Letters, 4(5):40– 44, March/April 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Buxton:1985:KAF

[Bux85b] John N. Buxton. Keynote address, future APSE workshop. ACM SIGADA Ada Letters, 4(5):40–44, March/April 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Buzdalov:2016:SAM

[Buz16] Denis Buzdalov. Simulation of AADL models with software-in-the-loop execution. ACM SIGADA Ada Letters, 36(2):49–53, December 2016. CODEN AALEE5. ISSN 0736-721X.

Burns:2003:RSG

[BV03] Alan Burns and Tullio Vardanega. Report of session: generating new AIs. ACM SIGADA Ada Letters, 23(4): 93–95, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2013:SSLa

[BV13] Alan Burns and Tullio Vardanega. Session summary: language profile and application frameworks. ACM SIGADA Ada Letters, 33(1): 146–149, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1987:RTA

[BW87] A. Burns and A. J. Wellings. Real-Time Ada issues. ACM SIGADA Ada Letters, 7(6): 43–46, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1989:PAA

[BW89] A. Burns and A. J. Wellings. Programming atomic actions in Ada. ACM SIGADA Ada Letters, 9(6):67–79, September/October 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bums:1990:RTA

[BW90a] A. Bums and A. J. Wellings.

Real-time Ada: outstanding problem areas. ACM SIGADA Ada Letters, 10(4): 5–14, Spring 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bums:1990:UAT

[BW90b] A. Bums and A. J. Wellings.

Usability of the Ada taskingmodel. ACM SIGADA Ada Letters, 10(4):49–56, Spring 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1990:RTA

[BW90c] A. Burns and A. J. Wellings.

Real-Time Ada: Outstanding problem areas. ACM SIGADA Ada Letters, 10(4): 5–14, Spring 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1990:UAT

[BW90d] A. Burns and A. J. Wellings.

Usability of the Ada tasking model. ACM SIG-ADA Ada Letters, 10(4): 49–56, Spring 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Berry:1991:MC

[BW91] R. H. Berry and G. H. Wedberg. Metrics for competitiveness. In ACM [ACM91b], pages 119–123. ISBN 0-89791-393-0. LCCN ????

Burns:1992:SAR

[BW92] A. Burns and A. J. Wellings. In support of the Ada 9X real-time facilities. ACM SIGADA Ada Letters, 12 (1):53–64, January/February 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1993:MME

[BW93a] A. Burns and A. J. Wellings. Measuring, monitoring and enforcing CPU execution time usage. ACM SIG-ADA Ada Letters, 13(2):54– 64, March/April 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1993:SIW

[BW93b] Alan Burns and Andy Wellings. Summary of the 6th International Workshop on real-time Ada issues. ACM SIGADA Ada Letters, 13 (2):21–36, March/April 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1994:IAH

[BW94] A. Burns and A. J. Wellings.

Implementing analysable hard real-time sporadic tasks in Ada 9X. ACM SIGADA Ada Letters, 14(1):38–49, January/February 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1997:FID

[BW97a] A. Burns and A. J. Wellings.

Feature interactions with dynamic priorities. ACM SIG-ADA Ada Letters, 17(5):24– 26, September/October 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1997:RTM

[BW97b] A. Burns and A. J. Wellings.

Restricted tasking models. ACM SIGADA Ada Letters, 17(5):27–32, September/October 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:1999:HVC

[BW99] A. Burns and A. J. Wellings.

How to verify concurrent Ada programs: the application of model checking. ACM SIGADA Ada Letters, 19(2): 78–83, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2002:ADQ

[BW02] A. Burns and A. J. Wellings. Accessing delay queues. ACM SIGADA Ada Letters, 22(4): 72–76, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2003:TAB

[BW03] A. Burns and A. J. Wellings. Task attribute-based scheduling: extending Ada’s support for scheduling. ACM SIG-ADA Ada Letters, 23(4):36– 41, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2007:PET

[BW07a] A. Burns and A. J. Wellings. Programming execution-time servers in Ada 2005. ACM SIGADA Ada Letters, 27(2): 48–52, August 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2007:IEA

[BW07b] Alan Burns and Andy Wellings. Implementation experience with Ada 2005: Introduction. ACM SIG-ADA Ada Letters, 27(2): 59–60, August 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2010:LVL

[BW10a] A. Burns and A. J. Wellings.

Language vulnerabilities: let’s not forget concurrency. ACM SIGADA Ada Letters, 30(1):26–32, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2010:MSS

[BW10b] A. Burns and A. J. Wellings.

Multiprocessor systems session summary. ACM SIG-ADA Ada Letters, 30(1): 147–151, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2010:SEM

[BW10c] A. Burns and A. J. Wellings.

Supporting execution on multiprocessor platforms. ACM SIGADA Ada Letters, 30(1): 16–25, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2013:LPM

[BW13a] A. Burns and A. J. Wellings.

Locking policies for multiprocessor Ada. ACM SIG-ADA Ada Letters, 33(2):59– 65, August 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2013:SMP

[BW13b] A. Burns and A. J. Wellings.

Support for multiprocessor platforms. ACM SIG-ADA Ada Letters, 33(1): 9–14, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2013:SSLb

[BW13c] Alan Burns and Andy Wellings. Session summary: locking protocols. ACM SIG-ADA Ada Letters, 33(2):123– 125, August 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2015:TCR

[BW15] A. Burns and A. J. Wellings. Testing conformity to the real-time annex. ACM SIG-ADA Ada Letters, 35(1): 17–25, April 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2016:STC

[BW16a] A. Burns and A. J. Wellings. Synchronous task control and synchronous barriers. ACM SIGADA Ada Letters, 36(1): 35–38, June 2016. CODEN AALEE5. ISSN 0736-721X.

Burns:2016:DFP

[BW16b] Alan Burns and Andy Wellings. The Deadline Floor Protocol and Ada. ACM SIGADA Ada Letters, 36(1): 29–34, June 2016. CODEN AALEE5. ISSN 0736-721X.

Burns:2016:SSD

[BW16c] Alan Burns and Andy Wellings. Session summary: Deadline floor protocol. ACM SIGADA Ada Letters, 36(1): 91–93, June 2016. CODEN AALEE5. ISSN 0736-721X.

Burns:1990:ATC

[BWD90] A. Burns, A. J. Wellings, and G. L. Davies. Asynchronous transfer of control in Ada 9X. ACM SIGADA Ada Letters, 10(9):75–84, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2001:DVD

[BWK^{+}01] A. Burns, A. J. Wellings, A. M. Koelmans, M. Koutny, A. Romanovsky, and A. Yakovlev. On developing and verifying design abstractions for reliable concurrent programming in Ada. ACM SIG-ADA Ada Letters, 21(1): 48–55, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2013:TRP

[BWM13] A. Burns, A. J. Wellings, and A. H. Malik. TTF-Ravenscar: a profile to support reliable high-integrity multiprocessor Ada applications. ACM SIGADA Ada Letters, 33(1): 15–23, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Burns:2003:RSF

[BWV03] Alan Burns, Andy Wellings, and Tullio Vardanega. Report of session: flexible scheduling in Ada. ACM SIGADA Ada Letters, 23(4): 32–35, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Yue:1993:ASG

[bY93] Kwok bun Yue. An Ada solution to the general mutual exclusion problem. ACM SIG-ADA Ada Letters, 13(4):37– 43, July/August 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Yue:1994:SA

[bY94] Kwok bun Yue. Semaphores in Ada-94. ACM SIG-ADA Ada Letters, 14(5):71– 79, September/October 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Berry:1986:RUP

[BYY86] Daniel M. Berry, Nancy Yavne, and Moshe Yavne. On the requirements for and the use of a program design language: Parameterization, abstract data typing, strong typing. ACM SIG-ADA Ada Letters, 6(1):82– 89, January/February 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carlsson:1989:DAI

[CA89] Mats Carlsson and Lars Asplund. A data acquisition and information handling system in Ada for electron spectroscopy. ACM SIG-ADA Ada Letters, 9(5):89– 100, July/August 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Courtieu:2013:TFS

[CAC^{+}13] Pierre Courtieu, Maria Virginia Aponte, Tristan Crolard, Zhi Zhang, Fnu Robby, Jason Belt, John Hatcliff, Jerome Guitton, and Trevor Jennings. Towards the formalization of SPARK 2014 semantics with explicit runtime checks using Coq. ACM SIGADA Ada Letters, 33(3): 21–22, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Campbell:1992:CSL

[Cam92] John A. Campbell. Creating structure from linearity in non-Ada interfaces. ACM SIGADA Ada Letters, 12 (4):20–23, July/August 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carter:1988:MSDa

[Car88a] J. R. Carter. MMAIM: a software development method for Ada. I. Description. ACM SIGADA Ada Letters, 8(3): 107–114, May/June 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carter:1988:MSDb

[Car88b] J. R. Carter. MMAIM: a software development method for Ada, part II — Example. ACM SIGADA Ada Letters, 8(5):47–60, September/October 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carter:1989:MSD

[Car89a] J. Carter. MMAIM: a software development method for Ada. ACM SIGADA Ada Letters, 8(3):107–114, May/June 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carter:1989:VLS

[Car89b] Jeffrey R. Carter. Variablelength string input in Ada. ACM SIGADA Ada Letters, 9(4):103–104, May/June 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carter:1990:FRA

[Car90] Jeffrey R. Carter. The form of reusable Ada components for concurrent use. ACM SIG-ADA Ada Letters, 10(1):118– 121, January/February 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carter:1991:CRA

[Car91] Jeffrey R. Carter. Concurrent reusable abstract data types. ACM SIGADA Ada Letters, 11(1):96–101, January/February 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carter:1992:ARC

[Car92] Jeffrey R. Carter. Ada 9X reusable components. ACM SIGADA Ada Letters, 12 (2):91–96, March/April 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carter:1994:ADG

[Car94] Jeffrey R. Carter. Ada’s design goals and object-oriented programming. ACM SIG-ADA Ada Letters, 14(6): 57–61, November/December 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carter:1996:BAP

[Car96] Jeffrey R. Carter. Breaking the Ada Privacy Act. ACM SIGADA Ada Letters, 16(3): 52–55, May/June 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carter:1997:OVR

[Car97] Jeffrey R. Carter. OOP vs. readability. ACM SIG-ADA Ada Letters, 17(2):63– 66, March/April 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carlisle:1998:GF

[Car98] Martin C. Carlisle. Graphics for free. ACM SIG-ADA Ada Letters, 18(5):47– 50, September/October 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carlisle:1999:TII

[Car99a] Martin C. Carlisle. A truly implementation independent GUI development tool. ACM SIGADA Ada Letters, 19(3): 47–52, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carpenter:1999:VRS

[Car99b] Paul B. Carpenter. Verification of requirements for safety-critical software. ACM SIGADA Ada Letters, 19(3): 23–29, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carlisle:2000:AOO

[Car00] Martin C. Carlisle. An automatic object-oriented parser generator for Ada. ACM SIGADA Ada Letters, 20 (2):57–63, June 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.acm.org/sigada/ada_
letters/june2000/adagoop.
pdf.

Carlisle:2001:KAC

[Car01] Martin Carlisle. Keynote address: confessions of an academic Ada zealot. ACM SIG-ADA Ada Letters, 21(4):71– 72, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carlisle:2002:EP

[Car02]  Martin Carlisle.  Edito-

rial policy. ACM SIG-ADA Ada Letters, 22(1): 3–10, March 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carter:2004:PRC

[Car04] Jeffrey R. Carter. The PragmAda reusable components. ACM SIGADA Ada Letters, 24(3):44–46, September 2004. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carlisle:2006:AOP

[Car06a] Martin C. Carlisle. Automatic OO parser generation using visitors for Ada 2005. ACM SIGADA Ada Letters, 26(3):3–8, December 2006.

CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carlisle:2006:HAI

[Car06b] Martin C. Carlisle. How Ada 2005 impacts CS1/2. ACM SIGADA Ada Letters, 26(1): 18–24, April 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carlisle:2011:WCB

[Car11] Martin C. Carlisle. Why I came back to Ada. ACM SIGADA Ada Letters, 31(3): 37–38, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carrez:2017:INS

[Car17] Stéphane Carrez. IP network stack in Ada 2012 and the Ravenscar profile. ACM SIG-ADA Ada Letters, 37(2):51– 58, December 2017. CODEN AALEE5. ISSN 0736-721X.

Carrez:2022:IBM

[Car22a] Stéphane Carrez. Implementing a build manager in Ada. ACM SIGADA Ada Letters, 42(1):68–75, June 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3577949.3577960.

Carter:2022:OAG

[Car22b] Jeffrey R. Carter. Overview of Ada GUI. ACM SIG-ADA Ada Letters, 42(1): 61–64, June 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3577949.3577958.

Casinghino:2020:LPH

[Cas20] Chris Casinghino. A language for programmable hardware security. ACM SIGADA Ada Letters, 39(1):71, January 2020. ISSN 0736-721X.

URL https://dl.acm.org/
doi/abs/10.1145/3379106.
3379115.

Cheng:1988:TCD

[CAU88] J. Cheng, K. Araki, and K. Ushijima. Tasking communication deadlocks in concurrent Ada programs. ACM SIGADA Ada Letters, 8(5):61–70, September/October 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carlisle:2007:TNN

[CB07] Martin C. Carlisle and Leemon C. Baird III. Timing neural networks in C and Ada. ACM SIG-ADA Ada Letters, 27(3):71– 74, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Colket:1997:AAT

[CBB^{+}97] Currie Colket, Gary Barnes, Steve Blake, Dan Cooper, Jesper Jørgensen, Clyde Roby, Dan Rittersdorf, Sergey Ryben, Alfred Strohmeier, and Bill Thomas. Architecture of ASIS: a tool to support code analysis of complex systems. ACM SIG-ADA Ada Letters, 17(1):35– 40, January/February 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chapman:1994:SWC

[CBW94] Roderick Chapman, Alan Burns, and Andy Wellings. Static worst-case timing analysis of Ada. ACM SIG-ADA Ada Letters, 14(5):88– 91, September/October 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chard:2021:EAP

[CBW^{+}21] Kyle Chard, Yadu Babuji, Anna Woodard, Ben Clifford, Zhuozhao Li, Mihael Hategan, Ian Foster, Mike Wilde, and Daniel S. Katz. Extended abstract: Productive parallel programming with Parsl. ACM SIGADA Ada Letters, 40(2):73–75, April 2021. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3463478.3463486.

Carlisle:1998:AFI

[CC98] Martin C. Carlisle and A. T. Chamillard. AdaGIDE: a friendly introductory programming environment for a freshman computer science course. ACM SIG-ADA Ada Letters, 18(2): 42–52, March 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carletto:2018:REF

[CC18] Paolo Carletto and Tullio Carletto. Ravenscar-EDF: Further results from improved comparative benchmarking. ACM SIGADA Ada Letters, 38(1):40, June 2018. CODEN AALEE5. ISSN 0736-721X.

Castagna:2021:AAN

[CCC21] Christian Castagna, Daniela Cancila, and Antonio Cammi. Adoption of ACPS in nuclear reactor analysis. ACM SIGADA Ada Letters, 41 (1):69–73, June 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3570315.3570320.

Costa:2023:CAD

[CCC23] Gonçalo Costa, José Cecílio, and António Casimiro. Cooperative autonomous driving in simulation. ACM SIGADA Ada Letters, 43 (1):78–82, June 2023. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3631483.3631496.

Comar:1997:TGJ

[CDG97] C. Comar, G. Dismukes, and F. Gasperoni. Targeting GNAT to the Java Virtual Machine. In ACM [ACM97], pages 149–164. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

CrespiReghizzi:1987:DAP

[CDM87] S. Crespi Reghizzi, A. Di Maio, and F. Maderna. Distributable Ada programs. ACM SIGADA Ada Letters, 7(6):67–69, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chaki:2016:CBV

[CdN16] Sagar Chaki and Dionisio de Niz. Contract-based verification of timing enforcers: [extended abstract]. ACM SIGADA Ada Letters, 36(2): 27–30, December 2016. CO-DEN AALEE5. ISSN 0736-721X.

Creuse:2023:ATV

[CEG23] L. Creuse, M. Eyraud, and V. Garèse. Automatic test value generation for Ada. ACM SIGADA Ada Letters, 43(1):100–105, June 2023. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3631483.3631500.

Celier:1997:MUD

[Cel97] V. Celier. Managing usage of dynamic structures with Ada controlled objects. In ACM [ACM97], pages 165– 172. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Charles:1982:LGA

[CF82] Philippe Charles and Gerald Fisher. A LALR(1) grammar for ’82 Ada. ACM SIG-ADA Ada Letters, 2(2):34– 45, September/October 1982. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carter:2013:SSA

[CFH^{+}13] Kyle Carter, Adam Foltzer, Joe Hendrix, Brian Huffman, and Aaron Tomb. SAW: the software analysis workbench. ACM SIGADA Ada Letters, 33(3):15–18, December 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chase:1982:CFA

[CG82] Anna I. Chase and Mark S.

Gerhardt. The case for full Ada as a design language. ACM SIGADA Ada Letters, 2(3):51–59, November/December 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cook:1987:NAA

[CG87a] David A. Cook and Dean W. Gonzalez. Notes on Ada abstraction. ACM SIG-ADA Ada Letters, 7(5):93– 95, September/October 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cook:1987:NDA

[CG87b] David A. Cook and Dean W. Gonzalez. Notes on data abstraction. ACM SIG-ADA Ada Letters, 7(5):93– 95, September/October 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Collard:1988:KBS

[CG88] Philippe Collard and Andre Goforth. Knowledge based systems and Ada: An overview of the issues. ACM SIGADA Ada Letters, 8(6):72–81, November/December 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chamillard:1997:TAI

[CH97] A. T. Chamillard and W. C. Hobart. Transitioning to Ada in an introductory course for non-majors. In ACM [ACM97], pages 37–40. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Chapman:2004:ESS

[CH04] Roderick Chapman and Adrian Hilton. Enforcing security and safety models with an information flow analysis tool. ACM SIG-ADA Ada Letters, 24(4):39– 46, December 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carlisle:2006:IAV

[CH06] Martin C. Carlisle and J. A.

Hamilton, Jr. Integrating Ada 2005 into visual studio 2005. ACM SIG-ADA Ada Letters, 26(3):15– 20, December 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chambers:1982:EAL

[Cha82] John M. Chambers. Extending Ada legally via preprocessors. ACM SIG-ADA Ada Letters, 1(4):55– 58, May/June 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chapman:2000:IES

[Cha00] Roderick Chapman. Industrial experience with SPARK. ACM SIGADA Ada Letters, 20(4):64–68, December 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

                                        /www.acm.org/sigada/ada_
                                        letters/dec2000/chapman-

paper.pdf. Special Issue: Presentations from SIGAda 2000.

Chapman:2007:CCP

[Cha07a] Rod Chapman. Correctness by construction: putting engineering (back) into software. ACM SIGADA Ada Letters, 27(3):100, December 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chapman:2007:MSC

[Cha07b] Rod Chapman. MF1: security by construction. ACM SIGADA Ada Letters, 27(3): 5–6, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Charlet:2009:GGA

[Cha09] Arnaud Charlet. Gem #59: generating Ada bindings for C headers. ACM SIG-ADA Ada Letters, 29(2):56– 60, August 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chapman:2011:GSS

[Cha11] Rod Chapman. Gem #80: speedy shift and rotate in SPARK. ACM SIG-ADA Ada Letters, 31(2): 30–32, August 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chaki:2013:BMC

[Cha13]  Sagar Chaki.  Bounded

model checking of highintegrity software. ACM SIG-ADA Ada Letters, 33(3):9– 10, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cheng:1990:CTD

[Che90]  Jingde Cheng.  A clas-

sification of tasking deadlocks. ACM SIGADA Ada Letters, 10(5):110–127, May/June 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cheng:1991:STD

[Che91a] Jingde Cheng. A survey of tasking deadlock detection methods. ACM SIG-ADA Ada Letters, 11(1):82– 91, January/February 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cherry:1991:SRM

[Che91b] George W. Cherry. Stimulusresponse machines: An Adabased graphic notation for specifying, designing, and implementing reactive or interactive systems. ACM SIG-ADA Ada Letters, 11(5):30– 46, July/August 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cheng:1992:TDN

[Che92] Jingde Cheng. The Task Dependence Net in Ada software development. ACM SIG-ADA Ada Letters, 12(4):24– 35, July/August 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cheng:1997:TDN

[Che97] J. Cheng. Task dependence nets for concurrent systems with Ada 95 and its applications. In ACM [ACM97], pages 67–78. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Chelini:2009:WTD

[Che09]  James Chelini.  Working

towards DO-178C/ED-12C, DO-248C/ED-94C, and DO-278A/ED109A. ACM SIG-ADA Ada Letters, 29(3):103– 104, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Creuse:2019:SEI

[CHGH19] Léo Creuse, Joffrey Huguet, Christophe Garion, and Jérôme Hugues. SPARK by example: an introduction to formal verification through the Standard C++ library. ACM SIGADA Ada Letters, 38(2):89–96, December 2019. ISSN 0736-721X.

URL https://dl.acm.org/
doi/abs/10.1145/3375408.
3375415.

Chelini:1990:EEDa

[CHHB90a] James V. Chelini, Donna D.

Hughes, Leonard J. Hoffman, and Denise M. Brunelle. An example of event-driven asynchronous scheduling with Ada. ACM SIGADA Ada Letters, 10(6):84–96, July/ August 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chelini:1990:EEDb

[CHHB90b] James V. Chelini, Donna D.

Hughes, Leonard J. Hoffman, and Denise M. Brunelle. An example of event-driven asynchronous scheduling with Ada. ACM SIGADA Ada Letters, 10(8):130–144, November/December 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chong:2019:PLS

[Cho19] Stephen Chong. Programming languages for security. ACM SIGADA Ada Letters, 38(2):69–88, December 2019. ISSN 0736-721X.

URL https://dl.acm.org/
doi/abs/10.1145/3375408.
3375414.

Christensen:1987:AFR

[Chr87a] Elisabeth Broe Christensen. Ada features and real-time embedded applications. ACM SIGADA Ada Letters, 7(6): 116–118, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Christiansen:1987:AFR

[Chr87b] Elisabeth Broe Christiansen. Ada features and real-time embedded applications. ACM SIGADA Ada Letters, 7(6): 116–118, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cross:1990:DC

[CKF90] Joe Cross, Mike Kamrad, and Sylvester Fernandez. Distributed communications. ACM SIGADA Ada Letters, 10(9):85–93, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clark:1987:DCO

[Cla87a] Robert G. Clark. Designing concurrent objects. ACM SIGADA Ada Letters, 7(6): 107–109, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clarson:1987:AIH

[Cla87b] Donald R. Clarson. Ada information hiding — additional notes. ACM SIGADA Ada Letters, 7(1):89–93, January/February 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clarson:1987:PAD

[Cla87c] Donald R. Clarson. Proposal for adding discriminants for Ada task types. ACM SIG-ADA Ada Letters, 7(5):96– 99, September/October 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clarke:1997:OCO

[Cla97] David Clarke. The OMG, CORBA, Orbix and Ada. ACM SIGADA Ada Letters, 17(3):97–108, May/June 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clemmensen:1982:FMD

[Cle82] G. B. Clemmensen. A formal model of distributed Ada tasking. In ACM [ACM82], pages 224–237. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Clemmensen:1986:RRD

[Cle86] Geert. B. Clemmensen. Retargeting and rehosting the DDC Ada compiler system: a case study — the Honeywell DPS 6. ACM SIGADA Ada Letters, 6(1):22–28, January/February 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chamillard:1998:UAN

[CLY98] A. T. Chamillard, Ronald J. Lisowski, and Richard R. Young. Using Ada in non-CS majors. ACM SIGADA Ada Letters, 18(6):61–67, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clapp:1989:AH

[CM89] Russell M. Clapp and Trevor Mudge. Ada on a hypercube. ACM SIGADA Ada Letters, 9(2):118–128, March/April 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clapp:1990:ISI

[CM90a] Russell M. Clapp and Trevor Mudge. Introduction to the special issue on Ada performance issues. ACM SIG-ADA Ada Letters, 10(3):10– 13, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clapp:1990:O

[CM90b] Russell M. Clapp and Trevor Mudge. Optimization. ACM SIGADA Ada Letters, 10 (3):59, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clapp:1990:PDI

[CM90c] Russell M. Clapp and Trevor Mudge. Parallel and distributed issues. ACM SIG-ADA Ada Letters, 10(3):33– 37, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clapp:1990:RDI

[CM90d] Russell M. Clapp and Trevor Mudge. A rationale for the design and implementation of Ada benchmark programs. ACM SIGADA Ada Letters, 10(3):8–13, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clapp:1990:SP

[CM90e] Russell M. Clapp and Trevor Mudge. The space problem. ACM SIGADA Ada Letters, 10(3):29–32, Winter 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clapp:1990:TB

[CM90f] Russell M. Clapp and Trevor Mudge. Taxonomy of benchmarks. ACM SIG-ADA Ada Letters, 10(3): 14–19, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clapp:1990:TP

[CM90g] Russell M. Clapp and Trevor Mudge. The time problem.

ACM SIGADA Ada Letters, 10(3):20–28, Winter 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.eecs.umich.edu/~tnm/
papers/adaTime.pdf.

Choi:1994:UIS

[CM94] Sungwoon Choi and Toshimi Minoura. User interface system based on active objects. ACM SIGADA Ada Letters, 14(Special Issue): 16–25, Fall 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carlisle:1998:RFP

[CM98] Martin C. Carlisle and Patrick Maes. RAPID: a free, portable GUI design tool. ACM SIGADA Ada Letters, 18(6):158–164, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Describes an interface between Ada and Tcl/ Tk.

Clapp:1990:RFT

[CMR90] Russell M. Clapp, Trevor Mudge, and Daniel Roy. Recommendations and future trends. ACM SIG-ADA Ada Letters, 10(3):98– 100, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chard:2021:LSP

[CMWT21] Kyle Chard, James Muns, Richard Wai, and S. Tucker Taft. Language support for parallel and distributed computing. ACM SIG-ADA Ada Letters, 40(2): 51–54, April 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463478.3463481.

Castellano:1996:SOM

[CN96] G. Vincent Castellano and Steven W. North. System Object Model (SOM) and Ada: an example of CORBA at work. ACM SIG-ADA Ada Letters, 16(3):39– 51, May/June 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cohen:1981:HAA

[Coh81]  Paul M. Cohen.  From

HOLWG to AJPO — Ada in transition. ACM SIG-ADA Ada Letters, 1(1):22– 25, July/August 1981. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cohen:1982:PQE

[Coh82] Norman H. Cohen. Parallel quicksort: An exploitation of concurrent programming in Ada. ACM SIG-ADA Ada Letters, 2(2):61– 68, September/October 1982.

CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cohen:1985:TAM

[Coh85] Norman H. Cohen. Tasks as abstraction mechanisms. ACM SIGADA Ada Letters, 5(3–6):30–44, November/December 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cohen:1986:UEC

[Coh86]  Ellis S. Cohen.  Updat-

ing elements of a collection in place. ACM SIGADA Ada Letters, 6(1):55–62, January/February 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cohen:1988:DAT

[Coh88] Norman H. Cohen. Dependence on Ada task scheduling is not “erroneous”. ACM SIGADA Ada Letters, 8(2): 77–83, March/April 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cohen:1994:EIR

[Coh94] Norman H. Cohen. Endianindependent record representation clauses. ACM SIG-ADA Ada Letters, 14(1):27– 29, January/February 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Collingbourne:1987:PAD

[Col87] L. R. Collingbourne. A practical approach to developing Real-Time Ada programs for embedded systems. ACM SIGADA Ada Letters, 7(6):15–17, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Collard:1989:OOP

[Col89] P. Collard. Object-oriented programming techniques with Ada — an example. ACM SIGADA Ada Letters, 9(6): 119–126, September/October 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Colbert:1990:S

[Col90] Edward Colbert. SigAda.

ACM SIGADA Ada Letters, 10(6):5, July/August 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Colket:1995:ASI

[Col95a]  Currie Colket.  Ada Se-

mantic Interface Specification (ASIS): frequently asked questions. ACM SIGADA Ada Letters, 15(4):50–63, July/August 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Colket:1995:HJA

[Col95b] Currie Colket. Highlights of the June 1995 ASISWG/ ASISRG meeting. ACM SIG-ADA Ada Letters, 15(5):32– 33, September/October 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Collins:1998:TSS

[Col98] W. Robert Collins. Tasking solutions to the Sieve of Eratosthenes. ACM SIG-ADA Ada Letters, 18(4): 107–110, July 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cole:1999:CAA

[Col99a] Oliver Cole. Converting an Ada 83 application to Ada 95. ACM SIGADA Ada Letters, 19(4):19–21, December 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Colket:1999:CAS

[Col99b] William Currie Colket. Code analysis of safety-critical and real-time software using ASIS. ACM SIG-ADA Ada Letters, 19(3):67– 76, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Colket:2001:MC

[Col01]  Currie Colket.  Message

from the Chair. ACM SIG-ADA Ada Letters, 21(3):1– 2, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Colket:2002:MC

[Col02]  Currie Colket.  Message

from the Chair. ACM SIG-ADA Ada Letters, 22(1): 1–2, March 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Command:1990:ACE

[Com90] Air Force Systems Command. Ada Compiler Evaluation Capability (ACEC) data analysis: an overview. ACM SIGADA Ada Letters, 10(3): 111–125, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Condic:1990:JFS

[Con90] Marin David Condic. Junk facts and the SlowSort. ACM SIGADA Ada Letters, 10(1):104–110, January/February 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Conn:1997:SCA

[Con97a] R. Conn. The Source Code Analysis Tool Construction Project. In ACM [ACM97], pages 141–148. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Conn:1997:DEE

[Con97b]  Richard Conn.  Defining

and exploring an efficient distributed process for the reuse of Ada software components and tools in a global theater — the Public Ada Library. ACM SIGADA Ada Letters, 17(4):59–65, July/August 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Conn:1997:TWC

[Con97c] Richard Conn. Tour of Walnut Creek Ada CDROM. ACM SIGADA Ada Letters, 17(4):31–58, July/August 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Conn:1997:WUS

[Con97d] Richard Conn. What users should know about the Public Ada Library (PAL). ACM SIGADA Ada Letters, 17 (4):17–30, July/August 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Conn:1998:RTP

[Con98] Richard Conn. The Reuse Tapestry Project. ACM SIG-ADA Ada Letters, 18(1):65– 69, January/February 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Condic:2003:PPC

[Con03a] Marin D. Condic. A plan for producing a conventional Ada library. ACM SIG-ADA Ada Letters, 23(3):16– 31, September 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Conn:2003:ACL

[Con03b] Richard Conn. Ada, CMM level 4, and the C-130J aircraft. ACM SIGADA Ada Letters, 23(1):10, March 2003. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cooper:1997:ABC

[Coo97] C. Daniel Cooper. ASISbased code analysis automation. ACM SIGADA Ada Letters, 17(6):65–69, November/December 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cornhill:1983:SDC

[Cor83] Dennis Cornhill. A survivable distributed computing system for embedded applications programs writtein in Ada. ACM SIGADA Ada Letters, 3(3):79–87, November/December 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cousins:2021:OA

[Cou21] Jeff Cousins. An overview of Ada 202x. ACM SIG-ADA Ada Letters, 41(1): 44–63, June 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3570315.3570317.

Colket:1997:ASI

[CR97] Currie Colket and Clyde Roby. Ada semantic interface specification (ASIS) frequently asked questions. ACM SIGADA Ada Letters, 17(2):26–28, March/April 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Comar:2005:DPL

[CR05] Cyrille Comar and Pat Rogers. On dynamic plugin loading with Ada 95 and Ada 2005. ACM SIG-ADA Ada Letters, 25(2): 31–41, June 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cheng:2007:IPC

[CR07] Albert M. K. Cheng and James Ras. The implementation of the Priority Ceiling Protocol in Ada-2005. ACM SIGADA Ada Letters, 27(1): 24–39, April 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carletto:2018:SSD

[CR18] Andy Carletto and Jorge Real. Session summary: Deadline floor protocol. ACM SIGADA Ada Letters, 38(1): 70–73, June 2018. CODEN AALEE5. ISSN 0736-721X.

Crafts:1982:CAS

[Cra82a] Ralph E. Crafts. Commercial applications software in Ada: a reality. ACM SIG-ADA Ada Letters, 1(4):46– 54, May/June 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cranc:1982:CLA

[Cra82b]  M. E. Cranc.  A com-

mand language for the Ada environment. In ACM [ACM82], pages 181–186. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Crawford:1995:PIA

[Cra95] Bard S. Crawford. Proposed icons for Ada 95. ACM SIG-ADA Ada Letters, 15(4):36– 45, July/August 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Crafts:1997:RNR

[Cra97]  Ralph Crafts.  Reaction

to NRC recommendations. ACM SIGADA Ada Letters, 17(1):18–20, January/ February 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Crawford:1998:AAS

[Cra98] Bard S. Crawford. Algorithm animation with symbol processing robots. ACM SIG-ADA Ada Letters, 18(6):217– 218, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Craeynest:2022:ADR

[Cra22] Dirk Craeynest. 11th Ada Developer Room at FOS-DEM 2022. ACM SIG-ADA Ada Letters, 42(1): 39–40, June 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3577949.3577951.

Criley:2001:SBM

[Cri01] Marc A. Criley. A socketbased manifestation of streams. ACM SIGADA Ada Letters, 21(2):53–64, June 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cross:1990:OCS

[Cro90] Joseph K. Cross. Other compiler support working group. ACM SIGADA Ada Letters, 10(4):144–158, Spring 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cronin:1995:IRM

[Cro95] Kevin J. Cronin. Integrating rate monotonic analysis into the preliminary Ada design process. ACM SIG-ADA Ada Letters, 15(2):40– 45, March/April 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Crocker:2014:CCM

[Cro14] David Crocker. Can C++ be made as safe as SPARK? ACM SIGADA Ada Letters, 34(3):5–12, December 2014. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cornhill:1987:PIA

[CS87] Dennis Cornhill and Lui Sha.

Priority inversion in Ada — or — what should be the priority of an Ada server task? ACM SIGADA Ada Letters, 7(7):30–32, November/December 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Celarier:1991:AML

[CS91] Donald A. Celarier and Donald W. Sando. An Ada math library for real-time avionics. ACM SIGADA Ada Letters, 11(7):274–284, Fall 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carter:1994:ADN

[CS94] Jeffrey R. Carter and Bo I. Sanden. Ada design of a neural network. ACM SIG-ADA Ada Letters, 14(3):61– 73, May/June 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carlisle:2002:AVG

[CS02] Martin C. Carlisle and Ricky E. Sward. An automatic “visitor” generator for Ada. ACM SIG-ADA Ada Letters, 22(3):42– 47, September 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carr:1987:IPC

[CSA^{+}87] P. Carr, R. Stevenson, J. Alea, J. Berthold, G. Groucher, M. Davis, G. Dobbins, D. Law, V. Szarek, and W. Webster. Implementation of a prototype CAIS environment. ACM SIG-ADA Ada Letters, 7(2):58– 72, March/April 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carlisle:2003:WAN

[CSH03] Martin C. Carlisle, Ricky E. Sward, and Jeffrey W. Humphries. Weaving Ada 95 into the .NET environment. ACM SIGADA Ada Letters, 23(1):22–26, March 2003.

CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cornhill:1987:LAR

[CSL^{+}87] Dennis Cornhill, Lui Sha, John P. Lehoczky, Ragunathan Rajkumar, and Hide Tokuda. Limitations of Ada for real-time scheduling. ACM SIGADA Ada Letters, 7(6):33–39, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cicalese:2009:USA

[CSSW09] Cynthia Cicalese, Joel Sherrill, Ricky E. Sward, and Richard Weatherly. Unmanned systems and Ada. ACM SIGADA Ada Letters, 29(3):11–12, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cicalese:2010:USA

[CSSW10] Cindy Cicalese, Joel Sherill, Ricky Sward, and Richard Weatherly. Unmanned systems with Ada and RTEMS. ACM SIGADA Ada Letters, 30(3):9–10, December 2010. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cheng:1989:NAT

[CU89] Jingde Cheng and Kazuo Ushijima. Naming Ada tasks at run-time. ACM SIG-ADA Ada Letters, 9(2):52– 61, March/April 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Carey:2003:NIF

[CVW03] Robert W. Carey, Paul J. Van Arsdall, and John P. Woodruff. The National Ignition Facility: early operational experience with a large Ada control system. ACM SIGADA Ada Letters, 23(1):11, March 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Clarke:1980:NAB

[CWW80] Lori A. Clarke, Jack C. Wileden, and Alexander L. Wolf. Nesting in Ada is for the birds. In ACM [ACM80], pages 139–145. CODEN SIN-ODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Chen:2001:DCE

[CXY01] Zhenqiang Chen, Baowen Xu, and Huiming Yu. Detecting concurrently executed pairs of statements using an adapted MHP algorithm. ACM SIGADA Ada Letters, 21(4):107–114, December 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Doran:2013:RMD

[DA13] Steven Doran and Stephanie August. Reddo: a model driven engineering toolset for embedded software development. ACM SIG-ADA Ada Letters, 33(3):47– 48, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dausmann:1987:LSR

[Dau87] Manfred Dausmann. Library structures for reusable components. In ACM [ACM87a], pages 226–336. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Davis:1982:COA

[Dav82] Mark Davis. Classical optimizations in Ada. ACM SIGADA Ada Letters, 1(2): 11–14, September 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Davis:2004:ISS

[Dav04] James F. Davis. Information systems security engineering: a critical component of the systems engineering lifecycle. ACM SIGADA Ada Letters, 24(4):13–18, December 2004. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Davis:2005:AAF

[Dav05] James F. Davis. The affordable application of formal methods to software engineering. ACM SIG-ADA Ada Letters, 25(4):57– 62, December 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

deBondeli:1997:AFR

[dB97a] Patrick de Bondeli. Annex: a fully reusable class of objects for synchronization and communication in Ada 95. ACM SIGADA Ada Letters, 17(5):35–39, September/October 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

deBondeli:1997:DRM

[dB97b] Patrick de Bondeli. Developing reusable multi-tasking components using objectoriented techniques in Ada 95. ACM SIGADA Ada Letters, 17(5):33–34, September/October 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dobbing:1998:RTP

[DB98] Brian Dobbing and Alan Burns. The Ravenscar tasking profile for high-integrity real-time programs. ACM SIGADA Ada Letters, 18(6): 1–6, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

deBondeli:1999:FRC

[dB99] Patrick de Bondeli. A fully reusable class of objects for synchronization and communication in Ada 95. ACM SIGADA Ada Letters, 19(1): 66–96, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dinh:2009:DCD

[DB09] Tong Dinh and Shan Barkataki.

Distributed container: a design pattern for fault tolerance and high speed data exchange. ACM SIGADA Ada Letters, 29(3):115–118, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

DeanHendrix:1997:VCS

[DCBM97] T. Dean Hendrix, J. H. Cross, L. A. Barowski, and K. S. Mathias. Visualization of control structure and complexity in Ada 95. In ACM [ACM97], pages 135– 140. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

DiMaio:1985:EMD

[DCC85] A. Di Maio, S. Ceri, and S. Crespi Reghizzi. Execution monitoring and debugging tool for Ada using relational algebra. ACM SIGADA Ada Letters, 5(2):109–123, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Donaho:1987:AES

[DD87] Jane E. D. Donaho and Genell K. Davis. Adaembedded SQL: the options. ACM SIGADA Ada Letters, 7(3):60–72, May/June 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dousette:1998:CCU

[DDJ98] Patricia J. Dousette, Ari Danesh, and Matthew Jones. Command and control using World Wide Web technology. ACM SIGADA Ada Letters, 18(6):212–214, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dobbing:2003:SSF

[DdlP03] Brian Dobbing and Juan Antonio de la Puente. Session: status and future of the Ravenscar profile. ACM SIGADA Ada Letters, 23(4): 55–57, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

deBondeli:1987:RTA

[de 87] Patrick de Bondeli. Real-Time Ada systems: Development methodology and realtime performance. ACM SIGADA Ada Letters, 7(6): 119–120, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

deBondeli:1988:ATC

[de 88] Patrick de Bondeli. Asynchronous transfer of control and scheduling problems. ACM SIGADA Ada Letters, 8(7):57–60, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Debest:1983:UFS

[Deb83] X. Debest. A user-friendly I/O system for Ada. ACM SIGADA Ada Letters, 2(4): 101–112, January/February 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

DeLoach:1988:IAP

[DeL88a]  Scott A. DeLoach.  An

interface-based Ada programming support environment. ACM SIGADA Ada Letters, 8(4):70–82, July/ August 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

DeLoach:1988:IBA

[DeL88b]  Scott A. DeLoach.  An

interface-based Ada programming support environment. ACM SIGADA Ada Letters, 8(4):70–82, July/ August 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

developer:2017:GMCa

[dev17a] An Unknown AdaCore developer. Gem #138: Master the command line — Part 1. ACM SIGADA Ada Letters, 37(1):39–42, June 2017. CO-DEN AALEE5. ISSN 0736-721X.

developer:2017:GMCb

[dev17b] An Unknown AdaCore developer. Gem #139: Master the command line — Part 2. ACM SIGADA Ada Letters, 37(1):43–45, June 2017. CO-DEN AALEE5. ISSN 0736-721X.

Developer:2017:GCF

[Dev17c] An Unknown AdaCore Developer. Gem #141: Configure it out. ACM SIG-ADA Ada Letters, 37(1):50– 53, June 2017. CODEN AALEE5. ISSN 0736-721X.

Dewar:1984:ALM

[Dew84] Robert B. K. Dewar. Ada language maintenance, a look at what is going on. ACM SIGADA Ada Letters, 4(2):65–76, September/October 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

DeWeese:1986:ALL

[DeW86] Keith Patrick DeWeese. Ada:

a life and legacy: Dorothy Stein book review. ACM SIGADA Ada Letters, 6(2): 13–14, March/April 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dewar:2001:KAF

[Dew01] Robert Dewar. Keynote address: future development of the Ada language. ACM SIGADA Ada Letters, 21(4): 1–2, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dewar:2006:AHI

[Dew06] Robert Dewar. Ada 2005 & high integrity systems. ACM SIGADA Ada Letters, 26(3): 43, December 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dewar:2007:BFW

[Dew07a] Robert Dewar. Birds-of-a-feather: where would you like to see GNAT go? ACM SIGADA Ada Letters, 27(3): 97–98, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dewar:2007:CSA

[Dew07b] Robert B. K. Dewar. The compiler as a static analysis tool. ACM SIG-ADA Ada Letters, 27(3):83– 88, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dewar:2009:GCDa

[Dew09a] Robert Dewar. Gem #27: changing data representation (part 1). ACM SIG-ADA Ada Letters, 29(1): 35–37, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dewar:2009:GCDb

[Dew09b] Robert Dewar. Gem #28: changing data representation (part 2). ACM SIG-ADA Ada Letters, 29(1): 38–40, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dewar:2009:GPP

[Dew09c] Robert Dewar. Gem #31: preconditions/postconditions. ACM SIGADA Ada Letters, 29(1):48–49, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dewar:2009:GIB

[Dew09d] Robert Dewar. Gem #46: incompatibilities between Ada 83 and Ada 95. ACM SIGADA Ada Letters, 29(1): 75–76, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dewar:2017:GU

[Dew17] Robert Dewar. Gem # #150:

Out and uninitialized. ACM SIGADA Ada Letters, 37(2): 37–39, December 2017. CO-DEN AALEE5. ISSN 0736-721X.

Daily:1984:APS

[DF84] Paulan D. Daily and John T.

Foreman. Ada programming standards and guidelines. ACM SIGADA Ada Letters, 3(6):79–94, May/June 1984. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dewar:2009:GAS

[DFGZ09] Robert Dewar, Vasiliy Fofanov, Franco Gasperoni, and Yang Zhang. Gem #22: Ada speaks many languages. ACM SIGADA Ada Letters, 29(1):23–24, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dewar:1980:NAT

[DFS^{+}80] Robert B. K. Dewar, Gerald A. Fisher, Jr., Edmond Schonberg, Robert Froelich, Stephen Bryant, Clinton F. Goss, and Michael Burke. The NYU Ada translator and interpreter. In ACM [ACM80], pages 194– 201. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Dorchak:1997:PIS

[DG97] Susan Fife Dorchak and S. Rollins Guild. Protecting internal state variables from subclasses. ACM SIGADA Ada Letters, 17(6):70–77, November/December 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

DelasHeras-Quiros:1997:PDF

[DGBMCG97] P. De las Heras-Quiros, J. Gonzalez-Barahona, M., and J. Centeno-Gonzalez. Programming distributed fault tolerant systems: The replicAda approach. In ACM [ACM97], pages 21–30. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Dapra:1984:UAA

[DGCR^{+}84] A. Dapra, S. Gatti, S. Crespi-Reghizzi, F. Maderna, D. Belcredi, A. Natali, R. A. Stammers, and M. D. Tedd. Using Ada and APSE to support distributed multimicroprocessor targets. ACM SIG-ADA Ada Letters, 3(6):57– 65, May/June 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Donzeau-Gouge:1985:TAP

[DGLM85] V. Donzeau-Gouge, B. Lang, and B. Me’le’se. A tool for Ada program manipulations: Mentor-Ada. ACM SIGADA Ada Letters, 5(2):297–308, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Duncan:1980:UAI

[DH80] A. G. Duncan and J. S.

Hutchison. Using Ada for industrial embedded microprocessor applications. In ACM [ACM80], pages 26– 35. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Duncan:1982:UAI

[DH82] A. G. Duncan and J. S.

Hutchison. Using Ada for industrial embedded microprocessor applications, II. In ACM [ACM82], pages 152– 161. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Dismukes:2009:GEP

[Dis09] Gary Dismukes. Gem #63: the effect of pragma suppress. ACM SIGADA Ada Letters, 29(2):65–67, August 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

delaPuente:2013:SSC

[dlPM13] Juan Antonio de la Puente and Stephen Michell. Session summary: concurrency issues. ACM SIG-ADA Ada Letters, 33(1): 150–156, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

delaPuente:2002:SSS

[dlPP02] Juan Antonio de la Puente and Luís Miguel Pinho. Session summary: safety improvements for consideration. ACM SIGADA Ada Letters, 22(4):120–122, December 2002. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

delaPuente:1999:RTP

[dlPRGB99] Juan A. de la Puente, José F. Ruiz, and Jesús M. González-Barahona. Real-time programming with GNAT: specialised kernels versus POSIX threads. ACM SIGADA Ada Letters, 19(2):73–77, June 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

delaPuente:2007:CPN

[dlPU07] Juan A. de la Puente and Santiago Urueña. Conclusions and plans for next IRTAW: Summary. ACM SIGADA Ada Letters, 27(2): 96–97, August 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

delaPuente:2003:ETC

[dlPZ03] Juan Antonio de la Puente and Juan Zamorano. Executiontime clocks and Ravenscar kernels. ACM SIG-ADA Ada Letters, 23(4):82– 86, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

delaPuente:2001:DIO

[dlPZR^{+}01] Juan A. de la Puente, Juan Zamorano, José Ruiz, Ramón Fernández, and Rodrigo García. The design and implementation of the open Ravenscar kernel. ACM SIGADA Ada Letters, 21(1): 85–90, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Delrio:1991:RDR

[DM91] P. Delrio and F. Mazzanti. The risk of destructive runtime errors. ACM SIGADA Ada Letters, 11(1):102–113, January/February 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

deNiz:2023:SRC

[dNW23] Dionisio de Niz and Lutz Wrage. Symbolic refinement for CPS. ACM SIG-ADA Ada Letters, 43(1): 88–93, June 2023. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3631483.3631498.

Dobbs:1983:AEA

[Dob83] Paul Dobbs. Ada experience on the Ada capability study. ACM SIGADA Ada Letters, 2(6):59–62, May/June 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dobbing:1990:DAS

[Dob90] B. Dobbing. Distributed Ada: a suggested solution for Ada 9X. ACM SIG-ADA Ada Letters, 10(9): 94–102, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dobbing:1993:EPM

[Dob93]  Brian Dobbing.  Experi-

ences with the partitions model. ACM SIGADA Ada Letters, 13(2):65–77, March/April 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dobbing:2000:BPA

[Dob00] Brian Dobbing. Building partitioned architectures based on the Ravenscar profile. ACM SIGADA Ada Letters, 20(4):29–31, December 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/dec2000/dobbing-

paper.pdf. Special Issue: Presentations from SIGAda 2000.

Dobbing:2001:OSJ

[Dob01a] Brian Dobbing. Overview of the Sun Java Community Process’s Real-Time Expert Group specification of RT-Java: Session Summary. ACM SIGADA Ada Letters, 21(1):18–19, March 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dobbing:2001:RPH

[Dob01b] Brian Dobbing. The Ravenscar profile for high-integrity Java programs? ACM SIGADA Ada Letters, 21(1):

56–61, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

DoD:1987:DDC

[DoD87a] U. S. DoD. DoD directive 3405.1: Computer programming languages policy. ACM SIGADA Ada Letters, 7(4): 42–44, July/August 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

DoD:1987:DDU

[DoD87b] U. S. DoD. DoD directive 3405.2: Use of Ada in weapon systems. ACM SIG-ADA Ada Letters, 7(4):45– 53, July/August 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Domitz:1987:RTA

[Dom87] R. O. Domitz. Real-Time Ada debugging. ACM SIG-ADA Ada Letters, 7(6): 18–20, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Donaldson:1990:LE

[Don90] Cameron Donaldson. Letter from the editor. ACM SIG-ADA Ada Letters, 10(8):12, November/December 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Doran:1999:ILL

[Dor99] Steven Doran. Interfacing low-level C device drivers with Ada 95. ACM SIG-ADA Ada Letters, 19(3):133– 143, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dowson:1994:PWG

[Dow94] Mark Dowson. “process”

working group summary SETA2. ACM SIGADA Ada Letters, 14(Special Issue): 104–108, Fall 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dissaux:1997:CDT

[DPB^{+}97] Pierre Dissaux, Laurent Pautet, Lars Björnfot, Yvon Kermarrec, and Dominique LeCampion. Communication and distribution tools for embedded distributed applications: a case study with Ada 95 and its distributed systems annex. ACM SIG-ADA Ada Letters, 17(5):40– 44, September/October 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Delange:2009:VSI

[DPP^{+}09] Julien Delange, Laurent Pautet, Alain Plantec, Mickael Kerboeuf, Frank Singhoff, and Fabrice Kordon. Validate, simulate, and implement ARINC653 systems using the AADL. ACM SIG-ADA Ada Letters, 29(3):31– 44, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dobbing:1999:RTP

[DR99] Brian Dobbing and George Romanski. The Ravenscar Tasking Profile — experience reporting. ACM SIG-ADA Ada Letters, 19(2): 28–32, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dobbing:1997:STS

[DRF97] Brian Dobbing and Marc Richard-Foy. T-SMART — task-safe, minimal Ada realtime toolset. ACM SIG-ADA Ada Letters, 17(5):45– 50, September/October 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Davis:1998:TCN

[DRH98] Noël Davis, Scot Ransbottom, and Drew Hamilton. Teaching computer networks through modeling. ACM SIGADA Ada Letters, 18(5):104–110, September/October 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dritc:1989:PHS

[Dri89a]  K. Dritc.  Plugging the

holes in the Sieve of Eratosthenes. ACM SIG-ADA Ada Letters, 9(2):72– 77, March/April 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dritz:1989:PHS

[Dri89b] Kenneth W. Dritz. Plugging the holes in the Sieve of Eratosthenes. ACM SIG-ADA Ada Letters, 9(2):72– 77, March/April 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dritz:1991:PSGa

[Dri91a] K. W. Dritz. Proposed standard for a generic package of elementary functions for Ada. ACM SIGADA Ada Letters, 11(7):9–46, Fall 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dritz:1991:PSGb

[Dri91b] K. W. Dritz. Proposed standard for a generic package of primitive functions for Ada. ACM SIGADA Ada Letters, 11(7):66–82, Fall 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dritz:1991:IPS

[Dri91c] Kenneth W. Dritz. Introduction to the proposed standard for the elementary functions in Ada. ACM SIGADA Ada Letters, 11(7):3–8, Fall 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dritz:1991:RPSa

[Dri91d] Kenneth W. Dritz. Rationale for the proposed standard for a generic package of elementary functions for Ada. ACM SIGADA Ada Letters, 11(7):47–65, Fall 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dritz:1991:RPSb

[Dri91e] Kenneth W. Dritz. Rationale for the proposed standard for a generic package of primitive functions for Ada. ACM SIGADA Ada Letters, 11(7):83–90, Fall 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dross:2022:CSS

[Dro22] Claire Dross. Containers for specification in SPARK. ACM SIGADA Ada Letters, 42(2):62–68, December 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3591335.3591341.

Denzler:2023:EMO

[DRSK23] P. Denzler, D. Ramsauer, D. Scheuchenstuhl, and W. Kastner. Experiences modeling a OPC UA /DDS gateway in AADL in the context of fog computing. ACM SIGADA Ada Letters, 43(1):58, June 2023. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3631483.3631490.

Druffel:1982:NPD

[Dru82] Larry E. Druffel. THe need for a programming discipline to support the APSE: Where does the APSE path lead? ACM SIGADA Ada Letters, 1(4):21–23, May/June 1982. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Drury:1999:UAD

[Dru99] Pace Drury. Using ASIS for data base insulation. ACM SIGADA Ada Letters, 19(1): 64–65, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Das:1987:ALI

[DS87] Souripriya Das and Stephen R.

Schach. An Ada-LISP interface generator. ACM SIG-ADA Ada Letters, 7(4):88– 97, July/August 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Duerinckx:1997:CRC

[Due97] Guido Duerinckx. Cyclic redundancy checks in Ada95. ACM SIGADA Ada Letters, 17(1):41–53, January/ February 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Duff:2008:GLTc

[Duf08a] Bob Duff. Gem # 3: Limited types in Ada 2005 — constructor functions. ACM SIGADA Ada Letters, 28(1): 36–37, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Duff:2008:GLTa

[Duf08b] Bob Duff. Gem #1: Limited types in Ada 2005 — limited aggregates. ACM SIGADA Ada Letters, 28(1): 31–33, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Duff:2008:GLTb

[Duf08c] Bob Duff. Gem #2: Limited types in Ada 2005 — notation in aggregates. ACM SIGADA Ada Letters, 28(1): 34–35, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Duff:2009:GNCa

[Duf09a] Bob Duff. Gem #23: null considered harmful. ACM SIGADA Ada Letters, 29(1): 25–26, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Duff:2009:GNCb

[Duf09b] Bob Duff. Gem #24: null considered harmful (part 2 – efficiency). ACM SIG-ADA Ada Letters, 29(1): 27–28, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Duff:2009:GMA

[Duf09c] Bob Duff. Gem #26: the mod attribute. ACM SIG-ADA Ada Letters, 29(1): 33–34, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Duff:2009:GAC

[Duf09d] Bob Duff. Gem #44: accessibility checks (part III). ACM SIGADA Ada Letters, 29(1):71–73, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Duff:2009:GOR

[Duf09e] Bob Duff. Gem #50: overload resolution. ACM SIG-ADA Ada Letters, 29(1): 81–83, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dulman:2003:VAD

[Dul03]  Leonid Dulman.  Visual

Ada developer. ACM SIG-ADA Ada Letters, 23(1): 30–34, March 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Duncan:1998:RAL

[Dun98] Arthur G. Duncan. Reusable Ada libraries supporting infinite data structures. ACM SIGADA Ada Letters, 18(6): 89–103, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dobbing:2001:RSA

[DV01] Brian Dobbing and Tullio Vardanega. Report of session: analysis of the J consortium real-time Java proposal. ACM SIGADA Ada Letters, 21(1):17–18, March 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

DiGrazia:1987:ADM

[DZM87] Joseph C. DiGrazia, Jehuda Ziegler, and Richard Mueller. An Ada distributed multiprocessor executive: From conceptualization to implementation. In ACM [ACM87a], pages 147–156. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Early:1992:ART

[Ear92] Marvin Early. An Ada realtime executive rate scheduler. ACM SIGADA Ada Letters, 12(2):62–75, March/April 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Edgerton:2001:ERA

[Edg01] Scott Edgerton. Experience report: architecture-based software development on the Crusader program. ACM SIGADA Ada Letters, 21 (4):127–128, December 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ehresman:2001:EMB

[EF01] Kenneth L. Ehresman and Joey L. Frantzen. Electronic maneuvering board and dead reckoning tracer decision aid for the officer of the deck. ACM SIGADA Ada Letters, 21(4):61–70, December 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ekiba:2013:NTT

[EGC13] Takeo Ekiba, Yuichi Goto, and Jingde Cheng. New types of tasking deadlocks in Ada 2012 programs. ACM SIGADA Ada Letters, 33(1):

169–179, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Efstathopoulos:2013:OVE

[EH13] Pavlos Efstathopoulos and Andrew Hawthorn. Optimizing verification effort with SPARK 2014. ACM SIG-ADA Ada Letters, 33(3):19– 20, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Eventoff:1980:RMC

[EHP80] W. Eventoff, D. Harvey, and R. J. Price. The rendezvous and monitor concepts; is there an efficiency difference? In ACM [ACM80], pages 156–165. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Ehrenfried:1994:SAA

[Ehr94] Daniel H. Ehrenfried. Static analysis of Ada programs. ACM SIGADA Ada Letters, 14(4):28–35, July/August 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Etienne:2016:SHP

[EJ16] Jean-Frédéric Etienne and Eric Juppeaux. SafeProver: a high-performance verification tool. ACM SIGADA Ada Letters, 36(2):47–48, December 2016. CODEN AALEE5. ISSN 0736-721X.

Eisenhauer:1989:TTC

[EJK89] Greg Eisenhauer, Rakesh Jha, and J. Michael Kamrad, II. Targeting a traditional compiler to a distributed environment. ACM SIGADA Ada Letters, 9(2): 45–51, March/April 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Eilers:2011:MNE

[EK11] Dan R. Eilers and Tero Koskinen. Making the non-executable ACATS tests executable. ACM SIG-ADA Ada Letters, 31(3):75– 80, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Eilers:2012:AAU

[EK12] Dan R. Eilers and Tero Koskinen. Adapting ACATS for use with run-time checks suppressed. ACM SIG-ADA Ada Letters, 32(3):97– 102, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Evangelista:2004:VLT

[EKPPR04] S. Evangelista, C. Kaiser, J. F. Pradat-Peyre, and P. Rousseau. Verifying linear time temporal logic properties of concurrent Ada programs with Quasar3. ACM SIGADA Ada Letters, 24(1): 17–24, March 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Elliott:1983:RAW

[Ell83] Jon K. Elliott. The ROLM Ada work center. ACM SIGADA Ada Letters, 2(4): 97–100A, January/February 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Elrad:1988:CSC

[Elr88] Tzilla Elrad. Comprehensive scheduling controls for Ada tasking. ACM SIGADA Ada Letters, 8(7):12–19, Fall 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Elrad:1989:IMC

[Elr89] Tzilla Elrad. The issue of mutual control: synchronization and decision making control for Ada. ACM SIG-ADA Ada Letters, 9(4):105– 112, May/June 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Elsom:1990:PAA

[Els90a] K. C. Elsom. Prioritized asynchronism in Ada 9X. ACM SIGADA Ada Letters, 10(9):103–110, Fall 1990.

CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Elsom:1990:SV

[Els90b] Ken C. Elsom. Shared variables. ACM SIGADA Ada Letters, 10(9):29–30, Fall 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Elsom:1990:ACA

[Els90c]  Kenneth Elsom.  Asyn-

chronous communication in Ada. ACM SIGADA Ada Letters, 10(4):57–65, Spring 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Elsom:1991:OOP

[Els91] K. Elsom. Object oriented programming facilities in Ada 9X. ACM SIGADA Ada Letters, 11(6):64–65, September/October 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Emery:1983:DDS

[Eme83]  David Emery.  The De-

partment of Defense Software Initiative, a summary. ACM SIGADA Ada Letters, 2(4):84–87, January/ February 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Emery:1986:TUT

[ER86] Dave Emery and Steve Rosen. Two UNIX tools supporting a common style. ACM SIGADA Ada Letters, 6(2):84, March/April 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Frankel:1982:LAC

[FA82] G. Frankel and R. Arnold. Linkage of Ada components — theme and variations. In ACM [ACM82], pages 201– 211. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Fagin:2000:AIL

[Fag00a] Barry Fagin. An Ada interface to Lego Mindstorms. ACM SIGADA Ada Letters, 20(3):20–40, September 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

                                        /www.acm.org/sigada/ada_
                                        letters/sept2000/mindstorms.
                                        pdf.

Fagin:2000:AMU

[Fag00b] Barry Fagin. Ada/Mindstorms 1.0 user’s guide and reference manual. ACM SIG-ADA Ada Letters, 20(3):32– 40, September 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fairley:1980:ADT

[Fai80] Richard E. Fairley. Ada debugging and testing support environments. In ACM [ACM80], pages 16– 25. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

FDU:1994:RSC

[Fai94] Fairleigh Dickinson University, Teaneck, NJ. Reusable software components. ACM SIGADA Ada Letters, 14 (2):24–49, March/April 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Falis:1982:DIA

[Fal82] Ed Falis. Design and implementation in Ada of a runtime task supervisor. In ACM [ACM82], pages 1– 9. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Falcone:1991:ACE

[Fal91] Marco Falcone. Ada compiler evaluation on the Columbus Software Development Environment Project. ACM SIG-ADA Ada Letters, 11(2):107– 114, March/April 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fantechi:1984:IRE

[Fan84]  A. Fantechi.  Interfacing

with real environments from Ada programs. ACM SIG-ADA Ada Letters, 4(2):35– 43, September/October 1984. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Farkas:1982:ABA

[Far82]  E. Farkas.  Annoying

bagatelles in Ada. ACM SIG-ADA Ada Letters, 1(4):24– 26, May/June 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fassbender:2001:RAP

[Faß01] Heinz Faßbender. Reengineering an Ada95-programmed command and control information system by using UML. ACM SIG-ADA Ada Letters, 21(4):53– 60, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ferrandez:2023:WCE

[FAT^{+}23] Ivan Rodriguez Ferrandez, Alvaro Jover Alvarez, Matina Maria Trompouki, Leonidas Kosmidis, and Francisco J. Cazorla. Worst case execution time and power estimation of multicore and GPU software: a pedestrian detection use case. ACM SIG-ADA Ada Letters, 43(1):

111–117, June 2023. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3631483.3631502.

Favaro:1991:WPR

[Fav91] John Favaro. What price reusability? A case study. ACM SIGADA Ada Letters, 11(3):115–124, Spring 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fong:2010:WIN

[FBL^{+}10] Elizabeth Fong, Paul E.

Black, Richard F. Leslie, Simson Garfinkel, Larry Wagoner, Gary McGraw, and Jeff Williams. Wouldn’t it be nice to have software labels. ACM SIGADA Ada Letters, 30(3): 31–32, December 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ford:1991:AGP

[FC91] Ray Ford and Hong Chew.

AWING: a general purpose command interface generator (and an exercise in software reuse). ACM SIG-ADA Ada Letters, 11(3): 73–82, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fernandez:1983:EMM

[FCS83] John D. Fernandez, Homer Carlisle, and Sallie Sheppard. Experience with matrix multiplication using Ada tasks. ACM SIG-ADA Ada Letters, 2(5):76– 84, March/April 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Feiler:2016:AFT

[FD16] Peter Feiler and Julien Delange. Automated fault tree analysis from AADL models. ACM SIGADA Ada Letters, 36(2):39–46, December 2016. CODEN AALEE5. ISSN 0736-721X.

Feiler:2014:AMB

[Fei14]  Peter H. Feiler.  AADL

and model-based engineering. ACM SIGADA Ada Letters, 34(3):17–18, December 2014. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Feller:1986:SE

[Fel86] Peter H. Feller. The SEI environment. ACM SIG-ADA Ada Letters, 6(2):83, March/April 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Feldman:2009:IA

[Fel09] Michael B. Feldman. Introduction to Ada. ACM SIGADA Ada Letters, 29(3): 1–2, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Feldman:2011:IA

[Fel11] Michael Feldman. Introduction to Ada. ACM SIG-ADA Ada Letters, 31(3):9– 10, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fernandez:1997:TCM

[Fer97] José L. Fernandez. A taxonomy of coordination mechanisms used by real-time processes. ACM SIGADA Ada Letters, 17(2):29–54, March/April 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fantechi:1982:PAP

[FG82] A. Fantechi and F. Gallo.

Portable Ada programming system: a proposed runtime architecture. In ACM [ACM82], pages 48– 56. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Firesmith:1986:RAR

[FG86] Donald G. Firesmith and Colin B. Gilyeat. Resolution of Ada-related concerns in DoD-STD-2167, revision A. ACM SIGADA Ada Letters, 6(5):29–33, September/October 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Falquet:1985:STL

[FGN85] G. Falquet, J. Guyot, and L. Nerima. Simple tools to learn Ada. ACM SIG-ADA Ada Letters, 4(6):44– 48, May/June 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Faasch:1983:AMN

[FHN83] M. Faasch, V. Haarslev, and H.-H. Nagel. Ada on a minicomputer-network for image sequence analysis: An investigative implementation. ACM SIGADA Ada Letters, 2(4):92–96, January/February 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Firesmith:1986:SCL

[Fir86] Donald G. Firesmith. SD-SAWG chairperson’s letter. ACM SIGADA Ada Letters, 6(2):59, March/April 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Firesmith:1987:TIP

[Fir87a] Donald G. Firesmith. Two Impediments to the proper use of Ada. ACM SIG-ADA Ada Letters, 7(5):104, September/October 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Firth:1987:PAA

[Fir87b] Robert Firth. A pragmatic approach to Ada insertion. ACM SIGADA Ada Letters, 7(6):24–26, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Firesmith:1988:MAO

[Fir88] Donald G. Firesmith. Mixing apples and oranges: or what is an Ada line of code anyway? ACM SIGADA Ada Letters, 8(5):110–112, September/October 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Firesmith:1990:OAB

[Fir90] D. G. Firesmith. OOD and Ada bibliography. ACM SIG-ADA Ada Letters, 10(6):114– 128, July/August 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Firesmith:1991:OOG

[Fir91a] Donald Firesmith. Objectoriented graphics for requirements analysis and logical design. ACM SIGADA Ada Letters, 11(9):100–115, November/December 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Firesmith:1991:SAO

[Fir91b] Donald Firesmith. Structured analysis and objectoriented development are not compatible. ACM SIGADA Ada Letters, 11(9):56–66, November/December 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fischer:1983:STI

[Fis83]  Herman Fischer.  Soft-

ware Technology Initiative Raleigh Workshop: An editorial report. ACM SIG-ADA Ada Letters, 2(6):45– 50, May/June 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fisher:1984:LGA

[Fis84a] Gerry Fisher. A LALR(1) grammar for ANSI Ada. ACM SIGADA Ada Letters, 3(4):37–50, January/ February 1984. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fisher:1984:UAP

[Fis84b]  Gerry Fisher.  Universal

arithmetic packages. ACM SIGADA Ada Letters, 3(6): 30–47, May/June 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). See erratum [Fro15].

Fisher:2012:HHA

[Fis12] Kathleen Fisher. HACMS: high assurance cyber military systems. ACM SIG-ADA Ada Letters, 32(3):51– 52, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Fleck:1986:SAM

[Fle86] Thomas J. Fleck. A specification for Ada machine code insertions. ACM SIGADA Ada Letters, 6(6):54–60, November/December 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Flint:1998:UJA

[Fli98] Shayne Flint. Using Java APIs with native Ada compilers. ACM SIGADA Ada Letters, 18(6):193–203, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fernandez-Marina:2009:GACa

[FM09a] Ramón Fernández-Marina.

Gem # 33: accessibility checks (part I: Ada95). ACM SIGADA Ada Letters, 29(1): 51–52, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fernandez-Marina:2009:GACb

[FM09b] Ramón Fernández-Marina.

Gem #41: accessibility checks (part II: Ada2005). ACM SIGADA Ada Letters, 29(1):66–68, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fagin:2001:TCS

[FME01] Barry S. Fagin, Laurence D. Merkle, and Thomas W. Eggers. Teaching computer science with robotics using Ada/Mindstorms 2.0. ACM SIGADA Ada Letters, 21(4): 73–78, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Freitas:1990:OOR

[FMG90] Maria Manuel Freitas, Ana Moreira, and Pedro Guerreiro. Object oriented requirements analysis in an Ada project. ACM SIG-ADA Ada Letters, 10(6):97– 109, July/August 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Filipski:1980:AST

[FMN80] Gary L. Filipski, Donald R. Moore, and Major John E. Newton. Ada as a software transition tool. In ACM [ACM80], pages 176– 182. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Fleener:1998:RLE

[FMS98] Nathan Fleener, Laura Moody, and Mary Stewart. A reusable lightweight executive for command and control systems. ACM SIGADA Ada Letters, 18(6):81–88, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fox:1985:AKD

[FNS^{+}85] Stephen Fox, Anil Nori, John M. Smith, Arvola Chan, and Sy Danberg. Atool kit for database programming in Ada. ACM SIG-ADA Ada Letters, 5(2):41– 57, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fukuyama:1987:EGU

[FOFY87] Shunichi Fukuyama, Naoi Okuse, Matsuto Fujimaru, and Seiichi Yamaski. Empirical guidelines to use Ada effectively. In ACM [ACM87a], pages 25–30. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Foster:2020:WDW

[Fos20] Jeff Foster. Who decides what is allowed?: User interactions and permissions use on Android. ACM SIGADA Ada Letters, 39(1):71, January 2020. ISSN 0736-721X.

URL https://dl.acm.org/

                                        doi/abs/10.1145/3379106.
                                        3379114.

Ferreira:2023:SBS

[FOSC23] José Ferreira, Alan Oliveira, André Souto, and José Cecílio. Software-based security approach for networked embedded devices. ACM SIGADA Ada Letters, 43 (1):73–77, June 2023. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3631483.3631495.

Francl:1987:PMS

[Fra87a]  Fred Francl.  Pioneering

mission-critical software. In ACM [ACM87a], pages 31– 35. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Frankel:1987:IAT

[Fra87b] Gary Frankel. Improving Ada tasking performance. ACM SIGADA Ada Letters, 7(6): 47–48, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

French:1986:API

[Fre86a] Stewart French. AIM project introduction. ACM SIG-ADA Ada Letters, 6(2):85– 86, March/April 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

French:1986:TAS

[Fre86b] Stewart French. Transporting an Ada software tool: a case study. ACM SIG-ADA Ada Letters, 6(2):90– 91, March/April 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fritz:1983:AUD

[Fri83]  Robert Fritz.  The Ada

user and the DoD software initiative. ACM SIG-ADA Ada Letters, 2(5):85– 88, March/April 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Frigo:1987:EVA

[Fri87] G. Vittorio Frigo. Evaluation of the VAX^{T M} Ada(R) compiler and APSE by means of a real program. ACM SIG-ADA Ada Letters, 7(3):84– 93, May/June 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Frisberg:1998:AGF

[Fri98a] Bo Frisberg. Ada in the Gripen flight control system. ACM SIGADA Ada Letters, 18(6):140–141, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Frisberg:1998:UAG

[Fri98b] Bo Frisberg. Usage of Ada in the Gripen flight control system. ACM SIGADA Ada Letters, 18(6):140–141, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Froggatt:1987:FPC

[Fro87]  Terry Froggatt.  Fixed-

point conversion, multiplication, and division in Ada. ACM SIGADA Ada Letters, 7(1):71–81, January/ February 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Froggatt:2015:EAU

[Fro15]  Terry Froggatt.  An er-

ror in the Ada universal arithmetic package. ACM SIGADA Ada Letters, 35 (2):14, August 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). See [Fis84b]. The 32-year-old error is a test with digit t that has if (t > BASE), but the operator should instead be >=.

Fofanov:1997:AID

[FRS97] V. Fofanov, S. Rybin, and A. Strohmeier. ASIStint: An interpreter for debugging and testing ASIS implementations. In ACM [ACM97], pages 205–212. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Flynn:1987:ETA

[FSS87] Susan Flynn, Edith Schonberg, and Edmond Schonberg. The efficient termination of Ada tasks in a multi-processor environment. ACM SIGADA Ada Letters, 7(7):55–76, November/December 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Fujita:1987:SDO

[Fuj87] Shohei Fujita. Self-organizing distributed operating system — implementation and problem using Ada. In ACM [ACM87a], pages 157–158. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Fussichen:1991:AIS

[Fus91] K. Fussichen. Ada in information systems. ACM SIG-ADA Ada Letters, 11(6):77– 79, September/October 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Goldsack:1990:OOA

[GA90] S. J. Goldsack and C. Atkinson. An object oriented approach to virtual nodes: Are package types an answer? ACM SIGADA Ada Letters, 10(4):78–84, Spring 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Galvin:2020:UGR

[Gal20] David S. Galvin. Using genericity and reflection to create a declarative elaboration logger for large scale realtime embedded Ada systems. ACM SIGADA Ada Letters, 40(1):45–52, October 2020. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3431235.3431237.

Galeotti:2022:SLA

[Gal22] Gabriele Galeotti. SweetAda: a lightweight Adabased framework. ACM SIGADA Ada Letters, 42 (1):50–53, June 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3577949.3577955.

Gantsou:2001:TAD

[Gan01] Dhavy Gantsou. Targeting Ada95/DSA for distributed simulation of multiprotocol communication networks. ACM SIGADA Ada Letters, 21(4):91–96, December 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gantsou:2003:AFS

[Gan03] Dhavy Gantsou. An architectural framework for supporting distributed object based routing. ACM SIG-ADA Ada Letters, 23(1): 27–29, March 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gantsou:2004:DMD

[Gan04] Dhavy Gantsou. A DSA model for data access in selforganizing systems. ACM SIGADA Ada Letters, 24(1): 25–28, March 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gardner:1983:UAC

[Gar83] Michael R. Gardner. Using Ada for commercial software. ACM SIGADA Ada Letters, 2(5):56–59, March/April 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gardner:1984:WUP

[Gar84] Michael R. Gardner. When to use private types. ACM SIGADA Ada Letters, 3(6): 66–78, May/June 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gargaro:1990:VND

[Gar90]  Anthony Gargaro.  Vir-

tual nodes/distributed systems working group. ACM SIGADA Ada Letters, 10(4): 66–77, Spring 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gardinier:2009:OSD

[Gar09] Mark Gardinier. Open source development of a safety critical dual redundant (ada95/ C++) signal control program environment (SCOPE). ACM SIGADA Ada Letters, 29(3): 23–30, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gasperoni:2008:GBN

[Gas08]  Franco Gasperoni.  Gem

#7: The beauty of numeric literals in Ada. ACM SIGADA Ada Letters, 28(1): 45–47, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gaumer:1990:PTR

[Gau90a]  Dale Gaumer.  PIWG

test results. ACM SIG-ADA Ada Letters, 10(3):146– 210, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gaumer:1990:RPT

[Gau90b] Dale Gaumer. Reporting PIWG test results. ACM SIGADA Ada Letters, 10(3): 211–216, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gauthier:1995:EHA

[Gau95] Michel Gauthier. Exception Handling in Ada-94: Initial Users’ Requests and Final Features. ACM SIG-ADA Ada Letters, 15(1):70– 82, January/February 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gauthier:1996:WNS

[Gau96] Michel Gauthier. What’s new for scanning with Ada-95? ACM SIGADA Ada Letters, 16(4):57–72, July/August 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gargaro:1987:IWR

[GB87] Anthony Gargaro and Benjamin Brosgol. International workshop on Real-Time Ada issues. ACM SIGADA Ada Letters, 7(6):??, Fall 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Giering:1994:TDS

[GB94] E. W. Giering, III and T. P. Baker. A tool for the deterministic scheduling of real-time programs implemented as periodic Ada tasks. ACM SIGADA Ada Letters, 14(Special Issue): 54–73, Fall 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gacek:2014:RAC

[GBC^{+}14] Andrew Gacek, John Backes, Darren Cofer, Konrad Slind, and Mike Whalen. Resolute: an assurance case language for architecture models. ACM SIGADA Ada Letters, 34(3): 19–28, December 2014. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gonzalez-Barahona:1997:TNP

[GBCGDBC97] J. M. Gonzalez-Barahona, J. Centeno-Gonzalez, P. De las Heras-Quiros, and F. J. Ballesteros-Camara. Teaching network programming with Ada and LowerLayer. In ACM [ACM97], pages 105– 112. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Goforth:1990:PMP

[GCM90] Andre Goforth, Philippe Collard, and Matthew Marquardt. Performance measurement of parallel Ada: An applications based approach. ACM SIGADA Ada Letters, 10(3):38–58, Winter 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gasperoni:2000:MPJ

[GD00] Franco Gasperoni and Gary Dismukes. Multilanguage programming on the JVM: The Ada 95 benefits. ACM SIGADA Ada Letters, 20 (4):3–28, December 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/dec2000/ada-on-

jvm.pdf. Special Issue: Presentations from SIGAda 2000.

GonzalezHarbour:1997:IRC

[GDAG97] M. Gonzalez Harbour, J. M.

Drake Moyano, M. Aldea Rivas, and J. Garcia Fernandez. Implementing robot controllers under real-time POSIX and Ada. ACM SIG-ADA Ada Letters, 17(5):57– 64, September/October 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gutierrez:2002:MSA

[GDHM02] J. Javier Gutiérrez, José M.

Drake, Michael González Harbour, and Julio L. Medina. Modeling and schedulability analysis in the development of real-time distributed Ada systems. ACM SIGADA Ada Letters, 22(4):

58–65, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

GonzalezHarbour:2002:SRT

[GdlP02] Michael González Harbour and Juan Antonio de la Puente. Session on realtime, fault tolerance, and distribution. ACM SIG-ADA Ada Letters, 22(4):123– 124, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Genillard:1991:SML

[Gen91] Christian Genillard. SYN-TAX ANALYSER G: a multilanguage syntax analysis package. ACM SIGADA Ada Letters, 11(1):57–70, January/February 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Genillard:1989:RDR

[GES89] C. Genillard, N. Ebel, and A. Strohmeier. Rationale for the design of reusable abstract data types implemented in Ada. ACM SIG-ADA Ada Letters, 9(2):62– 71, March/April 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Grau:1987:CMA

[GG87] J. Kaye Grau and Kathleen A. Gilroy. Compliant mappings of Ada programs to the DoD-STD-2167 static structure. ACM SIG-ADA Ada Letters, 7(2):73– 84, March/April 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

GutierrezGarcia:1999:PRP

[GG99] J. J. GutiérrezGarcía and M. GonzálezHarbour. Prioritizing remote procedure calls in Ada distributed systems. ACM SIGADA Ada Letters, 19(2):67–72, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gaucher:2016:DES

[GG16] Fabien Gaucher and Yves Génevaux. Debugging embedded systems requirements before the design begins: “The beginning is the most important part of the work”

— Plato. ACM SIGADA Ada Letters, 36(2):58–59, December 2016. CODEN AALEE5. ISSN 0736-721X.

Gargaro:1990:AAD

[GGP^{+}90] A. B. Gargaro, S. J. Goldsack, R. K. Power, R. A. Volz, and A. J. Wellings. Adapting Ada for distribution and fault tolerance. ACM SIG-ADA Ada Letters, 10(9): 111–117, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Garcia:1999:PRP

[GH99] J. J. Gutiérrez García and M. González Harbour. Prioritizing remote procedure calls in Ada distributed systems. ACM SIGADA Ada Letters, 19(2):67–72, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Garcia:2001:TRT

[GH01] José Javier Guitiérrez García and Michael González Harbour. Towards a realtime distributed systems annex in Ada. ACM SIG-ADA Ada Letters, 21(1): 62–66, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

German:1982:MDA

[GHL82] S. M. German, D. P. Helmbold, and D. C. Luckham. Monitoring for deadlocks in Ada tasking. In ACM [ACM82], pages 11– 25. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Gonzalez-Harbour:2003:RSC

[GHV03] Michael Gonzalez-Harbour and Tullio Vardanega. Report of session: current real-time AIs. ACM SIG-ADA Ada Letters, 23(4):22– 23, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Goldsack:1993:TAP

[GHVVW93] S. J. Goldsack, A. A.

Holzbacher-Valero, R. Volz, and R. Waldrop. Translating an AdaPT partition to Ada9X. ACM SIG-ADA Ada Letters, 13(2):78– 90, March/April 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Goldsack:1994:AA

[GHVVW94] S. J. Goldsack, A. A.

Holzbacher-Valero, R. Volz, and R. Waldrop. AdaPT and Ada 9X. ACM SIG-ADA Ada Letters, 14(2):80– 92, March/April 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gibson:2000:TAT

[Gib00] David S. Gibson. Two approaches to teaching software components using Ada 95. ACM SIGADA Ada Letters, 20(1):38–57, March 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gicca:1990:SSA

[Gic90]  Greg Gicca.  A simple

standardized Ada command line interface. ACM SIG-ADA Ada Letters, 10(5):88– 100, May/June 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gicca:1991:RSR

[Gic91] Greg Gicca. Reuse System: software repository tool concepts. ACM SIGADA Ada Letters, 11(1):70–81, January/February 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Giddings:1996:DSU

[Gid96] Victor Giddings. Distributed systems using CORBA and Ada. ACM SIGADA Ada Letters, 16(5):59–69, September/October 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gilroy:1984:EAG

[Gil84] Kathleen Gilroy. Experience with Ada for the graphical kernal system. ACM SIG-ADA Ada Letters, 4(2):54– 64, September/October 1984. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gilroy:1992:RSa

[Gil92a]  K. Gilroy.  Rendezvous

summary. ACM SIGADA Ada Letters, 12(4):12–??, July/August 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gilroy:1992:RSb

[Gil92b] K. Gilroy. Rendezvous summary. ACM SIGADA Ada Letters, 12(5):15–??, September/October 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gilroy:1992:RSc

[Gil92c] K. Gilroy. Rendezvous summary. ACM SIGADA Ada Letters, 12(6):16–??, November/December 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gilroy:1993:RSa

[Gil93a]  K. Gilroy.  Rendezvous

summary. ACM SIGADA Ada Letters, 13(2):12–??, March/April 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gilroy:1993:RSb

[Gil93b]  K. Gilroy.  Rendezvous

summary. ACM SIGADA Ada Letters, 13(3):15–??, May/June 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gilroy:1993:RSc

[Gil93c] K. Gilroy. Rendezvous summary. ACM SIGADA Ada Letters, 13(5):12–??, September/October 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gilroy:1993:RSd

[Gil93d] K. Gilroy. Rendezvous summary. ACM SIGADA Ada Letters, 13(6):28–??, November/December 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gilroy:1994:RSa

[Gil94a]  K. Gilroy.  Rendezvous

summary. ACM SIGADA Ada Letters, 14(2):16–??, March/April 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gilroy:1994:RSb

[Gil94b]  K. Gilroy.  Rendezvous

summary. ACM SIGADA Ada Letters, 14(3):14–??, May/June 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gilchrist:1999:AAM

[Gil99a]  Ian Gilchrist.  Attitudes

to Ada — a market survey. ACM SIGADA Ada Letters, 19(3):229–242, September 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gilchrist:1999:AAU

[Gil99b]  Ian Gilchrist.  Attitudes

to Ada in the UK highreliability software sector (plenary session). ACM SIG-ADA Ada Letters, 19(3):

221, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Goldenberg:1989:AAS

[GL89] Joanne Goldenberg and Gertrude Levine. Ada’s abort statement: license to kill. ACM SIGADA Ada Letters, 9(6):97–103, September/October 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gluch:2009:ESE

[Glu09] David Gluch. Embedded systems engineering with the AADL: modeling & analysis. ACM SIGADA Ada Letters, 29(3):7–8, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gargaro:1997:DFT

[GLV97] Anthony Gargaro, Douglass Locke, and Richard Volz. Distributed and fault tolerant systems (session summary). ACM SIGADA Ada Letters, 17(5):8–10, September/October 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Garrido:2016:SER

[GLZdlP16] Jorge Garrido, Beatriz Lacruz, Juan Zamorano, and Juan A. de la Puente. In support of extending the Ravenscar profile. ACM SIGADA Ada Letters, 36(1):63–67, June 2016. CODEN AALEE5. ISSN 0736-721X.

Gray:1992:RSS

[GMO92] Lewis Gray, David S. Maibor, and Jim O’Day. Report from the SIGAda Software Development Standards and Ada Working Group (SDSAWG). ACM SIG-ADA Ada Letters, 12(2):31– 32, March/April 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Goldfedder:1993:CIP

[Gol93] Brandon Goldfedder. Counterintuitive programming. ACM SIGADA Ada Letters, 13 (4):63–70, July/August 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gonzalez:1988:ATD

[Gon88] D. W. Gonzalez. An Ada tasking demo. ACM SIG-ADA Ada Letters, 8(5):87– 91, September/October 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gonzalez:1990:MSC

[Gon90] Dean W. Gonzalez. Multitasking software components. ACM SIGADA Ada Letters, 10(1):92–96, January/February 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gonzalez:1991:CHA

[Gon91a]  D. W. Gonzalez.  Con-

sidered harmful (Ada private types). ACM SIG-ADA Ada Letters, 11(2):56– 59, March/April 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gonzalez:1991:CH

[Gon91b] Dean W. Gonzalez. “=” considered harmful. ACM SIG-ADA Ada Letters, 11(2):56– 59, March/April 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Goodenough:1980:ACV

[Goo80] John B. Goodenough. The Ada compiler validation capability. In ACM [ACM80], pages 1–8. CODEN SIN-ODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Goodenough:1985:DA

[Goo85] John B. Goodenough. On defining “=” in Ada. ACM SIGADA Ada Letters, 4 (4):27–31, January/February 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Goodenough:1990:RTT

[Goo90] John Goodenough. Realtime tasking semantics working group. ACM SIG-ADA Ada Letters, 10(4): 32–48, Spring 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Goodenough:2013:BCS

[Goo13] John B. Goodenough. Building confidence in system behavior. ACM SIGADA Ada Letters, 33(3):49–50, December 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gordon:1983:BPD

[Gor83] Michael Gordon. The Byron program design language -1-. ACM SIGADA Ada Letters, 2(4):76–83, January/February 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gorski:2020:UEB

[Gór20]  Janusz Górski.  Using

evidence-based arguments to support dependability assurance — experiences and challenges. ACM SIG-ADA Ada Letters, 40(1):53– 59, October 2020. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3431235.3431238.

Gonzalez:1993:ADA

[GP93] Dean W. Gonzalez and Tim Peart. Applying dimensional analysis. ACM SIG-ADA Ada Letters, 13(4):77– 86, July/August 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gregertsen:2018:SSC

[GP18] Kristoffer Nyborg Gregertsen and Luis Miguel Pinho. Session summary: Clock issues. ACM SIGADA Ada Letters, 38(1):77–78, June 2018. CO-DEN AALEE5. ISSN 0736-721X.

Garrido:2021:VCD

[GPZdlP21] Jorge Garrido, David Pisonero, Juan Zamorano, and Juan A. de la Puente. Vectorization challenges in digital signal processing. ACM SIG-ADA Ada Letters, 40(2): 92–95, April 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463478.3463490.

Groves:1980:DVM

[GR80] L. J. Groves and W. J.

Rogers. The design of a virtual machine for Ada. In ACM [ACM80], pages 223– 234. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Gaumer:1990:RTR

[GR90] Dale Gaumer and Daniel Roy. Reporting test results. ACM SIGADA Ada Letters, 10(3): 211–216, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Grabber:1983:MWA

[Gra83] Eran Grabber. The middle way approach for Ada based PDL syntax. ACM SIGADA Ada Letters, 2(4):64–67, January/February 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Green:1990:AVP

[Gre90] Geir Green. Access values pointing to any object. ACM SIGADA Ada Letters, 10 (5):101–109, May/June 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Grein:1999:AF

[Gre99a] Christoph Grein. Add finalization. ACM SIG-ADA Ada Letters, 19(4):24– 31, December 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Grein:1999:SP

[Gre99b] Christoph Grein. Safe pointers. ACM SIGADA Ada Letters, 19(4):44–48, December 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Grein:2005:DLL

[Gre05] Christoph Grein. Dead live longer: a dramoletto. ACM SIGADA Ada Letters, 25(3): 28–31, September 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gregertsen:2013:ERP

[Gre13] Kristoffer Nyborg Gregertsen. An extended Ravenscar profile for execution time control. ACM SIG-ADA Ada Letters, 33(2):109– 114, August 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gregertsen:2016:RAT

[Gre16] Kristoffer Nyborg Gregertsen. Revising the Ada timers and group budgets to support execution time control for interrupt handling. ACM SIGADA Ada Letters, 36(1): 39–50, June 2016. CODEN AALEE5. ISSN 0736-721X.

Gregertsen:2018:PPC

[Gre18] Kristoffer Nyborg Gregertsen. Position paper: Clock support in Ada. ACM SIG-ADA Ada Letters, 38(1):55– 57, June 2018. CODEN AALEE5. ISSN 0736-721X.

Gregertsen:2021:EER

[Gre21] Kristoffer Nyborg Gregertsen. Ember: an embedded robotics library in SPARK. ACM SIGADA Ada Letters, 41(2):61–65, December 2021. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3530801.3530803.

Griffin:1995:ASA

[Gri95] Michael D. Griffin. 1995 ACM/SIGAda Awards Program. ACM SIGADA Ada Letters, 15(5):16–??, September 1, 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Grier:1998:EPU

[Gri98] Samuel Grier. Early projects using Ada at the United States Air Force Academy. ACM SIGADA Ada Letters, 18(1):92–109, January/ February 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Grover:1986:EMI

[Gro86] Vinod Grover. On expressing module interconnections in Ada. ACM SIGADA Ada Letters, 6(1):90–93, January/February 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Grosman:2007:HEA

[Gro07]  Tom Grosman.  Hibachi:

the Eclipse Ada Development Toolset. ACM SIG-ADA Ada Letters, 27(3): 99, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gupta:1985:ESM

[GS85] Rajiv Gupta and Mary Lou Soffa. The efficiency of storage management schemes for Ada programs. ACM SIGADA Ada Letters, 5(2): 164–172, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14–16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Goodenough:1988:PCP

[GS88] John B. Goodenough and Lui Sha. The priority ceiling protocol: a method for minimizing the blocking of high priority Ada tasks. ACM SIGADA Ada Letters, 8(7): 20–31, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Garcia:2002:ERI

[GS02] Rodrigo García García and Alfred Strohmeier. Experiences report on the implementation of EPTs for GNAT. ACM SIGADA Ada Letters, 22(4):22–27, December 2002. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gregertsen:2010:ETC

[GS10] Kristoffer Nyborg Gregertsen and Amund Skavhaug. Execution-time control for interrupt handling. ACM SIGADA Ada Letters, 30(1): 33–44, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gregertsen:2013:ETT

[GS13] Kristoffer Nyborg Gregertsen and Amund Skavhaug. Execution time timers for interrupt handling. ACM SIGADA Ada Letters, 33(2): 87–96, August 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gaudel:2011:ADP

[GSP^{+}11] Vincent Gaudel, Frank Singhoff, Alain Plantec, Stephane Rubini, Pierre Dissaux, and Jerome Legrand. An Ada design pattern recognition tool for AADL performance analysis. ACM SIGADA Ada Letters, 31(3):61–68, December 2011. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gargaro:1997:FDA

[GST^{+}97] Anthony Gargaro, Gary Smith, Ronald J. Theriault, Richard A. Volz, and Raymond Waldrop. Future directions in Ada — distributed execution and heterogeneous language interoperability toolsets. ACM SIG-ADA Ada Letters, 17(5):51– 56, September/October 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gargaro:1997:ACA

[GSTV97] A. Gargaro, G. Smith, R. J.

Theriault, and R. A. Volz. Aria-Java communication in ADEPT. In ACM [ACM97], pages 231–246. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Gedela:1999:FMS

[GSX99] Ravi K. Gedela, Sol M. Shatz, and Haiping Xu. Formal modeling of synchronization methods for concurrent objects in Ada 95. ACM SIG-ADA Ada Letters, 19(3):211– 220, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Goos:1980:TCF

[GW80] Gerhard Goos and Georg Winterstein. Towards a compiler front-end for Ada. In ACM [ACM80], pages 36–46. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Garrido:2015:AIP

[GZdlP15] Jorge Garrido, Juan Zamorano, and Juan A. de la Puente. ARINC-653 inter-partition communications and the Ravenscar profile. ACM SIGADA Ada Letters, 35(1): 38–45, April 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Garrido:2018:PAP

[GZdlP18] Jorge Garrido, Juan Zamorano, and Juan A. de la Puente. On protocols for accessing protected objects on multiprocessors. ACM SIGADA Ada Letters, 38(1):29–33, June 2018. CODEN AALEE5. ISSN 0736-721X.

Haden:1990:LML

[Had90]  Steven Haden.  LEXI-

CAL ANALYZER G: a multilanguage lexical analysis package. ACM SIGADA Ada Letters, 10(1):131–139, January/February 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hagihara:1991:AJ

[Hag91] T. Hagihara. Ada in Japan. In ACM [ACM91a], pages 367–375. ISBN 0-89791-445-7. LCCN ????

Hait:2000:AOP

[Hai00] Férial Benachour Hait. Agent oriented programming with Ada 95: Application to financial markets. ACM SIG-ADA Ada Letters, 20(1): 67–80, March 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hall:1983:ADM

[Hal83] Patrick A. V. Hall. Adding database management to Ada. ACM SIGADA Ada Letters, 2(4):88–91, January/February 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hart:1982:ADA

[Har82]  Hal Hart.  Ada for de-

sign: An approach for transitioning industry software developers. ACM SIG-ADA Ada Letters, 2(1):50– 57, July/August 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Harbaugh:1985:XEA

[Har85] Sam Harbaugh. XAda — an executable Ada design language methodology. ACM SIGADA Ada Letters, 4(6): 27–31, May/June 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Harkleroad:1987:AAC

[Har87] Joseph Harkleroad. Analyzing Ada concurrent algorithms. ACM SIGADA Ada Letters, 7(2):118–134, March/April 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Harbaugh:1988:CRM

[Har88]  Sam Harbaugh.  Com-

ments and recommendation on MOSI from an Ada point of view. ACM SIGADA Ada Letters, 8(2):107–109, March/April 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hart:1994:LCC

[Har94a]  Hal Hart.  Letter from

the Chair: Changes in the Ada world. ACM SIG-ADA Ada Letters, 14(2):13– ??, March/April 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hart:1994:MC

[Har94b] Hal Hart. Message from the Chair. ACM SIG-ADA Ada Letters, 14(3):12– ??, May/June 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hart:1994:SBG

[Har94c] Hal Hart. SIGAda being a good citizen within ACM and helping Ada too! ACM SIG-ADA Ada Letters, 14(4):12– 15, July/August 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hart:1997:SEP

[Har97] H. Hart. Software engineering plan reviews: Better or worse for Ada than the mandate. In ACM [ACM97], pages 305– 307. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Harbour:1999:DAR

[Har99a] Michael Gonzalez Harbour.

Distributed Ada and realtime (session summary). ACM SIGADA Ada Letters, 19(2):15–18, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hart:1999:SAW

[Har99b] Hal Hart. 1998 SIGAda awards winners and 1999 nominations. ACM SIG-ADA Ada Letters, 19(1): 19–??, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hart:2000:SAW

[Har00] Hal Hart. 1999 SIGAda awards winners and 2000 nominations. ACM SIG-ADA Ada Letters, 20(1):

12–15, March 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hart:2001:SAN

[Har01] Hal Hart. SIGAda 2000 awards and 2001 nomination information. ACM SIGADA Ada Letters, 21 (2):89, June 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hardin:2022:HSC

[Har22] David Hardin. Hardware/ software co-assurance for the Rust programming language applied to zero trust architecture development. ACM SIGADA Ada Letters, 42 (2):55–61, December 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3591335.3591340.

Harmon:1988:AIM

[HB88] Marion G. Harmon and Ted P. Baker. An Ada implementation of Marsaglia’s “universal” random number generator. ACM SIGADA Ada Letters, 8(2):110–112, March/April 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hagar:1996:UFS

[HB96] Jon Hagar and James M. Bieman. Using formal specifications as test oracles for system-critical software. ACM SIGADA Ada Letters, 16(6):55–72, November/December 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hart:1999:WHI

[HBTW99] Hal Hart, Barry Boehm, S. Tucker Taft, and Tony Wasserman. What happened to integrated environments? (panel session). ACM SIG-ADA Ada Letters, 19(3):225– 226, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hendrix:1998:GSE

[HCBM98a] T. Dean Hendrix, James H.

Cross, II, Larry A. Barowski, and Karl S. Mathias. GRASP: software engineering with Ada 95 for Windows 95 and NT. ACM SIGADA Ada Letters, 18(1):70–77, January/February 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hendrix:1998:VSI

[HCBM98b] T. Dean Hendrix, James H.

Cross, II, Larry A. Barowski, and Karl S. Mathias. Visual support for incremental abstraction and refinement in Ada 95. ACM SIGADA Ada Letters, 18(6):142–147, November/December 1998.

CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Also mistakenly reprinted on pp. 153–157.

Hendrix:1998:AGU

[HCT^{+}98] T. Dean Hendrix, James H. Cross, II, Joe C. Teate, Larry A. Barowski, and Karl S. Mathias. Assessing GRASP utilization through instrumentation. ACM SIG-ADA Ada Letters, 18(5):51– 56, September/October 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Humphries:2004:MPA

[HCW04] Jeffrey W. Humphries, Martin C. Carlisle, and Terry A. Wilson. Multilanguage programming with Ada in the .NET environment. ACM SIGADA Ada Letters, 24(1): 1–3, March 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hammons:1985:CCP

[HD85] Charles Hammons and Paul Dobbs. Coupling, cohesion, and package unity in Ada. ACM SIGADA Ada Letters, 4(6):49–59, May/June 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hopper:1998:UAD

[HDHH98] Jim Hopper, Jennifer DeVilbiss, Harry Heaton, and Tom Haberlandt. Use of Ada 95 in Digital Radar Landmass Simulation (DRLMS). ACM SIGADA Ada Letters, 18(6):137–139, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Heaney:2004:CSA

[Hea04] Matthew J. Heaney. Charles:

an STL for Ada95. ACM SIGADA Ada Letters, 24(3): 23–30, September 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Heaney:2008:GKB

[Hea08a]  Matthew Heaney.  Gem

#5: Key-based searching in set containers. ACM SIG-ADA Ada Letters, 28(1): 38–40, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Heaney:2008:GAM

[Hea08b] Matthew Heaney. Gem #6:

The Ada95 multiple views idiom vs. Ada05 interfaces. ACM SIGADA Ada Letters, 28(1):41–44, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Heaney:2008:GFF

[Hea08c] Matthew Heaney. Gem #8:

Factory functions. ACM SIGADA Ada Letters, 28(1):

48–51, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Heaney:2008:GCO

[Hea08d] Matthew Heaney. Gem #9: Classwide operations, iterators, and generic algorithms. ACM SIGADA Ada Letters, 28(1):52–58, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Heker:1983:SCE

[Hek83] Wolf-Dieter Heker. Some comments on “experiences with matrix multiplication using Ada tasks”. ACM SIG-ADA Ada Letters, 3(2):76– ??, September/October 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Heker:1989:SER

[Hek89] Wolf-Dieter Heker. Sieve of Eratosthenes revisited. ACM SIGADA Ada Letters, 9(5): 83, July/August 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hulse:1999:RMC

[HEUV99] Christine Hulse, Scott Edgerton, Michael Ubnoske, and Louis Vazquez. Reducing maintenance costs through the application of modern software architecture principles. ACM SIGADA Ada Letters, 19(3):101–110, September 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Harbaugh:1984:TSU

[HF84] Sam Harbaugh and John A.

Forakis. Timing studies using a synthetic Whetstone benchmark. ACM SIGADA Ada Letters, 4(2):23–35, September/October 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Harbour:2007:PPL

[HG07] Michael González Harbour and J. Javier Gutiérrez. Programming patterns and libraries: Introduction. ACM SIGADA Ada Letters, 27(2): 37–40, August 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hugues:2014:LAS

[HG14] Jérôme Hugues and Christophe Garion. Leveraging Ada 2012 and SPARK 2014 for assessing generated code from AADL models. ACM SIG-ADA Ada Letters, 34(3):39– 46, December 2014. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hughes:1990:EED

[HHBC90] D. Hughes, L. Hoffman, D. Brundelle, and J. Chelini. An example of eventdriven asynchronous scheduling with Ada. ACM SIG-ADA Ada Letters, 10(9): 130–144, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hibbard:1986:SAS

[HHR^{+}86] Peter Hibbard, Andy Hisgen, Jonathan Rosenberg, Mary Shaw, and Mark Sherman. Studies in Ada style. ACM SIGADA Ada Letters, 6(2): 103, March/April 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hilfinger:1982:ISA

[Hil82]  P. N. Hilfinger.  Im-

plementation strategies for Ada tasking idioms. In ACM [ACM82], pages 26– 30. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Hild:2022:ALG

[Hil22] Stefan Hild. Ada looks good, now program a game without knowing anything. ACM SIGADA Ada Letters, 42 (1):41–42, June 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3577949.3577952.

Hirasuna:1992:UIP

[Hir92] Michael Hirasuna. Using inheritance and polymorphism with Ada in government sponsored contracts. ACM SIGADA Ada Letters, 12 (2):43–56, March/April 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hirasuna:1994:ASIa

[Hir94a] Michael Hirasuna. An Ada 9X subset for inheritancebased reuse and its translation to Ada 83 (part 1). ACM SIGADA Ada Letters, 14(1):50–60, January/February 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hirasuna:1994:ASIb

[Hir94b] Michael Hirasuna. An Ada 9X subset for inheritancebased reuse and its translation to Ada 83 (part 2). ACM SIGADA Ada Letters, 14 (2):58–67, March/April 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hirasuna:1994:BSS

[Hir94c] Michael Hirasuna. BATCES solution #2: a simplified SA/ OOD approach. ACM SIG-ADA Ada Letters, 14(3):39– 60, May/June 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Helbold:1985:RDD

[HL85a] D. Helbold and D. C. Luckham. Runtime detection and description of deadness errors in Ada tasking. ACM SIG-ADA Ada Letters, 4(6):60– 72, May/June 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Helmbold:1985:RDD

[HL85b] D. Helmbold and D. C. Luckham. Runtime detection and description of deadness in Ada tasking. ACM SIG-ADA Ada Letters, 4(6):60– 72, May/June 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Helmbold:1985:TTS

[HL85c] David Helmbold and David C. Luckham. TSL: Task sequencing language. ACM SIGADA Ada Letters, 5(2): 255–274, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14–16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Harrison:1986:GIA

[HL86] George C. Harrison and Dar-Biau Liu. Generic implementations via analogies in the Ada programming language. ACM SIG-ADA Ada Letters, 6(4):34– 43, July/August 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hisgen:1980:RRA

[HLRS80] Andy Hisgen, David Alex Lamb, Jonathan Rosenberg, and Mark Sherman. A runtime representation for Ada variables and types. In ACM [ACM80], pages 82–90. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Howell:1991:EHL

[HM91] C. Howell and D. Mularz. Exception handling in large Ada systems. In ACM [ACM91b], pages 90–101. ISBN 0-89791-393-0. LCCN ????

Howe:2003:AFV

[HM03] Douglas J. Howe and Stephen Michell. An approach to formal verification of real time concurrent Ada programs. ACM SIGADA Ada Letters, 23(4):87–92, December 2003. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Herr:1988:CVR

[HMC88] C. S. Herr, D. G. McNicholl, and S. G. Cohen. Compiler validation and reusable Ada parts for real-time, embedded applications. ACM SIG-ADA Ada Letters, 8(5):75– 86, September/October 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Harbour:1997:IRC

[HMRF97] M. González Harbour, J. M. Drake Moyano, M. Aldea Rivas, and J. García Fernández. Implementing robot controllers under real-time POSIX and Ada. ACM SIGADA Ada Letters, 17(5):57–64, September/October 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hamilton:2000:PLI

[HMZ00] J. A. Drew Hamilton, Jr., Jeanne L. Murtagh, and Richard G. Zoller. Programming language impacts on learning. ACM SIG-ADA Ada Letters, 20(3):12– 19, September 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.acm.org/sigada/ada_
letters/sept2000/pascal.
pdf.

Hoffman:1998:TGA

[HNS98] Daniel Hoffman, Jayakrishnan Nair, and Paul Strooper. Testing generic Ada packages with APE. ACM SIGADA Ada Letters, 18(6):255–262, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hodgson:1991:PSP

[Hod91a] Graham S. Hodgson. Proposed standard for packages of real and complex type declarations and basic operations for Ada (including vector and matrix types). ACM SIGADA Ada Letters, 11(7): 91–130, Fall 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hodgson:1991:RPS

[Hod91b] Graham S. Hodgson. Rationale for the proposed standard for packages of real and complex type declarations and basic operations for Ada (including vector and matrix types). ACM SIGADA Ada Letters, 11(7): 131–139, Fall 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hoffmann:1986:ADT

[Hof86] K. E. Hoffmann. Appropriate data-types in Ada (apparently not a simple subject). ACM SIGADA Ada Letters, 6(1):20–21, January/February 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hoskins:1988:DIK

[Hos88]  Rose Hoskins.  The de-

sign and implementation of a Karel compiler and interpreter. ACM SIG-ADA Ada Letters, 8(4):83– 96, July/August 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hosch:1989:MPA

[Hos89] Frederick A. Hosch. Message passing and administrators in Ada. ACM SIG-ADA Ada Letters, 9(2):106– 117, March/April 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hosch:1990:GIC

[Hos90] Frederick A. Hosch. Generic instantiations as closures. ACM SIGADA Ada Letters, 10(1):122–130, January/February 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Houghton:1983:TTF

[Hou83] Raymond C. Houghton. A taxonomy of tool features for the Ada programming support environment (APSE). ACM SIGADA Ada Letters, 3(3):63–78, November/December 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hovater:2000:DGU

[Hov00] Steven V. Hovater. Document generation using ASIS tools. ACM SIGADA Ada Letters, 20(4):40–49, December 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.acm.org/sigada/ada_
letters/dec2000/hovater-

paper.pdf. Special Issue: Presentations from SIGAda 2000.

Howell:1986:MCI

[How86] Chuck Howell. Minutes of CAIS implementor’s group meeting. ACM SIGADA Ada Letters, 6(2):75–76, March/April 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Harbour:2001:SSD

[HP01] Michael González Harbour and Luis Miguel Pinho. Session summary: distribution and real-time. ACM SIG-ADA Ada Letters, 21(1): 14–16, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Haertig:1981:TST

[HPT81] Herman Haertig, Andreas Pfitzmann, and Leo Treff. Task state transitions in Ada. ACM SIG-ADA Ada Letters, 1(1):31– 41, July/August 1981. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Harbour:2003:MME

[HR03] Michael González Harbour and Mario Aldea Rivas. Managing multiple executiontime timers from a single task. ACM SIGADA Ada Letters, 23(4):28–31, December 2003. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hallmark:2007:PEG

[HR07] Tyler B. Hallmark and Eugene K. Ressler. Parallel evolution of game evaluation functions in Ada. ACM SIG-ADA Ada Letters, 27(3):59– 62, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Harbaugh:1987:GPM

[HS87] Sam Harbaugh and Greg Saunders. GKS/Ada post mortem, a cost analysis. In ACM [ACM87a], pages 14– 24. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Heinfeld:1998:SET

[HS98] Blaine W. Heinfeld and James L. Silver. A software engineering training program for non-software engineers. ACM SIGADA Ada Letters, 18(5):39–46, September/October 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hatcliff:2022:ACL

[HSB^{+}22] John Hatcliff, Danielle Stewart, Jason Belt, Robby ., and August Schwerdfeger. An AADL contract language supporting integrated modeland code-level verification. ACM SIGADA Ada Letters, 42(2):45–54, December 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3591335.3591339.

Hutcheon:1987:PDD

[HSW87] A. D. Hutcheon, D. S. Snowden, and A. J. Wellings. Programming and debugging distributed real-time applications in Ada. ACM SIGADA Ada Letters, 7(6): 73–76, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hardin:2012:DCD

[HSWP12] David S. Hardin, Konrad L. Slind, Michael W. Whalen, and Tuang-Hung Pham. A DSL for crossdomain security. ACM SIG-ADA Ada Letters, 32(3):53– 62, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Huff:1982:FQA

[Huf82] Edward Huff. FIFO queues in Ada: An exercise. ACM SIGADA Ada Letters, 1(4): 32–33, May/June 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hugues:2023:MRP

[Hug23] J. Hugues. Mechanization of the Ravenscar profile in Coq. ACM SIGADA Ada Letters, 43(1):106–110, June 2023. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3631483.3631501.

Humphrey:2022:BFV

[Hum22] Laura Humphrey. Basic formal verification of aWaypoint manager for unmanned air vehicles in SPARK. ACM SIGADA Ada Letters, 42 (2):79–88, December 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3591335.3591344.

Hunt:1988:IA

[Hun88] J. R. Hunt. Interrupts and Ada. ACM SIGADA Ada Letters, 8(7):61–64, Fall 1988.

CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Huijsman:1987:TAP

[HvKPT87] R. D. Huijsman, J. van Katwijk, C. Pronk, and W. J. Toetenel. Translating Algol 60 programs into Ada: Report on a feasability study. ACM SIGADA Ada Letters, 7(5):42–50, September/October 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hucheon:1988:SAD

[HW88a] A. D. Hucheon and A. J.

Wellings. Supporting Ada in a distributed environment. ACM SIGADA Ada Letters, 8(7):113–117, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hutcheon:1988:SAD

[HW88b] A. D. Hutcheon and A. J.

Wellings. Supporting Ada in a distributed environment. ACM SIGADA Ada Letters, 8(7):113–117, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Imbruglia:2022:CSC

[ICS22] Antonio Imbruglia, Daniela Cancila, and Marina Settembre. 5G communication and security in connected vehicles. ACM SIG-ADA Ada Letters, 42(2):109– 113, December 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3591335.3591351.

Inverardi:1983:DKA

[ILMV83] P. Inverardi, G. Levi, U. Montanari, and G. N. Vallario. A distributed KAPSE architecture. ACM SIGADA Ada Letters, 3(2):55–61, September/October 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Inverardi:1985:UAD

[IMM85] P. Inverardi, F. Mazzanti, and C. Montangero. The use of Ada in the design of distributed systems. ACM SIG-ADA Ada Letters, 5(2):85– 96, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Irwin:1996:CLM

[Irw96] Jess Irwin. Choosing a language for maintainable software. ACM SIGADA Ada Letters, 16(1):54–57, January/February 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

ISO-IEC-JTC1-SC22-WG9:1991:PSGa

[ISO91a] ISO-IEC and JTC1 and SC22 and WG9 (Ada) Numerics Rapporteur Group. Proposed standard for a generic package of elementary functions for Ada. ACM SIGADA Ada Letters, 11(7):9–46, Fall 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

ISO-IEC-JTC1-SC22-WG9:1991:PSGb

[ISO91b] ISO-IEC and JTC1 and SC22 and WG9 (Ada) Numerics Rapporteur Group. Proposed standard for a generic package of primitive functions for Ada. ACM SIG-ADA Ada Letters, 11(7): 66–82, Fall 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Jones:1982:CED

[JA82] A. Jones and A. Ardo. Comparative efficiency of different implementations of the Ada rendezvous. In ACM [ACM82], pages 212– 223. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Jackson:2013:EDS

[Jac13] Ethan K. Jackson. Engineering domain-specific languages with formula 2.0. ACM SIG-ADA Ada Letters, 33(3):3– 4, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

James:1998:DMU

[Jam98a] Scott James. A dataflow model using protected types in a distributed system. ACM SIGADA Ada Letters, 18(6):39–44, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

James:1998:EDD

[Jam98b] Scott James. The evolution of a distributed dataflow processing model using Ada. ACM SIGADA Ada Letters, 18(6):39–44, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

James:1999:RDA

[Jam99] Scott James. Redistribution in distributed Ada. ACM SIGADA Ada Letters, 19(3): 3–8, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Jansohn:1988:ADS

[Jan88] Hans-Stephan Jansohn. Ada for distributed systems. ACM SIGADA Ada Letters, 8(7): 101–103, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Jarzombek:2007:WSA

[Jar07] Joe Jarzombek. Wanted:

software with assurance built-in. ACM SIGADA Ada Letters, 27(3):9–10, December 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Jover-Alvarez:2022:SCA

[JARKS22] Alvaro Jover-Alvarez, Ivan Rodriguez, Leonidas Kosmidis, and David Steenari. Space compression algorithms acceleration on embedded multi-core and GPU platforms. ACM SIG-ADA Ada Letters, 42(1): 100–104, June 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3577949.3577969.

John:2022:APA

[JBT^{+}22] Wolfgang John, Ali Balador, Jalil Taghia, Andreas Johnsson, Johan Sjöberg, Ian Marsh, Jonas Gustafsson, Federico Tonini, Paolo Monti, Pontus Sköldström, and Jim Dowling. ANIARA Project — automation of network edge infrastructure and applications with artificial intelligence. ACM SIGADA Ada Letters, 42 (2):92–95, December 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3591335.3591347.

Jha:1989:ISD

[JEKC89] Rakesh Jha, Greg Eisenhauer, J. Michael Kamrad, II, and Dennis Cornhill. An implementation supporting distributed execution of partitioned Ada programs. ACM SIGADA Ada Letters, 9(1):147–160, January/February 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Jennings:2009:SLL

[Jen09] Trevor J. Jennings. SPARK: the Libre language and toolset for high-assurance software engineering. ACM SIGADA Ada Letters, 29(3): 9–10, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Jarc:1998:ESW

[JF98a] Duane J. Jarc and Michael B. Feldman. An empirical study of Web-based algorithm animation courseware in an Ada data structure course. ACM SIGADA Ada Letters, 18(6):68–74, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Jarc:1998:SES

[JF98b] Duane J. Jarc and Michael B. Feldman. A [sic] empirical study of Web-based algorithm animation courseware in an Ada data structure course. ACM SIGADA Ada Letters, 18(6):68–74, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Jha:1990:PAI

[Jha90] Rakesh Jha. Parallel Ada:

Issues in programming and implementation. ACM SIG-ADA Ada Letters, 10(9): 126–132, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Jones:1985:ISR

[JLM^{+}85] Bill Jones, Steve Litvintchouk, Jerry Mungle, Herb Krasner, John Melby, and Herb Willman. Issues in software reusability. ACM SIG-ADA Ada Letters, 4(5):97– 99, March/April 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Future Ada Environment Workshop.

Johansson:1993:OOP

[Joh93]  Henrik Johansson.  Ob-

ject oriented programming and virtual functions in conventional languages (an extended abstract). ACM SIG-ADA Ada Letters, 13(4):44– 48, July/August 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Johns:1994:AAI

[Joh94] Janet Faye Johns. Activities of the artificial intelligence working group. ACM SIG-ADA Ada Letters, 14(2):50– 53, March/April 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Jemli:2010:MAK

[JR10] Mamdouh Jemli and Jean-Pierre Rosen. A methodology for avoiding known compiler problems using static analysis. ACM SIG-ADA Ada Letters, 30(3):23– 30, December 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kamrad:1983:ROA

[Kam83] J. Michael Kamrad. Runtime organization for the Ada language system programs. ACM SIGADA Ada Letters, 3(3):58–68, November/December 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kamrad:1991:PRA

[Kam91] Mike Kamrad. Protected records in Ada 9X. ACM SIGADA Ada Letters, 11(6): 49–53, September/October 1991. CODEN AALEE5.

ISSN 1094-3641 (print), 1557-9476 (electronic).

Kamrad:1995:SAW

[Kam95] Mike Kamrad. Summary of ARTEWG workshop on distributed systems. ACM SIG-ADA Ada Letters, 15(5):34– 45, September/October 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kamrad:1998:AER

[Kam98] Mike Kamrad. Ada experience report for BlazeNet, Inc. ACM SIGADA Ada Letters, 18(6):215–216, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kamrad:1999:FTS

[Kam99] Mike Kamrad. Fault tolerance (session summary). ACM SIGADA Ada Letters, 19(2):10–11, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kanig:2012:GGC

[Kan12a] Johannes Kanig. Gem #104:

Gprbuild and configuration files — part 1. ACM SIG-ADA Ada Letters, 32(2):43– 44, August 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kanig:2012:LEA

[Kan12b] Johannes Kanig. Leadingedge Ada verification technologies: combining testing and verification with GNAT-Test and GNATProve — the Hi-Lite Project. ACM SIG-ADA Ada Letters, 32(3):5– 6, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Krieg-Brueckner:1983:CCA

[KB83] Berndt Krieg-Brueckner. Consistency checking in Ada and Anna: a transformational approach. ACM SIGADA Ada Letters, 3(2):46–54, September/October 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Karam:1987:EAT

[KB87] Gerald M. Karam and Raymond J. A. Buhr. Experience with the automatic temporal analysis of multitasking Ada designs. In ACM [ACM87a], pages 36–44. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Kim:1997:CSD

[KB97a] Hyoseob Kim and Cornelia Boldyreff. A case study on design pattern discovery in Ada. ACM SIGADA Ada Letters, 17(6):98–107, November/December 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kim:1997:SRI

[KB97b] Hyoseob Kim and Cornelia Boldyreff. Software reusability issues in code and design. ACM SIGADA Ada Letters, 17(6):91–97, November/December 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Krieg-Bruekner:1980:ATL

[KBL80] Bernd Krieg-Bruekner and David C. Luckham. ANNA: Towards a language for annotating Ada programs. In ACM [ACM80], pages 128– 138. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Kirkham:1984:USS

[KBT84] J. A. Kirkham, A. Burns, and R. J. Thomas. The use of structured systems analysis in the rapid creation of information management systems prototypes written in Ada. ACM SIG-ADA Ada Letters, 4(1):74– 87, July/August 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kamrad:1990:DC

[KC90] M. Kamrad and J. Cross. Distributed communications. ACM SIGADA Ada Letters, 10(9):85–93, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1982:SPA

[Ker82]  Judith Kerner.  Should

PDL/Ada be compilable? ACM SIGADA Ada Letters, 2(2):49–50, September/October 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1986:ADD

[Ker86] Judy Kerner. Ada DL developers matrix update. ACM SIGADA Ada Letters, 6(2): 57–58, March/April 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1988:ADL

[Ker88a]  J. Kerner.  Ada design

language developers matrix. ACM SIGADA Ada Letters, 8(6):35–48, November/December 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1988:DMC

[Ker88b]  J. Kerner.  Development

methodology committee — ADL developers matrix. ACM SIGADA Ada Letters, 8(3):69–80, May/June 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1989:ADL

[Ker89]  J. Kerner.  Ada design

language developers matrix. ACM SIGADA Ada Letters, 9(4):30–42, May/June 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1990:ADLa

[Ker90a] Judy Kerner. Ada design language developers matrix. ACM SIGADA Ada Letters, 10(5):48–61, May/June 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1990:ADLb

[Ker90b] Judy Kerner. Ada Design Language Developers Matrixr. ACM SIGADA Ada Letters, 10(8):34, November/December 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1992:ADLa

[Ker92a]  Judy Kerner.  Ada De-

sign Language/CASE developers matrix. ACM SIG-ADA Ada Letters, 12(3):67– 83, May/June 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1992:ADLb

[Ker92b] Judy Kerner. Ada design language/CASE developers matrix. ACM SIGADA Ada Letters, 12(6):29–45, November/December 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1993:ADLa

[Ker93a]  Judy Kerner.  Ada de-

sign language/CASE developer matrix. ACM SIG-ADA Ada Letters, 13(3):21– ??, May/June 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1993:ADLb

[Ker93b] Judy Kerner. Ada design language/CASE developers matrix. ACM SIGADA Ada Letters, 13(6):37–55, November/December 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1994:ADLa

[Ker94a]  Judy Kerner.  Ada de-

sign language/CASE developers matrix. ACM SIG-ADA Ada Letters, 14(3):20– 38, May/June 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1994:ADLb

[Ker94b] Judy Kerner. Ada design language/CASE developers matrix. ACM SIGADA Ada Letters, 14(6):19–40, November/December 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1995:ADL

[Ker95] Judy Kerner. Ada design language/CASE developers matrix. ACM SIGADA Ada Letters, 15(6):22–43, November/December 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1996:ADLa

[Ker96a]  Judy Kerner.  Ada de-

sign language/CASE developers matrix. ACM SIG-ADA Ada Letters, 16(3): 19, May/June 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1996:ADLb

[Ker96b] Judy Kerner. Ada design language/CASE matrix — updates only. ACM SIGADA Ada Letters, 16(6):40–50, November/December 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1997:ADL

[Ker97] Judy Kerner. Ada design language/CASE matrix — updates only. ACM SIG-ADA Ada Letters, 17(4):74– 87, July/August 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kerner:1998:CAA

[Ker98]  Judy Kerner.  Commer-

cially available Ada design language/CASE productsupdates only. ACM SIG-ADA Ada Letters, 18(4): 22–31, July 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kermarrec:1999:CVA

[Ker99] Yvon Kermarrec. CORBA vs. Ada 95 DSA: a programmer’s view. ACM SIG-ADA Ada Letters, 19(3):39– 46, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kruchten:1996:ATI

[KETT96] Philippe Kruchten, Dan Ehrenfried, Kim Thompson, and Chris Thompson. Ada type interchange — moving data between platforms. ACM SIGADA Ada Letters, 16(1):46–53, January/February 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kaisler:1998:OOC

[KF98] Stephen H. Kaisler and Michael B. Feldman. Objectoriented and concurrent program design issues in Ada 95. ACM SIGADA Ada Letters, 18(6):246–254, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kann:1997:EPA

[KFS97] Charles W. Kann, Michael B.

Feldman, and John Sibert. Experience programming applets with Ada95. ACM SIG-ADA Ada Letters, 17(3):17– 29, May/June 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Uses an early beta release of AppletMagic for compiling Ada95 programs into code for the Java Virtual Machine.

Kuang:1998:IEH

[KGL98] Shan Kuang, K. M. George, and Lan Li. Implementation of event handling in GNA95GP. ACM SIG-ADA Ada Letters, 18(2): 53–66, March 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kamrad:1985:ART

[KGW^{+}85] Mike Kamrad, Kathleen Gilroy, Daryl Winters, Dock Allen, and Charles Mckay. Ada run-time environments working group (ARTEWG) report. ACM SIGADA Ada Letters, 5(3–6):63, November/December 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Khrabrov:1995:ALS

[Khr95] Alexy V. Khrabrov. An Ada-like separate compilation style in C. ACM SIG-ADA Ada Letters, 15(2):23– 30, March/April 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kiem:1989:KSD

[Kie89] Eric Kiem. The KEYSTONE system design methodology. ACM SIGADA Ada Letters, 9(5):101–108, July/August 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kienzle:1997:NAA

[Kie97] J. Kienzle. Network application in Ada 95. In ACM [ACM97], pages 3–10. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Kienzle:1999:CTT

[Kie99]  Jörg Kienzle.  Combin-

ing tasking and transaction. ACM SIGADA Ada Letters, 19(2):49–53, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kienzle:2001:EC

[Kie01] Jörg Kienzle. Exceptions and concurrency. ACM SIG-ADA Ada Letters, 21(3):13– 15, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kimmet:2021:AGC

[Kim21] James S. Kimmet. Autogenerated coherent data store for concurrent modular embedded systems. ACM SIG-ADA Ada Letters, 41(1): 74–77, June 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3570315.3570321.

Kirtchev:2012:NRE

[Kir12] Hristian Hristov Kirtchev. A new robust and efficient implementation of controlled types in the GNAT compiler. ACM SIGADA Ada Letters, 32(3):43–50, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Kamrad:1987:DA

[KJEC87] Mike Kamrad, Rakesh Jha, Greg Eisenhauer, and Dennis Cornhill. Distributed Ada. ACM SIGADA Ada Letters, 7(6):113–115, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Korochkin:2003:EPA

[KK03] Dmitry Korochkin and Sergey Korochkin. Experimental performance analysis of the Ada95 and Java parallel program on SMP systems. ACM SIGADA Ada Letters, 23(1): 53–56, March 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Klem:1989:KSD

[Kle89] E. Klem. The KEYSTONE system design methodology. ACM SIGADA Ada Letters, 9(5):101–108, July/August 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Klein:2006:UAL

[Kle06] Judith Klein. Use of Ada in Lockheed Martin for air traffic management and beyond. ACM SIGADA Ada Letters, 26(3):1, December 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kleinke:2021:CLL

[Kle21] Brian Kleinke. Challenges and lessons learned introducing an evolving open source technology into an established legacy Ada and C++ program. ACM SIG-ADA Ada Letters, 40(2): 70–72, April 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463478.3463485.

Klumpp:1987:ALA

[Klu87] Allan R. Klumpp. An Ada linear algebra package modeled after HAL/S. In ACM [ACM87a], pages 101–110. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Knapper:1981:RC

[KM81] Robert J. Knapper and Robert F. Mathis. Roberts Corporation. ACM SIG-ADA Ada Letters, 1(1):29– 30, July/August 1981. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kordon:1998:FAF

[KM98] Fabrice Kordon and Jean-Luc Mounier. FrameKit, an Ada framework for a fast implementation of CASE enviroments. ACM SIGADA Ada Letters, 18(5):57–66, September/October 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kini:1982:TIA

[KMS82] V. Kini, D. F. Martin, and A. Stoughton. Testing the INRIA Ada formal definition: The USC-ISI formal semantics project. In ACM [ACM82], pages 120– 128. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Krishnan:2008:SAT

[KNB08] R. Krishnan, Margaret Nadworny, and Nishil Bharill. Static analysis tools for security checking in code at Motorola. ACM SIG-ADA Ada Letters, 28(1): 76–82, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Knight:1987:AFT

[Kni87] John C. Knight. Ada on fault-tolerant distributed systems. ACM SIGADA Ada Letters, 7(6):61–63, Fall 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Knight:1990:AAP

[Kni90] John C. Knight. On the assessment of Ada performance. ACM SIGADA Ada Letters, 10(3):1–6, Winter 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Knight:2009:ENA

[Kni09] John Knight. Echo: a new approach to formal verification based on Ada. ACM SIGADA Ada Letters, 29(3):

85–86, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kurbel:1986:PAIb

[KP86a] K. Kurbel and W. Pietsch. A portable Ada implementation of index sequential inputoutput, part 2. ACM SIG-ADA Ada Letters, 6(3):31– 42, May/June 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kurbel:1986:PAIa

[KP86b] Karl Kurbel and Wolfram Pietsch. A portable Ada implementation of index sequential input-output, Part 1. ACM SIGADA Ada Letters, 6(2):29–40, March/ April 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kaiser:1997:CRP

[KPP97] C. Kaiser and J. F. Pradat-Peyre. Comparing the reliability provided by tasks or protected objects for implementing a resource allocation service: a case study. In ACM [ACM97], pages 51–66. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Kaiser:2006:CJC

[KPPÉR06] Claude Kaiser, Jean-François Pradat-Peyre, Sami Évangelista, and Pierre Rousseau. Comparing Java, C# and Ada monitors queuing policies: a case study and its Ada refinement. ACM SIG-ADA Ada Letters, 26(2): 23–37, August 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kaufman:1993:TAC

[KPR93] Vitali Sh. Kaufman, Mikhail V. Pavlov, and Sergei I. Rybin. Testing of Ada compiler diagnostics. ACM SIG-ADA Ada Letters, 13(4):71– 76, July/August 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Klemm:2021:OAH

[KQT^{+}21] Michael Klemm, Eduardo Quiñones, Tucker Taft, Dirk Ziegenbein, and Sara Royuela. The OpenMP API for high integrity systems: Moving responsibility from users to vendors. ACM SIG-ADA Ada Letters, 40(2): 48–50, April 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463478.3463480.

Knight:1988:NAF

[KR88] John C. Knight and Marc E.

Rouleau. A new approach to fault tolerance in distributed Ada programs. ACM SIGADA Ada Letters, 8(7): 123–126, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kienzle:2001:CTT

[KR01a] Jörg Kienzle and Alexander Romanovsky. Combining tasking and transactions, part II: open multithreaded transactions. ACM SIGADA Ada Letters, 21(1): 67–74, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kienzle:2001:IEO

[KR01b] Jörg Kienzle and Alexander Romanovsky. Implementing exceptions in open multithreaded transactions based on Ada 95 exceptions. ACM SIGADA Ada Letters, 21(3): 57–63, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kruchten:1990:EHL

[Kru90] Philippe Kruchten. Error handling in large, objectbased Ada systems. ACM SIGADA Ada Letters, 10(7): 91–103, September/October 1990. CODEN AALEE5.

ISSN 1094-3641 (print), 1557-9476 (electronic).

Kok:1984:PSB

[KS84] J. Kok and G. T. Symm. A proposal for standard basic functions in Ada. ACM SIGADA Ada Letters, 4(3): 44–52, November/December 1984. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kallberg:2001:SSS

[KS01] Björn Källberg and Rei Stråhle. Ship system 2000, a stable architecture under continuous evolution. ACM SIGADA Ada Letters, 21(4): 47–52, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Klein:2006:PFP

[KS06] Judith Klein and Drasko Sotirovski. Publisher Frame-Work (PFW). ACM SIG-ADA Ada Letters, 26(2):12– 22, August 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kanig:2012:HLC

[KSD12] Johannes Kanig, Edmond Schonberg, and Claire Dross. Hi-Lite: the convergence of compiler technology and program verification. ACM SIG-ADA Ada Letters, 32(3):27– 34, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Kownacki:1987:PED

[KT87] Ron Kownacki and S. Tucker Taft. Portable and efficient dynamic storage management in Ada. In ACM [ACM87a], pages 190–198. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Knight:1984:IUA

[KU84] John C. Knight and John I. A. Urquhart. On the implementation and use of Ada on fault-tolerant distributed systems. ACM SIGADA Ada Letters, 4(3):53–64, November/December 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kirchgassner:1983:OA

[KUP^{+}83] Walter Kirchgassner, Jurgen Uhl, Guido Perch, Manfred Dausmann, Sophia Drossopoulou, Hans-Stephan Jansohn, and Rudolph Landwehr. Optimization in Ada. ACM SIGADA Ada Letters, 3(3): 45–57, November/December 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Krishnam:1988:ITT

[KVT88a] P. Krishnam, R. A. Volz, and R. J. Theriault. Implementation of task types in distributed Ada. ACM SIGADA Ada Letters, 8(7): 104–107, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Krishnan:1988:ITT

[KVT88b] P. Krishnan, R. A. Volz, and R. J. Theriault. Implementation of task types in distributed Ada. ACM SIGADA Ada Letters, 8(7): 104–107, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kenward:1991:AUI

[KW91] P. D. Kenward and B. A. Wichmann. Approved uniformity issues. ACM SIGADA Ada Letters, 11(1):114–120, January/February 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kiddle:1998:EPT

[KW98] O. P. Kiddle and A. J. Wellings. Extensible protected types. ACM SIGADA Ada Letters, 18(6):229–239, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kuo:2011:GTDa

[KW11a] Dean Kuo and Angela Wallenburg. Gem #71: tokeneer discovery — lesson 1. ACM SIGADA Ada Letters, 31(1):32–36, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kuo:2011:GTDb

[KW11b] Dean Kuo and Angela Wallenburg. Gem #71: tokeneer discovery — lesson 2. ACM SIGADA Ada Letters, 31(1):37–38, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kuo:2011:GTDc

[KW11c] Dean Kuo and Angela Wallenburg. Gem #73: tokeneer discovery — lesson 3. ACM SIGADA Ada Letters, 31(1):39–42, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kuo:2011:GTDd

[KW11d] Dean Kuo and Angela Wallenburg. Gem #73: tokeneer discovery — lesson 4. ACM SIGADA Ada Letters, 31(1):43–46, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kuo:2011:GTDe

[KW11e] Dean Kuo and Angela Wallenburg. Gem #73: tokeneer discovery — lesson 5. ACM SIGADA Ada Letters, 31(1):47–48, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kuo:2011:GTDf

[KW11f] Dean Kuo and Angela Wallenburg. Gem #73: tokeneer discovery — lesson 6. ACM SIGADA Ada Letters, 31(1):49–52, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lundqvist:1999:FMA

[LA99] Kristina Lundqvist and Lars Asplund. A formal model of the Ada Ravenscar tasking profile; delay until. ACM SIGADA Ada Letters, 19(3): 15–21, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ladden:1989:SIC

[Lad89] Richard M. Ladden. A survey of issues to be considered in the development of an object-oriented development methodology for Ada. ACM SIGADA Ada Letters, 9(2): 78–89, March/April 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lahtinen:1982:MAA

[Lah82] Pekka Lahtinen. A machine architecture for Ada.

ACM SIGADA Ada Letters, 2(2):28–33, September/October 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lane:2010:SSI

[Lan10] Chris Lane. Systems software integrity assurance. ACM SIGADA Ada Letters, 30(3): 11–12, December 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lapping:2004:MDD

[Lap04] Andy Lapping. Model driven development with Ada. ACM SIGADA Ada Letters, 24(4): 19–22, December 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Larson:2014:FSP

[Lar14] Brian R. Larson. Formal semantics for the PACE-MAKER system specification. ACM SIGADA Ada Letters, 34(3):47–60, December 2014. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Larraz:2022:FLS

[Lar22] Daniel Larraz. Finding locally smallest cut sets using Max-SMT. ACM SIG-ADA Ada Letters, 42(2):32– 39, December 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3591335.3591337.

Latour:1991:MDR

[Lat91] Larry Latour. A methodology for the design of reuse engineered Ada components. ACM SIGADA Ada Letters, 11(3):103–113, Spring 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lathrop:2009:DAB

[Lat09] Steven M. Lathrop. Dynamic analysis of branch mispredictions in Ada. ACM SIG-ADA Ada Letters, 29(3):79– 84, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lau:2007:USB

[Lau07]  Kung-Kiu Lau.  Using

SPARK for a beginner’s course on reasoning about imperative programs. ACM SIGADA Ada Letters, 27(3): 75–78, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Laval:1995:ISR

[Lav95] Philippe Laval. Implementing self-reproducing artificial organisms with Ada. ACM SIGADA Ada Letters, 15 (2):46–53, March/April 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lawlis:1997:AAA

[Law97] P. K. Lawlis. Is the answer always Ada? In ACM [ACM97], pages 297– 304. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Lovengreen:1980:FMT

[LB80] Hans Henrik Lovengreen and Dines Bjorner. On a formal model of the tasking concept in Ada. In ACM [ACM80], pages 213– 222. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Llamosi:1984:UTR

[LBO84] Albert Llamosi, Pere Botella, and Fernando Orejas. On unlimited types and reliability of Ada programs. ACM SIG-ADA Ada Letters, 4(1):50– 60, July/August 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Little:1986:CSE

[LC86] Joyce C. Little and Lillian N. Cassel, editors. Computer science education: Papers of the seventeenth SIGCSE technical symposium (Cincinnati, Ohio, February 6–7, 1986). ACM Press, New York, NY, USA, 1986. ISBN 0-89791-178-4. LCCN QA76.27.A79 v.18 no.1. US$28. ACM Order No 457860. Published as ACM SIGCSE Bull. 18, Feb. 6–7, 1986.

Loureiro:2022:DLR

[LC22] J. Loureiro and J. Cecílio. Deep learning for reliable communication optimization on autonomous vehicles. ACM SIGADA Ada Letters, 42(1):90–94, June 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3577949.3577967.

Loseby:2009:USR

[LCB09] Chad Loseby, Peter Chapin, and Carl Brandon. Use of SPARK in a resource constrained embedded system. ACM SIGADA Ada Letters, 29(3):87–90, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lee:1991:RAA

[LCN91] Pen-Nan Lee, Chi-Hua Chin, and W. Nehman. A reselect alternative for Ada’s selective wait statement. ACM SIG-ADA Ada Letters, 11(2):72– 85, March/April 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lucas:1987:RAD

[LD87] L. Lucas and D. Dent. Real-Time Ada demonstration. In ACM [ACM87a], pages 159– 163. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Leach:1987:ETC

[Lea87a] Ronald J. Leach. Experiences teaching concurrency in Ada. ACM SIGADA Ada Letters, 7(5):40–41, September/October 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Leavitt:1987:APF

[Lea87b] Randal Leavitt. Adjustable precision floating point arithmetic in Ada. ACM SIG-ADA Ada Letters, 7(5):63– 78, September/October 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Leake:2004:ISA

[Lea04] Stephen Leake. Introduction to Stephe’s Ada library. ACM SIGADA Ada Letters, 24(3): 31–43, September 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Leblang:1982:ASB

[Leb82] D. B. Leblang. Abstract syntax based programming environments. In ACM [ACM82], pages 187–200. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Lederman:1992:DEB

[Led92] M. Lederman. The Difference Engine book review. ACM SIGADA Ada Letters, 12 (4):42–??, July/August 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ledru:1995:PTE

[Led95a]  Pascal Ledru.  Protected

types with entry barriers depending on parameters of the entries: some practical examples. ACM SIG-ADA Ada Letters, 15(4):46– 49, July/August 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ledru:1995:TPT

[Led95b] Pascal Ledru. Translation of the protected type mechanism in Ada 83. ACM SIG-ADA Ada Letters, 15(1):64– 69, January/February 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lefebvre:1987:RMA

[Lef87] Phillip J. Lefebvre. Reclamation of memory allocated for dynamic Ada tasking. In ACM [ACM87a], pages 199– 207. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Leif:1996:CA

[Lei96] Robert C. Leif. Commercializing Ada. ACM SIG-ADA Ada Letters, 16(1):44– 45, January/February 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Leif:1999:ADC

[Lei99a] Robert C. Leif. Ada developers cooperative license: (draft) version 0.3. ACM SIGADA Ada Letters, 19(1): 97–107, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Leif:1999:SWH

[Lei99b] Robert C. Leif. SIGAda ’98 workshop: How do we expedite the commercial use of Ada? ACM SIG-ADA Ada Letters, 19(1): 28–39, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Leif:2000:SWH

[Lei00] Robert C. Leif. SIGAda 99, workshop: how do we expedite the commercial use of Ada? ACM SIG-ADA Ada Letters, 20(2): 19–26, June 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

                                        www.acm.org/sigada/ada_
                                        letters/june2000/leif.pdf.

Leif:2002:SWC

[Lei02] Robert C. Leif. SIGAda 2001 workshop, “Creating a symbiotic relationship between XML and Ada”. ACM SIG-ADA Ada Letters, 22(3):24– 41, September 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Leif:2006:WCA

[Lei06] Robert C. Leif. Workshop, commercializing Ada. ACM SIGADA Ada Letters, 26(1): 16–17, April 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Leino:2012:DVP

[Lei12a] K. Rustan M. Leino. Developing verified programs with Dafny. ACM SIG-ADA Ada Letters, 32(3):9– 10, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Leino:2012:PPU

[Lei12b] K. Rustan M. Leino. Program proving using intermediate verification languages (IVLs) like Boogie and Why3. ACM SIGADA Ada Letters, 32(3):25–26, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Leonard:1985:AGK

[Leo85] Thomas M. Leonard. Ada and the Graphical Kernel System. ACM SIGADA Ada Letters, 5(2):136–150, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Leroy:2001:ET

[Ler01] Pascal Leroy. Exceptions as types. ACM SIG-ADA Ada Letters, 21(3):33– 34, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Leroy:2003:IA

[Ler03] Pascal Leroy. An invitation to Ada 2005. ACM SIG-ADA Ada Letters, 23(3):33– 55, September 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levy:1982:AAS

[Lev82a] Arnold J. Levy. The Ada atom system environment. ACM SIGADA Ada Letters, 1(4):34–45, May/June 1982. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levy:1982:MBD

[Lev82b] Arnold J. Levy. Motivation behind the design of the Ada atom system environment. ACM SIG-ADA Ada Letters, 1(3):62– 63, March/April 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1988:CPI

[Lev88] Gertrude Levine. The control of priority inversion in Ada. ACM SIGADA Ada Letters, 8(6):53–56, November/December 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1989:CDA

[Lev89] Gertrude Levine. Controlling deadlock in Ada. ACM SIGADA Ada Letters, 9(4): 87–91, May/June 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1990:RSC

[Lev90] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 10(5):62– 65, May/June 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1991:SWI

[Lev91] G. Levine. Signaling from within interrupt handlers reconsidered. ACM SIG-ADA Ada Letters, 11(2):53– 55, March/April 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1992:RSCa

[Lev92a] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 12(3):84– 91, May/June 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1992:RSCb

[Lev92b] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 12(5):43– ??, September/October 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1993:RSCa

[Lev93a] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 13(1):60– 62, January/February 1993.

CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1993:RSCb

[Lev93b] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 13(3):62– 73, May/June 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1993:RSCc

[Lev93c] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 13(4):23– 28, July/August 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1993:RSCd

[Lev93d] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 13(5):17– 19, September/October 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1993:RSCe

[Lev93e] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 13(6): 56–60, November/December 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1994:RSCa

[Lev94a] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 14(4):23– 27, July/August 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1994:RSCb

[Lev94b] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 14(5):47– 63, September/October 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1994:RSCc

[Lev94c] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 14(6): 41–52, November/December 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1995:RSCa

[Lev95a] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 15(1):24– 27, January/February 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1995:RSCb

[Lev95b] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 15(3):50– 70, May/June 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1995:RSCc

[Lev95c] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 15(5):26– 31, September/October 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1995:RSCd

[Lev95d] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 15(6): 44–45, November/December 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1996:RSCa

[Lev96a] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 16(1):25– 35, January/February 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1996:RSCb

[Lev96b] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 16(4):20– 44, July/August 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1997:GLA

[Lev97a] Gertrude Levine. The Game of Life with Ada tasks. ACM SIGADA Ada Letters, 17(6):19–31, November/December 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1997:RSCa

[Lev97b] Trucy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 17(1):25– 34, January/February 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1997:RSCb

[Lev97c] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 17(4):66– 73, July/August 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1998:DCA

[Lev98a]  Trudy Levine.  Deadlock

control with Ada95. ACM SIGADA Ada Letters, 18(2): 67–80, March 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1998:RSCa

[Lev98b] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 18(1):33– 39, January/February 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1998:RSCb

[Lev98c] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 18(4): 32–46, July 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1999:RSCa

[Lev99a] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 19(1): 22–27, March 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:1999:RSCb

[Lev99b] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 19(4):11– 12, December 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2000:RSC

[Lev00] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 20 (2):27–37, June 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http://

www.acm.org/sigada/ada_
letters/june2000/levine.
pdf.

Levine:2001:CRR

[Lev01a]  Gertrude Levine.  Con-

flict resolution for readers and writers. ACM SIG-ADA Ada Letters, 21(2): 81–88, June 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2001:RSC

[Lev01b] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 21(2): 17–25, June 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2002:RSCa

[Lev02a] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 22(1): 29–38, March 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2002:RSCb

[Lev02b] Trudy Levine, Jr. Reusable software components. ACM SIGADA Ada Letters, 22(3): 20–23, September 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2004:RSC

[Lev04] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 24(3):47– 48, September 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2005:ACI

[Lev05a] Gertrude Levine. Ada and the control of intrusion. ACM SIGADA Ada Letters, 25(3):

32–39, September 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2005:RSCa

[Lev05b] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 25(1): 57–65, March 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2005:RSCb

[Lev05c] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 25(2): 45–53, June 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2005:RSC

[Lev05d] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 25(3):40– 48, September 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2006:RSC

[Lev06] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 26(2): 75–83, August 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2008:RSC

[Lev08] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 28(1): 59–70, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2009:ACD

[Lev09a] Gertrude Levine. Ada for the control of degradation of service. ACM SIG-ADA Ada Letters, 29(2): 20–27, August 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2009:RSC

[Lev09b] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 29(1): 84–97, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2010:RSC

[Lev10] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 30(2): 67–78, August 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2011:PIF

[Lev11a]  Gertrude Levine.  Prior-

ity inversion with fungible resources. ACM SIG-ADA Ada Letters, 31(2):

9–14, August 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2011:RSCa

[Lev11b] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 31(1): 53–63, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2011:RSCb

[Lev11c] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 31(2): 59–69, August 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2013:RSC

[Lev13] Trudy Levine. Reusable software components. ACM SIG-ADA Ada Letters, 33(2):133– 140, August 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levine:2015:RSC

[Lev15a] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 35(2): 15–21, August 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Levy:2015:ITD

[Lev15b] David C. Levy. Illustrating timing drift. ACM SIG-ADA Ada Letters, 35(2):9– 13, August 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lewis:2002:SPG

[Lew02] Bruce Lewis. Software portability gains realized with METAH and Ada95. ACM SIGADA Ada Letters, 22(4): 37–46, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Leveson:2012:SES

[LFT12] Nancy Leveson, Cody Harrison Fleming, and John Thomas. Safety of embedded software. ACM SIG-ADA Ada Letters, 32(3):7– 8, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Locke:1988:PAC

[LG88] C. D. Locke and J. B. Goodenough. A practical application of the ceiling protocol in a real-time system. ACM SIGADA Ada Letters, 8(7):35–38, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Landwehr:1987:MPA

[LHBK87] Rudolf Landwehr, Peter Hensel, Rami Bayan, and Antonio Kung. A model for a portable Ada run-time library. ACM SIGADA Ada Letters, 7(6):93–96, Fall 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Larson:2013:IAE

[LHFD13] Brian Larson, John Hatcliff, Kim Fowler, and Julien Delange. Illustrating the AADL error modeling annex (v.2) using a simple safety-critical medical device. ACM SIG-ADA Ada Letters, 33(3):65– 84, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Li:1982:OSM

[Li82] W. Li. An operational semantics of multitasking and exception handling in Ada. In ACM [ACM82], pages 138–151. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Lindley:1982:APD

[Lin82] Lawrence M. Lindley. Ada program design language survey. ACM SIGADA Ada Letters, 2(3):32–33, November/December 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lindley:1983:APD

[Lin83] Lawrence M. Lindley. Ada program design language survey update. ACM SIGADA Ada Letters, 2(4):61–63, January/February 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Liskov:2012:KPP

[Lis12] Barbara Liskov. Keynote presentation: Programming the Turing machine. ACM SIG-ADA Ada Letters, 32(3):23– 24, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Littlefield:1997:OOA

[Lit97] Arthur Irving Littlefield, III.

An object-oriented approach to automated generation of challenge examinations using Ada 95. ACM SIG-ADA Ada Letters, 17(1):54– 68, January/February 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Liebrenz:2016:AAA

[LKH16] Timm Liebrenz, Verena Klös, and Paula Herber. Automatic analysis and abstraction for model checking HW/SW codesigns modeled in SystemC. ACM SIGADA Ada Letters, 36(2):9–17, December 2016. CODEN AALEE5. ISSN 0736-721X.

Loeper:1997:COA

[LKN97] Hans Loeper, Amro Khattab, and Peter Neubert. Concurrent objects in Ada 95. ACM SIGADA Ada Letters, 17(6):47–64, November/December 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ly:2019:SDA

[LKSL19] Dara Ly, Nikolai Kosmatov, Julien Signoles, and Frédéric Loulergue. Soundness of a dataflow analysis for memory monitoring. ACM SIG-ADA Ada Letters, 38(2):97– 108, December 2019. ISSN 0736-721X. URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3375408.3375416.

Liu:1988:MPF

[LL88] Jane W. S. Liu and Kwei-Jay J. Lin. On means to provide flexibility in scheduling. ACM SIGADA Ada Letters, 8(7):32–34, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Leif:1998:AEB

[LL98] Robert C. Leif and Suzanne B. Leif. Ada in embedded boards for scientific and medical instruments. ACM SIGADA Ada Letters, 18(6):114–120, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Llamosi:1992:APT

[Lla92] Albert Llamosí. On Ada packages, types and task types. ACM SIGADA Ada Letters, 12(5):47–58, September/October 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Leif:2003:XAC

[LLL03] Robert C. Leif, Suzanne B.

Leif, and Stephanie H. Leif. XML and Ada complement each other. ACM SIG-ADA Ada Letters, 23(1): 44, March 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Litvintchouk:1983:AARa

[LM83a] Steven D. Litvintchouk and A. S. Matsumoto. An algebraic approach to reusable Ada components. ACM SIG-ADA Ada Letters, 3(1):51– 54, July/August 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Litvintchouk:1983:AARb

[LM83b] Steven D. Litvintchouk and A. S. Matsumoto. An algebraic approach to reusable Ada components. ACM SIG-ADA Ada Letters, 3(2):89– 92, July/August 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lindquist:1994:HDY

[LM94] Timothy E. Lindquist and Robert G. Munck. How do you pronounce OO-ERA-RDBMS-OMS? ACM SIG-ADA Ada Letters, 14(Special Issue):93–98, Fall 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Leeson:1994:IAV

[LMA94] David Leeson, Glenn MacEwen, and David Andrews. Interfacing Ada with verification languages. ACM SIGADA Ada Letters, 14(Special Issue):74–81, Fall 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lander:1990:DPI

[LMP90] Leslie C. Lander, Sandeep Mitra, and Thomas F. Piatkowski. Deterministic priority inversion in Ada selective waits. ACM SIG-ADA Ada Letters, 10(7):55– 62, September/October 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Locke:1993:RPT

[LMV93] C. Douglass Locke, Thomas J. Mesler, and David R. Vogel. Replacing passive tasks with Ada 9X protected records. ACM SIG-ADA Ada Letters, 13(2):91– 96, March/April 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lee:1991:ORT

[LN91] Pen-Nan Lee and William Nehman. An overview of real-time issues and Ada. ACM SIGADA Ada Letters, 11(9):83–95, November/December 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Luckham:1987:EAS

[LNR87] David C. Luckham, Randall Neff, and David S. Rosenblum. An environment for Ada software development based on formal specification. ACM SIGADA Ada Letters, 7(3):94–106, May/June 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Locke:1991:SIA

[Loc91] C. Douglass Locke. Scheduling issues in Ada 9X. ACM SIGADA Ada Letters, 11(6): 69–74, September/October 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Loftus:1993:AY

[Lof93] C. Loftus, editor. Ada yearbook 1993. IOS Press, Amsterdam, The Netherlands, 1993. xvi + 431 pp.

Logozzo:2013:PSV

[Log13a] Francesco Logozzo. Practical specification and verification with code contracts. ACM SIGADA Ada Letters, 33(3): 7–8, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Logozzo:2013:TIC

[Log13b] Francesco Logozzo. Technology for inferring contracts from code. ACM SIG-ADA Ada Letters, 33(3):13– 14, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lomuto:1983:SRA

[Lom83] Nico Lomuto. Self-reproducing Ada tasks. ACM SIG-ADA Ada Letters, 2(5):62– 75, March/April 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lopes:1999:ASO

[Lop99] Arthur V. Lopes. Ada + SQL — an overview. ACM SIG-ADA Ada Letters, 19(3):157– 162, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lowe:1999:EAA

[Low99a] Tony Lowe. Extending Ada to assist multiprocessor embedded development. ACM SIGADA Ada Letters, 19 (3):125–132, September 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lowe:1999:PPW

[Low99b] Tony Lowe. Pinching pennies while losing dollars. ACM SIGADA Ada Letters, 19 (3):183–193, September 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Luckham:1980:PMD

[LP80] David C. Luckham and Wolfgang Polak. A practical method of documenting and verifying Ada programs with packages. In ACM [ACM80], pages 113– 122. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

LeDoux:1985:STA

[LP85] Carol H. LeDoux and D. Stott Parker, Jr. Saving traces for Ada debugging. ACM SIGADA Ada Letters, 5(2): 97–108, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14–16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Ludwig:2006:DDE

[LP06] Luke Ludwig and Paul Pukite. DEGAS: discrete event Gnu advanced scheduler. ACM SIGADA Ada Letters, 26(3):35–42, December 2006. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

laPuente:2018:SSM

[lPB18] Juan Ade la Puente and Alan Burns. Session summary: Multiprocessor locking. ACM SIGADA Ada Letters, 38(1): 61, June 2018. CODEN AALEE5. ISSN 0736-721X.

Liang:2009:APG

[LRS09] Sheldon X. Liang, Lyle Reibling, and Samuel Sambasivam. ‘Automatic Prototype Generating’ restated with re-ADA: perspective-bridged architecture for documentdriven systems transitioning. ACM SIGADA Ada Letters, 29(3):45–60, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lupton:1998:SII

[LS98] William Lupton and Vojislav Stojkovic. Solving incomplete and incorrect information problems using conditional planning, execution monitoring, and situated planning agents. ACM SIG-ADA Ada Letters, 18(5):87– 96, September/October 1998.

CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Laski:1998:DAA

[LSH98] Janusz Laski, William Stanley, and Jim Hurst. Dependency analysis of Ada programs. ACM SIGADA Ada Letters, 18(6):263–275, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Laski:2001:BAP

[LSP01] Janusz Laski, William Stanley, and Pawel Podgorski. Beyond ASIS: program data bases and tool-oriented queries. ACM SIGADA Ada Letters, 21(4):81–90, December 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Locke:1988:PIC

[LSR^{+}88] Douglass Locke, Lui Sha, Ragunathan Rajkumar, John Lehoczky, and Greg Burns. Priority inversion and its control: An experimental investigation. ACM SIGADA Ada Letters, 8(7):39–42, Fall 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Li:2012:ART

[LSRM12] Shuai Li, Frank Singhoff, Stéphane Rubini, and Bourdellès Michel. Applicability of real-time schedulability analysis on a software radio protocol. ACM SIG-ADA Ada Letters, 32(3):81– 94, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Lortz:1999:RDR

[LT99] Henry A. Lortz and Timothy A. Tibbetts. The role of distributed, real-time Ada & C++ on the Airborne Surveillance Testbed (AST) program. ACM SIG-ADA Ada Letters, 19(3):181– 182, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Locke:1987:PAR

[LV87] C. Douglass Locke and David R. Vogel. Problems in Ada runtime task scheduling. ACM SIGADA Ada Letters, 7(6):51–53, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lucchetti:2023:ACF

[LV23] Federico Lucchetti and Marcus Voelp. Achieving crash fault tolerance in autonomous vehicle autopilot software stacks through safety-critical module rejuvenation. ACM SIG-ADA Ada Letters, 43(1): 83–87, June 2023. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3631483.3631497.

Locke:1990:PRT

[LVM90] C. D. Locke, D. R. Vogel, and T. J. Mesler. Predictable real-time avionics design using Ada tasks and rendezvous: a case study. ACM SIGADA Ada Letters, 10(9): 118–125, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Larraz:2022:BMC

[LVTL22] Daniel Larraz, Arjun Viswanathan, Cesare Tinelli, and Mickaël Laurent. Beyond model checking of idealized Lustre in Kind 2. ACM SIG-ADA Ada Letters, 42(2):40– 44, December 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3591335.3591338.

Lundqvist:1997:RL

[LW97] Kristina Lundqvist and Göran Wall. A rendezvous with Linda. ACM SIGADA Ada Letters, 17(3):87–96, May/June 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Liang:2001:OUO

[LW01] Xianzhong Liang and Zhenyu Wang. Omega: a uniform object model easy to gain Ada’s ends. ACM SIGADA Ada Letters, 21(2): 65–80, June 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Liang:2002:EBI

[LW02] Xianzhong Liang and Zhenyu Wang. Event-based implicit invocation decentralized in Ada. ACM SIG-ADA Ada Letters, 22(1): 11–16, March 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lau:2007:VCB

[LW07] Kung-Kiu Lau and Zheng Wang. Verified componentbased software in SPARK: experimental results for a missile guidance system. ACM SIGADA Ada Letters, 27(3):51–58, December 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lin:2013:ARS

[LWB13] S. Lin, A. J. Wellings, and A. Burns. Ada 2012: resource sharing and multiprocessors. ACM SIGADA Ada Letters, 33(1):32–44, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Latour:1991:DPA

[LWF91] Larry Latour, Tom Wheeler, and Bill Frakes. Descriptive and predictive aspects of the 3Cs model, SETA1 working group summary. ACM SIGADA Ada Letters, 11(3): 9–17, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Li:1998:TAS

[LXY98] Bangqing Li, Baowen Xu, and Huiming Yu. Transforming Ada serving tasks into protected objects. ACM SIG-ADA Ada Letters, 18(6):240– 245, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Li:2010:EAS

[LYB^{+}10] You Li, Lu Yang, Lei Bu, Linzhang Wang, Jianhua Zhao, and Xuandong Li. Extending Ada to support multi-core based monitoring and fault tolerance. ACM SIGADA Ada Letters, 30(3): 53–62, December 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Liang:2003:APG

[LZL03] Sheldon X. Liang, Lynn Zhang, and Luqi. Automatic prototype generating via optimized object model. ACM SIGADA Ada Letters, 23(2):22–31, June 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

MacLaren:1980:ETA

[Mac80] Lee MacLaren. Evolving toward Ada in real-time systems. In ACM [ACM80], pages 146–155. CODEN SIN-ODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

MacanAirchinnign:1984:APU

[Mac84] Mícheál Mac an Airchinnign. Ada packages and the user’s conceptual model. ACM SIG-ADA Ada Letters, 3(4):70– 77, January/February 1984. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

MacanAirchinnigh:1986:RIA

[Mac86] M. Mac an Airchinnigh. The real issues in Ada education/training. ACM SIG-ADA Ada Letters, 6(5):86– 93, September/October 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Macpherson:1987:WUW

[Mac87] George W. Macpherson. We’re using the wrong name. ACM SIGADA Ada Letters, 7(1):94–96, January/ February 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Macpherson:1996:RAP

[Mac96] George W. Macpherson. A reusable Ada package for scientific dimensional integrity. ACM SIGADA Ada Letters, 16(3):56–69, May/June 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mahani:2011:MAR

[Mah11]  Negin Mahani.  Making

alive register transfer level and transaction level modeling in Ada. ACM SIG-ADA Ada Letters, 31(2):15– 22, August 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mahani:2012:MAR

[Mah12a]  Negin Mahani.  Making

alive register transfer level and transaction level modeling in Ada. ACM SIG-ADA Ada Letters, 32(2):9– 16, August 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mahani:2012:TRR

[Mah12b] Negin Mahani. TLM request response channel in SystemAda. ACM SIG-ADA Ada Letters, 32(1): 13–18, April 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mahani:2013:IST

[Mah13] Negin Mahani. Investigating SystemAda: TLM FIFO detailed characteristics proof, TLM2.0 interfaces implementation, simulation time comparison to SystemC. ACM SIGADA Ada Letters, 33(1): 157–168, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Maloney:1988:UVV

[Mal88] James J. Maloney. Using the VAX/VMS lock manager with Ada tasks. ACM SIG-ADA Ada Letters, 8(2):84– 95, March/April 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Martin:1986:NAA

[Mar86] Donald G. Martin. Non-Ada to Ada conversion. ACM SIGADA Ada Letters, 6(1):72–81, January/ February 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mardis:1999:ESR

[Mar99] Mike Mardis. Endian-safe record representation clauses for Ada programs. ACM SIG-ADA Ada Letters, 19(4):13– 18, December 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mark:2005:DSB

[Mar05] Matt Mark. Data sharing between Ada and C/ C++. ACM SIGADA Ada Letters, 25(4):93–102, December 2005. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Martin:2019:CVE

[Mar19]  Bob Martin.  Common

Vulnerabilities Enumeration (CVE), Common Weakness Enumeration (CWE), and Common Quality Enumeration (CQE): Attempting to systematically catalog the safety and security challenges for modern, networked, software-intensive systems. ACM SIGADA Ada Letters, 38(2):9–42, December 2019. ISSN 0736-721X.

URL https://dl.acm.org/
doi/abs/10.1145/3375408.
3375410.

Martignano:2020:C

[Mar20] Maurizio Martignano. A:

the compiler. ACM SIG-ADA Ada Letters, 39(2): 25–28, April 2020. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3394514.3394516.

Martignano:2021:SAA

[Mar21] Maurizio Martignano. Static analysis for Ada, C/C++ and Python: Different languages, different needs. ACM SIGADA Ada Letters, 41 (2):77–80, December 2021. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3530801.3530807.

Mathis:1987:EFP

[Mat87a] Robert F. Mathis. Elementary functions package for Ada. In ACM [ACM87a], pages 95–100. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Matthews:1987:OPA

[Mat87b] Edmund R. Matthews. Observations on the portability of Ada I/O. ACM SIGADA Ada Letters, 7(5):100–103, September/October 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mattini:1991:HTE

[Mat91] M. Mattini. HP/Telegen2 encapsulation: an integration project of the Telesoft Ada environment with HP CASE and OSF/Motif. ACM SIG-ADA Ada Letters, 11(2):98– 106, March/April 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mathis:1996:CAQ

[Mat96] Robert Mathis. Commonly asked questions about Ada: the standardized development language. ACM SIG-ADA Ada Letters, 16(6): 51–54, November/December 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Maurer:2007:UMI

[Mau07]  Ward D. Maurer.  Us-

ing mathematics to improve Ada compiled code, part 2: the proof. ACM SIG-ADA Ada Letters, 27(3):11– 26, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mazzanti:1989:AE

[Maz89a] Franco Mazzanti. The AIDA experiment. ACM SIG-ADA Ada Letters, 9(5):109– 114, July/August 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mazzanti:1989:RUA

[Maz89b] Franco Mazzanti. Reducing unpredictability in Ada executions. ACM SIGADA Ada Letters, 9(6):90–96, September/October 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Matthews:1991:VAI

[MB91] Ed Matthews and Greg Burns. VADS APSE: An integrated Ada programming support environment. ACM SIGADA Ada Letters, 11(3): 61–72, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Martin:2008:CWE

[MB08] Robert A. Martin and Sean Barnum. Common weakness enumeration (CWE) status update. ACM SIG-ADA Ada Letters, 28(1): 88–91, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mitchell:2001:ME

[MBW01] S. E. Mitchell, A. Burns, and A. J. Wellings. MOPping up exceptions. ACM SIG-ADA Ada Letters, 21(3):80– 92, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Matthews:1990:LE

[MC90] John Matthews and Jeffrey R. Carter. Letters to the editor. ACM SIG-ADA Ada Letters, 10(5):9– 14, May/June 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mathisen:2005:OSN

[MC05] David G. Mathisen and Robert W. Carey. Orchestrating shots for the National Ignition Facility. ACM SIG-ADA Ada Letters, 25(4):69– 78, December 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Miranda:2009:GIC

[MC09a] Javier Miranda and Arnaud Charlet. Gem #61: interfacing with C++ constructors. ACM SIGADA Ada Letters, 29(2):61–62, August 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Miranda:2009:GCC

[MC09b] Javier Miranda and Arnaud Charlet. Gem #62: C++ constructors and Ada 2005. ACM SIGADA Ada Letters, 29(2):63–64, August 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mosteo:2022:AU

[MC22] Alejandro R. Mosteo and Fabien Chouteau. Alire 2022 update. ACM SIG-ADA Ada Letters, 42(1): 46–49, June 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3577949.3577954.

McCormick:1987:SDA

[McC87a] Frank McCormick. Scheduling difficulties of Ada in the hard real-time environment. ACM SIGADA Ada Letters, 7(6):49–50, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McCoy:1987:IAR

[McC87b] L. Scott McCoy. Interfacing Ada and relational databases. ACM SIGADA Ada Letters, 7(3):50–59, May/June 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McCoy:1990:BAa

[McC90a] L. Scott McCoy. Bindings and Ada. ACM SIGADA Ada Letters, 10(8):156–160, November/December 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McCoy:1990:BAb

[McC90b]  L. Scott McCoy.  Bind-

ings and Ada. ACM SIG-ADA Ada Letters, 10(9): 156–160, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McCormick:1999:AMR

[McC99]  John McCormick.  Ada,

model railroading, and realtime software engineering education (keynote address). ACM SIGADA Ada Letters, 19(3):111–112, September 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McCormick:2000:SEE

[McC00] John W. McCormick. Software engineering education: On the right track with Ada. ACM SIGADA Ada Letters, 20(3):41–49, September 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/sept2000/right_
track.pdf.

McCormick:2006:SAA

[McC06a] John W. McCormick. 2005 SIGAda awards awarded at SIGAda 2005 in Atlanta Georgia on November 16. ACM SIGADA Ada Letters, 26(1):12–15, April 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McCormick:2006:SRS

[McC06b] John W. McCormick. Special report by SIGAda Chair. ACM SIGADA Ada Letters, 26(1):7–11, April 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McCormick:2007:MRT

[McC07] John W. McCormick. MA1: real-time and parallel processing in Ada. ACM SIG-ADA Ada Letters, 27(3): 7, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McCormick:2009:ART

[McC09] John W. McCormick. Ada for real-time and parallel processing. ACM SIG-ADA Ada Letters, 29(3):13– 14, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McCormick:2010:APE

[McC10] John W. McCormick. Ada for parallel, embedded, and real-time applications. ACM SIGADA Ada Letters, 30(3): 5–6, December 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McDonald:1988:AAT

[McD88a] C. McDonald. The Ada ASEET team. ACM SIG-ADA Ada Letters, 8(3):115– 122, May/June 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McDonald:1988:ASE

[McD88b]  C. W. McDonald.  The

Ada Software Engineering Education and Training (ASEET) team. ACM SIG-ADA Ada Letters, 8(3):115– 122, May/June 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McDonald:1989:AAT

[McD89] C. McDonald. The Ada ASEET team. ACM SIG-ADA Ada Letters, 8(3):115– 122, May/June 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McEvilley:2003:EIA

[McE03]  Michael McEvilley.  The

essence of information assurance and its implications for the Ada community. ACM SIGADA Ada Letters, 23(1): 35–39, March 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Michell:1997:UAA

[MCS97] Stephen Michell, Dan Craigen, and Mark Saaltink. Using analytical approaches for high integrity Ada95 systems. ACM SIGADA Ada Letters, 17(5):65–70, September/October 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Maymir-Ducharme:1990:DPP

[MD90] Fred A. Maymir-Ducharme.

Dynamic priorities, priority scheduling and priority inheritance. ACM SIG-ADA Ada Letters, 10(9): 39–45, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moy:2022:PCG

[MD22] Yannick Moy and Claire Dross. Proving the correctness of GNAT light runtime library. ACM SIG-ADA Ada Letters, 42(1): 65–67, June 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3577949.3577959.

Michell:2016:SST

[MdlP16] Stepehen Michell and Juan Antonio de la Puente. Session summary: Time vulnerabilities. ACM SIGADA Ada Letters, 36(1):103–106, June 2016. CODEN AALEE5. ISSN 0736-721X.

Maymir-Ducharme:1994:RHS

[MDPK94] Fred Maymir-Ducharme, Teri Payton, and Judy Kerner. “reuse” and “hybrid systems” working groups summary — SETA2 working groups 2 and 6. ACM SIGADA Ada Letters, 14(Special Issue):109–112, Fall 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mearns:1987:DRT

[Mea87] Ian Mearns. Developing Real-Time Ada systems. ACM SIGADA Ada Letters, 7(6): 124–126, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Medley:1991:TQM

[Med91] J. S. Medley. Total quality management manifested through Ada. In ACM [ACM91b], pages 24–39. ISBN 0-89791-393-0. LCCN ????

Mendal:1987:SRM

[Men87] Geoffrey O. Mendal. Storage reclamation models for Ada programs. In ACM [ACM87a], pages 180–189. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Mendal:1988:TRA

[Men88] Geoff Mendal. Three reasons to avoid the use clause. ACM SIGADA Ada Letters, 8(1):52–57, January/ February 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mentis:2009:RAD

[Men09] Alexander S. Mentis. A robotics API dialect for type-safe robots: translating Myro to Ada. ACM SIG-ADA Ada Letters, 29(3):91– 102, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mundie:1991:OOR

[MF91] David A. Mundie and David A. Fisher. Optimized overload resolution and type matching for Ada. ACM SIG-ADA Ada Letters, 11(3): 83–90, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Marco:2004:FDI

[MF04] Jordi Marco and Xavier Franch. A framework for designing and implementing the Ada Standard Container Library. ACM SIG-ADA Ada Letters, 24(1): 49–61, March 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moore:1985:PAA

[MFD85] Melody Moore, John Foreman, and Paulan Daily. Planning an AdaTEC/AdaJUG meeting. ACM SIGADA Ada Letters, 5(1):32–41, July/August 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Melde:1987:LSS

[MG87] John E. Melde and Phillip G.

Gage. Large system simulation using Ada. In ACM [ACM87a], pages 126–132. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Malaquias:2023:TMD

[MGB^{+}23] Felipe Lisboa Malaquias, Georgios Giantamidis, Stylianos Basagiannis, Simone Fulvio Rollini, and Isaac Amundson. Towards a methodology to design provably secure cyber-physical systems. ACM SIGADA Ada Letters, 43(1):94–99, June 2023. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3631483.3631499.

McGregor:2016:ADS

[MGF16] John D. McGregor, David P.

Gluch, and Peter H. Feiler. Analysis and design of safetycritical, cyber-physical systems. ACM SIGADA Ada Letters, 36(2):31–38, December 2016. CODEN AALEE5. ISSN 0736-721X.

Munck:1997:AJW

[MH97] Robert G. Munck and Richard F. Hilliard II. Ada and Java on the WWW. ACM SIGADA Ada Letters, 17(3): 3–16, May/June 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Murtagh:1998:CAP

[MH98] Jeanne L. Murtagh and John A. Hamilton, Jr. A comparison of Ada and Pascal in an introductory computer science course. ACM SIGADA Ada Letters, 18(6):75–80, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Murtagh:2009:HAO

[MH09] Jeanne Murtagh and Drew Hamilton. How Ada object orientation works. ACM SIGADA Ada Letters, 29(3): 5–6, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moore:2020:WSA

[MH20] Brad Moore and John A. Hamilton. Winners of 2018 SIGAda Awards. ACM SIG-ADA Ada Letters, 39(1):100, January 2020. ISSN 0736-

                                        721X. URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3379106.3379118.

Michell:2001:PPC

[Mic01] Stephen Michell. Position paper: completing the Ravenscar profile. ACM SIG-ADA Ada Letters, 21(1):

75–78, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Michell:2002:PIE

[Mic02] Stephen Michell. Practical implementations of embedded software using the Ravenscar Profile. ACM SIG-ADA Ada Letters, 22(4):28– 36, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Michell:2007:IAO

[Mic07] Stephen Michell. Interfacing Ada to operating systems. ACM SIGADA Ada Letters, 27(2):90–95, August 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Michell:2013:PLV

[Mic13] Stephen Michell. Programming language vulnerabilities: proposals to include concurrency paradigms. ACM SIGADA Ada Letters, 33(1): 101–115, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Michell:2016:TIP

[Mic16] Stephen Michell. Time issues in programs vulnerabilities for programming languages or systems. ACM SIGADA Ada Letters, 36(1):

77–82, June 2016. CODEN AALEE5. ISSN 0736-721X.

Middlemas:1987:AAE

[Mid87] Michael R. Middlemas. Ada applications on embedded targets. In ACM [ACM87a], pages 170–179. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Masters:1983:SDP

[MK83] Michael W. Masters and Michael J. Kuchinski. Software design prototyping using Ada. ACM SIGADA Ada Letters, 2(4):68–75, January/February 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Maarek:1987:UCC

[MK87] Yoelle S. Maarek and Gail E. Kaiser. Using conceptual clustering for classifying reusable Ada code. In ACM [ACM87a], pages 216–225. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Micallef:1991:EMS

[MK91] Josephine Micallef and Gail E. Kaiser. Extending the MER-CURY system to support teams of Ada programmers. ACM SIGADA Ada Letters, 11(3):49–60, Spring 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Matsakis:2014:RL

[MK14] Nicholas D. Matsakis and Felix S. Klock II. The Rust language. ACM SIG-ADA Ada Letters, 34(3):103– 104, December 2014. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moody:1999:STT

[MKK99] Scott Arthur Moody, Samuel Kwok, and Dale Karr. SimpleGraphics: Tcl/Tk visualization of real-time multithreaded and distributed applications. ACM SIG-ADA Ada Letters, 19(2): 60–66, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Micallef:1991:ALC

[MKP91a] Josephine Micallef, Gail E.

Kaiser, and Dewayne E. Perry. Ada libraries, configuration management, and version control. ACM SIG-ADA Ada Letters, 11(3): 29–??, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Micallef:1991:SWG

[MKP91b] Josephine Micallef, Gail E. Kaiser, and Dewayne E. Perry. SETA1 working group on Ada libraries, configuration management, and version control. ACM SIG-ADA Ada Letters, 11(3): 29–31, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Matthews:1986:AEE

[ML86] E. R. Matthews and W. Lively. The ATMAda environment: an enhanced APSE. ACM SIGADA Ada Letters, 6(3): 61–64, May/June 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Marr:1991:ADR

[ML91] J. S. Marr and P. K. Lawlis. Automatic determination of recommended test combinations for Ada compilers. In ACM [ACM91b], pages 77– 89. ISBN 0-89791-393-0. LCCN ????

Mignon:1995:AUB

[ML95a] Marie-France Mignon and Florence Lescroart. Ada used for on-board flight control. ACM SIGADA Ada Letters, 15(4):17–18, July/August 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mignon:1995:AUD

[ML95b] Marie-France Mignon and Florence Lescroart. Ada used to develop a simulator run by robots. ACM SIG-ADA Ada Letters, 15(4):15– 16, July/August 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Michell:1999:ESD

[ML99] Stephen Michell and Kristina Lundqvist. Extendable [sic], dispatchable task communication mechanisms. ACM SIGADA Ada Letters, 19(2): 54–59, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Macos:1998:RDL

[MM98] Dragan Macos and Frank Mueller. The rendezvous is dead — long live the protected object. ACM SIGADA Ada Letters, 18(6):287–293, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Marriott:2017:UGP

[MM17] Ahlan Marriott and Urs Maurer. Using GtkAda in practice. ACM SIGADA Ada Letters, 37(2):59–67, December 2017. CODEN AALEE5. ISSN 0736-721X.

Marriott:2021:MAN

[MM21] A. Marriott and U. Maurer. More Ada in non-Ada systems. ACM SIG-ADA Ada Letters, 41(2):71– 76, December 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3530801.3530806.

Maia:2003:VVM

[MMB^{+}03] R. Maia, F. Moreira, R. Barbosa, D. Costa, Kjeld Hjortaes, Patricia Rodriguez, and Luis Miguel Pinho. Verifying, validating and monitoring the open Ravenscar real time kernel. ACM SIG-ADA Ada Letters, 23(4):74– 81, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mahani:2009:SLH

[MMN09] Negin Mahani, Parnian Mokri, and Zainalabedin Navabi. System level hardware design and simulation with SystemAda. ACM SIG-ADA Ada Letters, 29(1): 19–22, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Michell:2013:RTP

[MMP13a] Stephen Michell, Brad Moore, and Luis Miguel Pinho. Realtime programming on accelerator many-core processors.

ACM SIGADA Ada Letters, 33(3):23–36, December 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moore:2013:PAG

[MMP13b] Brad Moore, Stephen Michell, and Luis Miguel Pinho. Parallelism in Ada: general model and Ravenscar. ACM SIGADA Ada Letters, 33(2): 14–32, August 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Michell:2016:CUE

[MMPT16] Stephen Michell, Brad Moore, Luis Miguel Pinho, and Tucker Taft. Constraints on the use of executors in realtime systems. ACM SIG-ADA Ada Letters, 36(1):25– 28, June 2016. CODEN AALEE5. ISSN 0736-721X.

Mahani:2009:SAB

[MMSN09] Negin Mahani, Parnian Mokri, Mahshid Sedghi, and Zainalabedin Navabi. SystemAda: an Ada based system-level hardware description language. ACM SIGADA Ada Letters, 29(2): 15–19, August 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Muller:2016:DRI

[MNG16] Josef Müller, Prashanth Lakshmi Narasimhan, and Swaminathan Gopalswamy. Design Requirements Iterative Process (DRIP) tool demonstration concurrent engineering of design, requirements and knowledge. ACM SIG-ADA Ada Letters, 36(2):60– 68, December 2016. CODEN AALEE5. ISSN 0736-721X.

Mogilensky:1991:PMG

[Mog91] J. Mogilensky. Process maturity as a guide to phased Ada adoption. In ACM [ACM91b], pages 16–23. ISBN 0-89791-393-0. LCCN ????

Molich:1983:ACQ

[Mol83] Rolf Molich. Ada compiler quality assurance. ACM SIG-ADA Ada Letters, 3(2):72– 75, September/October 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moore:1985:RWA

[Moo85] John I. Moore, Jr. Report on the 1985 Washington Ada Symposium. ACM SIGADA Ada Letters, 5(3–6):16–18, November/December 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moore:1991:ABS

[Moo91] James W. Moore. The ANSI binding of SQL to Ada. ACM SIGADA Ada Letters, 11 (5):47–61, July/August 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moore:1993:IAI

[Moo93] J. W. Moore. The impact of Ada 9X incompatibilities on projects converting from Ada 83. ACM SIG-ADA Ada Letters, 13(4):29– 36, July/August 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moore:1994:SDS

[Moo94] James W. Moore. A structure for a defense software reuse marketplace. ACM SIGADA Ada Letters, 14(3): 80–90, May/June 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moore:1996:FIS

[Moo96] James W. Moore. Future of IEEE standard for Ada PDL to be considered. ACM SIG-ADA Ada Letters, 16(2):35– 38, March/April 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moody:1997:OOR

[Moo97] Scott Arthur Moody. Objectoriented real-time systems using a hybrid distributed model of Ada 95’s builtin DSA capability (Distributed Systems Annex-E) and CORBA. ACM SIG-ADA Ada Letters, 17(5):71– 76, September/October 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moore:1998:OAS

[Moo98] James W. Moore. Overview of Ada standardization. ACM SIGADA Ada Letters, 18(3): 18–19, May 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moore:2010:PGA

[Moo10]  Brad J. Moore.  Paral-

lelism generics for Ada 2005 and beyond. ACM SIG-ADA Ada Letters, 30(3):41– 52, December 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moore:2011:SSP

[Moo11] Bradley J. Moore. Stack safe parallel recursion with Paraffin. ACM SIGADA Ada Letters, 31(3):27–34, December 2011. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moore:2018:SSA

[Moo18] Brad Moore. Synchronous signals: an abstraction for interleaving sequential and parallel code. ACM SIG-ADA Ada Letters, 38(1):19– 28, June 2018. CODEN AALEE5. ISSN 0736-721X.

Moreton:1987:PAL

[Mor87]  Trevor Moreton.  Parti-

tioned Ada libraries as a basis for variant control. In ACM [ACM87a], pages 60– 64. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Morrone:1995:DWE

[Mor95a] George Morrone. Did we ever really want to be liberated from the von Neumann architecture?: or, assignment statement considered a nuisance. ACM SIG-ADA Ada Letters, 15(5):52– 53, September/October 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Morrone:1995:RBF

[Mor95b]  George Morrone.  Re-

cursion: beyond factorial. ACM SIGADA Ada Letters, 15(6):70–72, November/December 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Morrone:1996:DAa

[Mor96a] George Morrone. Dr. Ada 95.

ACM SIGADA Ada Letters, 16(2):70, March/April 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Morrone:1996:DAb

[Mor96b] George Morrone. Dr. Ada 95. ACM SIGADA Ada Letters, 16(3):70–71, May/June 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mosley:2006:WML

[Mos06] David Mosley. When to migrate legacy embedded applications. ACM SIG-ADA Ada Letters, 26(3):77– 80, December 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mosteo:2020:RBA

[Mos20] Alejandro R. Mosteo. RCLAda, or bringing Ada to the Robot Operating System. ACM SIGADA Ada Letters, 39(2): 35–40, April 2020. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3394514.3394518.

Mosteo:2022:UAA

[Mos22] Alejandro R. Mosteo. Use (and abuse) of Ada 2022 features in designing a JSONlike data structure. ACM SIGADA Ada Letters, 42 (1):54–57, June 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3577949.3577956.

Moy:2011:GLSa

[Moy11a] Yannick Moy. Gem #68: let’s SPARK! — part 1. ACM SIGADA Ada Letters, 31(1): 19–23, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moy:2011:GLSb

[Moy11b] Yannick Moy. Gem #69: let’s SPARK! — part 2. ACM SIGADA Ada Letters, 31(1): 24–27, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moy:2011:GTBa

[Moy11c] Yannick Moy. Gem #82:

type-based security 1: handling tainted data. ACM SIGADA Ada Letters, 31(2): 36–39, August 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moy:2011:GTBb

[Moy11d] Yannick Moy. Gem #83:

type-based security 2: validating the input. ACM SIGADA Ada Letters, 31(2): 40–43, August 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moy:2017:GPBa

[Moy17a] Yannick Moy. Gem #146:

Su(per)btypes in Ada 2012 — Part 1. ACM SIGADA Ada Letters, 37(2):27–29, December 2017. CODEN AALEE5. ISSN 0736-721X.

Moy:2017:GPBb

[Moy17b] Yannick Moy. Gem #147: Su(per)btypes in Ada 2012 — Part 2. ACM SIGADA Ada Letters, 37(2):30–31, December 2017. CODEN AALEE5. ISSN 0736-721X.

Moy:2017:GPBc

[Moy17c] Yannick Moy. Gem #148: Su(per)btypes in Ada 2012 — Part 3. ACM SIGADA Ada Letters, 37(2):32–33, December 2017. CODEN AALEE5. ISSN 0736-721X.

Moy:2017:GAT

[Moy17d] Yannick Moy. Gem #149: Asserting the truth, but (possibly) not the whole truth. ACM SIGADA Ada Letters, 37(2):34–36, December 2017. CODEN AALEE5. ISSN 0736-721X.

Moy:2017:GSM

[Moy17e] Yannick Moy. Gem #151: Specifying mathematical properties of programs. ACM SIGADA Ada Letters, 37(2): 40–42, December 2017. CO-DEN AALEE5. ISSN 0736-721X.

Meiling:1984:CSC

[MP84] Erik Meiling and Steen U. Palm. A comparative study of CHILL and Ada on the basis of denotational descriptions. ACM SIGADA Ada Letters, 3(4):78–91, January/February 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mauger:1985:EDD

[MP85] Claude Mauger and Kevin Pammett. An event-driven debugger for Ada. ACM SIGADA Ada Letters, 5(2): 124–135, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14–16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Mysior:1989:EBC

[MP89] Jerzy Mysior and Andrzej Paprocki. An eight-bit character set in Ada programs. ACM SIGADA Ada Letters, 9(7):85–90, November/December 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Moore:1991:LBT

[MP91] M. Moore and A. Pruitt. A logic base tool set for realtime Ada software development. In ACM [ACM91b], pages 102–118. ISBN 0-89791-393-0. LCCN ????

Mills:1998:HSC

[MP98] Mike Mills and Greg Peterson. Hardware/software co-design: VHDL and Ada 95 code migration and integrated analysis. ACM SIG-ADA Ada Letters, 18(6): 18–27, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mezzetti:2010:TIR

[MPV10] Enrico Mezzetti, Marco Panunzio, and Tullio Vardanega. Temporal isolation with the Ravenscar profile and Ada 2005. ACM SIGADA Ada Letters, 30(1): 45–55, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McDermid:1983:LCS

[MR83] John McDermid and Knut Ripken. Life cycle support in the Ada environment. ACM SIGADA Ada Letters, 3(1): 57–62, July/August 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Maxted:1987:AGT

[MR87a] Amelia Maxted and John C. Rowe. An Ada graphical tool. In ACM [ACM87a], pages 87–94. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

McNickle:1987:EUA

[MR87b] Mark McNickle and Ann Reedy. Experiences in using Ada with DBMS applications. ACM SIG-ADA Ada Letters, 7(3):40– 49, May/June 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Michell:2010:CIR

[MR10] Stephen Michell and Jorge Real. Conclusions of the 14th International Real-Time Ada Workshop. ACM SIG-ADA Ada Letters, 30(1): 162–164, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Markow:2006:CST

[MRB06] Tanya Markow, Eugene Ressler, and Jean Blair. Catch that speeding turtle: latching onto fun graphics in CS1. ACM SIG-ADA Ada Letters, 26(3):29– 34, December 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Musser:1987:LGA

[MS87] David R. Musser and Alexander A. Stepanov. A library of generic algorithms in Ada. In ACM [ACM87a], pages 216–225. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Miranda:2004:GRA

[MS04] Javier Miranda and Edmond Schonberg. GNAT: on the road to Ada 2005. ACM SIGADA Ada Letters, 24(4): 51–60, December 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

McCormick:2011:BER

[MS11] John W. McCormick and Frank Singhoff. Building embedded real-time applications. ACM SIGADA Ada Letters, 31(3):15–16, December 2011. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Miranda:2005:IAS

[MSK05] Javier Miranda, Edmond Schonberg, and Hristian Kirtchev. The implementation of Ada 2005 synchronized interfaces in the GNAT compiler. ACM SIG-ADA Ada Letters, 25(4):41– 48, December 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Miranda:2003:DCP

[MSM^{+}03] Javier Miranda, Edmond Schonberg, Miguel Masmano, Jorge Real, and Alfons Crespo. Dynamic ceiling priorities in GNAT implementation report. ACM SIG-ADA Ada Letters, 23(4):24– 27, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Marmor-Squires:1985:MER

[MSW85] Ann Marmor-Squires and Jack Wileden. Methodology and environment relationships. ACM SIG-ADA Ada Letters, 4(5):79– 83, March/April 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Future Ada Environment Workshop.

Michell:1998:LSH

[MSW98a] Stephen Michell, Mark Saaltink, and Brian Wichmann. Looking into safety with the High-Integrity Rapporteur Group (HRG). ACM SIGADA Ada Letters, 18(6):7–11, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Michell:1998:LSS

[MSW98b] Stephen Michell, Mark Saaltink, and Brian Wichmann. Looking into safety with the safety and security Rapporteur group. ACM SIGADA Ada Letters, 18(6):7–11, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Michell:2001:TOO

[MT01] Stephen Michell and Joyce L. Tokar. Tasking and object orientation. ACM SIG-ADA Ada Letters, 21(1): 9–10, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mudge:1987:UDD

[Mud87] Trevor Mudge. Units of distribution for distributed Ada. ACM SIGADA Ada Letters, 7(6):64–66, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mundie:1991:IMS

[Mun91a] David Mundie. Integration mechanism subgroup. ACM SIGADA Ada Letters, 11(3): 33–??, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mundie:1991:RIM

[Mun91b] David Mundie. Report of the integration mechanisms working group. ACM SIG-ADA Ada Letters, 11(3): 33–35, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Munck:1996:AJM

[Mun96] Bob Munck. Ada95 and Java: a major opportunity for the Ada community. ACM SIG-ADA Ada Letters, 16(1):18– 20, January/February 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). New mailing list

web_ada@acm.org created for

discussion of Ada-Java issues. Send subscription requests to

mailserv@acm.org with no

subject line and a body consisting of the lines subscribe

web ada and help.

Murray:1987:LOA

[Mur87] L. E. Murray. A life-cycle oriented Ada design language. In ACM [ACM87a], pages 81–86. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Murray:1990:ATT

[Mur90] A. G. Murray. Ada tasking as a tool for ecological modelling. ACM SIG-ADA Ada Letters, 10(7):85– 90, September/October 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Martin:1999:BTT

[MVG99] Juan Carlos Díaz Martín, Isidro Irala Veloso, and José Manuel Rodríguez García. Building Tcl-Tk GUIs for HRT–HOOD systems. ACM SIGADA Ada Letters, 19 (3):113–123, September 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Michell:2010:RPN

[MWM10] Stephen Michell, Luke Wong, and Brad Moore. Realtime paradigms needed post Ada 2005. ACM SIG-ADA Ada Letters, 30(1): 62–67, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Murugesan:2013:CVM

[MWRH13] Anitha Murugesan, Michael W. Whalen, Sanjai Rayadurgam, and Mats P. E. Heimdahl. Compositional verification of a medical device system. ACM SIGADA Ada Letters, 33(3):51–64, December 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Monroe:1998:SEU

[MY98] Joseph Monroe and H. Yu. A software engineering using Ada 95 course. ACM SIG-ADA Ada Letters, 18(1):86– 91, January/February 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Myers:1985:SEA

[Mye85] Gil Myers. Software Engineering Automation for Tactical Embedded Systems (SEATECS). ACM SIG-ADA Ada Letters, 4(5):45– 48, March/April 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Naeser:2005:PIM

[Nae05]  Gustaf Naeser.  Priority

inversion in multi processor systems due to protected actions. ACM SIG-ADA Ada Letters, 25(1): 43–47, March 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Naeser:2005:STM

[NAF05] Gustaf Naeser, Lars Asplund, and Johan Furunäs. SafetyChip: a time monitoring and policing device. ACM SIGADA Ada Letters, 25(4): 63–68, December 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Naks:2020:USE

[NAT20] T. Naks, M. A. Aiello, and S. T. Taft. Using SPARK to ensure system to software integrity: a case study. ACM SIGADA Ada Letters, 40(1):74–78, October 2020. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3431235.3431241.

Nogueira:2020:NFR

[NBZ^{+}20] Luis Nogueira, António Barros, Cristina Zubia, David Faura, Daniel Gracia Pérez, and Luis Miguel Pinho. Nonfunctional requirements in the ELASTIC architecture. ACM SIGADA Ada Letters, 40(1):85–90, October 2020. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3431235.3431243.

Needham:1998:COO

[NDM98] Donald M. Needham, Steven A. Demurjian, Sr., and Margaret M. McMahon. Concurrency in object-oriented propagation modeling using Ada95. ACM SIGADA Ada Letters, 18(5):97–103, September/October 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Needham:1997:ABP

[NDP97] D. M. Needham, S. A. Demurjian, and T. J. Peters. An Ada95 basis for propagation modeling. In ACM [ACM97], pages 263–272. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Needham:1999:TDO

[NDP99] Donald M. Needham, Steven A. Demurjian, Sr., and Thomas J. Peters. Towards a distributed object-oriented propagation model using Ada95. ACM SIGADA Ada Letters, 19(3): 203–210, September 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Needham:2000:IAM

[NDP00] D. Needham, S. Demurjian, and T. Peters. An IDL to Ada95 mapping to support propagation modeling. ACM SIGADA Ada Letters, 20(1): 58–66, March 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Newport:1995:PMR

[New95] John R. Newport. A performance model for realtime systems. ACM SIG-ADA Ada Letters, 15(2):59– 73, March/April 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Newport:1999:RTP

[New99] John R. Newport. A realtime, path guidance cue. ACM SIGADA Ada Letters, 19(1):59–63, March 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Nielsen:1986:TCC

[Nie86]  Kjell W. Nielsen.  Task

coupling and cohesion in Ada. ACM SIGADA Ada Letters, 6(4):44–52, July/ August 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Nilsen:2012:RTJ

[Nil12a] Kelvin Nilsen. Real-time Java in modernization of the Aegis Weapon System. ACM SIG-ADA Ada Letters, 32(3):63– 70, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Nilsen:2012:TOU

[Nil12b]  Kelvin Nilsen.  Tutorial

overview: understanding dynamic memory management in safety critical Java. ACM SIGADA Ada Letters, 32(3): 15–22, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Nettleton:2007:TDC

[NIM07] Chris Nettleton, Wilson Ifill, and Colin Marsh. Towards a demonstrably-correct Ada compiler. ACM SIG-ADA Ada Letters, 27(3):89– 96, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Nakao:1993:ACD

[NKN93] Zensho Nakao, Masaya Kinjo, and Masahiro Nakama. Ada and C: differences as the language for system programming. ACM SIGADA Ada Letters, 13(5):22–31, September/October 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Naeser:2005:TSV

[NLA05] Gustaf Naeser, Kristina Lundqvist, and Lars Asplund. Temporal skeletons for verifying time. ACM SIG-ADA Ada Letters, 25(4):49– 56, December 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Nelson:1992:OOP

[NM92] Michael L. Nelson and Gilberto F. Mota. Objectoriented programming in Classic-Ada. ACM SIG-ADA Ada Letters, 12(2):102– 110, March/April 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Nelson:1992:COO

[NMT92] Michael L. Nelson, Gilberto F.

Mota, and Vassilios Theologitis. Concurrent objectoriented programming in Classic Ada. ACM SIG-ADA Ada Letters, 12(5):77– 83, September/October 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Notkin:1980:EPA

[Not80] David S. Notkin. An experience with parallelism in Ada. In ACM [ACM80], pages 9–15. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Neri:1997:DDA

[NPT97] D. Neri, L. Pautet, and S. Tardieu. Debugging distributed applications with replay capabilities. In ACM [ACM97], pages 189–196. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Narfelt:1985:ESP

[NS85] Kjell-Hakan Narfelt and Dick Schefstrom. Extending the scope of the program library. ACM SIGADA Ada Letters, 5(2):25–40, September/October 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Neville:2003:DGG

[NS03] Melvin Neville and Anaika Sibley. Developing a generic genetic algorithm. ACM SIGADA Ada Letters, 23(1): 45–52, March 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Nissen:1983:AEG

[NW83] J. C. D. Nissen and B. A.

Wichmann. Ada-Europe guidelines for Ada compiler specification and selection. ACM SIGADA Ada Letters, 3(1):37–50, July/ August 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Nissen:1984:AEG

[NW^{+}84] J. C. D. Nissen, P. A.

Wichmann, et al. Ada-Europe guidelines for Ada compiler specification and selection. ACM SIGADA Ada Letters, 3(5):50–62, March/April 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Nissen:1982:AEG

[NWW82] J. C. D. Nissen, P. Wallis, and B. A. Wichmann. Ada-Europe guidelines for the portability of Ada programs. ACM SIGADA Ada Letters, 1(3):44–61, March/April 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Nyberg:1987:URC

[Nyb87] Karl A. Nyberg. Using representation clauses as an operating system interface. ACM SIGADA Ada Letters, 7(4): 98–101, July/August 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Nyberg:2005:WDD

[Nyb05] Karl Nyberg. Windows disk drive recovery with Ada95: an application note. ACM SIGADA Ada Letters, 25(2): 42–44, June 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Nyberg:2007:MCM

[Nyb07] Karl Nyberg. Multi-core + multi-tasking = multiopportunity? ACM SIG-ADA Ada Letters, 27(3):79– 82, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Nyberg:2010:AGD

[Nyb10a] Karl Nyberg. Automatically generating DTD-specific XML parsers. ACM SIGADA Ada Letters, 30(2):13–18, August 2010. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Nyberg:2010:PHD

[Nyb10b]  Karl Nyberg.  Parsing

Hierarchical Data Format (HDF) files. ACM SIG-ADA Ada Letters, 30(2): 19–24, August 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Oh:1997:OAT

[OB97] D.-I. Oh and T. P. Baker.

Optimization of Ada’95 tasking constructs. In ACM [ACM97], pages 79–90. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Oberndorf:1985:SCR

[Obe85] T. Oberndorf. Second CAIS review meeting. ACM SIG-ADA Ada Letters, 4(6):35– 43, May/June 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Oberndorf:1994:PSI

[Obe94] Patricia A. Oberndorf, editor. Proceedings of the Second International Symposium on Environments and Tools for Ada (SETA2), volume 14 (Special Issue) of ACM SIGADA Ada Letters. ACM Press, New York, NY, USA, Fall 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Obry:2009:GIA

[Obr09] Pascal Obry. Gem #29: introduction to the Ada Web Server (AWS). ACM SIG-ADA Ada Letters, 29(1): 41–44, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Obry:2012:GSWa

[Obr12a] Pascal Obry. Gem #101: SOAP/WSDL server part. ACM SIGADA Ada Letters, 32(2):35–36, August 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Obry:2012:GSWb

[Obr12b] Pascal Obry. Gem #102: SOAP/WSDL client part. ACM SIGADA Ada Letters, 32(2):37–38, August 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2009:GEI

[Och09a]  Quentin Ochem.  Gem

#48: extending interfaces in Ada 2005. ACM SIG-ADA Ada Letters, 29(1): 78–79, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2009:GIA

[Och09b] Quentin Ochem. Gem #55: introduction to Ada /Java interfacing. ACM SIG-ADA Ada Letters, 29(2): 43–45, August 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2009:GCA

[Och09c] Quentin Ochem. Gem #56:

creating Ada to Java calls using GNAT-AJIS. ACM SIGADA Ada Letters, 29(2): 46–49, August 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2009:GASa

[Och09d] Quentin Ochem. Gem #57:

Ada /Java cross dispatching. ACM SIGADA Ada Letters, 29(2):50–52, August 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2009:GASb

[Och09e]  Quentin Ochem.  Gem

#58: Ada /Java exception handling. ACM SIG-ADA Ada Letters, 29(2): 53–55, August 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2009:MLP

[Och09f]  Quentin Ochem.  Multi-

language programming with Ada. ACM SIGADA Ada Letters, 29(3):19–20, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2011:GAQ

[Och11] Quentin Ochem. Gem #86: Ada quiz 1 — basic types. ACM SIGADA Ada Letters, 31(2):52–55, August 2011. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2012:GGS

[Och12a]  Quentin Ochem.  Gem

#88 GPS: smart completion (part 1 of 2). ACM SIGADA Ada Letters, 32(1): 19–21, April 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2012:GSC

[Och12b] Quentin Ochem. Gem #91: smart completion (part 2 of 2). ACM SIGADA Ada Letters, 32(1):30–31, April 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2012:GDS

[Och12c]  Quentin Ochem.  Gem

#95: dynamic stack analysis in GNAT. ACM SIG-ADA Ada Letters, 32(1): 46–48, April 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Office:1987:ACV

[Off87] Ada Joint Program Office. Ada compiler validation procedures and guidelines, version 1.1. ACM SIG-ADA Ada Letters, 7(2):28– 57, March/April 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

OUSDA:1988:ABR

[Off88a] Office of the Under Secretary of Defense for Acquisition. Ada Board response to the Report of the Defense Science Board Task Force on Military Software. ACM SIG-ADA Ada Letters, 8(4):47– 68, July/August 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

OUSDA:1988:EFR

[Off88b] Office of the Under Secretary of Defense for Acquisition. Excerpts from Fall 1987 report of the Defense Science Board Task Force on military software. ACM SIG-ADA Ada Letters, 8(4):35– 46, July/August 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

OUSDA:1988:RDS

[Off88c] Office of the Under Secretary of Defense for Acquisition. Report of the Defense Science Board Task Force on military software: September 1987. ACM SIG-ADA Ada Letters, 8(4):35– 46, July/August 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

OLeary:2007:FAA

[O’L07] Jeff O’Leary. Federal Aviation Administration and Ada. ACM SIGADA Ada Letters, 27(3):69–70, December 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Oliver:1994:PIB

[Oli94] S. Ron Oliver. Of pyramids and igloos: a brief cultural perspective. ACM SIG-ADA Ada Letters, 14(4):36– 42, July/August 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Oberndorf:1985:PD

[OP85a] P. A. Oberndorf and M. H. Penedo. Project databases. ACM SIGADA Ada Letters, 4(5):65–78, March/ April 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Orberndorf:1985:PDW

[OP85b] Patricia A. Orberndorf and Maria H. Penedo. Project database working group. ACM SIGADA Ada Letters, 4(5):65–78, March/ April 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Future Ada Environment Workshop.

Orberndorf:1985:SCR

[Orb85] T. Orberndorf. The second CAIS review meeting. ACM SIGADA Ada Letters, 4(6): 35–43, May/June 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

OLeary:2012:FCP

[OS12] Jeffrey O’Leary and Alok Srivastava. FAA’s controller pilot automatic data communication (data comm) system software development. ACM SIGADA Ada Letters, 32(3): 71–72, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Olsen:1982:ATD

[OW82] Eric W. Olsen and Stephen B.

Whitehill. Ada technology development at Irvine Computer Sciences Corporation. ACM SIGADA Ada Letters, 1(3):77–85, March/April 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

OLeary:2008:AST

[OWSB08] Jeff O’Leary, Frederick Woodard, Alok Srivastava, and Denise S. Beidleman. Assessment of string tests strategy for an en route air traffic control system. ACM SIGADA Ada Letters, 28(1):24–30, April 2008.

CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pagan:1982:TAI

[Pag82]  Frank G. Pagan.  Tam-

ing Ada for introductory teaching purposes — an approximation. ACM SIG-ADA Ada Letters, 1(4):27– 31, May/June 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Panunzio:2012:GCAd

[Pan12a] Marco Panunzio. Gem #103: code archetypes for real-time programming — part 5. ACM SIGADA Ada Letters, 32(2): 39–42, August 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Panunzio:2012:G

[Pan12b] Marco Panunzio. Gem #89. ACM SIGADA Ada Letters, 32(1):22–26, April 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Panunzio:2012:GCAa

[Pan12c] Marco Panunzio. Gem #92: code archetypes for real-time programming — part 2. ACM SIGADA Ada Letters, 32(1): 32–36, April 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Panunzio:2012:GCAb

[Pan12d] Marco Panunzio. Gem #94:

code archetypes for real-time programming — part 3. ACM SIGADA Ada Letters, 32(1): 39–45, April 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Panunzio:2012:GCAc

[Pan12e] Marco Panunzio. Gem #96:

code archetypes for real-time programming — part 4. ACM SIGADA Ada Letters, 32(2): 17–23, August 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Papay:1989:FCA

[Pap89]  David Papay.  Forcing

the completion of abnormal tasks. ACM SIGADA Ada Letters, 9(6):104–107, September/October 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Paulk:1986:MD

[Pau86] Mark C. Paulk. Minutes of the DAWG. ACM SIG-ADA Ada Letters, 6(2):76, March/April 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Paulk:1987:RTP

[Pau87] Mark C. Paulk. Real-time performance of distributed Ada programs. ACM SIG-ADA Ada Letters, 7(6): 77–78, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Paulkovich:1993:AOR

[Pau93] Michael Paulkovich. Ada overhead reconsidered. ACM SIGADA Ada Letters, 13(3): 86–87, May/June 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pazy:1990:PPA

[Paz90] Offer Pazy. Problems with Pthreads and Ada. ACM SIGADA Ada Letters, 10(9): 133–140, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Petren:1998:RWW

[PB98] John Petren and John Beidler. ReUse/Web: Webbased Ada reuse. ACM SIGADA Ada Letters, 18(2): 81–88, March 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Parsian:1988:ATT

[PBB^{+}88] Mahmoud Parsian, Brayan Basdell, Yusuf Bhayat, Ian Caldwell, Neva Garland, Bruce Jubanowsky, and Jeanne Robinette. Ada translation tools development: Automatic translation of FORTRAN to Ada. ACM SIGADA Ada Letters, 8(6):57–71, November/December 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pollack:1990:CRP

[PC90] Robert H. Pollack and David J. Campbell. Clock resolution and the PIWG benchmark suite. ACM SIG-ADA Ada Letters, 10(3): 91–97, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pedersen:2005:AAO

[PC05] Knut H. Pedersen and Constantinos Constantinides. AspectAda: aspect oriented programming for Ada95. ACM SIGADA Ada Letters, 25(4):79–92, December 2005. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pneuli:1982:RAP

[PD82] A. Pneuli and W. P. DeRoever.

Rendezvous with Ada — a proof theoretical view. In ACM [ACM82], pages 128– 137. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Persch:1983:EEP

[PDG83] Guido Persch, Manfred Dausmann, and Gerhard Goos. Early experience with the programming language Ada. ACM SIGADA Ada Letters, 3(1):63–70, July/August 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pulido:2007:ACP

[PdlPH^{+}07] José A. Pulido, Juan A. de la Puente, Jérome Hugues, Matteo Bordin, and Tullio Vardanega. Ada 2005 code patterns for metamodelbased code generation. ACM SIGADA Ada Letters, 27(2): 53–58, August 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Price:1997:RMF

[PDN97] M. W. Price, S. A. Demurjian, and D. M. Needham. A reusability measurement framework and tool for Ada 95. In ACM [ACM97], pages 125–134. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Plinta:1998:SCG

[PDV98] Charles Plinta, Richard D’Ippolito, and Roger Van Scoy. A specification and code generation tool for message translation and validation. ACM SIGADA Ada Letters, 18(6):276–286, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Penedo:1991:SRM

[Pen91] Maria H. Penedo. SEE reference model working group — summary. ACM SIG-ADA Ada Letters, 11(3): 37–46, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Perez:1988:SIA

[Per88] E. Perez. Simulating inheritance with Ada. ACM SIG-ADA Ada Letters, 8(5):37– 46, September/October 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pettit:2010:DRT

[Pet10] Robert G. Pettit, IV. Designing real-time, concurrent, and embedded software systems using UML and Ada. ACM SIGADA Ada Letters, 30(3): 7–8, December 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Procter:2020:AEL

[PF20] Sam Procter and Peter Feiler.

The AADL error library: an operationalized taxonomy of system errors. ACM SIG-ADA Ada Letters, 39(1):63– 70, January 2020. ISSN 0736-

721X. URL https://dl.
acm.org/doi/abs/10.1145/
3379106.3379113.

Purser:1991:AAL

[PG91] Lynn Purser and Robin Graham. Analysis of Ada-LINPACK benchmark results. ACM SIGADA Ada Letters, 11(4):91–98, May/June 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Paul:1994:HRE

[PG94] Michael J. Paul and John E. Gochenouer. A high resolution event timer Ada package for DOS environments. ACM SIGADA Ada Letters, 14(1):61–67, January/February 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Popov:1992:PS

[PGRZ92] Arcady Popov, Ilia Gindysh, Vadim Rupp, and Vasily Ziabkin. Pallada system. ACM SIGADA Ada Letters, 12 (3):117–125, May/June 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pierce:1985:AEP

[Pie85] R. H. Pierce. Ada in the ECLIPSE project support environment. ACM SIGADA Ada Letters, 5(2):309–320, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Pierce:1987:UPT

[Pie87] R. H. Pierce. On the use of passive tasks in Ada. ACM SIGADA Ada Letters, 7(6): 121–123, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pierpoint:1990:MMA

[Pie90] Tom Pierpoint. Making music with Ada. ACM SIG-ADA Ada Letters, 10(7):63– 69, September/October 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Piotrowski:1986:AIH

[Pio86] W. G. Piotrowski. Ada information hiding — a design goal missing? ACM SIG-ADA Ada Letters, 6(3):43– 55, May/June 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

PhD:2011:SVP

[PJPD11] Joyce L. Tokar PhD, F. David Jones, Paul E. Black PhD, and Chris E. Dupilka. Software vulnerabilities precluded by SPARK. ACM SIGADA Ada Letters, 31(3): 39–46, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pazy:1997:OLS

[PK97] Offer Pazy and Mike Kamrad. Outstanding language (session summary). ACM SIGADA Ada Letters, 17(5):11–15, September/October 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pukite:2007:GDE

[PL07] Paul Pukite and Luke Ludwig. Generic discrete event simulations using DEGAS: application to logic design and digital signal processing. ACM SIGADA Ada Letters, 27(3):27–40, December 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Platek:1986:CLF

[Pla86] Richard Platek. Chairperson’s letter: Formal methods committee. ACM SIG-ADA Ada Letters, 6(2):51– 52, March/April 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ploedereder:1984:PS

[Plo84] Erhard Ploedereder. Project SPERBER. ACM SIGADA Ada Letters, 3(4):92–99, January/February 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ploedereder:1992:HPA

[Plo92] Erhard Ploedereder. How to program in Ada 9X, using Ada 83. ACM SIGADA Ada Letters, 12(6):50–58, November/December 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ploedereder:1998:RGA

[Plo98] Erhard Ploedereder. A readers’ guide to the Ada issues. ACM SIGADA Ada Letters, 18(3):20–112, May 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ploedereder:2001:PMI

[Plo01] Erhard Ploedereder. Panel:

the making of ISO/IEC 8652: Ada 2005. ACM SIG-ADA Ada Letters, 21(4):129– 130, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pinho:2016:SSP

[PM16] Luis Miguel Pinho and Stephen Michell. Session summary: Parallel and multicore systems. ACM SIG-ADA Ada Letters, 36(1):83– 90, June 2016. CODEN AALEE5. ISSN 0736-721X.

Patino-Martinez:2001:ITU

[PMJPA01] M. Patiño-Martínez, R. Jiménez-Peris, and S. Arévalo. Implementing transactions using Ada exceptions: which features are missing? ACM SIGADA Ada Letters, 21(3): 64–75, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pinho:2013:AMC

[PMM13a] Luis Miguel Pinho, Stephen Michell, and Brad Moore. Ada and many-core platforms. ACM SIGADA Ada Letters, 33(2):40–48, August 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pinho:2013:SSP

[PMM13b] Luis Miguel Pinho, Stephen Michell, and Brad Moore. Session summary: parallel and multicore systems. ACM SIGADA Ada Letters, 33(2): 115–122, August 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pinho:2015:SSF

[PMM15] Luis Miguel Pinho, Stephen Michell, and Brad Moore. Session summary: Finegrained parallelism. ACM SIGADA Ada Letters, 35(1): 97–101, April 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pinho:2015:RTF

[PMMT15] Luís Miguel Pinho, Brad Moore, Stephen Michell, and S. Tucker Taft. Real-time fine-grained parallelism in Ada. ACM SIGADA Ada Letters, 35(1):46–58, April 2015. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Potratz:2004:PCB

[Pot04] Eric Potratz. A practical comparison between Java and Ada in implementing a realtime embedded system. ACM SIGADA Ada Letters, 24(1): 71–83, March 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Powers:1990:ASA

[Pow90]  Richard Powers.  Asyn-

chronous and stand-alone entries. ACM SIGADA Ada Letters, 10(9):31–34, Fall 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Powers:1997:ICU

[Pow97] Richard D. Powers. Implementing CIFO using Ada 95 and POSIX. ACM SIG-ADA Ada Letters, 17(5):77– 82, September/October 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pinho:2018:CTM

[PQR18] Luis Miguel Pinho, Eduardo Quiñones, and Sara Royuela. Combining the tasklet model with OpenMP. ACM SIG-ADA Ada Letters, 38(1):14– 18, June 2018. CODEN AALEE5. ISSN 0736-721X.

Pautet:1999:CCS

[PQT99] Laurent Pautet, Thomas Quinot, and Samuel Tardieu. CORBA and CORBA services for DSA. ACM SIG-ADA Ada Letters, 19(3):31– 38, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Payton:1986:CL

[PR86] Teri Payton and Ann Reedy. Chairperson’s letter. ACM SIGADA Ada Letters, 6(2): 73–74, March/April 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Powers:1990:ASR

[PR90] Richard D. Powers and Chuck Roark. Ada support for real-time systems. ACM SIGADA Ada Letters, 10(4): 114–118, Spring 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pritchett:1998:ABS

[PR98] William W. Pritchett, IV and John D. Riley. An ASIS-based static analysis tool for high-integrity systems. ACM SIGADA Ada Letters, 18(6):12–17, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Privitera:1982:ADL

[Pri82] J. P. Privitera. Ada design language for the structured design methodology. In ACM [ACM82], pages 76–90. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Pritchett:1996:AOO

[Pri96] William W. Pritchett IV. Applying object-oriented metrics to Ada 95. ACM SIG-ADA Ada Letters, 16(5):48– 58, September/October 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pritchett:2001:OOM

[Pri01] William W. Pritchett IV. An object-oriented metrics suite for Ada 95. ACM SIG-ADA Ada Letters, 21(4):117– 126, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Procter:2020:ALS

[Pro20] Sam Procter. Architecturelevel security concerns in a safety critical system. ACM SIGADA Ada Letters, 39(1):

50–62, January 2020. ISSN 0736-721X. URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3379106.3379112.

Pinho:2021:RTI

[PRQ21] Luis Miguel Pinho, Sara Royuela, and Eduardo Quiñones. Real-time issues in the Ada parallel model with OpenMP. ACM SIGADA Ada Letters, 40(2):96–102, April 2021. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3463478.3463491.

Phillips:1984:RAR

[PS84] Stephen P. Phillips and Peter R. Stevenson. The role of Ada in real time embedded applications. ACM SIG-ADA Ada Letters, 3(4):99– 111, January/February 1984. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Plantec:2006:RAL

[PS06] Alain Plantec and Frank Singhoff. Refactoring of an Ada 95 library with a Meta CASE tool. ACM SIG-ADA Ada Letters, 26(3):61– 70, December 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pautet:1999:WFD

[PT99] Laurent Pautet and Samuel Tardieu. What future for the distributed systems annex? ACM SIGADA Ada Letters, 19(3):77–82, September 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pucci:2017:GHT

[Puc17] Vincent Pucci. Gem #136:

How tall is a kilogram? ACM SIGADA Ada Letters, 37(1): 26–30, June 2017. CODEN AALEE5. ISSN 0736-721X.

Puk:1988:RMI

[Puk88] Richard F. Puk. Report on the meeting of ISO/IEC JTC1/SC24/WG4 computer graphics language bindings held in Tokyo, Japan, April 12–20, 1988. ACM SIG-ADA Ada Letters, 8(4):97– 103, July/August 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pukite:1993:AIC

[Puk93]  Paul R. Pukite.  Auto-

mated interface code generation from Ada specifications. ACM SIGADA Ada Letters, 13(3):74–85, May/June 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pukite:1994:AMW

[Puk94]  Paul R. Pukite.  Ada

for MS-Windows applications. ACM SIGADA Ada Letters, 14(1):30–37, January/February 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pullan:1995:PAS

[Pul95] Wayne Pullan. A pragmatic Ada software design/development methodology. ACM SIGADA Ada Letters, 15(2):31–39, March/April 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pinho:1998:MAB

[PV98] Luís Miguel Pinho and Francisco Vasques. Multi-μ: an Ada 95 based architecture for fault tolerance support of real-time systems. ACM SIGADA Ada Letters, 18(6):52–60, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pinho:1999:RMR

[PV99a] Luís Miguel Pinho and Francisco Vasques. Replica management in real-time Ada 95 applications. ACM SIG-ADA Ada Letters, 19(2): 21–27, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pinho:1999:AAA

[PV99b] Luís Miguel Pinho and Francisco Vasques. To Ada or not to Ada: Adaing vs. Javaing in real-time systems. ACM SIGADA Ada Letters, 19(4): 37–43, December 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pinho:2002:URS

[PV02] Luís Miguel Pinho and Francisco Vasques. Using Ravenscar to support fault-tolerant real-time applications. ACM SIGADA Ada Letters, 22(4): 47–52, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Panunzio:2013:CEA

[PV13] Marco Panunzio and Tullio Vardanega. Charting the evolution of the Ada Ravenscar code archetypes. ACM SIGADA Ada Letters, 33(1): 64–83, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pinho:2018:SSP

[PV18] Luis Miguel Pinho and Tullio Vardanega. Session summary: Parallel programming. ACM SIGADA Ada Letters, 38(1):58–60, June 2018. CO-DEN AALEE5. ISSN 0736-721X.

Pinho:2001:PAM

[PVF01] Luís Miguel Pinho, Francisco Vasques, and Luis Ferreira. Programming atomic multicast in CAN. ACM SIGADA Ada Letters, 21(1): 79–84, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Poutanen:1985:NBR

[PVV85] Olavi Poutanen, Kari-Matti Varanki, and Tapio Välimäki. Notes on building a relational database management system in Ada. ACM SIG-ADA Ada Letters, 5(2):14– 24, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Pautet:1997:TFS

[PW97] L. Pautet and T. Wolf. Transparent filtering of streams in GLADE. In ACM [ACM97], pages 11–20. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Pritchett:2001:VTT

[PW01] William Pritchett and Brian Wood. Vetronics technology testbed: experience report. ACM SIGADA Ada Letters, 21(4):115–116, December 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Persch:1980:OPA

[PWDD80] Guido Persch, Georg Winterstein, Manfred Dausmann, and Sophia Drossopoulou. Overloading in preliminary Ada. In ACM [ACM80], pages 47–56. CODEN SIN-ODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Pyle:1984:PSA

[Pyl84]  I. C. Pyle.  A pack-

age for specifying Ada programs. ACM SIGADA Ada Letters, 3(5):63–68, March/April 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Pyster:1985:EEE

[Pys85]  Arthur Pyster.  Experi-

ence with existing environments. ACM SIG-ADA Ada Letters, 4(5):59– 64, March/April 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Future Ada Environment Workshop.

Paprzycki:1997:ADS

[PZ97a] Marcin Paprzycki and Janusz Zalewski. Ada in distributed systems: an overview. ACM SIGADA Ada Letters, 17 (2):67–81, March/April 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Paprzycki:1997:PCA

[PZ97b] Marcin Paprzycki and Janusz Zalewski. Parallel computing in Ada: an overview and critique. ACM SIG-ADA Ada Letters, 17(2):55– 62, March/April 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Quinot:2001:DTG

[QKP01] Thomas Quinot, Fabrice Kordon, and Laurent Pautet. DROOPI: Towards a generic middleware. ACM SIG-ADA Ada Letters, 21(2): 26–52, June 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Quiggle:1990:ATCb

[Qui90a] Thomas J. Quiggle. Asynchronous transfer of control and interrupt handling. ACM SIGADA Ada Letters, 10 (9):46–49, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Quiggle:1990:ATCa

[Qui90b] Thomas J. Quiggle. Asynchronous transfer of control working group. ACM SIGADA Ada Letters, 10(4): 15–24, Spring 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Quiggle:1990:EPE

[Qui90c] Thomas J. Quiggle. Efficient periodic execution of Ada tasks. ACM SIG-ADA Ada Letters, 10(9): 141–146, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Quiggle:1990:RRI

[Qui90d] Thomas J. Quiggle. Ramifications of re-introducing asynchronous exceptions to the Ada language. ACM SIGADA Ada Letters, 10(4): 25–31, Spring 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Quinot:2011:GDSa

[Qui11a] Thomas Quinot. Gem #84:

the distributed systems annex 1 — simple client/server. ACM SIGADA Ada Letters, 31(2):44–47, August 2011. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Quinot:2011:GDSb

[Qui11b] Thomas Quinot. Gem #85:

the distributed systems annex 2 — distributed objects. ACM SIGADA Ada Letters, 31(2):48–51, August 2011. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Quinot:2011:GDSc

[Qui11c] Thomas Quinot. Gem #87: the distributed systems annex, part 3 — mailboxes. ACM SIGADA Ada Letters, 31(2):56–58, August 2011. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Quinot:2012:GDS

[Qui12] Thomas Quinot. Gem #90: the distributed systems annex, part 4 — DSA and C. ACM SIGADA Ada Letters, 32(1):27–29, April 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Quinot:2017:GBE

[Qui17] Thomas Quinot. Gem #140: Bridging the endianness gap. ACM SIGADA Ada Letters, 37(1):46–49, June 2017. CO-DEN AALEE5. ISSN 0736-721X.

Rosenfeld:1991:ECP

[RA91] D. A. Rosenfeld and G. G. Allen. The EACM code performance anomaly detector. In ACM [ACM91b], pages 124–135. ISBN 0-89791-393-0. LCCN ????

Racine:1988:WUC

[Rac88] R. Racine. Why the use clause is beneficial (Ada). ACM SIGADA Ada Letters, 8(3):123–127, May/June 1988. CODEN AALEE5.

ISSN 1094-3641 (print), 1557-9476 (electronic).

Racine:1989:WUC

[Rac89] Roger Racine. Why the use clause is beneficial. ACM SIGADA Ada Letters, 8(3): 123–127, May/June 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Radi:1994:AIQ

[Rad94] Thomas S. Radi. Automating improvements to the quality of your code, software quality. ACM SIG-ADA Ada Letters, 14(4):58– 71, July/August 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Raiha:1994:DA

[Räi94] Liisa Räihä. Delegation with Ada 9x. ACM SIGADA Ada Letters, 14(6):53–56, November/December 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Riccardi:1985:RSS

[RB85] G. A. Riccardi and T. P.

Baker. A runtime supervisor to support Ada tasking: Rendezvous and delays. ACM SIGADA Ada Letters, 5(2):329–342, September/October 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Roby:2001:SAW

[RC01] Clyde Roby and Currie Colket. SIGAda 2000 ASIS Workshop report. ACM SIGADA Ada Letters, 21 (2):12–16, June 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ras:2010:DRT

[RC10a] Jim Ras and Albert M. K. Cheng. A deterministic runtime environment for Ada-05 on the ATmega16 microcontroller. ACM SIG-ADA Ada Letters, 30(3):13– 22, December 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Real:2010:IOM

[RC10b] Jorge Real and Alfons Crespo. Incorporating operating modes to an Ada realtime framework. ACM SIG-ADA Ada Letters, 30(1): 73–85, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Real:2002:PCC

[RCWB02] Jorge Real, Alfons Crespo, Andy Wellings, and Alan Burns. Protected ceiling changes. ACM SIG-ADA Ada Letters, 22(4):66– 71, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Roger:2023:AMS

[RD23] Jean-Charles Roger and Pierre Dissaux. AADL modelling with SysML v2. ACM SIGADA Ada Letters, 43 (1):42–45, June 2023. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3631483.3631486.

Real:2013:SSO

[RdlP13] Jorge Real and Juan Antonio de la Puente. Session summary: open issues. ACM SIGADA Ada Letters, 33(2): 131–132, August 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ruiz:2001:ESR

[RdlPZFM01] José Ruiz, Juan A. de la Puente, Juan Zamorano, and Ramón Fernández-Marina. Exception support for the Ravenscar Profile. ACM SIG-ADA Ada Letters, 21(3):76– 79, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Riley:1997:IAD

[RDP97] J. Riley, S. Dungrani, and W. Pritchett. An instance of the application download pattern: The SPAIDS software loader/verifier domain analysis and implementation. In ACM [ACM97], pages 273–278. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Reisner:1998:ASO

[RDS98] John A. Reisner, Steven A. Demurjian, and Sr. Addressing security for objectoriented design and Ada 95 development. ACM SIG-ADA Ada Letters, 18(2):89– 104, March 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Reboul:2017:GAQa

[Reb17a]  Valentine Reboul.  Gem

#137: Ada quiz 2— an heir and a spare? ACM SIG-ADA Ada Letters, 37(1):31– 38, June 2017. CODEN AALEE5. ISSN 0736-721X.

Reboul:2017:GAQb

[Reb17b]  Valentine Reboul.  Gem

#145: Ada quiz 3— statements. ACM SIGADA Ada Letters, 37(2):23–26, December 2017. CODEN AALEE5. ISSN 0736-721X.

Redwine:1985:EA

[Red85] Sam Redwine. Environment architectures. ACM SIG-ADA Ada Letters, 4(5):100– 104, March/April 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Future Ada Environment Workshop.

Reedy:1985:ACL

[Ree85] Ann Reedy. Ada contracts list. ACM SIGADA Ada Letters, 5(3–6):64–66, November/December 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Reedy:1986:ACL

[Ree86] Ann Reedy. Ada contracts list update. ACM SIG-ADA Ada Letters, 6(2):94, March/April 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Reedy:1988:CCR

[Ree88] Ann Reedy. CAIS comments and responses. ACM SIG-ADA Ada Letters, 8(2):28– 38, March/April 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rehmer:1987:DIM

[Reh87] Karl Rehmer. Development and implementation of the Magnavox generic Ada basic mathematics package. ACM SIGADA Ada Letters, 7(3): 73–83, May/June 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Reifer:1987:AIQ

[Rei87] Donald J. Reifer. Ada’s impact: a quantitative assessment. In ACM [ACM87a], pages 1–13. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Reznik:2022:GSA

[Rez22] Max Reznik. Getting started with AdaWebPack. ACM SIGADA Ada Letters, 42 (1):58–60, June 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3577949.3577957.

Roy:1990:PAM

[RG90] Daniel Roy and Lakshmi Gupta. PIWG analysis methodology. ACM SIG-ADA Ada Letters, 10(3):217– 229, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Raymond:1991:SRE

[RH91] G. E. Raymond and D. M. Hollis. Software reuse economics model. In ACM [ACM91b], pages 141–155. ISBN 0-89791-393-0. LCCN ????

Roberts-Hayden:1996:LSV

[RH96] Charlene Roberts-Hayden.

Letter from SIGAda Vice-Chair of Meetings. ACM SIGADA Ada Letters, 16(1): 16–??, January 1, 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rivas:2001:EAR

[RH01] Mario Aldea Rivas and Michael González Harbour. Extending Ada’s real-time systems annex with the POSIX scheduling services. ACM SIGADA Ada Letters, 21(1):20–26, March 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rivas:2002:ADS

[RH02] Mario Aldea Rivas and Michael González Harbour. Application-defined scheduling in Ada. ACM SIG-ADA Ada Letters, 22(4):77– 84, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rivas:2003:ADS

[RH03] Mario Aldea Rivas and Michael González Harbour. Application-defined scheduling in Ada. ACM SIG-ADA Ada Letters, 23(4):42– 51, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rivas:2007:OSS

[RH07] Mario Aldea Rivas and Michael González Harbour. Operating system support for execution time budgets for thread groups. ACM SIGADA Ada Letters, 27(2): 67–71, August 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rivas:2010:ETM

[RH10] Mario Aldea Rivas and Michael González Harbour. Execution time monitoring and interrupt handlers: position statement. ACM SIG-ADA Ada Letters, 30(1): 68–72, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rivas:2016:SSL

[RH16] Mario Aldea Rivas and Michael González Harbour. Session summary: Language issues. ACM SIGADA Ada Letters, 36(1):94–97, June 2016. CODEN AALEE5. ISSN 0736-721X.

Richards:2020:CPM

[Ric20] Ray Richards. CASE program: Motivation and challenges. ACM SIGADA Ada Letters, 39(1):9–16, January 2020. ISSN 0736-721X.

URL https://dl.acm.org/
doi/abs/10.1145/3379106.
3379108.

Riehle:1994:AC

[Rie94] Richard Riehle. Ada in China. ACM SIGADA Ada Letters, 14(4):72–75, July/August 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Riehle:1998:NIG

[Rie98] Richard Riehle. New ideas for generic components in Ada. ACM SIGADA Ada Letters, 18(5):67–86, September/October 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rivera:2017:HBD

[Riv17] J. Germán Rivera. Hardwarebased data protection/isolation at runtime in Ada code for microcontrollers. ACM SIG-ADA Ada Letters, 37(2):43– 50, December 2017. CODEN AALEE5. ISSN 0736-721X.

Rusanova:1999:SPP

[RK99] Olga Rusanova and Alexandr Korochkin. Scheduling problems for parallel and distributed systems. ACM SIG-ADA Ada Letters, 19(3):195– 201, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Romanovsky:2001:EC

[RK01] Alexander Romanovsky and Jörg Kienzle. Exceptions and concurrency. ACM SIG-ADA Ada Letters, 21(3):13– 15, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Real:2001:SDC

[RLC01] Jorge Real, Albert Llamosí, and Alfons Crespo. A semantics for dynamic ceiling priorities in Ada. ACM SIGADA Ada Letters, 21(1): 91–95, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosenberg:1980:CAC

[RLHS80] Jonathan Rosenberg, David Alex Lamb, Andy Hisgen, and Mark Sherman. The charrette Ada compiler. In ACM [ACM80], pages 72– 81. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Reisner:1998:ICS

[RLPD98] John A. Reisner, Zeenat Lainwala, Thomas J. Peters, and Steven Demurjian, Sr. Implementing a culling and self-intersection algorithm for stereo-lithography files in Ada 95. ACM SIGADA Ada Letters, 18(6):104–113, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Roark:1988:AAM

[RM88] Chuck Roark and Ron McAfee. The applicability of Ada to MIL-STD-1750A. ACM SIGADA Ada Letters, 8(3):84–86, May/June 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Real:2007:BAI

[RM07] Jorge Real and Stephen Michell. Beyond Ada 2005: Introduction. ACM SIG-ADA Ada Letters, 27(2): 72–74, August 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Real:2018:SST

[RM18] Jorge Real and Brad Moore.

Session summary: Time triggered scheduling in Ravenscar. ACM SIGADA Ada Letters, 38(1):66–69, June 2018. CODEN AALEE5. ISSN 0736-721X.

Rosen:2011:HMA

[RMT11] Jean-Pierre Rosen, Brad Moore, and Tucker Taft. How to make Ada go ‘viral’. ACM SIGADA Ada Letters, 31(3): 35–36, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Roast:1988:AAR

[Roa88] C. Roast. The applicability of Ada (R) to MIL-STD-1750A. ACM SIGADA Ada Letters, 8(3):84–86, May/June 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Roast:1989:AAM

[Roa89] C. Roast. The applicability of Ada to Mil-Std-1750A. ACM SIGADA Ada Letters, 8(3): 84–86, May/June 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Roby:1986:CCS

[Rob86] Clyde Roby. CAIS/CASWG/ SEI workshop summary. ACM SIGADA Ada Letters, 6(2):77–78, March/ April 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Roberts:1992:DDR

[Rob92] Steve Roberts. Difficulties in developing re-usable software components arising from the lack of user redefinition of standard assignment. ACM SIGADA Ada Letters, 12 (4):36–41, July/August 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Roby:1997:MDA

[Rob97] Clyde Roby. Minutes of 3 December 1996 ASISWG/ ASISRG meeting with Tri-Ada’96. ACM SIGADA Ada Letters, 17(2):18–25, March/April 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:1985:ICA

[Rog85] Mike W. Rogers. IT companies’ acceptance of and attitudes towards Ada. ACM SIGADA Ada Letters, 5(2):1– 13, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Rogers:1988:DAA

[Rog88] P. Rogers. Dimensional analysis in Ada. ACM SIGADA Ada Letters, 8(5):92–100, September/October 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:1997:BRC

[Rog97] Patrick Rogers. Book review: Concurrency In Ada, by Alan Burns and Andy Wellings. ACM SIGADA Ada Letters, 17(6):108, November/December 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2009:EHR

[Rog09a] Pat Rogers. Embedded, hard, real-time systems with Ada. ACM SIGADA Ada Letters, 29(3):17–18, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2009:GBBa

[Rog09b]  Pat Rogers.  Gem #35:

bounded buffer package in GNAT hierarchy (part 1). ACM SIGADA Ada Letters, 29(1):54–56, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2009:GBBb

[Rog09c]  Pat Rogers.  Gem #37:

bounded buffer package in GNAT hierarchy (part 2). ACM SIGADA Ada Letters, 29(1):58–60, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2009:GES

[Rog09d] Pat Rogers. Gem #39: efficient stream I/O for array types. ACM SIG-ADA Ada Letters, 29(1): 62–64, April 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2009:RBR

[Rog09e]  Pat Rogers.  Review of

the book: Real-Time Systems and Programming Languages (4th edition) by Alan Burns and Andy Wellings. ACM SIGADA Ada Letters, 29(2):71, August 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2011:LCS

[Rog11a]  James S. Rogers.  Lan-

guage choice for safety critical applications. ACM SIG-ADA Ada Letters, 31(3):81– 90, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2011:GSL

[Rog11b] Pat Rogers. Gem #70: the scope locks idiom. ACM SIGADA Ada Letters, 31(1): 28–31, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2011:GGS

[Rog11c]  Pat Rogers.  Gem #81:

GNAT semaphores. ACM SIGADA Ada Letters, 31(2): 33–35, August 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2011:RBB

[Rog11d] Patrick Rogers. Review of the book: Building parallel, embedded, and real-time applications with Ada, by John McCormick, Frank Singhoff, and Jérôme Hugues. ACM SIGADA Ada Letters, 31 (2):70, August 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2012:GHPa

[Rog12a] Pat Rogers. Gem #93: high performance multi-core programming — part 1. ACM SIGADA Ada Letters, 32(1): 37–38, April 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2012:GHPc

[Rog12b] Pat Rogers. Gem #98: high performance multi-core programming — part 2. ACM SIGADA Ada Letters, 32(2): 28–30, August 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rogers:2021:APS

[Rog21] Patrick Rogers. From Ada to Platinum SPARK: a case study. ACM SIG-ADA Ada Letters, 40(2): 76–91, April 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3463478.3463488.

Romanowsky:1986:AP

[Rom86] Helen Romanowsky. Ada publications. ACM SIG-ADA Ada Letters, 6(2):109– 110, March/April 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Romanowsky:1988:EPW

[Rom88] H. Romanowsky. Educational products working group Ada publications list. ACM SIG-ADA Ada Letters, 8(3):81– 83, May/June 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Romanovsky:2000:DDC

[Rom00] Alexander Romanovsky. Diversely designed classes for use by multiple tasks. ACM SIGADA Ada Letters, 20(1): 25–37, March 2000. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Romanovsky:2001:HEE

[Rom01] Alexander Romanovsky. How to evolve exception handling in Ada. ACM SIG-ADA Ada Letters, 21(3):16– 18, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Romanski:2005:AAI

[Rom05] George Romanski. Ada in the avionics industry. ACM SIG-ADA Ada Letters, 25(4):109– 114, December 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Roski:1986:DSD

[Ros86a] S. Roski. DoD-STD-2167 default Ada design and coding standard. ACM SIG-ADA Ada Letters, 6(5):34– 44, September/October 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Roski:1986:DSC

[Ros86b]  Steve Roski.  DoD-STD-

2167A coding standard (draft). ACM SIGADA Ada Letters, 6(5):34–44, September/October 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ross:1986:CAP

[Ros86c] Donald L. Ross. Classifying Ada packages. ACM SIG-ADA Ada Letters, 6(4):53– 65, July/August 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosen:1987:DUC

[Ros87a] J. P. Rosen. In defense of the “use” clause. ACM SIGADA Ada Letters, 7(7):77–81, November/December 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosen:1987:CDA

[Ros87b] Steven M. Rosen. Controlling dynamic Ada objects in large Ada systems. ACM SIG-ADA Ada Letters, 7(5):79– 92, September/October 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosen:1987:CDO

[Ros87c] Steven M. Rosen. Controlling dynamic objects in large Ada systems. ACM SIG-ADA Ada Letters, 7(5):79– 92, September/October 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosenblum:1987:ECK

[Ros87d] David S. Rosenblum. An efficient communication kernel for distributed Ada runtime tasking supervisors. ACM SIGADA Ada Letters, 7(2): 102–117, March/April 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ross:1989:FPI

[Ros89] Donald L. Ross. The form of a passive iterator. ACM SIG-ADA Ada Letters, 9(2):102– 105, March/April 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosen:1995:NCC

[Ros95]  J.-P. Rosen.  A naming

convention for classes in Ada 9X. ACM SIGADA Ada Letters, 15(2):54–58, March/April 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosen:1996:AAA

[Ros96] J.-P. Rosen. All aboard Ada 95! ACM SIGADA Ada Letters, 16(1):70, January/February 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosen:2004:EDT

[Ros04] J.-P. Rosen. Experiences in developing a typical Web/ database application. ACM SIGADA Ada Letters, 24(1): 38–48, March 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosen:2009:AP

[Ros09]  J.-P. Rosen.  The Ada

paradox(es). ACM SIG-ADA Ada Letters, 29(2):28– 35, August 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosen:2010:UOO

[Ros10] Jean-Pierre Rosen. Use of object oriented technologies in high reliability system. ACM SIGADA Ada Letters, 30(3): 3–4, December 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosen:2011:DCC

[Ros11a] Jean-Pierre Rosen. Designing and checking coding standards for Ada. ACM SIG-ADA Ada Letters, 31(3):13– 14, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosen:2011:DPU

[Ros11b] Jean-Pierre Rosen. Developing a profile for using object-oriented Ada in highintegrity systems. ACM SIGADA Ada Letters, 31(1): 9–10, April 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosen:2021:AVL

[Ros21]  J-P. Rosen.  ASIS vs.

LibAdalang: a comparative assessment. ACM SIG-ADA Ada Letters, 41(2):81– 85, December 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3530801.3530808.

Rosen:2022:ANM

[Ros22]  Jean-Pierre Rosen.  The

Ada numerics model. ACM SIGADA Ada Letters, 42 (1):43–45, June 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3577949.3577953.

Roubine:1985:PLF

[Rou85] O. Roubine. Programming large and flexible systems in Ada. ACM SIGADA Ada Letters, 5(2):197–209, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Roy:1990:PMM

[Roy90a] Daniel Roy. PIWG measurement methodology. ACM SIGADA Ada Letters, 10(3): 72–90, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Roy:1990:RI

[Roy90b] Daniel M. Roy. Results introduction. ACM SIG-ADA Ada Letters, 10(3): 138, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosenfeld:1990:IOA

[RR90] David Rosenfeld and Mike Ryer. Issues in optimizing Ada code. ACM SIG-ADA Ada Letters, 10(3):60– 71, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Real:2013:SSM

[RR13] Jorge Real and José F. Ruiz.

Session summary: multiprocessor issues, part 1. ACM SIGADA Ada Letters, 33(1): 134–137, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rathje:2014:FMC

[RR14] William Rathje and Brad Richards. A framework for model checking UDP network programs with Java Pathfinder. ACM SIG-ADA Ada Letters, 34(3):81– 86, December 2014. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Real:2016:SSE

[RR16] Jorge Real and Pat Rogers.

Session summary: “experience”. ACM SIGADA Ada Letters, 36(1):101–102, June 2016. CODEN AALEE5. ISSN 0736-721X.

Rogers:2015:TER

[RRG15] P. Rogers, J. Ruiz, and T. Gingold. Toward extensions to the Ravenscar profile. ACM SIGADA Ada Letters, 35(1):32–37, April 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rennels:1991:PAT

[RS91] Deborah Rennels and Edmond Schonberg. A program analysis tool for evaluating the Ada compiler validation suite. ACM SIG-ADA Ada Letters, 11(3):137– 146, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Romanovsky:2001:EEH

[RS01] Alexander Romanovsky and Bo Sandén. Except for exception handling . . .. ACM SIGADA Ada Letters, 21(3): 19–25, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Real:2016:CST

[RSC16] Jorge Real, Sergio Sáez, and Alfons Crespo. Combined scheduling of time-triggered plans and priority scheduled task sets. ACM SIGADA Ada Letters, 36(1):68–76, June 2016. CODEN AALEE5. ISSN 0736-721X.

Real:2018:RST

[RSC18] Jorge Real, Sergio Sáez, and Alfons Crespo. Ravenscar support for time-triggered scheduling. ACM SIG-ADA Ada Letters, 38(1):41– 54, June 2018. CODEN AALEE5. ISSN 0736-721X.

Runge:2019:TSC

[RSK^{+}19] Tobias Runge, Ina Schaefer, Alexander Knüppel, Loek Cleophas, Derrick Kourie, and Bruce W. Watson. Tool support for confidentialityby-construction. ACM SIG-ADA Ada Letters, 38(2):64– 68, December 2019. ISSN 0736-721X. URL https:

//dl.acm.org/doi/abs/10.
1145/3375408.3375413.

Rybin:1996:AGG

[RSZ96] Sergey Rybin, Alfred Strohmeier, and Eugene Zueff. ASIS for GNAT: goals, problems and implementation strategy. ACM SIGADA Ada Letters, 16(2):39–49, March/April 1996. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rosen:2009:NSM

[RT09] Jean-Pierre Rosen and Tucker Taft. The new semantic model in ASIS for Ada 2005. ACM SIGADA Ada Letters, 29(3):127–132, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rivas:2021:MAA

[RT21] Mario Aldea Rivas and Hector Perez Tijero. M2OS for Arduino Uno: Ada tasks and Arduino libraries working together. ACM SIG-ADA Ada Letters, 41(1): 78–82, June 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3570315.3570322.

Rivas:2015:MAP

[RTH15] Mario Aldea Rivas, Héctor Pérez Tijero, and Michael González Harbour. Multiprocessor Ada platform based on MaRTE OS and GNAT. ACM SIG-ADA Ada Letters, 35(1): 74–79, April 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Roubine:1982:LLL

[RTM82] O. Roubine, J. Teller, and O. Maurel. LOLITA — a low level intermediate language for Ada. In ACM [ACM82], pages 251– 260. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821. Rudolph:1983:ODA

[Rud83]  Bruce L. Rudolph.  An

overview of the design of an Ada ballistics system. ACM SIGADA Ada Letters, 2(5): 60–61, March/April 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ruiz:2010:TRE

[Rui10]  José F. Ruiz.  Towards

a Ravenscar extension for multi-processor systems. ACM SIGADA Ada Letters, 30(1): 86–90, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ruiz:2013:GRT

[Rui13] José F. Ruiz. Going real-time with Ada 2012 and GNAT. ACM SIGADA Ada Letters, 33(1):45–52, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ruocco:2005:EUS

[Ruo05] Anthony S. Ruocco. Experiences using SPARK in an undergraduate CS course. ACM SIGADA Ada Letters, 25(4): 37–40, December 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Real:1999:DCP

[RW99] Jorge Real and Andy Wellings.

Dynamic ceiling priorities and Ada 95. ACM SIG-ADA Ada Letters, 19(2): 41–48, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rybin:1994:ARO

[Ryb94] Sergei I. Rybin. Ada in Russia: an overview. ACM SIG-ADA Ada Letters, 14(3):74– 79, May/June 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rymer:1994:EAC

[Rym94] John Rymer. Evolving an Ada curriculum to 9X. ACM SIGADA Ada Letters, 14 (4):76–80, July/August 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Rymer:1998:RTA

[Rym98] John Rymer. Rethinking testing with Ada95. ACM SIGADA Ada Letters, 18 (1):40–47, January/February 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sacha:1989:AAR

[Sac89] Krzysztof M. Sacha. Ada: Adding reliability and efficiency to task communication in programming distributed control systems. ACM SIG-ADA Ada Letters, 9(6):80– 89, September/October 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sherrill:2001:IPL

[SAH01] Joel Sherrill, Jennifer Averett, and Glenn Humphrey. Implementing a product line-based architecture in Ada. ACM SIGADA Ada Letters, 21(4): 39–46, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Saidi:2008:LFS

[Saï08] Hassen Saïdi. Logical foundation for static analysis: application to binary static analysis for security. ACM SIGADA Ada Letters, 28(1):

96–102, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Salwin:1989:VV

[Sal89] Arthur E. Salwin. The variabilities are variable. ACM SIGADA Ada Letters, 9(4): 84–86, May/June 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Salwin:1992:UPE

[Sal92]  Arthur E. Salwin.  Us-

ing the proposed elementary functions standard to build a strongly typed trig package. ACM SIGADA Ada Letters, 12(5):59–63, September/October 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sankar:1989:AST

[San89] Sriram Sankar. APE — a set of TEX macros to format Ada programs. ACM SIGADA Ada Letters, 9(7):114–128, November/December 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sanden:1997:CDP

[San97] B. I. Sanden. Concurrent design patterns for resource sharing. In ACM [ACM97], pages 173–188. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Sanden:2000:ISM

[San00] Bo I. Sandén. Implementation of state machines with tasks and protected objects. ACM SIGADA Ada Letters, 20(2):38–56, June 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/june2000/sanden.
pdf.

Sanden:2001:EP

[San01a] Bo I. Sandén. Exception propagation. ACM SIG-ADA Ada Letters, 21(3):8– 10, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Santhanam:2001:ASM

[San01b] Usha Santhanam. Automating software module testing for FAA certification. ACM SIGADA Ada Letters, 21(4): 31–38, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sanden:2003:RTP

[San03a] Bo I. Sandén. Real-time programming safety in Java and Ada. ACM SIG-ADA Ada Letters, 23(2): 32–46, June 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Santhanam:2003:AFQ

[San03b] V. Santhanam. The anatomy of an FAA-qualifiable Ada subset compiler. ACM SIG-ADA Ada Letters, 23(1): 40–43, March 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sanden:2012:HTO

[San12] Bo I. Sandén. HILT’12 tutorial overview /design of multitask software: the entitylife modeling approach. ACM SIGADA Ada Letters, 32(3): 1–2, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Sautejeau:2005:MSS

[Sau05] Xavier Sautejeau. Modeling SPARK systems with UML. ACM SIGADA Ada Letters, 25(4):11–16, December 2005. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sherman:1980:FSA

[SB80] Mark S. Sherman and Martha S.

Borkan. A flexible semantic analyzer for Ada. In ACM [ACM80], pages 62– 71. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Shen:1999:LKM

[SB99] Hongfeng Shen and Theodore P. Baker. A Linux kernel module implementation of restricted Ada tasking. ACM SIGADA Ada Letters, 19(2): 96–103, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sward:2005:OSP

[SB05] Ricky E. Sward and Leemon C. Baird, III. Optimizing the SPARK program slicer. ACM SIGADA Ada Letters, 25(4): 17–22, December 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sward:2011:SOA

[SB11] Ricky E. Sward and Jeff Boleng. Service-oriented architecture (SOA)concepts and implementations. ACM SIGADA Ada Letters, 31(3): 3–4, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sward:2012:SOA

[SB12] Ricky E. Sward and Jeff Boleng. Service-oriented architecture (SOA) concepts and implementations. ACM SIGADA Ada Letters, 32(3): 11–12, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Senn:2023:MRB

[SB23] E. Senn and L. W. J. Bourdon. Modeling ROS based applications with AADL. ACM SIGADA Ada Letters, 43(1):64–68, June 2023. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3631483.3631492.

Shing:1998:MSS

[SBH^{+}98] M. Shing, V. Berzins, M. Holden, C. Eagle, and Luqi. Master of science in software engineering via distance learning. ACM SIGADA Ada Letters, 18(5):111–125, September/October 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Schultz:1987:ABA

[SC87] William L. Schultz and Asheem Chandna. An Ada based approach to factory scale MAP network simulation. In ACM [ACM87a], pages 116–125. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Shen:1992:GFP

[SC92] Jun Shen and Gordon V.

Cormack. On generic formal package parameters in Ada 9X. ACM SIGADA Ada Letters, 12(3):110–116, May/June 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Strohmeier:2001:SSC

[SC01] Alfred Strohmeier and Stanislav Chachkov. A side-by-side comparison of exception handling in Ada and Java. ACM SIGADA Ada Letters, 21(3): 41–56, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sward:2004:AAP

[SC04a] Ricky E. Sward and A. T. Chamillard. AdaSlicer: an Ada program slicer. ACM SIGADA Ada Letters, 24(1): 10–16, March 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sward:2004:REG

[SC04b] Ricky E. Sward and A. T. Chamillard. Re-engineering global variables in Ada. ACM SIGADA Ada Letters, 24(4): 29–34, December 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Shindi:2006:EPC

[SC06] Rajaa S. Shindi and Shaun Cooper. Evaluate the performance changes of processor simulator benchmarks When context switches are incorporated. ACM SIG-ADA Ada Letters, 26(3):9– 14, December 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Saez:2013:DSS

[SC13] Sergio Sáez and Alfons Crespo. Deferred setting of scheduling attributes in Ada 2012. ACM SIG-ADA Ada Letters, 33(1): 93–100, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sousa:2022:ANN

[SCC22] Inês Sousa Sousa, António Casimiro, and José Cecílio. Artificial neural networks for real-time data quality assurance. ACM SIG-ADA Ada Letters, 42(1): 86–89, June 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3577949.3577966.

Smith:1985:TKD

[SCD^{+}85] John M. Smith, Arvola Chan, Sy Danberg, Stephen Fox, and Anil Nori. A tool kit for database programming in Ada. ACM SIG-ADA Ada Letters, 5(2):41– 57, September/October 1985. Ada in Use: Proceedings of the Ada International Conference, Paris, 14–16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Shen:1992:LPI

[SCD92] Jun Shen, Gordon V. Cormack, and Dominic Duggan. Local package instances are not equivalent to generic formal package parameters. ACM SIGADA Ada Letters, 12(6):47–49, November/December 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sward:2004:CAU

[SCFG04] Ricky E. Sward, Martin C. Carlisle, Barry S. Fagin, and David S. Gibson. The case for Ada at the USAF Academy. ACM SIGADA Ada Letters, 24(1):68–70, March 2004. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Schacht:1987:APT

[Sch87a] Eric N. Schacht. Ada programming techniques, research and experiences on a fast control loop system. In ACM [ACM87a], pages 164– 169. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Schefstrom:1987:SET

[Sch87b]  Dick Schefstrom.  The

system-oriented editor — a tool for managing large software systems. In ACM [ACM87a], pages 56–59. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Schuler:1991:EOO

[Sch91] M. P. Schuler. Evolving object oriented design, a case study. In ACM [ACM91b], pages 50–61. ISBN 0-89791-393-0. LCCN ????

Schmidt:2009:ARD

[Sch09] Richard B. Schmidt. An Ada retrospective: developing large, mature, reliable systems. ACM SIG-ADA Ada Letters, 29(3):21– 22, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Schmidt:2010:ERA

[Sch10a] Richard B. Schmidt. Experience report: Ada & Java integration in the FAA’s ERAM SWIM program. ACM SIG-ADA Ada Letters, 30(3):33– 34, December 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Schonberg:2010:TAI

[Sch10b] Edmond Schonberg. Towards Ada 2012: an interim report. ACM SIGADA Ada Letters, 30(3):63–70, December 2010. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sebesta:1987:YAS

[Seb87] R. W. Sebesta. Yet another survey of Ada usage and Ada training. ACM SIG-ADA Ada Letters, 7(5):34– 39, September/October 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Seidewitz:1991:OOP

[Sei91]  Ed Seidewitz.  Object-

oriented programming through type extension in Ada 9X. ACM SIGADA Ada Letters, 11(2):86–97, March/April 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Seidewitz:1992:OOP

[Sei92]  Ed Seidewitz.  Object-

oriented programming with mixins in Ada. ACM SIG-ADA Ada Letters, 12(2):76– 90, March/April 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Seidewitz:2014:UME

[Sei14] Ed Seidewitz. UML with meaning: executable modeling in foundational UML and the Alf action language. ACM SIGADA Ada Letters, 34(3):61–68, December 2014. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Selic:1999:APC

[Sel99] Brian Selic. Architectural patterns for complex realtime systems (keynote address) (abstract only). ACM SIGADA Ada Letters, 19(3): 1, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Schonberg:1982:EMH

[SF82] E. Schonberg and G. A.

Fisher. An efficient method for handling operator overloading in Ada. In ACM [ACM82], pages 107–111. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Sward:2006:DSC

[SG06] Ricky E. Sward and Mark Gerken. Developing safety critical software for an unmanned aerial vehicle situational awareness tool. ACM SIGADA Ada Letters, 26(3): 45–50, December 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sterne:1989:SGN

[SGJP89] D. Sterne, A. Glendening, B. Jachowski, and G. Pretti.

A simplified graphic notation for Ada programs. ACM SIGADA Ada Letters, 9(6):108–118, September/October 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Saeed:1992:ICM

[SGS92] Faisel Saeed, K. M. George, and M. H. Samadzadeh. Implementation of classical mutual exclusion algorithms in Ada. ACM SIGADA Ada Letters, 12(1):73–84, January/February 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Strohmeier:1990:IBC

[SGW90a] Alfred Strohmeier, Christian Genillard, and Mats Weber. Implementation of 8-bit coded character sets in ADA. ACM SIGADA Ada Letters, 10(6):47–60, July/August 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Strohmeier:1990:OCS

[SGW90b] Alfred Strohmeier, Christian Genillard, and Mats Weber. Ordering of characters and strings. ACM SIG-ADA Ada Letters, 10(7):70– 84, September/October 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Shapiro:1993:ADA

[Sha93] Michael D. Shapiro. Another D 1 acronym. ACM SIG-ADA Ada Letters, 13(5):20– 21, September/October 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sherman:1980:ACG

[SHLR80] Mark Sherman, Andy Hisgen, David Alex Lamb, and Jonathan Rosenberg. An Ada code generator for VAX 11/780 with Unix. In ACM [ACM80], pages 91– 1?? CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Shore:1987:DES

[Sho87] R. W. Shore. Discrete-event simulation in Ada: Concepts. ACM SIGADA Ada Letters, 7(5):105–112, September/October 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sherman:1982:MPA

[SHR82] M. Sherman, A. Hisgen, and J. Rosenberg. A methodology for programming abstract data types in Ada. In ACM [ACM82], pages 66–75. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Singhoff:2023:AWS

[SHT^{+}23] Frank Singhoff, Jerome Hugues, Hai Nam Tran, Gianluca Bardaro, Dominique Blouin, Marco Bozzano, Patrick Denzler, Pierre Dissaux, Eric Senn, Xiong Xu, and Zhibin Yang. ADEPT 2022 workshop: a summary of strengths and weaknesses of the AADL ecosystem. ACM SIGADA Ada Letters, 43(1):37–41, June 2023. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3631483.3631485.

Shumate:1987:ECS

[Shu87] Ken Shumate. An example case study on Ada tasking. ACM SIGADA Ada Letters, 7(7):33–54, November/December 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Shumate:1991:SAO

[Shu91] Ken Shumate. Structured analysis and object-oriented design are compatible. ACM SIGADA Ada Letters, 11(4): 78–90, May/June 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Shumate:1993:BSO

[Shu93] Ken Shumate. BATCES solution #1: an objectoriented design from functional requirements analysis. ACM SIGADA Ada Letters, 13(6):133–161, November/December 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Siebert:2021:FST

[Sie21] Dr. Fridtjof Siebert. Fuzion — safety through simplicity. ACM SIGADA Ada Letters, 41(1):83–86, June 2021. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3570315.3570323.

Silberberg:1998:APS

[Sil98] David Silberberg. Applying the Personal Software Process (PSP)^{sm} with Ada. ACM SIGADA Ada Letters, 18(6):219–228, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Simpson:1982:ACF

[Sim82] R. T. Simpson. The ALS Ada compiler front end architecture. In ACM [ACM82], pages 98–106. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Singhoff:2007:MRT

[Sin07] Frank Singhoff. MP1: real time scheduling theory and its use with Ada. ACM SIGADA Ada Letters, 27(3): 8, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Solsi:1991:SYC

[SJ91] Swathi C. Solsi and Edward L. Jones. Simple yet complete heuristics for transforming data flow diagrams into Booch style diagrams. ACM SIGADA Ada Letters, 11(2):115–127, March/April 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sole:2022:CSA

[SK22] Marc Solé and Leonidas Kosmidis. Compiler support for an AI-oriented SIMD extension of a space processor. ACM SIGADA Ada Letters, 42(1):95–99, June 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3577949.3577968.

Slater:1995:OGP

[Sla95] Paul Slater. Output from generic packages. ACM SIG-ADA Ada Letters, 15(3):76– 79, May/June 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Singhoff:2004:CFR

[SLNM04] F. Singhoff, J. Legrand, L. Nana, and L. Marcé.

Cheddar: a flexible real time scheduling framework. ACM SIGADA Ada Letters, 24(4): 1–8, December 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Singhoff:2005:SMR

[SLNM05] F. Singhoff, J. Legrand, L. Nana, and L. Marcé. Scheduling and memory requirements analysis with AADL. ACM SIGADA Ada Letters, 25(4):1–10, December 2005. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sterrett:1992:PMA

[SM92] Anthony Sterrett and Marvin Minei. Performance measures of the Ada Rendezvous. ACM SIGADA Ada Letters, 12(2): 97–101, March/April 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Smart:2009:LAB

[Sma09] Jay C. Smart. A look at Ada from both sides now (a government, and a defense contractor perspective). ACM SIGADA Ada Letters, 29 (3):119–120, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Smith:1984:ASA

[Smi84]  David A. Smith.  ANSI

standard Ada — quick reference sheet. ACM SIG-ADA Ada Letters, 4(1):61– 66, July/August 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Smith:1997:W

[Smi97] D. Douglas Smith. WebAda. ACM SIGADA Ada Letters, 17(3):30–35, May/June 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Smith:2004:MEA

[Smi04] Geoff T. Smith. Measuring the effectiveness of ACATS. ACM SIGADA Ada Letters, 24(4):9–12, December 2004. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Shumate:1988:TAP

[SN88a] Ken Shumate and Kjell Nielsen. A taxonomy of Ada packages. ACM SIG-ADA Ada Letters, 8(2):55– 76, March/April 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sumate:1988:TAP

[SN88b] Ken Sumate and Kjell Nielsen. A taxonomy of Ada packages. ACM SIG-ADA Ada Letters, 8(2):55– 76, March/April 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Schilling:1994:ACR

[SN94] Jonathan L. Schilling and Johan Olmütz Nielsen. Automatic compiler recognition of monitor tasks. ACM SIG-ADA Ada Letters, 14(3):91– 104, May/June 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Soricone:2004:CAG

[SN04] Robert Soricone and Melvin Neville. Comparative analysis of genetic algorithm implementations. ACM SIG-ADA Ada Letters, 24(4):35– 38, December 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Snyder:1991:UAP

[Sny91] C. R. Snyder. Using Ada for PC-based software development. In ACM [ACM91b], pages 1–9. ISBN 0-89791-393-0. LCCN ????

SPSI:1988:NAC

[Sof88] Software Productivity Solutions, Inc. Naval Avionics Center Ada-Based Design Languages Workshop summary of events. ACM SIG-ADA Ada Letters, 8(4):104– 118, July/August 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Solderitsch:1991:LRS

[Sol91a] James Solderitsch. Library and representation subgroup. ACM SIGADA Ada Letters, 11(3):3–??, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Solderitsch:1991:WGR

[Sol91b] James Solderitsch. Working group report library and representation subgroup of methods and tools for design, specification, and reuse. ACM SIGADA Ada Letters, 11(3):3–7, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sotirovski:2006:THD

[Sot06] Drasko Sotirovski. Time horizon in distributed object societies. ACM SIG-ADA Ada Letters, 26(3):71– 74, December 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Singhoff:2007:AMA

[SP07] Frank Singhoff and Alain Plantec. AADL modeling and analysis of hierarchical schedulers. ACM SIG-ADA Ada Letters, 27(3):41– 50, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Schonberg:2012:ISD

[SP12] Edmond Schonberg and Vincent Pucci. Implementation of a simple dimensionality checking system in Ada 2012. ACM SIGADA Ada Letters, 32(3):35–42, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Spicer:2000:SEL

[Spi00] Kelly L. Spicer. A successful example of a layeredarchitecture based embedded development with Ada 83 for standard-missile control. ACM SIGADA Ada Letters, 20(4):50–63, December 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/dec2000/spicer-

paper.pdf. Special Issue: Presentations from SIGAda 2000.

SPS:1988:NAC

[SPS88] SPS, Inc. Naval Avionics Center Ada-Based Design Languages Workshop summary of events. ACM SIG-ADA Ada Letters, 8(4):103– 118, July/August 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Spurrier:1986:BAP

[Spu86] Tom Spurrier. Biography of an Ada project. ACM SIG-ADA Ada Letters, 6(1):49– 54, January/February 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Squire:1986:PCL

[Squ86] Jon Squire. PIWG chairperson’s letter. ACM SIG-ADA Ada Letters, 6(2):93, March/April 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Squire:1991:PSG

[Squ91a] Jon S. Squire. Proposed standard for a generic package of complex elementary functions (Ada). ACM SIG-ADA Ada Letters, 11(7): 140–165, Fall 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Squire:1991:RPS

[Squ91b] Jon S. Squire. Rationale for the proposed standard for a generic package of complex elementary functions (Ada). ACM SIGADA Ada Letters, 11(7):166–179, Fall 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Squire:1991:TVG

[Squ91c] Jon S. Squire. Towards validation of generic elementary functions and other standard Ada numerics packages. ACM SIGADA Ada Letters, 11(7):217–243, Fall 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Saez:2013:AMM

[SRC13a] Sergio Sáez, Jorge Real, and Alfons Crespo. Adding multiprocessor and mode change support to the Ada realtime framework. ACM SIG-ADA Ada Letters, 33(1): 116–127, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Saez:2013:DAS

[SRC13b] Sergio Sáez, Jorge Real, and Alfons Crespo. Deferred and atomic setting of scheduling attributes for Ada. ACM SIGADA Ada Letters, 33(2): 97–108, August 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Saez:2015:ITE

[SRC15] Sergio Sáez, Jorge Real, and Alfons Crespo. Implementation of timing-event affinities in Ada/Linux. ACM SIGADA Ada Letters, 35(1): 80–92, April 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Srivastava:2006:AIG

[Sri06a] Alok Srivastava. Ada issue 00354: group executiontime budgets. ACM SIG-ADA Ada Letters, 26(2):38– 47, August 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Srivastava:2006:AIP

[Sri06b] Alok Srivastava. Ada issue 00355: priority specific dispatching including round robin. ACM SIG-ADA Ada Letters, 26(2): 48–59, August 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Srivastava:2006:AIS

[Sri06c] Alok Srivastava. Ada issue 00357: support for deadlines and earliest deadline first scheduling. ACM SIG-ADA Ada Letters, 26(2):60– 68, August 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Srivastava:2006:AIR

[Sri06d] Alok Srivastava. Ada issue 00394: redundant restriction identifiers and completing Ravenscar definition. ACM SIGADA Ada Letters, 26(2):69–74, August 2006. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Srivastava:2006:EP

[Sri06e]  Alok Srivastava.  Edito-

rial policy. ACM SIG-ADA Ada Letters, 26(1): 2–3, April 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Srivastava:2006:ED

[Sri06f]  Alok Srivastava.  From

the Editor’s desk. ACM SIGADA Ada Letters, 26 (1):1, April 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sankar:1985:IA

[SRN85] Sriram Sankar, David Rosenblum, and Randall Neff. An implementation of Anna. ACM SIGADA Ada Letters, 5(2):285–296, September/October 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Schonberg:1985:HPA

[SS85] Edith Schonberg and Edmond Schonberg. Highly parallel Ada — Ada on an Ultracomputer. ACM SIG-ADA Ada Letters, 5(2):58– 71, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Seidewitz:1987:TGO

[SS87] Ed Seidewitz and Mike Stark. Towards a general objectoriented software development methodology. ACM SIGADA Ada Letters, 7(4): 54–67, July/August 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Schiper:1989:TUC

[SS89] Andre Schiper and Roland Simon. Traps using the COUNT attribute in the readers-writers problem. ACM SIGADA Ada Letters, 9(5): 123–128, July/August 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Seidewitz:1991:OAP

[SS91] E. Seidewitz and M. Stark. An object-oriented approach to parameterized software in Ada. In ACM [ACM91b], pages 62–76. ISBN 0-89791-393-0. LCCN ????

Smith:1994:MTS

[SS94] Milton Smith and Jag Sodhi. Marching towards a Software Reuse Future. ACM SIGADA Ada Letters, 14(6):

62–72, November/December 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Suchan:1997:UAT

[SS97] W. Suchan and T. L. Smith.

Using Ada 95 as a tool to teach problem solving to non-CS majors. In ACM [ACM97], pages 31–36. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Schranz:2020:MSI

[SS20] M. Schranz and M. Sende.

Modeling swarm intelligence algorithms for CPS swarms. ACM SIGADA Ada Letters, 40(1):64–73, October 2020. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3431235.3431240.

Schranz:2020:MCS

[SSB^{+}20] M. Schranz, M. Sende, A. Bagnato, E. Brosse, and A. Eckel. Modeling CPS swarms: an automotive use case. ACM SIG-ADA Ada Letters, 40(1):60– 63, October 2020. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3431235.3431239.

StDennis:1986:MCR

[SSFO86] R. St. Dennis, P. Stachour, E. Frankowski, and E. Onuegbe. Measurable characteristics of reusable Ada software. ACM SIG-ADA Ada Letters, 6(2):41– 50, March/April 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sousa:2022:MNF

[SSGH^{+}22] Rita Sousa, Eudald Sabate, Marco Gonzalez-Hierro, António Barros, Cristina Zubia, Luis Miguel Pinho, and Elli Kartsakli. Managing non-functional requirements in an ELASTIC edge-cloud continuum. ACM SIG-ADA Ada Letters, 42(2):114– 118, December 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3591335.3591352.

Schill:1985:CCC

[SSJ85] John Schill, Roger Smeaton, and Richard Jackman. The conversion of command & control software to Ada: Experiences and lessons learned. ACM SIGADA Ada Letters, 4(4):38–48, January/ February 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Standish:1983:IAA

[Sta83] Thomas A. Standish. Interactive Ada in the Arcturus environment. ACM SIG-ADA Ada Letters, 3(1):23– 36, July/August 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Stevenson:1980:ATA

[Ste80] David R. Stevenson. Algorithms for translating Ada tasking. In ACM [ACM80], pages 166–175. CODEN SIN-ODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Steele:2012:PLL

[Ste12] Guy L. Steele, Jr. Programming language life cycles. ACM SIGADA Ada Letters, 32(3):95–96, December 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Seidowitz:1998:PAS

[STF98] Ed Seidowitz, William Thomas, and Michael Feldman, editors. Proceedings: ACM SIG-Ada Annual International Conference (SIGAda ’98) (formerly TriAda), November 8–12, 1998, Omni Shoreham Hotel, Washington, DC, USA, volume 18(6) of ACM SIGADA Ada Letters. ACM Press, New York, NY, USA, 1998. ISBN 1-58113-033-3. Three papers in this volume were incorrectly printed, and a corrected supplement was issued in December 1998. Papers in that supplement have page numbers ending in ‘A’.

Spicer:1991:MMA

[SU91] Kelly L. Spicer and David A. Umphress. A method for mapping an analysis to a reusable design. ACM SIG-ADA Ada Letters, 11(9): 67–82, November/December 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sumate:1987:ECS

[Sum87] Ken Sumate. An example case study on Ada tasking. ACM SIGADA Ada Letters, 7(7):33–54, November/December 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Smith:1999:DPI

[SV99] Gary W. Smith and Richard A. Volz. Distributed programming with intermediate IDL. ACM SIG-ADA Ada Letters, 19(2): 90–95, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Szabo:2014:MEL

[SVK^{+}14] Tamás Szabó, Markus Voelter, Bernd Kolb, Daniel Ratiu, and Bernhard Schaetz. mbeddr: extensible languages for embedded software development. ACM SIG-ADA Ada Letters, 34(3):13– 16, December 2014. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sarkar:1987:IAF

[SW87] J. P. Sarkar and T. T. Wong.

Impacts of Ada features on real-time performance. ACM SIGADA Ada Letters, 7(6): 88–92, Fall 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sward:2007:SEA

[Swa07a] Ricky E. Sward. SP2: exposing Ada Web services using a service-oriented architecture (SOA). ACM SIG-ADA Ada Letters, 27(3): 4, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sward:2007:UAS

[Swa07b]  Ricky E. Sward.  Using

Ada in a service-oriented architecture. ACM SIG-ADA Ada Letters, 27(3):63– 68, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sward:2009:GIU

[Swa09a] Ricky E. Sward. Georegistration of imagery from unmanned aircraft systems using Ada. ACM SIG-ADA Ada Letters, 29(3):121– 126, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sward:2009:SOA

[Swa09b]  Ricky E. Sward.  Ser-

vice oriented architecture (SOA) concepts and implementations. ACM SIG-ADA Ada Letters, 29(3):15– 16, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sward:2010:RFP

[Swa10] Ricky E. Sward. The rise, fall and persistence of Ada. ACM SIGADA Ada Letters, 30(3): 71–74, December 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sammet:1982:PAD

[SWR82] Jean E. Sammet, Douglas W. Waugh, and Robert W. Reiter, Jr. PDL/Ada — a design language based on Ada. ACM SIGADA Ada Letters, 2(3):19–31, November/December 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Syiek:1995:CVA

[Syi95] David Syiek. C vs. Ada:

arguing performance religion. ACM SIGADA Ada Letters, 15(6):67–69, November/December 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Strom:1985:VAP

[SYW85] Rob Strom, Shaula Yemini, and Peter Wegner. Viewing Ada from a process model perspective. ACM SIGADA Ada Letters, 5(2):241–254, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Taft:1982:DIR

[Taf82] S. T. Taft. DIANA as an internal representation in an Ada-In-Ada compiler. In ACM [ACM82], pages 261– 265. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Taft:1991:BDT

[Taf91a] Tucker Taft. Building, debugging and testing realtime and distributed systems. ACM SIGADA Ada Letters, 11(3):19–??, Spring 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taft:1991:SWG

[Taf91b] Tucker Taft. SETA1 working group on building, debugging and testing real-time and distributed systems. ACM SIGADA Ada Letters, 11(3): 19–27, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taft:1997:SRN

[Taf97] Tucker Taft. Selected rationale for NRC recommendations. ACM SIGADA Ada Letters, 17(1):21–24, January/February 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taft:2001:EES

[Taf01a] S. Tucker Taft. Enhancing exception support in Ada 95: a workshop position paper. ACM SIGADA Ada Letters, 21(3):31–32, September 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taft:2001:KAF

[Taf01b] S. Tucker Taft. Keynote address: fixing software before it breaks. ACM SIG-ADA Ada Letters, 21(4):97– 98, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taft:2001:UAC

[Taf01c] S. Tucker Taft. Using Ada 95 in a compiler course. ACM SIGADA Ada Letters, 21(4): 79–80, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taft:2006:WYS

[Taf06] Tucker Taft. Why you should be using Ada 2005 now! ACM SIGADA Ada Letters, 26(3):75, December 2006. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taft:2011:EPP

[Taf11] S. Tucker Taft. Experimenting with ParaSail: parallel specification and implementation language. ACM SIG-ADA Ada Letters, 31(3):11– 12, December 2011. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taft:2012:TMP

[Taf12] S. Tucker Taft. Tutorial:

multicore programming using divide-and-conquer and work stealing. ACM SIG-ADA Ada Letters, 32(3):13– 14, December 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). HILT ’12 conference proceedings.

Taft:2013:BSD

[Taf13a] S. Tucker Taft. Bringing safe, dynamic parallel programming to the SPARK verifiable subset of Ada. ACM SIGADA Ada Letters, 33(3): 37–40, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taft:2013:TPS

[Taf13b] S. Tucker Taft. Tutorial: proving safety of parallel /multi-threaded programs. ACM SIGADA Ada Letters, 33(3):1–2, December 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taft:2020:SFV

[Taf20] Tucker Taft. SPARK formal verification for security. ACM SIGADA Ada Letters, 39(1): 83–99, January 2020. ISSN 0736-721X. URL https:

//dl.acm.org/doi/abs/10.
1145/3379106.3379117.

Taft:2021:LMA

[Taf21] S. Tucker Taft. A layered mapping of Ada 202X to OpenMP. ACM SIG-ADA Ada Letters, 40(2): 55–58, April 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463478.3463482.

Taft:2022:RPM

[Taf22a] S. Tucker Taft. Rigorous pattern matching as a language feature. ACM SIG-ADA Ada Letters, 42(2):69– 74, December 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3591335.3591342.

Taft:2022:WSS

[Taf22b] S. Tucker Taft. A workstealing scheduler for Ada 2022, in Ada. ACM SIG-ADA Ada Letters, 42(1): 80, June 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3577949.3577964.

Tai:1986:GND

[Tai86] Kuo-Chung Tai. A graphical notation for describing executions of concurrent Ada programs. ACM SIGADA Ada Letters, 6(1):94–103, January/February 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tanaka:1991:UAN

[Tan91a] Kiyoshi Tanaka. Using Ada at NTT. ACM SIGADA Ada Letters, 11(1):92–95, January/February 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tang:1991:PGE

[Tan91b] Ping Tak Peter Tang. A portable generic elementary function package in Ada and an accurate test suite. ACM SIGADA Ada Letters, 11(7): 180–216, Fall 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

TFMSDSB:1988:RDS

[Tas88] Task Force on Military Software Defense Science Board. Report of the Defense Science Board Task Force on Military Software. ACM SIG-ADA Ada Letters, 8(4):35– 46, July/August 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tokar:2002:SSS

[TB02] Joyce Tokar and Ben Brosgol. Session summary: summary and plans for next IRTAW. ACM SIGADA Ada Letters, 22(4):131, December 2002. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tzruya:1998:PID

[TBA98] Yoav Tzruya and Mordechai Ben-Ari. A portable implementation of the Distributed Systems Annex in Java. ACM SIGADA Ada Letters, 18(6):204–211, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taft:2022:DPM

[TBD22] S. Tucker Taft, Stephen Baird, and Claire Dross. Defining a pattern matching language feature for Ada. ACM SIGADA Ada Letters, 42(1):79, June 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3577949.3577963.

Tetewsky:1988:MAE

[TCRW88] Avram Tetewsky, Ann Clough, Roger Racine, and R. Whittredge. Mapping Ada onto embedded systems: Memory constraints. ACM SIGADA Ada Letters, 8(5):101–109, September/October 1988. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tokar:2003:SSN

[TD03] Joyce L. Tokar and Brian Dobbing. Session summary: new core language features. ACM SIGADA Ada Letters, 23(4):11–12, December 2003. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Temte:1984:OOD

[Tem84] Mark Temte. Object-oriented design and ballistics software. ACM SIGADA Ada Letters, 4(3):25–36, November/December 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ternes:1987:DSC

[Ter87] David Ternes. Development software configuration and integration in a large Ada project. In ACM [ACM87a], pages 65–74. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Texel:1986:CL

[Tex86] Putnam P. Texel. Chairperson’s letter. ACM SIG-ADA Ada Letters, 6(2):96– 99, March/April 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tijero:2009:EII

[TG09] Héctor Pérez Tijero and J. Javier Gutierrez. Experience in integrating interchangeable scheduling policies into a distribution middleware for Ada. ACM SIG-ADA Ada Letters, 29(3):73– 78, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tijero:2010:SRT

[TGH10] Héctor Pérez Tijero, J. Javier Gutiérrez, and Michael González Harbour. Support for a real-time transactional model in distributed Ada. ACM SIGADA Ada Letters, 30(1): 91–103, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tijero:2013:AEE

[TGH13] Héctor Pérez Tijero, J. Javier Gutiérrez, and Michael González Harbour. Adapting the endto-end flow model for distributed Ada to the Ravenscar profile. ACM SIG-ADA Ada Letters, 33(1): 53–63, April 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Thall:1982:KAL

[Tha82] R. M. Thall. The KAPSE for the Ada language system. In ACM [ACM82], pages 31–47. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Theriault:1990:STT

[The90] Ronald J. Theriault. A scheme for the translation of the Tucker Taft Select-And statement into Standard ANSI Ada. ACM SIG-ADA Ada Letters, 10(6):110– 113, July/August 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tichy:1982:ADA

[Tic82] W. F. Tichy. Adabase: a database for Ada programs. In ACM [ACM82], pages 57–65. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Tindell:1990:DCR

[Tin90]  Ken Tindell.  Dynamic

code replacement and Ada. ACM SIGADA Ada Letters, 10(7):47–54, September/October 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tischler:1983:NSA

[Tis83]  Ron Tischler.  Note on

scanning Ada. ACM SIG-ADA Ada Letters, 3(1):36– ??, July/August 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taft:2014:SPP

[TMPM14] S. Tucker Taft, Brad Moore, Luís Miguel Pinho, and Stephen Michell. Safe parallel programming in Ada with language extensions. ACM SIGADA Ada Letters, 34(3): 87–96, December 2014. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taft:2016:RPC

[TMPM16] Tucker Taft, Brad Moore, Luis Miguel Pinho, and Stephen Michell. Reduction of parallel computation in the parallel model for Ada. ACM SIGADA Ada Letters, 36(1): 9–24, June 2016. CODEN AALEE5. ISSN 0736-721X.

Tojo:2005:TDP

[TNGC05] Yasushi Tojo, Sinsuke Nara, Yuichi Goto, and Jingde Cheng. Tasking deadlocks in programs with the full Ada 95. ACM SIG-ADA Ada Letters, 25(1): 48–56, March 2005. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Toal:1996:UAC

[Toa96] Raymond J. Toal. Using Ada and C++ in computer science education. ACM SIG-ADA Ada Letters, 16(1):58– 69, January/February 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tokar:2003:STP

[Tok03] Joyce L. Tokar. Space & time partitioning with ARINC 653 and pragma profile. ACM SIGADA Ada Letters, 23(4): 52–54, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tokar:2015:UII

[Tok15] Joyce L. Tokar. Update of ISO/IEC technical reports on the use of the Ada programming language in high integrity systems. ACM SIGADA Ada Letters, 35(1): 93–94, April 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tokar:2016:CAO

[Tok16] Joyce L. Tokar. A comparison of avionics open system architectures. ACM SIGADA Ada Letters, 36(2):22–26, December 2016. CODEN AALEE5. ISSN 0736-721X.

Tombs:1997:UCN

[Tom97] D. J. Tombs. Using compliance notation to verify Ada tasking. ACM SIG-ADA Ada Letters, 17(5):83– 87, September/October 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tonndorf:1999:ACA

[Ton99]  Michael Tonndorf.  Ada

conformity assessments: a model for other programming languages? ACM SIG-ADA Ada Letters, 19(3):89– 99, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Toole:1991:AAM

[Too91] Betty Alexandra Toole. Ada, an analyst and a metaphysician. ACM SIGADA Ada Letters, 11(2):60–71, March/April 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tardieu:1998:BFT

[TP98] Samuel Tardieu and Laurent Pautet. Building fault tolerant distributed systems using IP multicast. ACM SIGADA Ada Letters, 18(6):45–51, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tardieu:2009:CAO

[TP09] Samuel Tardieu and Alexis Polti. Complementing Ada with other programming languages. ACM SIGADA Ada Letters, 29(3):105–114, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tijero:2021:FST

[TPG21] Héctor Pérez Tijero, Diego García Prieto, and J. Javier Gutiérrez. First steps towards an IEEE 802.1AS clock for EDF scheduling in distributed real-time systems. ACM SIGADA Ada Letters, 41 (1):87–91, June 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3570315.3570324.

Tetewsky:1987:ACS

[TR87] A. Tetewsky and R. Racine. Ada compiler selection for embedded targets. ACM SIGADA Ada Letters, 7(5):51–62, September/October 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tracz:1989:PCS

[Tra89] Will Tracz. Parameterization: a case study. ACM SIGADA Ada Letters, 9(4): 92–102, May/June 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Trono:2006:OTL

[Tro06] John A. Trono. Optimal table lookup for reserved words in Ada. ACM SIG-ADA Ada Letters, 26(1): 25–30, April 2006. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Trono:2012:UMW

[Tro12] John A. Trono. Updated MPHF weights for Ada 2012. ACM SIGADA Ada Letters, 32(1):9–12, April 2012. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Troiani:2020:ECR

[Tro20]  Mario Troiani.  Ensur-

ing cyber resilience through entropy-augmented replication. ACM SIGADA Ada Letters, 39(1):72, January 2020. ISSN 0736-721X.

URL https://dl.acm.org/
doi/abs/10.1145/3379106.
3379116.

Taft:2016:BTM

[TRT16] S. Tucker Taft, Elie Richa, and Andres Toom. Building trust in a model-based automatic code generator. ACM SIGADA Ada Letters, 36(2): 54–57, December 2016. CO-DEN AALEE5. ISSN 0736-721X.

Trub:1995:AUD

[Trü95] Ann Trüb. Ada used to develop a global positioning system for future spacecraft. ACM SIGADA Ada Letters, 15(4):22, July/August 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tomar:2020:MTV

[TS20] Ravi Tomar and Sarishma.

Maintaining trust in VANETs using blockchain. ACM SIG-ADA Ada Letters, 40(1):91– 96, October 2020. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3431235.3431244.

Thirion:2002:CPC

[TT02] Bernard Thirion and Laurent Thiry. Concurrent programming for the control of hexapod walking. ACM SIGADA Ada Letters, 22(1): 17–28, March 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taffs:1985:ACG

[TTRH85] D. A. Taffs, M. W. Taffs, J. C. Rienzo, and T. R. Hampson. The ALS Ada compiler global optimizer. ACM SIGADA Ada Letters, 5(2):355–366, September/October 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Tucker:1997:DHO

[Tuc97] K. Tucker. Debugging highly optimized Ada with code motion (DHACM). In ACM [ACM97], pages 197–204. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Toetenel:1988:ATC

[Tv88] W. J. Toetenel and J. van Katwijk. Asynchronous transfer of control in Ada.

ACM SIGADA Ada Letters, 8(7):65–79, Fall 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ujvary:1997:BHR

[UKDH97] Brian G. Ujvary, Nick I.

Kamenoff, and Jorge L. Diaz-Herrera. Benchmarking of hard real-time distributed systems with Ada 95. ACM SIGADA Ada Letters, 17(5):88–92, September/October 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Uruena:2007:INA

[UPRZ07] Santiago Urueña, José Pulido, José Redondo, and Juan Zamorano. Implementing the new Ada 2005 real-time features on a bare board kernel. ACM SIGADA Ada Letters, 27(2):61–66, August 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Uruena:2007:BHI

[UZ07] Santiago Urueña and Juan Zamorano. Building highintegrity distributed systems with Ravenscar restrictions. ACM SIGADA Ada Letters, 27(2):29–36, August 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

VanNeste:1986:ACS

[Van86] Karl F. VanNeste. Ada coding standards and conventions. ACM SIGADA Ada Letters, 6(1):41–48, January/February 1986. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

VanScoy:1990:CIW

[Van90] Roger Van Scoy. Communication issues working group. ACM SIGADA Ada Letters, 10(4):97–113, Spring 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

VanVlierberghe:1994:MMA

[Van94] Stef Van Vlierberghe. Memory management in Ada83 and Ada9X. ACM SIG-ADA Ada Letters, 14(4):43– 57, July/August 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vardanega:2001:CE

[Var01a] Tullio Vardanega. A case for exceptions. ACM SIG-ADA Ada Letters, 21(3):26– 30, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vardanega:2001:OOE

[Var01b] Tullio Vardanega. Object orientation and exception handling for Ada. ACM SIG-ADA Ada Letters, 21(3):11– 12, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vardanego:2001:CE

[Var01c]  T. Vardanego.  A case

for exceptions. ACM SIG-ADA Ada Letters, 21(3):26– 30, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vardanega:2003:RDP

[Var03] Tullio Vardanega. Ravenscar design patterns?: reflections on use of the Ravenscar profile. ACM SIG-ADA Ada Letters, 23(4):65– 73, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vasilescu:1991:UAR

[Vas91] E. Vasilescu. Using Ada for rapid prototyping of database applications. In ACM [ACM91b], pages 40– 49. ISBN 0-89791-393-0. LCCN ????

Vaughn:1998:ARY

[Vau98] Rayford B. Vaughn, Jr. The Ada recommendation — a year later. ACM SIG-ADA Ada Letters, 18(4): 95–100, July 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

VanScoy:1989:OD

[VBF89] Roger Van Scoy, Judy Bamberger, and Robert Firth. An overview of DARK. ACM SIGADA Ada Letters, 9(7):91–101, November/December 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

VanScoy:1990:DVD

[VBF90] Roger Van Scoy, Judy Bamberger, and Robert Firth. A Detailed view of DARK. ACM SIGADA Ada Letters, 10(6):68–83, July/August 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vardanega:2001:URP

[VC01] Tullio Vardanega and Gert Caspersen. Using the Ravenscar profile for space applications: the OBOSS case. ACM SIGADA Ada Letters, 21(1): 96–104, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

vanderLinden:1984:WDS

[vdL84] Peter van der Linden. Writing diagnostic software in Ada. ACM SIGADA Ada Letters, 4(2):44–53, September/October 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

vanderLinden:1985:LFA

[vdL85] Peter van der Linden. Looking forward with Ada. ACM SIGADA Ada Letters, 5(1): 49–54, July/August 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Verun:1992:CAM

[VE92] Ufuk Verün and Tzilla Elrad. A critique of the Ada 9X mutual control mechanism (requeue) and an alternative mapping (onlywhen). ACM SIGADA Ada Letters, 12(6):75–80, November/December 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Venet:2008:PAF

[Ven08] Arnaud Venet. A practical approach to formal software verification by static analysis. ACM SIGADA Ada Letters, 28(1):92–95, April 2008. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Verschelde:2021:PSO

[Ver21]  Jan Verschelde.  Parallel

software to offset the cost of higher precision. ACM SIGADA Ada Letters, 40(2): 59–64, April 2021. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463478.3463483.

Verschelde:2022:EAS

[Ver22]  Jan Verschelde.  Export-

ing Ada software to Python and Julia. ACM SIG-ADA Ada Letters, 42(1): 76–78, June 2022. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3577949.3577961.

Vestal:1989:MCP

[Ves89] Steve Vestal. Mixing coroutines and processes in an Ada tasking implementation. ACM SIGADA Ada Letters, 9(2):90–101, March/April 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vestal:1990:LBa

[Ves90a] Steve Vestal. Linear benchmarks. ACM SIGADA Ada Letters, 10(8):145–155, November/December 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vestal:1990:LBb

[Ves90b] Steve Vestal. Linear benchmarks. ACM SIGADA Ada Letters, 10(9):145–155, Fall 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vestal:1997:RMD

[VGD^{+}97] Steve Vestal, Laurent Guerby, Robert Dewar, David Mc-Connell, and Bruce Lewis.

Reimplementing a multiprocess distributed paradigm for real-time systems in Ada 95. ACM SIGADA Ada Letters, 17(5):93–99, September/October 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Valls:2020:SBV

[VGGS20] Joan J. Valls, Miguel García-Gordillo, and Sergio Sáez. Scenario-based validation & verification: The ENABLE-S3 approach. ACM SIG-ADA Ada Letters, 40(1):79– 84, October 2020. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3431235.3431242.

vonHenke:1985:SSA

[vHLKBO85] Friedrich W. von Henke, David Luckham, Bernd Krieg-Brueckner, and Olaf Owe. Semantic specification of Ada packages. ACM SIGADA Ada Letters, 5(2):185–196, September/October 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Vardanega:2010:SSL

[VHP10] Tullio Vardanega, Michael González Harbour, and Luís Miguel Pinho. Session summary: language and distribution issues. ACM SIGADA Ada Letters, 30(1):152–161, April 2010. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vladavsky:1993:AAS

[Vla93] Luba Vladavsky. Activities of the Ada semantic interface specification (ASIS) working group (ASISWG). ACM SIG-ADA Ada Letters, 13(3):39– 41, May/June 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vladavsky:1994:AAS

[Vla94] Luba Vladavsky. Activities of the Ada Semantic Interface Specification (ASIS) Working Group (ASISWG). ACM SIGADA Ada Letters, 14 (2):54–57, March/April 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Volz:1985:SPD

[VMNM85] Richard A. Volz, Trevor N. Mudge, Arch W. Naylor, and John H. Mayer. Some problems in distributing Real-Time Ada programs across machines. ACM SIG-ADA Ada Letters, 5(2):72– 84, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Voketaitis:1992:PRR

[Vok92] Arnold M. Voketaitis, Jr.

A portable and reusable RDBMS interface for Ada. ACM SIGADA Ada Letters, 12(5):64–76, September/October 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Volz:1987:DAE

[Vol87] Richard A. Volz. Distributed Ada execution: a definitional void. ACM SIGADA Ada Letters, 7(6):70–72, Fall 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Volz:1990:VNU

[Vol90] Richard A. Volz. Virtual nodes and units of distribution for distributed Ada. ACM SIGADA Ada Letters, 10(4):85–96, Spring 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vardanega:2003:SSF

[VP03] Tullio Vardanega and Luís Miguel Pinho. Session summary: the future of IRTAW. ACM SIGADA Ada Letters, 23(4): 96, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vardanega:2007:LII

[VR07] Tullio Vardanega and Josée F. Ruiz. Language issues: Introduction. ACM SIG-ADA Ada Letters, 27(2): 15–17, August 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vardanega:2016:SSA

[VR16] Tullio Vardanega and Pat Rogers. Session summary: Ada language profiles. ACM SIGADA Ada Letters, 36(1): 98–100, June 2016. CODEN AALEE5. ISSN 0736-721X.

Villaescusa:2021:QPM

[VRH21] David García Villaescusa, Mario Aldea Rivas, and Michael González Harbour. Queuing ports for mesh based many-core processors. ACM SIGADA Ada Letters, 41 (2):66–70, December 2021. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3530801.3530804.

Vardanega:2013:SSI

[VW13] Tullio Vardanega and Rod White. Session summary: improvements to Ada. ACM SIGADA Ada Letters, 33(2): 126–130, August 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Vardanega:2018:SSL

[VW18] Tullio Vardanega and Andy Andy Wellings. Session summary: Language issues. ACM SIG-ADA Ada Letters, 38(1):74– 76, June 2018. CODEN AALEE5. ISSN 0736-721X.

Ward:2002:LIC

[WA02] M. Ward and N. C. Audsley. Language issues of compiling Ada to hardware. ACM SIGADA Ada Letters, 22(4): 85–94, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ward:2007:SSB

[WA07] M. Ward and N. C. Audsley.

Suggestions for stream based parallel systems in Ada. ACM SIGADA Ada Letters, 27(2):82–87, August 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wade:1992:DRC

[Wad92] David M. Wade. Designing for reuse: a case study. ACM SIGADA Ada Letters, 12(3): 92–98, May/June 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wagreich:1985:MEE

[Wag85] Roberta G. Wagreich. Methodologies and environments for embedded systems lifecycle support. ACM SIGADA Ada Letters, 4(5):105–110, March/April 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Future Ada Environment Workshop.

Wagner:2020:FSA

[Wag20] Lucas Wagner. Formal specification and analysis of requirements using SpeAR. ACM SIGADA Ada Letters, 39(1):20–34, January 2020. ISSN 0736-721X.

URL https://dl.acm.org/
doi/abs/10.1145/3379106.
3379110.

Wainwright:1998:AEW

[Wai98] Ross H. Wainwright. An application engineering workbench for tailoring Ada flight components. ACM SIGADA Ada Letters, 18(6):165–174, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wai:2021:XA

[Wai21] Richard Wai. XERIS/APEX. ACM SIGADA Ada Letters, 40(2):65–69, April 2021. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3463478.3463484.

Walasek:1985:SLC

[Wal85a] Jan Walasek. Source listing with combs. ACM SIG-ADA Ada Letters, 4(6):32– 34, May/June 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wallis:1985:ALC

[Wal85b]  P. J. L. Wallis.  Auto-

matic language conversion and its place in the transition to Ada. ACM SIGADA Ada Letters, 5(2):275–284, September/October 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Ada in Use: Proceedings of the Ada International Conference, Paris, 14– 16 May, 1985, John G. P. Barnes and Gerald A. Fisher, Jr., eds.

Walters:1987:ESD

[Wal87] Michael D. Walters. Expert systems development in LISP and Ada. In ACM [ACM87a], pages 111–115. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Walters:1991:AOB

[Wal91] Neal L. Walters. An Ada object-based analysis and design approach. ACM SIG-ADA Ada Letters, 11(5):62– 78, July/August 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wallnau:1994:WSU

[Wal94] Kurt C. Wallnau. Workshop summary: user interface. ACM SIGADA Ada Letters, 14(Special Issue):99–103, Fall 1994. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wang:1990:UA

[Wan90] Y. E. Gail Wang. UNI-VERSAL FILE NAMES for Ada. ACM SIGADA Ada Letters, 10(1):111–117, January/February 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wang:1999:ISE

[Wan99]  Ming Wang.  Integrating

a software engineering approach into an Ada closed laboratory. ACM SIG-ADA Ada Letters, 19(3):163– 168, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Watson:1987:AM

[Wat87] S. E. Watson. Ada modules. ACM SIGADA Ada Letters, 7(4):79–84, July/ August 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Waugh:1983:ALP

[Wau83] Douglas W. Waugh. An Ada language programming course. ACM SIGADA Ada Letters, 2(5):34–41, March/April 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wu:1989:SCD

[WB89] Y. C. Wu and Ted P. Baker.

A source code documentation system for Ada. ACM SIG-ADA Ada Letters, 9(5):84– 88, July/August 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2007:BAA

[WB07a] A. J. Wellings and A. Burns.

Beyond Ada 2005: allocating tasks to processors in SMP systems. ACM SIG-ADA Ada Letters, 27(2):75– 81, August 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2007:FRT

[WB07b] A. J. Wellings and A. Burns.

A framework for real-time utilities for Ada 2005. ACM SIGADA Ada Letters, 27(2): 41–47, August 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2007:IOT

[WB07c] A. J. Wellings and A. Burns.

Integrating OOP and tasking: the missing requeue. ACM SIGADA Ada Letters, 27(2):23–28, August 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2010:GES

[WB10a] A. J. Wellings and A. Burns. Generalizing the EDF scheduling support in Ada 2005. ACM SIGADA Ada Letters, 30(1):116–124, April 2010. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2010:UDC

[WB10b] A. J. Wellings and A. Burns. User-defined clocks is it the right time now? ACM SIGADA Ada Letters, 30(1): 104–115, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2015:ITE

[WB15] A. J. Wellings and A. Burns. Interrupts, timing events and dispatching domains. ACM SIGADA Ada Letters, 35(1): 26–31, April 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2013:PSR

[WBCS13] A. J. Wellings, A. Burns, A. L. C. Cavalcanti, and N. K. Singh. Programming simple reactive systems in Ada: premature program termination. ACM SIGADA Ada Letters, 33(2):

75–86, August 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:1997:TTA

[WBP97] A. J. Wellings, A. Burns, and O. Pazy. Task termination and Ada 95. ACM SIGADA Ada Letters, 17(5):100–105, September/October 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Waligora:1997:IAO

[WBS97] Sharon Waligora, John Bailey, and Mike Stark. The impact of Ada and objectoriented design in NASA Goddard’s Flight Dynamics Division. ACM SIG-ADA Ada Letters, 17(3):67– 86, May/June 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Extensive study of a decade of large system software development in Ada, Fortran, C, and C++, with the finding that development is moving away from Ada to C and C++ on two main grounds: lack of adequate software development environments for Ada, and high cost (3 to 8 times per seat).

Wellings:2016:ISC

[WCB16] A. J. Wellings, V. Cholpanov, and A. Burns. Implementing safety-critical Java missions in Ada. ACM SIG-ADA Ada Letters, 36(1):51– 62, June 2016. CODEN AALEE5. ISSN 0736-721X.

Waterhouse:1993:RRT

[WD93] Daniel F. Waterhouse and Daniel L. Dyke. Rehost of a real-time interruptdriven simulation onto a DOS/PC/Ada environment using OOD. ACM SIG-ADA Ada Letters, 13(4):49– 62, July/August 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:1997:OOP

[WdlP97] Andy Wellings and Juan de la Puente. Object-oriented programming and real-time (session summary). ACM SIG-ADA Ada Letters, 17(5):16– 17, September/October 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Weatherly:2010:USA

[Wea10] Richard Weatherly. “unmanned systems and Ada”. ACM SIGADA Ada Letters, 30(3):35–36, December 2010. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Weber:1993:EOI

[Web93]  Mats Weber.  Elabora-

tion order issues in Ada 9X. ACM SIGADA Ada Letters, 13(1):63–75, January/February 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wegner:1982:AET

[Weg82] Peter Wegner. Ada education and technology transfer activities. ACM SIG-ADA Ada Letters, 2(2):51– 60, September/October 1982. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Weicker:1989:DBA

[Wei89] Reinhold P. Weicker. Dhrystone benchmark (Ada version 2): Rationale and measurement rules. ACM SIG-ADA Ada Letters, 9(5):60– 82, July/August 1989. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Weiderman:1990:HSB

[Wei90a] Nelson Weiderman. Hartstone: Synthetic benchmark requirements for hard realtime applications. ACM SIG-ADA Ada Letters, 10(3):126– 136, Winter 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Weidman:1990:MCA

[Wei90b] Henry Weidman. A method for converting abstract objects to discrete objects. ACM SIGADA Ada Letters, 10(2):52–63, March/April 1990. CODEN AALEE5.

ISSN 1094-3641 (print), 1557-9476 (electronic).

Weker:1990:CPP

[Wek90]  Mats Weker.  Comments

on the paper “Parameterization: a case study, by Will Tracz”. ACM SIG-ADA Ada Letters, 10(6):16– 17, July/August 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Welch:1985:STA

[Wel85] P. H. Welch. Structured tasking in Ada? ACM SIG-ADA Ada Letters, 5(1):17– 31, July/August 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:1990:RTR

[Wel90] Andy J. Wellings. Realtime requirements. ACM SIGADA Ada Letters, 10 (9):1–16, Fall 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:1991:SDS

[Wel91] A. J. Wellings. Support for distributed systems in Ada 9X. ACM SIGADA Ada Letters, 11(6):61–63, September/October 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Welch:1997:CCC

[Wel97a]  Lonnie R. Welch.  CO-

COON: Creator Of Concurrent Object OrieNted systems. ACM SIGADA Ada Letters, 17(6):32–38, November/December 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Welch:1997:PRE

[Wel97b] Lonnie R. Welch. PRISM:

a reverse engineering toolset. ACM SIGADA Ada Letters, 17(6):39–46, November/December 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:1999:NLF

[Wel99] Andy Wellings. New language features and other language issues (session summary). ACM SIGADA Ada Letters, 19(2):19–20, June 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2001:SFR

[Wel01] Andy Wellings. Status and future of the Ravenscar profile session summary. ACM SIGADA Ada Letters, 21(1): 5–8, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2003:JAR

[Wel03] Andy Wellings. Is Java augmented with the RTSJ a better real-time systems implementation technology than Ada 95? ACM SIG-ADA Ada Letters, 23(4):16– 21, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Westley:1997:TTA

[Wes97a] Terry J. Westley. TASH: Tcl Ada SHell, an Ada/ Tcl binding. ACM SIG-ADA Ada Letters, 17(2):82– 91, March/April 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Westly:1997:TTA

[Wes97b]  T. J. Westly.  TASH:

Tcl Ada SHell, an Ada/ Tcl binding. ACM SIG-ADA Ada Letters, 17(2):82– 91, March/April 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wand:1987:FFA

[WFF^{+}87] I. C. Wand, J. R. Firth, C. H. Forsyth, L. Tsao, and K. S. Walker. Facts and figures about the York Ada compiler. ACM SIG-ADA Ada Letters, 7(4):85– 87, July/August 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wagner:2020:ACO

[WG20] Lucas Wagner and Andrew Gacek. Automating certification objectives with SpeAR. ACM SIGADA Ada Letters, 39(1):35–49, January 2020. ISSN 0736-721X.

URL https://dl.acm.org/
doi/abs/10.1145/3379106.
3379111.

Wengelin:1990:AST

[WGA90a] Daniel Wengelin, Mats Carlsson Goethe, and Lars Asplund. Anonymous (special topic). ACM SIG-ADA Ada Letters, 10(1):97– 99, January/February 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wengelin:1990:ANT

[WGA90b] Daniel Wengelin, Mats Carlsson Göthe, and Lars Asplund. Anonymous (no title) [A portable Ada solution to the problem of suspending a caller on one node during a call to a remote node]. ACM SIGADA Ada Letters, 10(1):97–99, January/February 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wang:2017:NDU

[WGC17] Bo Wang, Hongbiao Gao, and Jingde Cheng. A new definition-use net generator for Ada 2012 programs. ACM SIGADA Ada Letters, 37(1):

9–25, June 2017. CODEN AALEE5. ISSN 0736-721X.

Whalen:2013:SFA

[Wha13] Michael W. Whalen. Up and out: scaling formal analysis using model-based development and architecture modeling. ACM SIG-ADA Ada Letters, 33(3):41– 42, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wheeler:1984:CIA

[Whe84] Thomas J. Wheeler. A command interpreter for Ada. ACM SIGADA Ada Letters, 3(4):51–61, January/ February 1984. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wheeler:1986:EDD

[Whe86] Thomas J. Wheeler. An example of the developer’s documentation for an embedded computer system written in Ada, Part 1. ACM SIGADA Ada Letters, 6(6):61–71, November/December 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wheeler:1987:EDD

[Whe87] Thomas J. Wheeler. An example of the developer’s documentation for an embedded computer system written in Ada, Part 2. ACM SIGADA Ada Letters, 7(1):40–48, January/February 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wheeler:1995:LAT

[Whe95] David A. Wheeler. Lovelace:

an Ada 95 tutorial. ACM SIGADA Ada Letters, 15(6): 57–66, November/December 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wheeler:1997:ACC

[Whe97] David A. Wheeler. Ada, C, C++, and Java vs. The Steelman. ACM SIGADA Ada Letters, 17(4):88–112, July/August 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wheeler:2019:ACR

[Whe19] David A. Wheeler. Approaches to cyber-resilience through language system design. ACM SIGADA Ada Letters, 38(2):43–57, December 2019. ISSN 0736-721X.

URL https://dl.acm.org/
doi/abs/10.1145/3375408.
3375411.

Whitaker:1981:FLF

[Whi81] Lt Col William A. Whitaker.

FORTRAN-like formatted output with Ada. ACM SIG-ADA Ada Letters, 1(1):26– 28, July/August 1981. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Whitehill:1982:AVO

[Whi82] S. B. Whitehill. An Ada virtual operating system. In ACM [ACM82], pages 238– 250. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

White:1985:ETS

[Whi85]  John R. White.  Ex-

tended terms for SIG officers. ACM SIGADA Ada Letters, 5(3–6):6–10, November/December 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Whitaker:1995:ADH

[Whi95]  William Whitaker.  Ac-

tivities of the DoD High Order Language Working Group. ACM SIGADA Ada Letters, 15(1):28–38, January/February 1995. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

White:1997:PIS

[Whi97] J. B. White. Performance issues of scientific programming in Ada 95. In ACM [ACM97], pages 279–296. ISBN 0-89791-981-5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

White:2010:PAR

[Whi10] Rod White. Providing additional real-time capability and flexibility for Ada 2005. ACM SIGADA Ada Letters, 30(1):135–146, April 2010. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Woodside:1991:CPA

[WHNB91] C. M. Woodside, E. M. Hagos, E. Neron, and R. J. A. Buhr. The CAEDE performance analysis tool. ACM SIGADA Ada Letters, 11(3): 125–136, Spring 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wichmann:1982:TMR

[Wic82] Brian A. Wichmann. Tutorial material on the real datatypes in Ada. ACM SIG-ADA Ada Letters, 1(2):15– 33, September 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wichmann:1986:AFA

[Wic86] B. A. Wichmann. Ackermann’s function in Ada. ACM SIGADA Ada Letters, 6(3):65–70, May/June 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wichmann:1993:BS

[Wic93]  B. A. Wichmann.  Are

Booleans safe? ACM SIG-ADA Ada Letters, 13(3):88– 90, May/June 1993. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wichmann:1998:GUA

[Wic98] B. A. Wichmann. Guidance for the use of the Ada programming language in high integrity systems. ACM SIGADA Ada Letters, 18(4): 47–94, July 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wilder:1983:MHK

[Wil83] William L. Wilder. Minimal host for the KAPSE. ACM SIGADA Ada Letters, 3(2):77–88, September/October 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wilder:1985:KIS

[Wil85] William L. Wilder. KAPSE implementation strategies. ACM SIGADA Ada Letters, 5(1):61–70, July/August 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Williams:1987:URR

[Wil87] Charles Williams. Use of the rational R1000 Ada development environment for an IBM based command and control system. In ACM [ACM87a], pages 45–55. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

Will:1991:SPE

[Wil91] C. A. Will. Software patents and economic competitiveness. In ACM [ACM91b], pages 136–140. ISBN 0-89791-393-0. LCCN ????

Winkler:1984:MBS

[Win84] J. F. H. Winkler. More on block structure: Using Ada. ACM SIGADA Ada Letters, 3(6):48–56, May/June 1984. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Winkler:1990:DLC

[Win90] Juergen F. H. Winkler. A definition of lines of code for Ada. ACM SIGADA Ada Letters, 10(2):89–94, March/April 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Winter:1991:FPA

[Win91] Dik T. Winter. Floating point attributes in Ada. ACM SIGADA Ada Letters, 11(7): 244–273, Fall 1991. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wing:2013:FMI

[Win13] Jeannette M. Wing. Formal methods: an industrial perspective. ACM SIG-ADA Ada Letters, 33(3):85– 86, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wisniewski:1999:TAA

[Wis99] Joseph R. Wisniewski. Transitioning an ASIS application: version 1 to Ada95 2.0. ACM SIGADA Ada Letters, 19(3):53–65, September 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2001:EPT

[WJS^{+}01] A. J. Wellings, B. Johnson, B. Sanden, J. Kienzle, T. Wolf, and S. Michell. Extensible protected types: proposal status. ACM SIG-ADA Ada Letters, 21(1):105– 110, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2002:IOO

[WJS^{+}02] A. J. Wellings, B. Johnson, B. Sanden, J. Kienzle, T. Wolf, and S. Michell. Integrating object-oriented programming and protected objects in Ada 95. ACM SIGADA Ada Letters, 22(2): 11–44, June 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:1984:PAR

[WKT84] A. J. Wellings, D. Keeffe, and G. M. Tomlinson. A problem with Ada and resource allocation. ACM SIGADA Ada Letters, 3(4):112–124, January/February 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wong:1998:KAU

[WL98] Sy Wong and Gertrude Levine. Kernel Ada to unify hardware and software design. ACM SIGADA Ada Letters, 18(6):28–38, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2010:ACN

[WMAB10] A. J. Wellings, A. H. Malik, N. C. Audsley, and A. Burns. Ada and cc-NUMA architectures what can be achieved with Ada 2005? ACM SIGADA Ada Letters, 30(1): 125–134, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wong:2010:NMP

[WMM10] Luke Wong, Stephen Michell, and Brad Moore. Named memory pool for Ada. ACM SIGADA Ada Letters, 30(1):

55–61, April 2010. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wolverton:1984:PHF

[Wol84] David Alan Wolverton. A perfect hash function for Ada reserved words. ACM SIG-ADA Ada Letters, 4(1):40– 44, July/August 1984. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wolfe:1985:AIC

[Wol85]  J. Wolfe.  Artificial in-

telligence and the CAIS. ACM SIGADA Ada Letters, 5(3–6):76–83, November/December 1985. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wolf:1997:FTD

[Wol97] Thomas Wolf. Fault tolerance in distributed Ada 95. ACM SIGADA Ada Letters, 17(5):106–110, September/October 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wolf:1999:TRF

[Wol99] Thomas Wolf. Transparent replication for fault tolerance in distributed Ada 95. ACM SIGADA Ada Letters, 19(2):33–40, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wolf:2001:EFC

[Wol01] Thomas Wolf. On exceptions as first-class objects in Ada 95. ACM SIG-ADA Ada Letters, 21(3):35– 40, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wong:1990:CAC

[Won90] Sy Wong. Considerations of Ada in Chinese. ACM SIG-ADA Ada Letters, 10(2):84– 88, March/April 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wong:1999:ATL

[Won99] Sy Wong. Ada as a teaching language. ACM SIG-ADA Ada Letters, 19(4):22– 23, December 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Woodger:1987:OAF

[Woo87] Michael Woodger. Origins of Ada features. ACM SIGADA Ada Letters, 7(1):59–70, January/February 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wood:1988:ACAa

[Woo88a] D. Wood. The algorithm capture approach to Ada transition. ACM SIGADA Ada Letters, 8(1):80–90, January/February 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wood:1988:ACAb

[Woo88b] David P. Wood. The algorithm capture approach to Ada transition. ACM SIG-ADA Ada Letters, 8(2):96– 106, March/April 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wood:1999:ACF

[Woo99] Dave Wood. Ada: a commercial flop and proud of it! -or-how to deal with Java envy. ACM SIGADA Ada Letters, 19(4):32–36, December 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Workman:1997:UGA

[Wor97] David A. Workman. Understanding generics in Ada95. ACM SIGADA Ada Letters, 17(6):78–90, November/December 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2013:SSM

[WP13] Andy Wellings and Luís Miguel Pinho. Session summary: multiprocessor issues, part 2 (resource control protocols). ACM SIGADA Ada Letters, 33(1):138–145, April 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wetherell:1983:ALT

[WQ83] Charles Wetherell and M. E.

Quinn. An Ada language type checking problem and two morals. ACM SIG-ADA Ada Letters, 3(1):55– 56, July/August 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2015:SS

[WR15] Andy Wellings and Jorge Real. Session summary. ACM SIGADA Ada Letters, 35(1): 102–104, April 2015. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wrege:1992:PKA

[Wre92] D. E. Wrege. Protected kernels and Ada 9X real-time facilities. ACM SIGADA Ada Letters, 12(6):81–87, November/December 1992. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ward:2013:AIC

[WRL13] Donald T. Ward, David A.

Redman, and Bruce A. Lewis. An approach to integration of complex systems: the SAVI virtual integration process. ACM SIGADA Ada Letters, 33(3):43–46, December 2013.

CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wood:1988:IFS

[WT88] David P. Wood and David Turcaso. Implementing a faster string search algorithm in Ada. ACM SIG-ADA Ada Letters, 8(3):87– 97, May/June 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wood:1989:IFS

[WT89] P. Wood and D. Turcaso. Implementing a faster string search algorithm in Ada. ACM SIGADA Ada Letters, 8(3):87–97, May/June 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2003:SSI

[WT03] Andy Wellings and Joyce L. Tokar. Session summary: integration versus orthogonality (RTSJ scheduling policies versus Ada’s). ACM SIG-ADA Ada Letters, 23(4):13– 15, December 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Woodruff:1998:LDC

[WV98] John P. Woodruff and Paul J. Van Arsdall. A large distributed control system using Ada in fusion research.

ACM SIGADA Ada Letters, 18(6):121–131, November/December 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wolf:2001:OOE

[WV01] Thomas Wolf and Tullio Vardanega. Object orientation and exception handling for Ada. ACM SIG-ADA Ada Letters, 21(3):11– 12, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2002:RSL

[WV02] Andy Wellings and Tullio Vardanega. Report of session: language changes for scheduling, modeling and analysis. ACM SIG-ADA Ada Letters, 22(4):125– 127, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

White:2001:DAL

[WW01] Laura J. White and Norman Wilde. Dynamic analysis for locating product features in Ada code. ACM SIG-ADA Ada Letters, 21(4):99– 106, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Walker:1999:ETE

[WWB99] W. M. Walker, P. T. Woolley, and A. Burns. An experimental testbed for embedded real time Ada 95. ACM SIGADA Ada Letters, 19(2):84–89, June 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Xu:2004:MCP

[XCZ04] Baowen Xu, Zhenqiang Chen, and Jianjun Zhao. Measuring cohesion of packages in Ada95. ACM SIG-ADA Ada Letters, 24(1): 62–67, March 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Xing:1988:IAP

[XRL^{+}88] Guo-Guang Xing, Hui Rao, Bin Liu, Jun Shen, and Ming-Yuan Zhu. An integrated Ada programming environment: AWA. ACM SIGADA Ada Letters, 8(6):82–91, November/December 1988. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Xu:2023:UGC

[XWZ^{+}23] Xiong Xu, Shuling Wang, Bohua Zhan, Xiangyu Jin, Naijun Zhan, and Jean-Pierre Talpin. Unified graphical co-modelling, analysis and verification of cyberphysical systems by combining AADL and Simulink/ Stateflow. ACM SIG-ADA Ada Letters, 43(1): 46–49, June 2023. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3631483.3631487.

Xianzhong:2002:EBI

[XZ02] Liang Xianzhong and Wang Zhenyu. Event-based implicit invocation decentralized in Ada. ACM SIG-ADA Ada Letters, 22(1): 11–16, March 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Yavne:1985:SAR

[Yav85] Nancy Linden Yavne. A simple approach to a relaxed syntax for an Ada PDL. ACM SIGADA Ada Letters, 5(1): 71–78, July/August 1985. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Yehudai:1982:DAT

[Yeh82] Amiram Yehudai. Data abstraction: Types vs. objects. ACM SIGADA Ada Letters, 2(2):46–48, September/October 1982. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Yemini:1982:SAM

[Yem82] S. Yemini. On the suitability of Ada multitasking for expressing parallel algorithms. In ACM [ACM82], pages 91–97. ISBN 0-89791-087-7. LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

Young:1980:GVA

[YG80] William D. Young and Donald I. Good. Generics and verification in Ada. In ACM [ACM80], pages 123–

127. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

Yang:2023:CMD

[YQZ^{+}23] Zhibin Yang, Zhikai Qiu, Yong Zhou, Zhiqiu Huang, Jean-Paul Bodeveix, and Mamoun Filali. C2AADL Reverse: a model-driven reverse engineering approach for development and verification of safety-critical software. ACM SIGADA Ada Letters, 43 (1):50–53, June 2023. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3631483.3631488.

Yu:1997:UOT

[Yu97]  H. Yu.  Using object-

oriented techniques to develop reusable components. In ACM [ACM97], pages 117–124. ISBN 0-89791-981-

5. LCCN ???? Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Yu:1998:CSR

[Yu98]  Huiming Yu.  A course

in software reuse with Ada 95. ACM SIGADA Ada Letters, 18(1):48–53, January/February 1998. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Zerzelidis:2007:CEP

[ZBW07] A. Zerzelidis, A. Burns, and A. J. Wellings. Correcting the EDF protocol in Ada 2005. ACM SIGADA Ada Letters, 27(2):18–22, August 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Zamorano:2002:PRT

[ZdlP02] Juan Zamorano and Juan Antonio de la Puente. Precise response time analysis for Ravenscar kernels. ACM SIGADA Ada Letters, 22(4): 53–57, December 2002. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Zamorano:2013:RTP

[ZdlP13] Juan Zamorano and Juan A.

de la Puente. On realtime partitioned multicore systems. ACM SIGADA Ada Letters, 33(2):33–39, August 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Zou:2022:RAM

[ZDM22] Jie Zou, Xiaotian Dai, and John A. McDermid. Resilience-aware mixed-criticality DAG scheduling on multicores for autonomous systems. ACM SIGADA Ada Letters, 42(1):81–85, June 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL

https://dl.acm.org/doi/
10.1145/3577949.3577965.

Zamorano:2013:ART

[ZEdlP13] Juan Zamorano, ’Angel Esquinas, and Juan A. de la Puente. Ada real-time services and virtualization. ACM SIGADA Ada Letters, 33(1):128–133, April 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Zalila:2006:IIC

[ZHP06] Bechir Zalila, Jérôme Hugues, and Laurent Pautet. An improved IDL compiler for optimizing CORBA applications. ACM SIGADA Ada Letters, 26(3):21–28, December 2006. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Zhu:1990:DTF

[Zhu90] Ming-Yuan Zhu. Design of a text formatter with AUTO STAR. ACM SIGADA Ada Letters, 10(1):140–159, January/February 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Zeigler:1983:ALS

[ZW83] Stephen F. Zeigler and Reinhold P. Weiker. Ada language statistics for the iMAX 432 operating system. ACM SIGADA Ada Letters, 2(6): 63–67, May/June 1983. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).